A Complete Bibliography of Publications in Marine Biology Research

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

21 August 2024 Version 1.16

Title word cross-reference

3 [149, 437]. ^{+} [1211]. ^{13} [53]. _{2} [679, 1344, 353, 1319, 795, 80, 1143]. _{3} [1344]. a [918, 112, 283, 402]. : [74]. D [483, 653]. F/R [898].

-loop [483, 653]. -NPs [1344]. -ratio [898].

/K [1211].

1 [540]. 1-year [322]. 16S [719]. 18 [1359]. 18S [657].

2 [541]. 2011 [391]. 2014 [694]. 20^{o}N [400]. 23^{o}S [193]. 25-year [615]. 25th [946].

3 [1077]. 30-year [1041]. 36^{o}N [23]. 36^{o}S [877].

41^{o}S [877]. 42^{o} [753]. 42^{o}S [1191]. 43^{o}S [1035]. 45^{o}S [753]. 47^{o} [632].

51^{o}N [632]. 51^{o}S [1035].

6^{o}S [400].

70 [452]. 71^{o}N [86].

9th [317].

=Caudofoveata [69].

A. [447, 622, 1160]. abalone [124, 456]. abbreviata [401]. abdominal [251]. abei [406]. ability [565, 410, 898]. abiotic [1178, 1067, 340]. Ablennes [1179]. abnormality [1001]. above [582]. abrupt [787]. absence [463, 457]. Abudefduf [1297]. Abundance [783, 717, 1349, 205, 892, 256, 1320, 648, 1180, 649, 388, 63, 590, 565, 886, 222, 980, 66, 126, 1150, 640, 1371, 413, 700, 1060, 788, 379, 801, 40, 1152, 1086, 31, 628, 1058, 836, 1287, 1281]. abundance-occupancy [1281]. abundances [934]. Abyssal [242, 1373, 843]. acanthias [511]. Acanthina [421]. Acari [526, 829]. Acartia [888, 565]. accelerometers [1068]. acclimated [787]. according [905, 115]. account [854]. accounts [885]. Accumulation [282, 1182]. acherusicum [835]. acid [1344, 1356, 1250, 463, 630, 252, 570, 658, 1040]. Acidification [338, 1090, 722, 285, 1249, 883]. acids [1000, 832, 1143]. acknowledgement [169, 232, 293, 370, 490, 586, 683, 805, 896, 948]. acoustic [328, 736]. acquisition [1113, 1111]. Acropora [1320, 1194, 509, 913, 629]. across [1338, 855, 1164, 60, 19, 184]. act [335]. Actiniaria [227, 407, 287, 126, 30, 1153, 464, 791]. actinomycetes [744]. Actinopterygii [1346, 802, 1109]. action [920]. active [1110]. activities [672, 1211, 744, 902]. activity [399, 53, 1244, 679, 1366, 1233, 761, 1239, 111, 43, 494, 160, 238]. aculeata [891]. acute [459]. acutorostrata [832, 50, 623, 600]. adaptation [458]. adaptive [229]. additional [546, 1029]. additions [1072]. additive [1058]. addressed [1292]. adds [495]. adequately [1292]. Adipicola [1240]. Adjacent [916, 588, 926, 928, 919, 551, 917, 927]. adjustment [530]. adjustments [115]. adopted [1052]. Adour [348]. adrenergic [755]. Adriatic [1251, 737, 518, 644, 511, 27, 150, 485, 661, 484, 521, 768, 215, 1041]. adrift [1051]. adspersus [520]. adult [983, 252]. adults [977]. Advances [1339, 1267]. Advantages [22]. adventitious [398]. adverse [199]. advice [699]. Aega [131]. Aegean [1269, 448, 126, 269, 930, 1105, 1189, 1374, 719, 57]. Aegidae [1087, 131]. Aegina [963]. Aequipecten [164]. aff. [1072]. affect [941, 773]. affected [1244, 556, 1030]. affecting [633]. affects [1009, 1023, 1325, 451]. affiliation [1171]. affinities [588, 770]. Africa [271, 482, 1185, 1275, 454, 407, 1234, 813, 1141, 988, 396, 1026]. African [1093, 855, 1354, 1242, 456]. Afropinnotheres [811]. after [558, 530, 347, 677, 422, 457]. Agardh [1146]. Agassiz [1199]. agassizii [59]. Agder [1165]. Age [512, 894, 1045, 1367, 345, 324, 871, 424, 625, 144, 236, 943, 768, 1189, 448, 1263, 175, 1074, 1148, 692, 485, 164, 174, 356]. age-dependent [164]. ageing [784, 388, 398]. Aggregation [1366, 500, 34]. aggregations [1366, 987, 190]. agility [390]. Aglou [1294]. Aglou-Sidi [1294]. Agrobacterium [240]. Aguado [292, 580]. aibuhitensis [864, 1055]. aided [414]. Aires [911]. Akashiwo [787, 1145]. al [694]. Alabama [1372, 1181]. Alaska [884, 287]. alatus [184]. albiflora [957]. albomaculata [709]. Alboran [550, 298, 1226, 266]. Albunea [145]. alcocki [1357]. Alcyonacea [443, 542]. Alcyoniidae [542]. Aleutian [287]. aleutiensis [287]. alfredi [265]. alga [458, 389, 868, 199]. algae [459, 1342, 416, 826, 857]. algal [316, 1070, 1024, 582, 1247, 827, 1145]. Algarve [236, 1277, 145]. Alicia [126]. Alien [1063, 88, 480, 1325, 889, 517, 153, 1374]. Alismatales [655]. alkalinized [223]. Alle [723]. allelopathic [1145]. allelopathy [800]. Allen [884]. Allman [965]. allocation [1111]. Allochthonous [1116]. Allostichaster [815]. Alloteuthis [732]. almansae [589]. Almirante [276]. along [1345, 304, 1180, 1312, 1251, 560, 855, 890, 93, 534, 1310, 1218, 793, 483, 1134, 831, 43, 416, 946, 1217, 1283, 1213, 591, 1370, 20, 957, 796, 1074]. alpha [874, 885]. alpha-taxonomic [885]. alpha-taxonomy [874]. Alpheidae [1066]. Alpheus [1066, 1326]. also [832]. alterations [954]. alternative [842, 360, 12, 1322]. alvarezii [910, 766]. always [711]. amansii [868]. Amazonian [1206]. ambereye [604]. amberjack [1068]. ambient [279]. Ambilimbus [23]. Ambush [390]. ameliae [210]. amendment [1276]. America [1214, 91, 374, 237, 552, 975, 412, 68, 110]. American [1266, 156, 496, 495, 554, 455, 527, 1219, 497, 498]. americanus [496, 1221]. Amiantis [510]. Ammodytes [1137]. Ammodytidae [1137]. Ammothea [1077, 122]. Ammotheidae [1077, 122]. amnesic [463]. among [1079, 1328, 1230, 460, 1376, 709, 1253, 271, 1043]. Amphan [1355]. Amphibalanus [1063, 660]. Amphicrossus [23]. Amphilepidida [1098]. amphipod [995, 319, 812, 1139, 309, 120]. Amphipoda [545, 688, 1333, 303, 1260, 1358, 51, 260, 760, 1184, 517, 911, 997]. amphipods [73, 835]. Amphiprion [631, 1153]. Amphisbetia [716]. Amphithallia [1097]. amphitrite [1063, 660]. amplicon [810]. amplification [1256]. Amplified [97, 84]. anadromous [615, 616, 613]. anaerobic [37]. anal [487]. analyses [483, 255, 579, 1365, 847]. Analysis [1136, 1028, 78, 1265, 852, 209, 458, 880, 1279, 62, 810, 760, 452, 418, 621, 1234, 467, 653, 822, 249, 831, 1153, 976, 1060, 665, 940, 1179, 525, 632, 849, 647, 702, 1040]. analytical [106]. Anarhichas [247]. Anasterias [105]. Anatoly [212]. anchialine [89, 27, 150]. anchovy [946]. and/or [450]. Anda [471]. Andaman [1173, 1154, 1250, 142, 1258]. andersoni [1296]. Andres [717]. Andrey [138]. Andriashev [212]. andromeda [1284]. anemone [287, 30, 464, 791]. anemonefish [1153]. anemones [1153].

Angiospermae [2]. anglerfish [506]. anglerfishes [1163]. Angola [102]. Anguilla [348]. Anguillid [939]. Anguilliformes [241, 1031]. anguina [1128]. Animal [640, 654]. Animal-sediment [640]. Anisomysis [67]. Anker [1066]. annandalei [1350]. annelid [119, 181]. Annelida [117, 26, 1206, 1079, 992, 1162, 771, 1154, 1002, 889, 1114, 1311, 186, 552, 976, 1037, 83, 668, 243, 292, 675, 24, 553, 81]. Announcement [627]. Annual [1338, 1330, 214, 1191, 1054, 1050, 174]. annulata [254]. annulus [324]. Anomalous [636]. Anomura [1084, 935, 145, 1062]. ANT [1077]. ANT-XXIX [1077]. ANT-XXIX/3 [1077]. Antarctic [419, 479, 1077, 1375, 1340, 561]. Antarctica [122]. ante [251]. ante-mortem [251]. Anthiines [585]. Anthoathecata [376, 1006]. Anthomedusae [76]. Anthoptilum [1315]. Anthozoa [287, 30, 1194, 691, 1260, 407, 423, 126, 133, 1153, 1252, 227]. anthropogenic [397, 1284, 1266, 1233, 1246]. anti [111]. anti-microbial [111]. antifouling [680]. antillarum [801, 1047]. antimicrobial [1246]. antimicrobial-resistant [1246]. antioxidant [1232, 1211, 334, 910, 335]. Antonio [1072]. Antygomonas [759]. apatite [495]. aper [1367, 1189]. Aphrocallistidae [541]. Apicomplexa [59]. Aplysina [62]. Apodida [1161]. Apostichopus [461]. apparatus [910]. apparent [820]. appearance [1049]. appendages [972, 49]. applicability [562]. application [1230, 343]. Applications [710, 393]. applied [295]. Appraisal [398]. appressum [1225]. Approach [137, 1356, 51, 953, 516, 664, 12, 363, 874, 1252, 355]. Approaches [415, 1365, 745]. Apristurus [559]. Apseudidae [1159]. Aqaba [558, 304, 961]. aquaculture [71, 949, 530, 336, 422, 614, 901]. Aquatic [200, 616]. aquiferous [1138]. Arabian [1335, 1308, 1288, 1204, 757, 1195, 402, 714, 1120, 403, 1217]. arabicum [1308]. Araeolaimida [667]. Arai [939]. Arbacia [1158]. Arc [152]. Archaeal [1117, 1060]. archigregarine [59]. Archipelago [942, 1201, 1154, 1152, 166, 1101, 991, 992, 709, 907]. Arcoida [32]. Arctic [952, 575, 10, 860, 141, 112, 172, 1375, 29, 488, 1029, 515, 573, 576, 101, 723, 283, 3, 202, 224, 297, 116, 728, 570, 76, 628, 254, 176, 356, 1036, 175]. Arctic-Atlantic [3]. Arctic-boreal [202]. Arctica [351, 391]. Arcticidae [351]. arcuata [782]. Area [537, 859, 869, 157, 650, 1209, 1034, 538, 324, 890, 610, 1295, 220, 640, 322, 832, 1001, 507, 547, 669, 634, 57, 991, 1163]. areas [689, 1312, 44, 357, 246, 949, 1362, 843, 471, 1100, 899]. arenaria [1023]. argentatus [488]. argentea [149]. argenteus [302]. Argentina [512, 1072, 867, 895, 911, 401, 332, 583, 909, 590, 1166, 1274, 777, 971]. Argentine [1191, 357, 531, 529, 1074]. Argentinian [751, 346]. Argentiniformes [1202, 151]. Argobuccinum [466]. argus [501]. Argyrosomus [730]. ariakensis [377]. Ariidae [1190]. armoured [656]. arrowhead [474]. arsenic [371]. artefact [106]. Artemia [1356]. arthritica [144, 195]. Arthropoda [122, 1134, 690]. Arthrotardigrada [1019, 942, 776]. arthrotardigrades [708]. articulation [1172]. Artificial [223, 558, 1079, 289, 1358, 155, 1168, 239, 1210]. Arve [179]. ascidian [334]. ascidians [1324]. Asexual [815, 378]. Asian [778, 622, 1100]. asiatica [1096]. Aspects [4, 194, 275, 687, 1112, 1340]. Aspidiophorus [198]. assay [822]. assemblage [479, 908, 1059, 551, 341]. assemblages [1101, 765, 731, 350, 995, 1011, 872, 926, 1009, 739, 64, 2, 373, 101, 556, 451, 924, 858, 470, 181, 857, 669, 632, 298, 550, 55]. assess [1209]. assessed [454, 810, 949, 621]. Assessing [785, 847, 557, 845]. Assessment [304, 1123, 923, 593, 902, 126, 953, 824, 643, 34, 1226, 160, 1210, 1257]. assessments [1132]. assimilation [925]. assistance [807]. associated [952, 1345, 352, 1288, 5, 765, 751, 1011, 333, 1373, 1201, 458, 565, 319, 72, 385, 1331, 73, 837, 824, 1305, 857, 1237, 550, 719, 1200, 1143]. Association [911, 1296, 998, 674, 1139, 544, 211]. associations [912, 961, 728]. Assortative [1102]. Asterias [1304]. asterocherid [90]. Asterocheridae [90]. Asteroid [134, 546]. Asteroidea [357, 105, 595]. Astrocaneum [934]. Astrodictyum [934]. Astroides [275, 323]. Astrophorida [276]. asymmetrical [902]. asymmetry [634]. asynchronous [473]. Ateret [694]. Athelges [712]. Atlantic [26, 90, 973, 1075, 140, 564, 912, 1315, 1342, 867, 618, 895, 1067, 542, 468, 135, 244, 1092, 753, 911, 1252, 529, 291, 1294, 231, 1049, 75, 689, 138, 194, 1020, 991, 888, 137, 1183, 1312, 667, 131, 506, 157, 242, 267, 1178, 838, 1198, 771, 134, 546, 886, 1176, 1229, 1076, 328, 692, 615, 127, 130, 537, 538, 1106, 933, 539, 871, 460, 1147, 1158, 1316, 545, 980, 408, 66, 23, 37, 813, 900, 299, 722, 188, 1224, 543, 431, 935, 1133, 104, 540, 1134, 640, 413, 1062, 962, 528, 924, 1130, 132, 284]. Atlantic [422, 128, 133, 865, 1309, 1039, 3, 225, 1013, 6, 433, 993, 709, 548, 927, 1071, 379, 547, 943, 835, 1168, 772, 1152, 69, 544, 833, 599, 616, 617, 252, 570, 129, 541, 591, 1140, 211, 931, 31, 870, 877, 235, 997, 612, 1370, 1095, 203, 721, 136, 1074, 1257, 641]. atlantica [379]. atlanticus [131, 266]. Atlantorchestoidea [1132]. atolls [1193]. ATP [143]. ATPase [1211, 359, 789]. atrovitta [1154]. attaching [1188]. attack [495, 164]. attacking [1290]. attracted [271, 482]. attributed [1071]. attributes [1267]. atun [1234]. Audouin [1157]. auk [723, 628]. auks [183]. Aulopiformes [326]. Aurelia [912, 209, 1053]. aurita [912, 209, 168]. Aust [1165]. Aust-Agder [1165]. austral [975]. Australasian [897]. Australia [1275, 648, 597, 856, 154, 384, 503, 19, 273, 829, 738]. Australian [226, 39, 695, 553, 362]. australis [357]. Austrodecidae [1172]. Authors [394]. auto [1122]. auto-correlated [1122]. autochthonous [1116, 1126, 813]. Autolytinae [771]. autolytines [771]. autotrophic [52, 1197]. autumn [445, 33, 1028]. Auxis [483]. availability [399, 1244, 228]. avoidance [270]. Axiannassidae [1310]. Axiidea [447, 790, 1085, 1147, 1310]. axillary [768]. Axinellidae [91]. Azadinium [673]. Azam [393]. Azores [194, 140, 543, 118, 1152, 599, 1370].

B [853, 978]. bachei [912]. bachus [446]. back [1036]. bacteria [1291, 565, 1356, 898, 1246]. Bacterial [1144, 912, 1200, 1345, 1288, 1321, 1195]. bacterioplankton [620]. Badijella [27]. Bahamas [89]. Bahía [583, 1072, 501]. bairdi [889, 367]. bait [271, 482]. bait-attracted [271, 482]. Baja [1104, 982]. Bakkhali [341]. Balaenoptera [832, 50, 623, 600]. Balanus [955]. Balearic [992]. Baleen [1233]. Baltic [747, 1095, 776, 78, 417, 305, 1065, 262, 520, 932, 486, 914, 1227, 819, 779]. baltica [389]. bamboo [1343]. Banc [1190]. banding [241]. bandtail [874]. Bangiaceae [84]. Bangladesh [876, 1350, 341]. Bank [31, 182, 361]. Banks [468, 355]. banlaensis [817]. barbatus [603, 602, 1346]. barcode [1142]. Barcoding [1123, 562, 726, 824, 9, 874, 1043]. Barents [1010, 574, 1220, 178, 177, 674, 325, 64, 173, 699, 567, 568, 1000, 573, 572, 575, 171, 577, 571, 983, 50, 569]. barnacle [1063, 680, 1113, 660, 1354, 955, 1169, 645]. barnacles [718]. Barnard [1205]. barreleyes [1202]. barrier [365]. basal [925, 149]. based [876, 1279, 864, 1344, 692, 1329, 436, 84, 530, 1062, 919, 422, 750, 579, 270, 736, 301, 940, 1179, 1276, 647, 874, 438, 497, 355, 719, 1263, 702]. baseline [1239]. Basin [667, 1025, 326, 154, 618, 646, 516, 102, 161, 104, 30, 669]. Basin-scale [646]. basins [1198, 310, 890, 10, 887]. basket [934]. bass [1109, 410]. bathyal [539, 3, 31, 994]. Bathylagus [151]. Bathymetric [312, 506, 793, 1105]. bathymetry [632, 497]. Bathymodiolinae [302, 1240]. bathypelagic [326]. bathyporum [1149]. Bathypterois [326]. Bathyrhombila [311]. Batillipedidae [708, 776]. Batillipes [708, 776]. Batoids [721]. Bay [315, 26, 1332, 70, 480, 1144, 1245, 775, 1238, 1372, 747, 384, 700, 707, 518, 1091, 1058, 1282, 373, 945, 1013, 1004, 36, 1111, 968, 1296, 1084, 1195, 1197, 714, 1156, 668, 411, 960, 205, 1300, 779, 620]. Bayesian [516]. bays [1237]. BC [746]. be [54, 22, 1219, 870]. beach [1244, 1116, 874]. beaches [1132, 340]. Beagle [419, 909]. beaked [177, 674]. beam [1306]. bearing [637]. bed [1304, 181, 857]. beds [1049, 158, 765, 707]. behave [1351]. behavior [1304, 143]. Behavioral [475]. behaviour [762, 326, 49, 42, 466, 1233, 893, 65, 766, 313, 405, 482, 1028]. Behavioural [1208, 791]. behaviours [1273]. belcheri [381]. Belize [374]. belli [1130]. belly [194, 267, 268]. belonging [671]. Belonidae [1348]. below [582]. Belt [1317]. beneath [685]. benefit [500]. Bengal [1321, 1296, 1195, 1197, 714, 1300]. Benguela [114, 730]. Benign [756]. Benthic [138, 127, 642, 1285, 543, 297, 1323, 114, 1180, 583, 137, 686, 53, 480, 185, 588, 880, 1306, 739, 539, 1325, 16, 452, 980, 1115, 747, 384, 1133, 548, 228, 139, 887, 570, 497]. Benthoctopus [75]. Benthophilinae [314]. Benthos [537, 538, 515]. Benthosema [230]. benzo [167]. Bergen [1142]. Bering [999, 375]. Bermuda [81]. Bernard [1017, 1018]. Bertrum [989, 990]. beta [1282]. better [1209, 1339]. between [1104, 1196, 428, 574, 972, 1262, 872, 812, 105, 1282, 890, 1234, 429, 509, 543, 913, 867, 1016, 551, 1153, 7, 633, 313, 450, 50, 605, 606, 513, 911, 685, 398, 1145, 498, 1151, 219]. Beyond [954, 339]. Bibliography [107, 212]. bicornutus [369]. bidentatus [693]. big [1291]. Bight [155, 1233]. bilayer [638]. billfish [984]. bio [1334, 980]. bio-collectors [980]. bio-optical [1334]. bioaccumulation [866, 986]. bioassay [425]. biochemical [825, 251, 1250, 336, 1221, 124, 1303, 282, 1249, 238]. biochemistry [40]. Biodiversity [233, 488, 769, 1192, 857, 887, 779, 1260, 597, 162]. Bioengineering [1325]. biofilm [1187]. biofouler [974]. biogeochemistry [330, 554]. biogeographic [590, 844, 1362, 963]. biogeographical [1340]. Biogeography [14, 887]. BIOICE [442]. Biological [194, 1112, 1168, 744, 970, 72, 9, 1269]. Biology [1, 294, 317, 804, 491, 596, 1359, 272, 82, 100, 201, 383, 694, 213, 514, 153, 939, 745, 685, 885, 595, 237, 1332, 492, 564, 666, 426, 295, 535, 447, 306, 687, 214, 1084, 1225, 1167, 413, 259, 351, 1039, 4, 391, 40, 1222, 772, 9, 409, 807, 1374]. Bioluminescence [1248]. biomarkers [220, 221, 335]. biomass [1335, 357, 205, 1165, 736]. Biometric [1221]. biometry [323]. biomineralization [1293]. biosorbent [1337]. Biosphere [387, 501]. Biota [916, 917, 883]. biotic [1067, 969]. biotopes [611]. biotransformation [335]. bipolarity [61]. Birds [926]. Biscay [26, 1370, 668]. Bispira [987]. bivalve [479, 603, 644, 1056, 1290, 633, 449]. bivalves [36]. Bivalvia [1218, 775, 192, 666, 785, 1325, 336, 32, 351, 510, 302, 518, 1283, 1240, 940, 1080, 184, 561]. Bizerta [290]. Bizerte [1346]. Black [869, 1099, 1180, 1352, 663, 1019, 1268, 1128, 1334, 682, 1106, 117, 165, 1307, 106, 1287, 1353]. blackbelly [1255]. Blainville [1030]. Blanca [583]. bleached [913]. bleaching [1341, 1194]. Bleeker [653, 342]. Bleik [788]. Bloch [1106, 1213]. blood [705]. bloom [1196, 112, 42, 58, 52, 402, 505, 1026]. bloom-forming [42, 58]. Blooms [682, 966, 920, 1217, 1247]. blubber [832, 252]. Blue [152, 86, 1057, 1316, 434, 1033, 431, 433, 420, 791, 380]. Bluefin [1131]. bluespotted [508]. Board [382, 684, 806, 915]. boarfish [1367, 1189]. bobtail [388, 564, 379, 40]. bocourti [447]. body [1176, 496, 1339]. Boeck [1333, 1115]. boggessi [219]. Bolinao [471]. Bolinao-Anda [471]. Bolinus [1290]. Bonattere [1361]. bone [467]. bonito [1106]. bonnellii [329]. Book [146, 989, 990, 56, 1017, 1018, 48, 38, 654, 853, 191, 884, 47, 46, 585, 1187]. bopyrid [1222]. Bopyridae [712, 291]. boreal [666, 202]. borealis [956, 1285, 7, 6, 307]. Boreogadus [573]. boring [938, 1080]. Bosphorus [365]. Botany [384]. Bothidae [280, 118]. Bothus [280, 118]. Bott [1085]. bottleneck [849]. bottlenose [700, 927, 1122]. bottom [1010, 242, 1025, 798, 10, 1069, 470, 724, 166]. bottoms [895, 1192, 298, 550]. Bougainvillia [895]. boulder [1070]. boulder-reefs [1070]. boundaries [729, 186]. boundary [590, 890, 3]. Bovallius [1139]. box [641]. boxcore [31]. brachiata [282]. brachiopods [1061, 1183, 136]. Brachyura [691, 936, 163, 622, 662, 1120, 1160, 891, 1242, 693, 933, 1357, 1016, 288, 455, 1365, 665, 752, 870, 641]. brackish [1056, 776]. branches [1097]. branchial [1188, 1258, 1222]. branchial-attaching [1188]. branching [509]. Branchiomma [889, 367]. Branchiostoma [381]. brandaris [1290]. brasileiro [1066]. brasiliensis [1351, 1132, 581, 1257]. Bray [1207]. Brazil [920, 919, 1094, 533, 783, 1206, 977, 968, 667, 156, 784, 1260, 650, 326, 441, 775, 926, 816, 598, 785, 1078, 928, 1059, 1121, 1202, 821, 922, 1069, 639, 1066, 1112, 420, 1169, 1122, 921, 580, 276, 1015]. Brazilian [691, 715, 1310, 312, 1225, 1161, 1121, 1192, 1301, 411, 276, 796, 1111]. break [357, 1035, 946]. breakdown [216]. bream [236, 1300, 850]. breeding [1103, 971]. bregazzii [1340]. brevidactylus [450]. brevirostris [332, 270]. Brevoortia [110]. Brewed [1354]. Brian [391]. brink [728]. Britain [632]. British [505, 1089]. brittle [1236]. brittlestar [1225]. broad [630, 764, 88]. broad-nosed [764]. broad-scale [88]. brocade [715]. Brodski [193]. bronchostylus [472]. Brooding [668, 275]. brown [615, 389, 626, 937, 521, 616, 613, 1073]. brownish [1283]. Bruce [1258]. brunnea [987]. Bryopsidales [1146]. Bryozoa [1086]. bryozoan [844, 782, 883, 254]. bubble [635]. Buccinanops [512, 894, 1045]. budgets [864]. Buenos [911]. Bulaceros [1193]. Buoyancy [530]. Burgess [978]. burial [318]. burrfish [786]. burrowing [1310, 347]. burrows [21]. burying [65]. butyltin [290]. butyltins [866]. by-catch [1027, 742]. bycatch [1277, 470]. bycaught [1155, 623].

C [53, 74]. C. [790, 95, 377]. Ca [260]. cabrilla [871]. cadmium [1284, 334, 371, 220, 221, 986]. caeciliae [759]. Caenogastropoda [777]. caerulea [354]. caespitosa [1041]. Caicos [89]. calanoid [583, 27]. Calanoida [89, 27, 578]. Calanus [141, 1076, 172, 193, 436, 52, 435, 437, 628]. Calcification [1249]. Calcinus [1062]. calcium [353]. Caledonia [907]. California [1104, 934, 470, 1272, 982, 1199, 858, 884, 846, 1270, 364]. californicus [1231]. Callanthias [830]. Callanthiidae [830]. callarias [305]. Callianassidae [1310, 447, 790, 1085, 1147]. Callichiridae [1310, 1222]. Callichirus [790, 1085, 1222]. Callinectes [783, 1316, 420]. Callyspongia [62]. calycularis [275, 323]. Calyptronema [581]. camouflage [820]. Campanulariidae [967]. Campos [667, 326, 104, 618]. camtschaticus [49, 325]. Can [469, 1219, 612, 9, 870]. Canada [505, 722, 1205, 640, 1046, 903]. Canadian [259]. Canal [1275, 472, 1203]. canaliculus [158, 944, 755]. Canaria [1246]. Canary [1246, 90, 1186, 203]. Cancer [933, 937, 752]. cancrid [752]. Cancridae [665, 752]. canicula [269, 930, 484]. canis [851]. cannibalism [372, 176, 1036]. cannonball [1199]. canopy [648, 1338, 41]. canopy-forming [648]. canosai [1224]. Canyon [26, 788]. canyons [597]. Cap [468]. capability [708]. capacity [37]. Capbreton [26]. Cape [91, 709]. capelin [173, 699, 573, 572, 462, 224]. Capitella [315]. Caprella [303, 51, 997]. Caprellidae [1358]. caprellids [260, 760, 517]. Capron [637]. Capros [1367, 1189]. captive [396]. captivity [677]. capture [216, 1364, 530, 422]. capture-based [530, 422]. carabus [145]. Caraguatatuba [775]. Carangidae [1229]. carapace [1187, 1265, 1174]. carapaces [70]. carbon [672, 985, 565, 1136, 864, 386, 467, 451, 1195, 94, 277, 363]. Carbonate [495, 883, 353, 569]. Carcharhiniformes [559]. Carcharhinus [881]. Carcharias [1208, 396]. carcini [1113, 65]. Carcinoscorpius [1243]. Carcinus [1113, 820, 65, 903]. care [1139]. Caretta [1265]. Carey [951].

Cariaco [516]. Caribbean [742, 373, 1253, 656, 501, 717, 243, 770, 1004, 98, 1297]. Caridea [1331, 1066, 248, 1119]. caridean [1192]. Caristiidae [1121]. caristiids [1121]. Carnivorous [618, 302]. carpet [1308]. carpogonial [1097]. carried [1359]. Carupella [817]. Carvalho [1017, 1018]. case [268, 1014, 902, 1136, 1230, 711, 963, 614, 1122, 1322, 254]. Caspian [1093, 1034]. Cassidinidea [1241]. Cassiopea [1284]. Catalan [1008]. Catalogue [721]. Catch [1307, 1027, 1306, 742, 663, 1073, 498]. Catenulida [374]. catfishes [1190]. catshark [930, 484, 559]. caudal [472]. caudatum [673]. caudatus [692]. caudicula [523]. caught [1266, 1361, 371, 724, 1226, 1028, 1151]. Caulerpa [1146, 181]. caused [1003, 347]. cave [1159, 27, 150]. caves [89, 226]. Cd [1182]. cefotaxime [240]. celata [1072]. cell [459]. celled [1097]. cells [1245]. cellular [463, 1318]. cellulase [672]. Celtic [632]. Central [1214, 374, 424, 1290, 203, 808, 671, 1014, 1166, 625, 687, 43, 1349, 313, 484, 521, 1240, 1235, 840, 1011, 552, 669]. centre [659]. Centrolophidae [671, 1302]. Centrophoridae [474]. Centrostephanus [1078, 507]. century [285]. Cephalopod [749, 206]. Cephalopoda [1071, 329, 388, 564, 279, 1167, 704, 994]. Cephalopods [235, 1370, 729]. cephalus [705]. Cerastoderma [220, 221, 347, 322]. Ceratioidei [1163]. Ceratoscopelus [1263]. Cereopsis [443]. Ceriantharia [1260, 1351]. Ceriantheomorphe [1351]. Cerro [859]. certain [132]. Cervimunida [1278]. cetacean [1169]. cetaceans [749, 599]. cf. [1194]. CGFZ [151]. Chaetodermomorpha [69]. chaetognath [714, 1092]. chaetognatha [142]. Chaetonotida [676, 198]. chain [984, 1292]. Chalinula [1082]. challenged [1356]. Challenges [958, 959]. chalumnae [414]. chambers [21]. Change [337, 319, 576, 1124, 1371, 577, 969, 527, 166, 600]. changed [1065]. Changes [574, 167, 162, 868, 916, 1232, 839, 178, 615, 1023, 928, 1158, 642, 609, 572, 787, 101, 171, 1056, 475, 1133, 80, 917, 493, 403, 630, 619, 164, 600, 569, 238, 658]. Changhai [822]. Channel [909, 441, 1021, 386, 3, 551, 419, 20]. Chaotic [1164]. character [1174, 266]. Characteristic [720]. characteristics [648, 1291, 998, 1334, 1126, 1138, 1084, 714, 314, 707, 462, 1247, 405, 752, 1269, 1151, 1323]. Characterization [1272, 761, 217, 241, 1337, 890, 1002, 1218, 1250, 1213, 652, 1125]. characterizations [1362]. Charlie [151, 133, 548, 547]. Charlie-Gibbs [151, 133]. Charybdis [936, 288, 772]. Chatham [1275, 446]. Chattonella [678]. Cheilostomata [1086, 254]. Chek [706]. chelar [496]. Chelidonichthys [1135]. chelipedal [1159]. chemical [1118, 70, 519, 8, 799]. chemistry [1295]. chemosynthetic [464]. chemotaxonomy [919]. Chile [688, 469, 1011, 466, 193, 1234, 206, 43, 659, 1240]. Chilean [320, 840, 908, 1035]. chilensis [909, 193, 908]. Chilomycterus [786]. Chiloscyllium [1308]. Chimaera [13]. Chimaeras [721]. Chimaeridae [13]. China [1061, 1157, 1245, 754, 316, 504, 400, 451, 913, 743, 121, 1263, 702, 825, 852, 744, 1279, 834, 1146, 95, 1096, 727, 957, 1058, 1323]. chinensis [336]. Chinese [882, 84]. Chionoecetes [325, 983]. Chirognathia [68]. Chiton [820, 189, 211]. Chlamys [670]. Chlorella [1303]. Chlorodendrophyceae [1052]. chlorophyll [918, 112, 283, 402]. chlorophyll- [283]. Chlorophyta [947, 1146, 795, 1052, 11]. Choanoflagellates [804]. choanosomal [17]. choice [499]. Chondracanthus [506]. chondrichthyans [877]. Chondrichthyes [332, 424, 626]. Chondrus [95]. Chorao [180]. Chordata [1324, 1170]. Christoffer [440]. chromodorid [619]. Chromodorididae [619]. chromosomal [241]. Chromosome [209, 1031]. chronic [459]. Chrysaora [659]. Chrysiptera [218, 399]. Chrysopetalidae [553]. Chthamalidae [1354]. Chthamaloidea [1354]. Chthamalus [982, 1354]. Chupa [601]. Cienfuegos [373]. Ciliary [229]. Ciliary-propelling [229]. ciliate [42, 58, 229, 646]. ciliates [535]. cinerascens [1182]. cinnamon [631]. circa [1014]. circa-littoral [1014]. circulation [193]. Cirolanidae [1259]. Cirrhipathes [1128]. Cirripedia [1220, 936, 1113, 645]. citizen [1071, 1170]. Cittarium [905]. civilization [70]. Cladocora [1041]. Cladorhizidae [618]. Claiborne [853]. clam [1003, 592, 180, 775, 1023, 336, 867, 895, 1277, 1226, 238]. clams [1332]. Clangula [486]. Clark [1078, 1331]. class [865]. classical [529]. Claus [770]. clausi [1139]. clava [334]. clavata [625, 165]. cleaner [949, 614]. Clibanarius [1206]. Climate [567, 568, 168, 576, 171, 920, 1371, 577, 969, 527, 600]. Climate-related [168]. Clinal [421]. Cliona [1072]. Clitellata [117]. clone [657]. Cloning [789]. close [918]. closely [465]. clownfish [631]. Clupea [431]. Clupeidae [1176, 834, 110]. Clytia [378, 967]. Cnidaria [1260, 1351, 1350, 441, 287, 1216, 1363, 963, 30, 962, 716, 863, 964, 748, 464, 409, 61, 791, 1252, 362]. cnidarian [739]. CO [679, 353, 795, 80, 1143]. Coast [916, 1294, 929, 1345, 304, 1204, 1206, 977, 1180, 825, 1266, 1312, 1328, 156, 1358, 588, 855, 1136, 718, 1176, 310, 715, 466, 324, 93, 534, 1310, 312, 1218, 206, 936, 793, 371, 687, 27, 483, 653, 1225, 919, 236, 1192, 1309, 416, 1277, 1349, 917, 946, 1248, 1050, 145, 1066, 486, 927, 1112, 1091, 1283, 40, 412, 914, 165, 1307, 396, 98, 1105, 1213, 68, 911, 1019, 380, 693, 1314, 957, 776, 388, 564]. Coastal [735, 193, 1191, 859, 884, 337, 918, 1094, 320, 419, 689, 765, 397, 25, 1296, 989, 990, 751, 346, 185, 1093, 838, 1110, 902, 842, 926, 882, 886, 1334, 598, 113, 928, 949, 93, 730, 1362, 1295, 1233, 747, 1181, 451, 183, 1371, 700, 1067, 1348, 945, 285, 1013, 1248, 526, 205, 139, 887, 108, 1089, 557, 529, 848, 362, 702]. Coastal-Marine [859]. coastline [1294]. coasts [292, 110]. cobbi [633]. cobble [980]. cobble-filled [980]. coccolithophore [1319]. cockle [221, 322]. cockles [347]. cod [574, 197, 886, 328, 324, 305, 534, 573, 446, 576, 530, 1065, 575, 37, 188, 182, 422, 571, 253, 50, 1314, 174, 176, 356, 1036, 175]. codfishes [405]. coelacanths [414]. Coelogorgia [1318]. coerulea [1053]. cohort [1023]. COI [1162]. Cold [744, 730, 900, 55, 1143, 1040]. cold-water [1143, 1040]. coli [1266]. Colin [393]. Collaboration [179, 170]. collapses [173]. collected [1061, 1077, 122, 1123, 718, 130, 1310, 368, 811, 690, 1081]. collectors [86, 980, 497]. Collett [4]. Collichthys [1291]. Colombian [797, 717, 742]. colonies [763, 716]. Colonization [155, 35, 1203]. colonized [507]. colonizers [1079]. Colony [323, 563]. Colour [1199, 300, 496, 563, 475]. Colpomenia [1024]. Columbia [505, 1089]. column [239]. comb [860, 1185]. comber [871]. combined [839, 810, 41, 1256]. combining [847]. Comesomatidae [667, 706]. Comments [694, 391, 237, 1149, 267, 1121, 276]. commercial [1027, 171]. commercially [665, 669]. commersonii [508]. commersonnii [834]. Common [233, 808, 388, 247, 1271, 1000, 45, 1304, 1290, 832, 1022, 927, 1122, 1082, 600, 636, 448, 364]. commonalities [1070]. communities [1345, 304, 1335, 941, 480, 1201, 588, 1123, 781, 798, 515, 386, 609, 900, 1372, 746, 1181, 1195, 913, 284, 285, 1060, 1117, 517, 1286, 843, 887, 846, 340, 796, 1281, 1040]. Community [1317, 1049, 1288, 1003, 242, 742, 598, 810, 1126, 642, 1285, 1133, 716, 1156, 1050, 906, 921, 1322, 505, 33, 779, 1119, 1100, 620, 719, 836, 702, 1323]. comparable [745]. Comparative [1044, 66, 745, 478, 940, 174]. compared [1000]. Comparing [267, 353, 894]. Comparison [347, 465, 160, 398, 1201, 241, 1176, 739, 1375, 354, 746, 1195, 266, 943, 914]. comparisons [1016, 551, 709, 738]. compensatory [267]. competence [660]. competition [450, 800]. Competitive [1145]. competitors [111]. complement [657]. Complete [196, 940]. Complex [471, 935, 571, 865]. component [631]. components [519]. Composition [1126, 12, 166, 203, 1118, 1335, 869, 1099, 1180, 1352, 825, 70, 1064, 1025, 1144, 644, 607, 1344, 1103, 663, 336, 1221, 717, 1124, 640, 851, 468, 124, 551, 659, 630, 1152, 1307, 252, 1165, 779, 215, 600, 238, 658, 1323, 1200, 1040]. Compositional [1060]. compositions [400]. compounds [974]. comprehensive [874, 885]. compromised [862]. Computer [414]. Computer-aided [414]. concentration [839, 678, 221, 253]. Concentrations [371, 1327, 1230, 290, 322, 582]. conceptacles [1275]. concepts [61]. concert [233]. concurrent [397]. Condition [696, 1101, 1176, 44, 563, 279, 65, 750, 335, 634]. conditions [303, 679, 787, 723, 199, 11, 1035, 224, 449, 356]. Conference [317, 627]. confirms [1093]. confocal [944]. confused [1096]. conger [485, 485]. connecting [1097, 1056]. connectivity [985, 1162, 460, 1234, 1348]. consequence [600]. Conservation [735, 769, 146, 923, 926, 489, 1316, 616]. conservatism [1178]. Conserve [1268]. conserving [502]. consider [1132]. considered [522]. consistency [593, 812]. consistent [840]. constricta [238]. constrictus [312, 1013]. constructing [21]. consumer [826]. consumers [74]. consuming [565]. consumption [1075, 1107, 1165, 437, 1327]. containing [1245]. contaminants [1266, 1361]. Contamination [200, 1284]. content [929, 644, 74, 847]. contents [726, 749, 579, 623]. contestability [1012]. contests [390]. context [344]. continental [597, 886, 1078, 1021, 757, 43, 13, 833, 103, 796]. continuum [1298]. contractility [965]. Contrasting [812, 230, 375, 121, 597, 1176, 1218, 967]. Contribution [210, 740]. contributions [729]. control [218, 747, 982, 516, 363]. controlling [139]. copepod [1204, 888, 1296, 63, 565, 1057, 66, 27, 150, 355]. Copepoda [237, 89, 754, 23, 1115, 770, 741, 578, 215]. copepodites [52]. copepods [320, 583, 1355, 33]. Cophylogenetic [1153]. copper [1125, 875]. copulation [214]. copy [188]. Coral [1010, 1127, 643, 996, 558, 387, 1320, 1101, 1090, 1194, 1259, 691, 797, 1193, 62, 295, 818, 275, 323, 306, 316, 504, 509, 900, 913, 1088, 618, 443, 1134, 1114, 893, 1318, 296, 1060, 1005, 335, 629, 1313, 1128, 1322, 1082, 31, 1041, 36, 1200, 1143, 1040]. coral-killing [1082]. Coralliidae [907]. Corallinales [1185, 1149, 860, 1097, 1275]. coralline [459]. coralliophaga [900]. coralliophaga-coral [900]. corals [1341, 1343, 717, 542, 767, 664, 863, 395, 1228, 471]. cordatus [349, 1339, 1083]. cordyla [1204]. core [1197]. coriacea [329, 1143]. cornetfish [508]. coronata [1080]. coronatus [707]. Corophium [77, 120]. Correction [1360]. correlated [1203, 1122]. Corrigendum [477, 780]. Corycaeidae [215]. Coryphaena [364]. cosmopolitan [1162]. Costa [1214]. costatum [800]. could [941]. Coulter [62]. counter [62]. coupled [436, 437]. coupling [139, 570]. cover [572]. coverage [686, 575]. crab [1187, 783, 1206, 349, 311, 369, 1211, 1368, 972, 1262, 1366, 855, 590, 49, 325, 715, 933, 281, 1113, 1261, 1084, 1357, 778, 935, 1016, 1062, 811, 65, 662, 937, 288, 1120, 1160, 1301, 1339, 1112, 665, 630, 420, 772, 153, 1242, 983, 1243, 870, 878, 848, 358, 641, 1083]. crabs [712, 1038, 1129, 1316, 820, 163, 903, 450]. Crambionella [1350]. Crangon [1073]. Crassostrea [1266, 1347, 785, 115, 217, 377]. creek [836]. creeled [44]. Cribrilina [254]. Crimean [1099]. criniticornis [450]. crispae [555]. crispus [95]. cristata [93]. criteria [1278]. Critical [337, 268, 462]. croaker [1291, 1007]. Croatian [27]. crocodilus [1348]. Croisettea [1295]. cross [377]. crowned [707]. crucis [243]. cruise [1077, 122]. Crustacea [1260, 1168, 536, 204, 1188, 1173, 533, 783, 303, 25, 311, 67, 286, 331, 649, 131, 102, 1259, 156, 972, 140, 856, 51, 260, 760, 545, 1184, 1159, 1205, 1241, 688, 214, 935, 1258, 662, 1174, 288, 1120, 1160, 455, 517, 835, 891, 98, 1140, 110, 342, 693]. Crustacean [536, 250, 1212, 598]. crustaceans [980, 961]. cryptic [1230, 1305, 1179, 591]. cryptobiotic [708]. Cryptocentrus [352]. Cryptochiridae [691]. cryptogenic [1114]. Ctenodiscus [357]. ctenophores [973, 911]. Cuapetes [1331]. Cuban [1101]. cucumber [1175, 259, 461, 901]. cucumbers [756, 1246]. cues [799]. cultivated [84]. culturable [1288]. culture [62]. Cumacea [140]. Cumacean [872]. cumaceans [237]. cupido [1238]. Current [1272, 1216, 1363, 1063, 819, 1223, 320, 730, 858]. Currie [853]. cuticle [495]. cuttlefish [1271, 125]. Cuvier [411, 1109, 828]. cyanea [218, 399]. cyanide [1313]. Cyanobacteria [205]. Cyanothece [1026]. cycle [966, 825, 592, 1175, 268, 775, 1330, 812, 511, 1158, 214, 336, 322, 493, 518, 777, 187, 1270]. cycles [975]. Cyclic [171]. cycling [360]. cyclone [1355]. cyclonic [1051].

Cyclopoida [23, 215]. Cyclorhagida [1029, 278, 406, 1042]. Cymodocea [1251, 1186]. cymothoid [1188]. Cymothoidae [1173, 1258, 110, 342]. cynomorium [473]. Cynoscion [751, 1111]. Cystophora [93]. cysts [16]. cytochrome [167, 1179]. Cytogenetic [241, 314]. cytometric [1375]. Cytophaga [121]. Cytotoxic [111].

D [853, 149, 437]. D. [406]. dactylopterus [344, 1255, 344]. Damage [1226, 1194, 1306, 1318, 1277]. damages [558]. damaging [603]. damicornis [797, 1088]. damselfish [218, 399]. Dana [1132, 565, 1128, 215]. danae [783, 420]. Danish [1113, 453]. Dardanus [715]. d’Arguin [1190]. Dark [904, 1143, 1052]. darkness [1319]. dart [1167]. Dasyatis [448]. data [289, 759, 1245, 694, 95, 540, 859, 953, 453, 1071, 1170, 624, 541, 1122, 1269, 276, 97, 657]. databases [963]. dating [1329]. David [841, 978]. day [812]. Daya [620]. Deania [474]. debris [21]. decade [685]. decades [457]. Decapod [250, 980, 493]. Decapoda [783, 311, 288, 1160, 1140, 693, 533, 656, 349, 691, 855, 933, 447, 790, 1085, 1147, 312, 936, 1331, 1084, 811, 163, 622, 1013, 1301, 145, 455, 1222, 1242, 931, 291, 641]. Decline [1215, 1090]. decompression [530, 422]. decorator [369]. Deep [132, 1121, 1202, 542, 468, 669, 1300, 791, 75, 533, 146, 102, 611, 151, 1373, 1025, 268, 703, 140, 1315, 810, 515, 386, 545, 385, 299, 1357, 618, 962, 302, 1120, 1160, 208, 668, 462, 464, 395, 360, 559, 189, 843, 931, 1080]. deep-living [962]. Deep-sea [1121, 1202, 542, 468, 669, 791, 75, 151, 703, 140, 1315, 515, 299, 1357, 618, 302, 1120, 208, 464, 395, 360, 189, 843, 1080]. deep-sequencing [810]. Deep-water [132, 533, 611, 268, 545, 1160, 668, 462, 559, 931]. defences [519]. deficiency [99]. defined [251]. defining [196]. defoliated [840]. Degradation [296]. degraded [471]. dehydration [766]. del [909]. Delaroche [280, 118, 1255, 344]. delays [764]. delimiting [462]. Delineation [95]. Demersal [284, 1105, 924, 858]. Demographic [859, 1047, 1191]. demosponge [1138, 1022, 898]. Demospongiae [1118, 91, 618, 276]. Dendraster [1293]. Dendrobranchiata [977]. Dendrochirotida [1130]. Dendrophylliidae [275, 323]. denitrification [1288]. Denmark [88]. dense [1046]. density [725, 369, 267, 933, 126, 867, 845, 1047, 736, 612, 498]. density-dependent [267]. dentatus [1354, 513]. deoxyribonuclease [761]. dependent [1298, 999, 267, 164]. deployment [558]. deposit [120]. deposition [318]. depressa [223, 1112]. deprived [454]. Depth [268, 101, 941, 319, 663, 34, 20, 796]. depth-integrated [941]. Depth-related [101]. depths [326, 688]. derived [1143]. dermis [252]. Dermochelidae [329]. Dermochelys [329]. desbarresi [584]. described [849]. Description [533, 131, 703, 407, 1002, 150, 1006, 552, 622, 1160, 493, 412, 741, 1019, 81, 204, 75, 1333, 67, 286, 331, 1148, 1010, 1077, 122, 992, 1260, 134, 798, 29, 1072, 584, 1238, 1331, 104, 618, 1130, 1161, 135, 581, 854, 698, 1137, 960, 548, 210, 406, 301, 1140, 110, 1081, 362]. Descriptions [163, 26, 1214, 25, 130, 132, 1037, 639, 243, 541, 907, 580]. descriptors [397, 465]. design [902, 1306]. designs [629]. Desmosomatidae [102]. despite [1242]. destructive [617]. detailed [646]. detect [397]. Detection [463, 720, 1267, 1256]. determinant [399]. determination [1102, 356, 175]. determine [1168]. determined [62, 328]. determines [826]. determining [74]. detritivore [73]. detritivorous [1273]. Deuterostomia [159]. Developing [558]. Development [315, 834, 218, 311, 234, 180, 1157, 63, 466, 1158, 281, 214, 368, 381, 811, 13, 258, 518, 944, 461, 505, 931, 1030, 1111, 654]. Developmental [807, 1299, 99, 1083]. developments [1127]. DGGE [719]. Diadema [1157, 801, 1047]. diadematid [1245]. Diadematidae [1157, 1078, 507]. Diadematoida [1157]. diagnosis [1259, 1276]. Diagnostic [752]. Diapterus [411]. diatom [1296]. Dicentrarchus [410]. dicyemid [729]. Diel [753, 678]. Diet [454, 602, 607, 1103, 385, 930, 851, 1107, 255, 445, 659, 772, 1165, 380, 1352, 1312, 1093, 1176, 408, 579, 258, 617, 525, 600, 1314, 971]. Dietary [565, 446, 760, 809, 1327]. Diets [93, 354, 431, 183]. differ [843]. difference [1022]. Differences [1262, 44, 1230, 549, 913, 766, 918, 1045, 1038, 644, 1195, 92, 246]. Different [773, 1104, 1320, 114, 425, 679, 947, 1230, 1347, 1212, 353, 1336, 37, 868, 795, 716, 1273, 629, 465, 1313, 34, 1246, 505, 340, 1030]. Differential [650, 615, 528, 1336, 278, 789]. differentially [358]. differentiate [1229]. differentiation [1275, 460, 946, 993, 240, 957]. Digenea [914]. digeneans [633]. digested [617, 525]. digestive [463]. digitalis [62]. digitata [857]. dilatata [51, 997]. dimorphism [1085, 1357, 1102]. dinoflagellate [904, 1032, 1287, 1125]. Dinophyceae [673]. Dinophyta [787]. Diodontidae [786]. Diogenidae [935, 1062]. dioica [753, 33]. diomedeae [1167]. Diplodus [321, 890]. dipstick [1256]. Dipturus [264]. Direct [94, 653]. directional [1242]. discard [663]. discarded [1306]. discards [1277, 1105]. discharge [185, 850]. Discodorididae [22]. discordance [1242]. discovered [1064]. discovery [146, 226, 1340]. discrepans [1185]. Discrete [186]. Dislodgement [905]. dispersal [635]. Dispersion [608, 812]. display [1273]. Disruption [1187]. dissolved [1126, 159, 363]. distance [126]. Distinct [1346, 790, 196]. Distinction [794]. distinctiveness [874]. distributed [1072, 1286]. Distribution [649, 1220, 388, 564, 866, 222, 1369, 373, 230, 1059, 821, 945, 468, 748, 266, 690, 40, 181, 1182, 591, 641, 21, 1264, 869, 1099, 783, 765, 977, 941, 1254, 592, 506, 972, 242, 1025, 1358, 268, 141, 1110, 1350, 1245, 1338, 590, 1209, 771, 1034, 357, 674, 444, 161, 970, 1325, 1002, 386, 312, 549, 434, 408, 206, 572, 66, 1261, 575, 604, 1250, 481, 1205, 1241, 1233, 601, 1150, 220, 463, 431, 867, 1124, 259, 737, 1197, 714, 922, 1349, 704, 455, 341, 1283, 975, 292, 833, 599]. distribution [892, 675, 1092, 753, 375, 34, 638, 12, 1042, 438, 430, 1119, 578, 215, 878, 203, 1058, 646, 57, 1074]. Distributional [1316, 560, 831, 817]. distributions [480, 793, 1056, 395]. disturbance [1003, 479, 347, 1069]. diurnal [998, 1248]. diver [497]. diver-based [497]. divergence [1300]. divergent [1043]. Diverse [611]. Diversity [1321, 844, 803, 504, 206, 747, 837, 724, 31, 36, 1099, 1020, 1328, 744, 882, 742, 785, 222, 1282, 687, 1250, 381, 960, 1089, 203, 957, 719, 121, 57]. Diving [873]. DMSP [459, 565]. DMSP-consuming [565]. DNA [562, 1142, 333, 1123, 84, 188, 483, 653, 381, 1062, 726, 824, 617, 109, 301, 355, 1043]. DNA-based [301]. do [9, 843, 92, 689, 520]. DOC [363]. dock [1123]. documenting [395]. Does [319, 1229, 37, 123, 1049, 1351, 820]. dofleini [999]. dogfish [272, 511, 851, 474]. Dokdo [245]. dokdoense [245]. dollar [1293]. dollfusi [936]. dolphin [674, 1122, 18]. dolphinfish [364]. dolphins [700, 927]. domed [1228]. dominant [320, 1056, 516]. dominate [816]. dominated [36]. domoic [463]. Donax [1277, 1226]. Dongdae [707]. donghaiense [1256, 800, 1125]. d’Orbigny [372]. Doridina [22]. Dorvilleidae [14, 119, 83]. Doryteuthis [840]. Dosidicus [953]. down [936]. Dracoderes [406]. Draconematidae [245]. draconis [656]. dracula [68]. dragonet [1269]. dredge [347, 1307]. dredging [747, 1277, 1226]. Dreissenidae [1325]. Drift [920, 93]. drilling [762]. drive [763]. drivers [1008, 1235]. drives [1090]. dropstone [515]. dropstone-induced [515]. drum [957]. ducks [486, 914]. due [402]. dufresnei [831, 777]. dugongs [489]. duration [1212]. during [651, 1194, 825, 1061, 432, 404, 1077, 122, 1110, 1123, 1375, 1272, 1106, 324, 463, 431, 670, 822, 322, 402, 124, 766, 313, 1247, 892, 505, 901, 430, 620, 176, 327, 358]. dust [1100]. duster [987]. dwelling [412, 1032]. dye [1290]. Dyer [271, 482]. dynamic [59]. Dynamics [428, 77, 1129, 758, 112, 694, 1271, 193, 161, 148, 17, 429, 889, 763, 967, 1067, 1024, 633, 835, 420, 624, 139, 363, 1042, 550, 997, 364]. Dynoides [1205].

E. [854]. Early [1299, 955, 1007, 1030, 923, 441, 251, 400, 125, 981, 944, 892]. East [194, 576, 575, 101, 556, 387, 234, 1136, 93, 3, 1076, 754, 188, 198, 1036, 121]. Eastern [423, 310, 1342, 1088, 203, 90, 25, 1148, 1025, 1358, 480, 729, 1085, 1002, 1065, 687, 1205, 1033, 930, 1059, 503, 1046, 7, 1066, 709, 1217, 165, 484, 521, 1169, 768, 375, 931, 168, 1269, 1374, 1151, 1180, 1330, 871, 299, 802, 1352, 1305, 1240]. Echinoderes [1029, 854, 1042]. Echinodermata [357, 130, 584, 1078, 1225, 1098, 1130, 259, 128, 1349, 1236, 258, 1270, 1081]. Echinoderms [756]. Echinoidea [1078, 1166, 1299, 128, 1349, 258, 1270]. Echinometra [1349, 1270]. Echinus [725]. Echiura [135]. echiuran [487]. ECO [138, 137, 127, 130, 548]. ECO-MAR [548]. Ecological [1101, 977, 734, 432, 1302, 1210]. Ecology [804, 489, 818, 555, 439, 182, 3, 695, 939, 595, 21, 146, 18, 339, 1190, 407, 393, 535, 621, 478, 467, 1033, 841, 809, 927, 1091, 847, 165, 103, 566, 1041]. ECOMAR [537, 538]. Economic [331, 1131]. Ecophenotypic [184]. Ecosystem [428, 1216, 1363, 173, 439, 577, 548, 605, 339, 429, 7, 205, 531, 847, 606, 1353]. Ecosystems [100, 1011, 378, 587]. ecotones [387]. Ecuador [423]. Ecuadorian [953]. eddy [1197, 1051]. edge [1021, 256]. edible [1045, 603, 866, 986]. edited [989, 990, 1017, 1018, 853, 939]. edition [853, 514]. editor [1359, 138]. Editorial [382, 684, 806, 915]. Editors [179]. edule [347]. edulis [86, 1064, 1057, 155, 453, 582, 96, 591, 1028, 1151]. Edward [903]. Edwards [1368, 447, 622, 1160]. Edwardsia [407]. edwardsii [1027, 897, 502, 665, 931]. eel [348, 241, 485, 661, 1137, 1031]. eelgrass [746, 707, 1048, 655, 875]. eelpout [703, 4]. Eels [939]. Effect [216, 1034, 1334, 1212, 334, 678, 228, 801, 261, 655, 800, 71, 839, 459, 1025, 718, 530, 1124, 65, 510, 229, 41, 1182, 849, 256, 1048]. effective [989, 990]. effectiveness [680]. Effects [428, 731, 1211, 1368, 762, 197, 479, 1053, 701, 1319, 515, 567, 934, 429, 631, 910, 1054, 195, 1355, 257, 1005, 494, 34, 461, 240, 1100, 55, 377, 875, 1010, 1244, 565, 864, 1347, 1344, 1356, 1177, 173, 1126, 568, 262, 722, 920, 811, 94, 750, 756, 527, 318, 983, 906, 1313, 1073, 1210]. efficiencies [62]. efficiency [814, 267, 904]. efficient [355]. efforts [723]. Egg [466, 1038, 1262, 1071, 931, 355]. Egg-laying [466]. eggs [583, 888, 834, 1044, 1095, 355]. Egypt [426]. Eilat [960]. Elasipodida [584, 116, 1081]. Elasmobranch [1152]. Elasmobranchii [991]. elassodon [375]. electrical [629]. electrophoretic [720]. electroreception [270]. elegans [816, 305, 262, 520]. element [343]. elemental [1135]. elements [1230]. elevated [1266, 1319, 795, 80]. Elizabeth [989, 990]. Elphidiidae [601]. Elpidiidae [116]. Elthusa [342]. elucidating [587]. emarginata [342]. embryo [513]. embryological [1157]. embryonic [828, 1326]. emerging [1280]. Emiliania [1319]. emphasis [1105]. empirical [593, 1151]. employing [149]. Enchelidiidae [581]. Enchytraeidae [117]. enclosed [1144, 373, 385, 1124, 737, 531]. Encounters [1170]. encrasicolus [946]. endangered [1361, 1292, 899]. endemic [860, 691, 275]. endogenous [832]. endolith [603]. endomyarian [464]. endosymbiont [58]. Energy [1111]. engelhartii’ [1185]. England [1123]. Engraulidae [834]. Engraulis [946]. enhance [996, 469]. Enhanced [53, 402]. Enhancing [525]. enrichment [620, 875]. ENSO [1054]. entering [457]. Enteroctopus [999]. Enteromorpha [11]. entomon [839, 143]. entoproct [412]. Entoprocta [210]. entry [498]. Environment [938, 975, 99, 686, 737, 340]. Environment-specific [938, 99]. Environmental [218, 864, 1191, 516, 1008, 337, 71, 1049, 765, 1338, 1034, 1282, 934, 747, 1133, 714, 1217, 1286, 375, 356]. Environments [200, 838, 1176, 1218, 1134, 664, 464, 187, 1042]. enzyme [1232, 1211]. enzymes [335]. epacteriscids [89]. ephemeral [416]. Epialtidae [662]. epibenthic [1064, 101, 846]. Epibiont [716]. epibiontic [1296]. epibionts [1265]. epibiosis [1063]. epibiota [1251]. epibiotic [718, 1243]. epifauna [12]. epifaunal [746]. Epinephelidae [1309]. Epinephelinae [687]. Epinephelus [1309]. epiphytes [319, 1024]. epiphytic [725]. Episesarma [1368]. epithelial [459]. epitypification [860]. Epizoic [1169, 1063]. Eptatretus [988]. Epulotheres [1140]. equations [525]. equilibra [997]. eradicated [1304]. Erebonasteridae [23]. Ericthonius [835]. Eriocheir [848]. erosa [180]. errata [1086]. Erratum [979]. Erylus [276]. erythrinus [808, 310]. escape [164]. escaping [1307].

Escherichia [1266]. esculentus [725]. Espeleonaushonia [656]. Espíritu [501]. established [481]. establishment [694, 11, 517, 243, 624]. esters [1327]. estimate [1148]. estimated [144, 1028]. Estimates [881, 501, 736, 1226]. Estimating [754, 437, 438]. estimation [894, 1074]. estimations [773]. estuaries [185]. Estuarine [566, 918, 1012, 838, 923, 1209, 1325, 823, 1056, 551, 121, 1257]. estuarine-coastal [918]. Estuary [583, 454, 407, 920, 1059, 922, 921, 1298, 888, 751, 348, 758, 928, 774, 919, 821, 1309, 551, 1069, 1355, 455, 341, 927, 420, 786, 1042, 926, 925, 916, 917]. estuary-associated [751]. estuary-dependent [1298]. Etmopterus [194, 267, 268]. Etneelva [615]. ETP [1292]. euglenids [154, 384]. eukalyptorhynch [1089]. eukaryotic [1100]. Eumonostilfera [1238]. Eunicida [72]. Euphausia [525]. euphausiids [638]. Euplectellidae [541]. Euretidae [104, 541]. Europe [782, 96]. European [274, 348, 1064, 1176, 792, 1057, 410, 485, 1001, 946, 263, 211]. euryops [151]. Eutardigrada [555]. Euthyneura [22]. eutrophic [786]. eutrophicated [945]. Evaluating [657, 1278]. Evaluation [397, 1073, 699, 1339, 898, 951, 1210]. Evans [853]. event [1341, 823]. events [1054, 1068]. Evidence [53, 66, 285, 613, 349, 1361, 264, 1025, 185, 481, 1131, 24, 522]. Evolution [804, 32, 456]. evolutionary [1012, 713, 1242]. evolved [827]. ex [1336]. Examining [283]. Example [576, 22, 1036]. examples [61]. excentricus [1293]. Exclusive [1131, 331]. exerts [1232]. exhibits [791]. exotic [824, 1168]. expanding [1103]. expansa [1332]. Expansion [1004, 520]. expedition [137, 1061, 127, 130, 1081, 138]. Experience [296, 197]. experiment [801, 55, 620]. experimental [1359, 1034, 705, 530, 363, 1073]. experiments [651, 114, 53, 1248]. explain [1049, 520, 612]. explained [450]. explains [1022]. explants [17]. exploitation [501, 450]. Exploited [696, 1083]. Exploration [100]. explore [51]. Exploring [877]. explosion [786]. export [277]. exposed [1284, 167, 868, 80]. Exposure [1320, 762, 905, 1318, 19, 1313, 591]. expressed [358]. Expression [670, 1232, 167, 458, 1293, 217, 359, 789, 652]. extension [1109, 587]. external [1376]. extraction [109, 355]. extractive [1339]. Extrapolating [1036]. Extreme [697, 282]. extrusion [1262]. exudation [651]. eyed [280, 118].

F [1017, 1018]. Fabricius [325, 1112]. Facetotecta [25]. facial [496]. facilitate [923, 1297]. facilitated [1064]. facilitation [1022]. facilities [750]. fact [106]. factor [631]. Factors [916, 934, 1067, 917, 1217, 1168, 139, 340]. Falcidens [69]. fall [1373, 588, 433]. false [1325, 163]. Family [1330, 976, 1037, 116, 829, 215]. fan [367]. FAO [721]. Far [198, 918, 1127]. Farming [605, 902, 610, 822, 606]. Faroe [63, 844, 781, 386, 182, 3, 20]. Faroese [408]. Farooq [393]. Farreidae [540]. farreri [670]. fasciatus [850]. fascicularis [316]. Faster [1065]. Fatty [832, 252, 1344, 1000, 1250, 630, 570, 658, 1143, 1040]. Fauchald [442]. Fauna [138, 1237, 672, 137, 686, 1373, 844, 134, 546, 319, 127, 10, 1331, 601, 3, 914, 244, 256, 20]. Fe [1344]. feather [987]. features [1272, 1021, 793, 1302, 195, 589]. Fecundity [838, 1271, 1215, 511, 269, 773]. feed [864]. Feeding [332, 671, 49, 280, 1106, 467, 1225, 13, 809, 165, 484, 521, 1264, 1312, 432, 197, 1367, 602, 247, 1055, 105, 52, 908, 485, 1139, 832, 411, 1273, 927, 120, 732, 1189, 578, 628, 1314]. feeds [159]. felix [1022]. Fell [713]. female [197, 594, 581]. females [1262, 870, 862]. ferruginea [899]. fertilized [1044]. ficiformis [719]. Fiddler [972]. Field [869, 453, 884, 140, 476, 801, 55]. fields [1208, 629, 843]. Fifteen [174]. Fifty [179, 170]. Fiji [1259]. filaments [1097]. filled [980]. Filling [1326, 1269, 1177]. filter [120]. filtration [582]. fin [610]. fin-fish [610]. find [1219]. Fine [1012, 348, 628]. fine- [628]. Fine-scale [1012]. fingerprinting [719]. finmarchicus [1076, 436, 52, 435, 437]. finned [865]. Finnmark [86]. FIPs [1292]. First [1333, 1296, 1361, 29, 1146, 226, 523, 1295, 126, 1239, 259, 854, 1137, 145, 245, 1255, 1283, 464, 292, 1109, 983, 1231, 580, 817, 776, 472, 1079, 1265, 1148, 488, 1309, 802, 581, 285, 1326, 412, 513, 1340]. first- [513]. Fish [428, 427, 60, 64, 551, 470, 1094, 1298, 1188, 1173, 814, 1075, 454, 985, 902, 1164, 1009, 1136, 742, 598, 834, 925, 1282, 949, 730, 549, 446, 1234, 610, 429, 1059, 1258, 571, 858, 1273, 341, 801, 724, 614, 34, 525, 632, 168, 437, 438, 430, 557, 343, 55, 971, 634]. fish-cod [571]. fisheri [366]. Fisheries [769, 415, 721, 185, 1014, 1190, 1364, 928, 502, 1152, 1307, 983]. fishery [1306, 774, 501, 953, 1292, 1210, 1292]. Fishes [803, 792, 200, 213, 696, 884, 5, 671, 991, 751, 973, 838, 923, 998, 842, 880, 818, 222, 488, 1203, 980, 15, 841, 1171, 978, 203, 853]. fishing [1027, 1010, 1291, 44, 347, 1277, 1313]. fissiparous [815]. Fistularia [508]. Fistulariidae [508]. fit [633]. Fitzhugh [976, 1037]. five [163]. fixation [451, 1143]. fixed [109]. fixing [1321]. fjord [297, 277, 505, 628, 612]. fjords [339, 112, 230]. Flabellina [709]. Flabellinidae [709]. flagging [1020]. flamingos [908]. flat [1064, 908, 1179, 184]. flatfish [720]. flathead [375]. flatworm [1276]. flatworms [374, 1089]. flavescens [860, 711]. flaviventris [1109]. Flavobacteria [121]. flecked [641]. Flemish [468, 468]. flesus [594, 1001]. floating [1223, 1123, 1071, 1255]. floats [435]. flood [642, 823]. flooding [554]. flora [743]. Florida [637, 330, 676, 978]. Florideophyceae [95]. flounder [167, 594, 280, 118, 1001, 375]. Flow [1375, 725, 762, 365, 1256]. fluctuating [634]. fluctuations [760]. Fluorescence [1194]. flux [386]. fluxes [672]. Focus [274, 1331, 589, 263]. Fold [1317]. follicle [594, 327]. follicle-stimulating [594, 327]. following [642]. fontaniana [372]. Food [399, 880, 626, 73, 661, 732, 320, 973, 1093, 607, 1177, 925, 305, 573, 640, 832, 39, 74, 297, 228, 529]. food-web [297]. forage [971]. foraging [1209, 621, 723]. foraminifer [8]. Foraminifera [710, 869, 539, 452, 1099]. foraminiferal [16, 601]. foraminiferans [28]. foraminifers [189]. force [905]. forceps [1365]. forcing [1009]. Forcipulata [105]. forest [929, 1011, 41]. forests [469, 1042]. foreword [393]. forgotten [443]. formalin [109]. formalin-fixed [109]. formation [1187, 324, 619, 638]. formerly [396]. forming [648, 731, 666, 458, 42, 58]. formulation [495]. forskali [901]. Forsskål [1330]. forsskali [1374]. fossil [351]. fouling [760, 65, 517, 1119]. found [1223, 1079]. Four [102, 1097, 25, 141, 63, 644, 130, 933, 1253, 243]. four-celled [1097]. four-year [933]. Fourmanoir [1374]. fourth [853]. fractioned [400]. fractions [1187]. Fracture [151, 133, 548, 547]. fragment [1228, 1004, 97]. fragments [1088]. framework [1058]. France [708]. Francesc [107]. freckled [413]. Free [154, 384, 1238, 504, 898, 1255, 843, 779]. free-floating [1255]. Free-living [154, 384, 504, 898, 779]. French [1113]. freshwater [454, 1001, 1026]. freshwater-deprived [454]. frigate [483]. frigida [237]. fringe [1298]. fringe-mudflat [1298]. fringing [717]. Front [537, 538, 548]. frontal [753, 57]. FRV [1370]. FSH [594]. Fucales [1223, 679, 1237]. fucata [1254]. Fucus [417, 524, 2, 389]. Fuego [909]. full [686, 657]. full-length [657]. fully [436, 437]. Function [487, 111]. Functional [965, 125, 666, 49, 351, 391]. functions [1200]. Fundy [1156]. fungi [837]. furnieri [784]. further [276]. fusca [486]. fyke [85].

G [989, 990, 884, 315]. G. [130]. Gabes [941, 424, 625, 626, 322, 950]. Gadidae [1020, 405]. Gadiformes [802, 1020]. gadoid [956]. gadolinium [1284]. Gadus [197, 886, 328, 324, 305, 530, 37, 422, 50, 1314, 176]. gahi [840]. Gaidropsarus [1020, 488]. gained [1118]. galateia [374]. Galaxea [316]. Galeocerdo [828]. Galeommatoidea [633, 561]. Galeus [266]. Galician [379]. galloprovincialis [1232, 365, 761]. Galveston [315]. gametogenesis [275, 1158]. Gametogenic [180, 336]. gammarids [1034]. gammarus [239, 494]. Ganges [1355]. ganglion [358]. gaps [1326, 1269]. garfish [740]. garricki [559]. garthi [790]. gastropod [894, 1045, 1244, 1129, 866, 905, 845, 658]. Gastropoda [22, 512, 562, 466, 368, 160, 398]. gastropods [1364, 986]. gastrotrich [472, 637]. Gastrotricha [472, 207, 676, 198]. gaudichaudii [525]. Gavin [1017, 1018]. GCLME [1216, 1363]. Gdańsk [305]. gear [1307]. Gebiidea [656, 1310]. Gebruk [138]. Gelatinous [973, 1075, 1139, 1071, 1255, 344, 1353]. Gelidiaceae [868]. Gelidium [868]. Geloina [1332]. Gempylidae [1234]. gen [1185, 1275, 662, 68]. gender [927]. gender-related [927]. gene [852, 458, 1279, 365, 1098, 670, 1293]. genera [1188, 1161, 132, 698]. general [4, 914]. generalist [1082]. generalized [1058]. generated [318]. generation [513, 931]. generic [1276]. generosa [592]. genes [876, 1232, 1137, 358, 456]. genesis [513]. Genetic [349, 879, 653, 946, 522, 957, 808, 1328, 348, 1012, 759, 855, 1164, 882, 310, 694, 785, 890, 730, 1218, 509, 1141, 483, 381, 778, 186, 1199, 993, 624, 1300, 849, 794, 850]. Genetics [92, 536, 230, 1348]. genomes [940, 343, 196]. Genomic [1267, 940]. Genus [1341, 952, 26, 1333, 1188, 1204, 533, 1097, 1079, 425, 1259, 91, 706, 140, 1350, 1193, 856, 942, 970, 119, 1147, 676, 1241, 1115, 27, 963, 443, 117, 713, 622, 288, 1160, 639, 245, 1170, 464, 210, 891, 1240, 1207, 741, 278, 406, 76, 673, 198, 342, 647, 738, 1019, 1080]. Geodiidae [276]. geoduck [592, 401]. geographic [1254, 1350, 417, 575, 1205, 1241, 1062, 168, 553]. geographical [481]. geography [497]. geometric [1365]. George [978]. Gephyrocapsa [1249]. Gerlach [581]. German [155]. germlings [947]. gerontophilic [351]. Gerridae [478]. geversianus [909]. Ghana [1345]. Ghar [1346]. ghost [447, 1085, 1147, 1222]. giant [876, 469, 999, 882]. Gibbes [1301, 891]. gibbesi [153]. Gibbs [548, 547, 151, 133]. Gibbula [562]. Gibraltar [855, 51]. Gibson [1207]. Gigartinales [95]. gigas [1187, 1347, 115, 778, 953, 217]. Gilbert [802]. gill [1211]. gills [1187, 1232]. Gippsland [154]. GIS [152]. Gjøsæter [179]. glaciale [230]. glacialis [604]. glacier [230, 634]. gland [463, 658]. glandula [955]. Glass [129, 540, 541]. glauca [1033]. glaucum [220, 221, 322]. Global [415, 969, 376]. Globicephala [865]. globicipitis [1169]. globulosus [512, 894, 1045]. Gnathorhynchidae [1089]. Gnathostomulida [68]. Goa [180, 288, 693]. goatfish [361, 647, 738, 874, 1374]. Goatfishes [108, 794]. gobies [352]. gobiid [5]. Gobiidae [314, 352]. goby [314, 932, 1095]. goes [1353]. Going [936]. Golfo [512]. goliath [1309]. gonad [124, 518]. gonadal [218, 1158, 214, 1111]. gonochoric [248]. Gonochorism [275]. Gonothyraea [965]. good [383, 845]. Gorgona [797]. Gorgonia [1022]. gorgonian [423, 1021, 993]. Gorgoniidae [1254, 423]. Gorgonocephalidae [934]. Gorringe [231]. Governance [769]. Gracilechinus [713]. gracilis [378]. gradient [43, 19, 184, 796]. gradients [432, 308, 737, 591]. Gran [1246]. Grand [468, 243]. grandimana [303]. Graneledone [994]. granularis [1299]. granulometry [308]. Grapsidae [855, 1301]. Grateloupia [852, 1279, 1096, 727]. Gray [443]. Grazing [417, 33, 389, 225, 1273, 41]. Great [1084, 1137, 632]. Greater [1213, 1068]. Greece [1189, 1269]. green [458, 859, 402, 820, 1248, 903]. Greenland [93, 101, 371, 773, 103]. gregarious [500]. grey [726, 1107, 1231]. Grid [1306]. Grid-net [1306]. gromiids [28]. Gross [79]. ground [1206, 1291, 602]. grounds [840, 1014, 1285, 468]. group [116, 794, 1119]. grouper [404, 1309]. Groupers [476, 687]. groups [1321]. Growth [679, 947, 1315, 45, 408, 1088, 1022, 118, 87, 582, 630, 1303, 96, 870, 471, 71, 996, 158, 512, 894, 1045, 999, 197, 1367, 784, 345, 268, 1190, 692, 615, 1023, 933, 324, 871, 1085, 1065, 279, 787, 424, 625, 1224, 631, 230, 1371, 965, 485, 94, 144, 236, 1046, 937, 815, 1066, 445, 257, 1005, 453, 943, 1228, 253, 768, 1086, 375, 461, 1189, 398, 261, 1048, 448, 634, 655, 1263, 875, 1074, 1015]. grypus [726, 1107]. guatucupa [751]. Guest [138, 179]. Guézé [1374]. Guide [803, 733, 884, 721, 476, 299, 392]. guidelines [249]. Guinea [1216, 1363, 662]. guinotae [1261]. guitarfish [332]. Gujarat [204]. Gulf [1087, 867, 895, 1131, 291, 401, 950, 558, 304, 1308, 387, 765, 1063, 941, 1266, 1361, 345, 209, 426, 222, 305, 418, 1158, 934, 424, 625, 626, 1167, 889, 322, 961, 1046, 1199, 470, 1222, 465, 1109, 846, 1270, 364, 1041]. Gulfs [1072]. gull [763]. Gullmarsfjord [638]. Gullmarsfjorden [12]. Gully [1163].

Gunnerus [1175]. Günther [1269]. gut [385, 487]. guttatus [499, 187, 361]. Gymnolaemata [254]. Gymnothorax [241].

H [989, 990, 853, 978]. H. [237, 1368, 1078, 1096, 375, 366]. haakonmosbiensis [675]. Haan [288]. Habitat [1122, 765, 648, 332, 77, 1010, 1220, 346, 972, 268, 771, 666, 113, 515, 897, 737, 224, 335, 405, 396, 92, 108]. habitat-forming [666]. Habitats [337, 1003, 1358, 1014, 842, 1070, 1114, 967, 284, 1046, 462, 1117, 1168, 616]. habits [332, 671, 280, 1106, 626, 485, 13, 732, 484, 521, 1189]. hadal [688]. haddock [574]. haemastoma [368, 718, 368]. haematopoietic [705]. haemocyanin [839, 493]. haemoglobin [37]. haemolymph [1211, 493]. Haemulidae [1330]. hagfishes [988]. Hainan [1136, 642, 743, 1323]. Haizhou [1058]. hake [531, 1074]. halanychi [69]. Halechiniscidae [708]. Haleciidae [970]. Halecium [970]. halibut [371, 773, 103]. Halicarcinus [870]. Halichoerus [726, 1107]. Halichondria [1138, 898]. Halichondrida [91]. Haliclona [71]. Haliotidae [124]. Haliotis [124]. Hallock [841]. Halobates [478]. Halobiotus [555]. Halophila [1004]. halophyte [282]. Halymeniaceae [852, 1279, 1096, 727]. hammerhead [1329, 1141, 943]. Hand [1277, 347]. handling [901]. hanging [1364]. hanging-net [1364]. Hanleya [211]. Hansen [237, 1241]. haploceros [708]. Haplosclerida [1118]. Haplosyllis [639]. Harald [179]. Harbour [457, 78, 467, 726, 1155, 1165, 997]. harbours [739, 58, 1324]. hard [816, 1230, 155]. Hardangerfjord [608, 611, 607, 609, 610, 616, 613, 605, 606, 614]. harengus [431]. harmful [1247]. Harmothoe [1162, 300, 255, 243]. harp [1000]. Harpacticoida [1115, 770, 741]. harpacticoids [1115, 747]. Harpinia [1333]. Harpiniinae [1333]. harshness [1049]. Hartman [442]. harvesting [1003, 949, 347]. hatchery [494]. hatchery-reared [494]. hatching [583, 1028]. Haug [179]. Hawaii [1114, 1311]. Hawaiian [837, 366, 907]. head [1291, 705]. healing [422]. Health [696, 750, 158, 643]. healthy [913]. heat [1336, 1125]. heavily [656]. heavy [389, 282]. hedgpethi [481]. height [995]. helena [661]. Helgoland [654, 239, 494]. Helicolenus [446, 1255, 344]. Helicosalpa [1170]. Hellenic [1361]. hellerii [772]. hemicellulase [672]. Hemiptera [478]. hemisphere [1333]. hemiuroid [1207]. Hemiuroidea [1207]. henslowianum [679]. hepatopancreas [1182, 658]. Hepatus [641]. herbivorous [1273]. herbivorous/detritivorous [1273]. Herbst [1211]. hermaphroditic [219]. hermatypic [1313]. hermit [1206, 712, 1129, 715, 1084, 935, 1062, 450]. hermits [845]. herring [574, 432, 444, 1177, 434, 431, 809, 736, 1095]. Herwig [1370]. Hess [1084]. Heterocapsa [904]. Heterogeneity [147, 855, 1282, 515]. Heteronemertea [92]. Heterosaccus [936]. Heterosigma [1145]. Heterotardigrada [942]. heterotrophic [1362, 52, 154, 384, 360]. heterozygote [99]. Hexacorals [227]. Hexactinellida [104, 540, 129, 541]. Hexactinosida [104]. Hexaplex [160, 398]. Hexapodidae [693]. Hexapus [693]. hians [1179]. hibernicum [1149]. hidden [1020]. hierarchical [516]. Higgins [278, 406]. High [1025, 460, 1001, 285, 993, 986, 689, 328, 353, 126, 283, 832, 1348, 123, 1060, 55, 36]. high-relief [55]. high-resolution [328]. higher [162]. highlights [1090]. highly [320, 500, 758, 694, 502, 659, 624, 899]. hilonis [366]. hinge [32]. Hippidea [145]. Hippocampus [707, 366]. Hippoglossoides [375, 371, 773, 103]. Hippolyte [248]. Hippolytidae [248]. hippurus [364]. Hirondellea [545, 688]. Hirondelleidae [545, 688]. Hirudinea [15]. Histamine [660]. Histioteuthidae [329]. Histioteuthis [329]. Histocompatibility [865]. Histological [1131]. histology [1238, 944]. histology-free [1238]. Historical [177, 609, 970]. History [61, 303, 77, 339, 852, 441, 1279, 715, 321, 881, 687, 1070, 713, 1326, 955, 1297, 1007, 175]. histrix [366]. Hokkaido [1314]. holdfasts [1305]. holding [750, 1073]. Holoplankton [1198]. Holothuria [1328, 901]. holothurian [544]. holothurians [1081, 130]. Holothuroidea [584, 1130, 1081, 130, 1161, 259, 116, 547]. Holthuis [533]. homalorhagid [589]. Homalorhagida [589]. homalos [553]. Homarus [496, 239, 494]. home [1209, 845]. Homoeoscelis [237]. homogeneity [808, 855, 653, 19]. Homolodromia [1160]. Homolodromiidae [1160]. Hong [712]. hooded [93]. hook [270]. Horizontal [455, 430, 941, 549, 1261, 438]. hormone [594, 631, 327]. hormone-like [327]. Horn [91]. horneri [1223]. horse [666, 1229]. horseshoe [1187, 778, 1243]. hospite [1336]. host [995, 936, 59, 1153, 826, 633]. hosts [767, 756]. hound [1191, 1348]. Hsp70 [452]. huanghaiensis [727]. Hubb [79]. Hubbs [802]. hubbsi [531, 1074]. hubbsorum [79]. Human [916, 902, 1133, 917]. human-mediated [1133]. Human-related [916, 917]. Humboldt [320]. hummelincki [967]. hustedtiana [1296]. Hutchings [976, 1037]. huttoni [257, 258]. huvei [966]. huxleyi [1319]. Huys [23]. hybrid [99, 377]. hybrids [720, 513]. Hydrachnidia [526, 829]. Hydractiniidae [968]. hydrocarbons [178]. Hydrodynamic [1138]. Hydrodynamism [369]. hydrography [549]. Hydroid [960, 966, 968, 378, 895, 716]. Hydroides [816]. hydroids [376]. hydrological [1272]. hydrology [920]. hydrolysing [1321]. hydromedusa [912]. hydromedusae [975]. hydroplasmic [965]. hydrostatic [594]. hydrothermal [140, 23, 30, 244, 227]. Hydrozoa [963, 964, 441, 1216, 1363, 970, 1006, 967, 716, 376, 863, 76]. Hydrozoan [958, 965, 967, 767, 969, 959, 958, 959]. hyemalis [486]. Hymedesmia [1143]. Hymenasterides [134]. Hymeniacidon [1072]. Hymenodora [604]. Hymenosomatidae [163, 870]. hyperborea [651, 929, 319, 74, 41]. hyperiid [1139]. Hyperiidae [911]. Hyperoche [911]. Hyperoplus [1137]. hypersaline [888, 557]. Hypo [1026]. Hypo-osmotic [1026]. Hyporhamphus [740]. hypotheses [885]. Hypoxia [1232, 764, 143].

Iberia [824]. Iberian [1367, 992, 589, 474, 724]. ice [93, 572]. Iceland [1264, 192, 1315, 161, 933, 604, 1285, 442, 189, 892]. Icelandic [45, 183, 1371, 1324, 809, 103, 600]. Ichthyofauna [689, 113, 231].

Ichthyoplankton [758, 1035]. Idas [302]. Identification [458, 358, 984, 1230, 84, 265, 726, 465, 474, 617, 414, 1289, 994, 343]. identifications [593]. identify [1224]. Identifying [1068, 638, 752]. identity [995, 496, 826]. Idotea [389]. Ifni [1294]. Iglesias [656]. II [761, 865, 335, 1370]. III [596]. imaging [944]. imbricata [1218, 1162, 300, 255]. immature [1209]. immatures [1268]. immigrant [1330]. immune [876, 631, 261]. immune-related [876]. immunochemistry [944]. immunocytochemical [149]. Impact [1057, 949, 823, 929, 1014, 1132, 353, 162, 610, 908, 33, 850, 620]. impacted [664]. Impacts [1266, 527, 1083, 397, 666, 1292]. Implications [71, 321, 642, 336, 763, 590, 926, 1316, 32, 955, 1047, 891, 632, 628]. Importance [1298, 74, 208, 462, 995, 480, 1093, 308, 418, 502, 737, 258]. important [973, 1209, 665, 669]. Imposex [290, 425, 718, 195, 954]. Improvement [1292]. inbreeding [1347]. Incidence [603, 1001]. Incipient [730, 14]. including [1260, 1054, 1317]. inclusion [443]. incorporate [1058]. increase [168]. Increased [500]. increasing [1284]. incrustans [1149]. incubation [651]. independence [1242]. India [204, 387, 1173, 1204, 1321, 67, 286, 331, 687, 1150, 653, 1258, 288, 1248, 887, 1213, 693, 817]. Indian [1296, 649, 681, 1154, 265, 142, 802, 669, 361, 1188, 1110, 1193, 460, 483, 1302, 1280, 1160, 1236, 1327, 1235]. indica [1052]. indicate [1348]. indicated [672]. Indication [1179]. indicator [1133, 1274]. Indicators [696, 640, 108]. indices [160]. indicus [278]. indigenous [1356, 782, 967, 1324]. Individual [409, 328, 265, 438, 355]. individual-based [436, 438]. Indo [1087, 698]. Indo-Pacific [1087]. Indo-West [698]. Indonesia [1332, 1148, 1295, 1060, 1082, 256]. Induced [916, 799, 515, 917, 613]. induces [425]. industrial [616]. Infauna [246]. Infaunal [77, 1069, 906]. infection [195]. infections [65]. inference [1276]. inferences [1062]. INFERNO [428, 429]. inferred [177, 95, 653, 1300]. Inferring [1302]. Infestation [291, 603]. infestations [814, 1169]. Influence [765, 787, 1261, 663, 1197, 629, 517, 1228, 725, 369, 70, 998, 193, 251, 1191, 556, 123, 975, 254, 702]. influenced [870, 634]. influences [635, 638, 1246]. information [989, 990, 692, 488, 1029, 657, 356]. information-based [692]. infranudis [802]. Ingolf [179]. inhabitants [1070]. inhabiting [838, 539, 1302, 927, 843, 1314]. inhalant [510]. inhibition [240]. Initial [289]. initiation [399]. Inlet [601]. inodorus [730]. Inorganic [363, 94]. input [1100, 1026]. inputs [528]. Inshore [840]. Insights [1118, 1020, 1162, 1033, 1283]. insignis [715]. instability [99]. Instantaneous [114]. Instructions [394]. instrumented [283]. insularity [321]. Integrated [491, 941, 492]. integrative [885, 1252]. intensities [679]. intensity [655]. intensive [610, 616]. inter [1099, 1191, 813, 1054, 1346, 738]. inter- [738]. inter-annual [1191, 1054]. inter-oceanic [813]. inter-regional [1099]. Interaction [519, 1104, 714]. Interactions [428, 330, 7, 605, 606, 1027, 429, 153, 271, 1227, 1145]. Interactive [722, 1244]. Interannual [1076, 403, 1287]. interface [989, 990]. interference [450]. Interindividual [496]. interlinked [387]. intermedia [509]. intermediate [633]. intermonsoon [402]. internal [84, 20]. International [317]. interspecific [51, 1376, 153, 1281]. interstitial [81]. Intertidal [1049, 842, 1310, 601, 1320, 5, 966, 369, 479, 985, 909, 222, 1186, 260, 1133, 981, 224, 19, 421, 728, 1323, 1281]. intestinal [1291]. Intra [1376, 51]. Intra- [1376, 51]. intracapsular [466, 368]. intracellular [459]. intrageneric [874]. intraguild [66]. intrapopulation [1346]. intraspecific [738]. introduced [933, 162, 199]. introgression [1242]. introvert [589]. invading [694, 624]. invasion [955]. invasions [1168]. invasive [508, 1064, 1012, 816, 1057, 974, 1218, 936, 262, 889, 835, 772, 367, 1119]. inverse [867]. invertebrate [469, 731, 746, 1070, 857]. Invertebrates [1223, 1156]. Investigating [1364, 1051]. Investigation [964, 944, 348, 310, 1375, 85, 1243]. investment [1262]. involvement [755]. Ionian [1189]. ions [974]. Iotrochota [111]. Iphitime [72]. Iran [222]. Ircinia [1022]. Irish [937]. iron [1344]. iron-based [1344]. irradiance [451, 123]. irroratus [933]. isaki [1184]. Ishii [712]. Iskenderun [205]. islagrande [1222]. Island [1173, 70, 797, 180, 717, 851, 903, 951, 271, 482, 743, 1323, 1150]. islandica [351, 391]. Islands [90, 287, 681, 265, 913, 1258, 1246, 203, 1094, 1331, 1121, 1202, 1192, 689, 844, 89, 1186, 153]. Isle [937]. Isognomon [184]. Isognomonidae [184]. Isohypsibiidae [555]. isolated [744, 1362, 931]. isolation [524, 730, 827, 219]. isopod [1188, 1173, 712, 389, 1258, 1222]. Isopoda [1173, 131, 1259, 856, 1205, 1241, 1258, 110, 342, 1087, 102, 712, 291]. isotope [78, 185, 880, 644, 621, 467, 826, 579]. isotopes [454, 985, 925, 1116, 1040]. isotopic [847, 570, 1257]. isozyme [720]. Israel [936]. Israeli [113]. Issue [138, 179]. Issues [1132, 536, 784]. isthmus [1214]. Istiophoridae [345]. Istiophorus [345, 532, 1039]. Isurus [1033]. itajara [1309]. Italy [25, 889]. item [329]. IX [1369].

J [1017, 1018, 853]. J. [1146]. Jacobia [135]. jalzici [27]. Janeiro [598]. Japan [70, 942, 563, 1238, 1261, 1159, 1084, 1056, 1174, 296, 1276, 1314, 1028]. Japanese [478, 214]. japonica [746]. japonicus [461, 327]. Jasus [1027, 897, 502]. Java [400]. Jawa [706]. Jellyfish [682, 1284, 882, 1371, 1199, 659, 1217, 1227]. johni [1278]. Jordanian [304]. Jorge [867, 895, 291]. journal [1]. Juarrero [656]. jumbo [953]. jumping [229]. Junction [140]. Just [874]. Juvenile [311, 671, 1312, 1366, 177, 897, 279, 1113, 1141, 859, 717, 32, 1139, 1309, 1001, 664, 922, 494, 471, 1058]. juveniles [977, 617, 636].

K-ATPase [789, 359]. Ka’an [501]. kaempferi [645]. Kagoshima [70]. Kalliapseudidae [1069]. kanamycin [240]. Kangerdlugssuaq [101]. Kappaphycus [910, 766]. Karadurmus [1359]. Karling [1340]. karyological [488]. karyotypes [1137]. Kasagia [1120]. Kastela [518]. Kattegat [78]. Kattegat/Skagerrak [78]. kD [452]. Kells [884]. kelp [651, 929, 469, 840, 1011, 319, 763, 416, 225, 728, 41]. Kemp [1209]. kempii [1209]. Kendari [1332]. Kerala [1204]. KEXUE [1081]. key [1188, 706, 63, 856, 834, 573, 935, 1292, 7, 198, 361, 1019, 1235, 362]. kidney [705]. killing [1082]. king [1262, 1366, 49, 325]. Kinorhynch [1372, 1181, 759]. Kinorhyncha [1029, 589, 278, 406, 1042]. kinorhynchs [854, 589]. knife [347]. knowledge [210]. Known [667, 1185, 472, 856, 1078, 539, 1029, 475, 673]. Koch [443]. Kochi [1248]. kofoidii [1032]. Koinocystididae [1089]. Kölliker [1315]. komatsui [1042]. Komodo [1082]. Kong [712]. Kongsfjorden [297]. Kononenko [131]. Korea [1362, 556, 1115, 707, 245, 741]. Korean [623]. Koslow [146]. krill [386, 892]. Kroyer [51]. Kuroshio [1223]. Kuruma [327]. Kussakin [1259]. Kyushu [70].

L [197, 1078, 45, 852, 1279, 86, 305]. L. [860, 1149, 44, 247, 328, 2, 246, 37, 625, 410, 485, 13, 1199, 518, 906, 768, 96, 1073, 862, 219]. Laboratory [16, 303, 679, 811, 1326, 752]. laboratory-reared [811, 752]. Labrador [7]. labrax [410]. Labridae [1376, 614]. labronica [14, 83]. Lacepede [483]. lack [1328, 882, 460]. lactic [1356]. lactuca [947, 795]. lacunosa [62]. laeve [860]. lageria [589]. Lagoon [926, 925, 920, 919, 922, 916, 917, 921, 765, 67, 751, 598, 1272, 290, 551, 557, 928]. Lagoons [337, 1067, 1346]. lake [823, 526, 1026]. lakes [1144, 1056, 1117]. Lakshadweep [67, 1331, 1150]. Lamarck [1232, 1299, 365, 510, 238]. Laminaria [651, 929, 319, 74, 857, 41]. lancelet [196]. lanceolatus [1137]. lances [1137]. Land [122, 349, 750]. land-based [750]. landing [326]. Langerhans [639]. lantern [194, 268]. lanternfish [230]. lanternshark [267]. Laomediidae [656]. Laptev [29]. Large [814, 1216, 1363, 416, 659, 1071, 634]. Large-scale [814, 416]. Larimichthys [1291, 1007]. Larry [884]. larvacean [753]. larvae [197, 1245, 590, 49, 1212, 444, 1177, 834, 549, 1150, 1293, 858, 258, 239, 377, 799, 1083]. Larval [234, 974, 1016, 811, 307, 634, 197, 1011, 1157, 1009, 60, 660, 258, 633, 891, 253, 461, 632, 755]. Last [1017, 1018]. late [431]. Lateral [200, 1256]. Latimeria [414]. latissima [416, 318]. Latitudinal [1358, 534, 316, 400, 846]. Latreille [1243]. Lawrence [1046]. layers [398]. laying [466]. LDH [720]. Leach [51]. lead [1284, 371]. leads [1320]. leaf [1338, 73]. Learned [270, 1127]. leatherback [329]. Leathesia [1024]. leaves [1251]. leeches [15]. leiachus [1111]. Leiosolenus [938]. lemon [270]. Length [1352, 707, 953, 525, 878, 97, 657]. lentiginosa [413]. lentil [564]. lentillifera [1146]. leopard [404]. Lepeophtheirus [814]. Lepidochelys [1209]. Lepidonotus [255]. Lepidophthalmus [447]. Lepidopus [692]. lepocreadiid [1171]. Leptochitonidae [189]. Leptophytum [860]. Leptothecata [376]. Lernaeenicus [1204]. lesions [79]. Lessepsian [1330, 113]. lessepsianus [849]. Lessonia [840, 1305]. lessoniana [279]. lessonii [1030]. Lessons [605, 606]. Lesueur [1204]. Letter [1359]. Leucosiidae [1357]. Leucozonia [650]. Leukophilia [404].

Levant [694, 624]. Levantine [480, 1025, 872]. Level [415, 1244, 1000, 115]. Levels [778, 679, 947, 594, 631, 451, 795, 582, 34]. LH [594]. lhcSR [458]. libraries [657]. lice [814, 608, 613]. lice-induced [613]. Life [303, 697, 715, 302, 1297, 966, 77, 852, 923, 441, 1279, 172, 881, 687, 1070, 1326, 955, 975, 892, 1007]. Life-history [715, 1297, 77, 1070, 955]. lifespan [1023]. light [1244, 679, 458, 904, 791, 655]. Ligia [1182]. like [322, 343, 327]. limestone [1159]. limit [592, 1074]. limitation [499]. limited [953]. limopsid [32]. Limopsis [32]. limpet [1030, 899]. lineage [1043]. lineages [808, 553]. linear [1365]. Linking [648]. Linnaeus [1204, 194, 839, 348, 1367, 310, 871, 251, 424, 269, 236, 124, 300, 1137, 1131, 493, 750, 1283, 165, 484, 521, 1297, 457, 1268, 655, 1041]. Linné [209]. lionfish [1087, 998]. Liparidae [157]. lipid [644, 630]. Lipids [172]. Lipophrys [985]. List [792]. Lithodes [1262, 1366, 291]. Lithodidae [291]. Lithognathus [236]. Lithophyllaceae [1149]. Lithophyllum [1149]. Litopenaeus [359, 652]. litter [73]. little [500, 1291, 723, 628, 717]. little-gregarious [500]. littoral [1014, 798, 113, 378, 298, 550]. littorea [97]. Littorina [97]. Live [750]. Living [567, 568, 452, 504, 220, 154, 384, 962, 577, 351, 863, 898, 779, 1119]. Livu [1259]. Lizardfish [1213, 849, 1058]. Lobster [317, 491, 492, 216, 1027, 500, 1278, 496, 251, 460, 495, 501, 502, 503, 499, 239, 497, 498]. lobsters [44, 897, 385, 494, 862]. Local [1264, 1338, 376, 36]. locality [247]. localization [1137]. located [289]. location [612]. locations [141, 312]. Loch [190]. locomotion [326]. locomotory [494]. loggerhead [1265]. loliginid [945]. Loliginidae [1230, 279, 1167]. Loligo [87]. Lolliguncula [1167]. Long [966, 1144, 923, 1023, 928, 199, 916, 927, 848, 918, 177, 865, 917, 1327, 1041, 851]. long-finned [865]. Long-term [966, 923, 1023, 928, 199, 916, 927, 848, 918, 177, 917, 1327]. longa [141, 578]. longevity [96]. longicollis [936]. longimanus [881]. Longipedia [770]. Longipediidae [770]. longispinulata [292]. longispinus [1078]. longtail [653]. Looking [496]. loop [483, 653]. López [580]. Lophelia [762]. lophii [506]. Lophiiformes [1163]. Lophius [506]. Loricifera [226, 29]. loss [705]. Lou [1096]. loveni [965]. Low [1328, 882, 1320, 459, 458, 910, 502, 904, 1060, 1047]. low-density [1047]. lower [91]. Loxosomella [412, 210]. loyola [968]. lucens [525]. lucerna [1135, 409]. Lucernariopsis [362]. Lucia [454, 407, 1026]. lucidus [1291]. Lucky [140]. lucunter [1349]. Ludwig [1130]. ludwigi [584]. Luiz [884]. lunar [1009, 812, 836]. lupus [247]. lusitanicus [854]. luteinizing [594]. lutjanid [475, 1253]. Lutjanus [465]. Lycenchelys [4]. Lychnorhiza [409]. Lysekil [166]. Lysianassoidea [545]. Lysmata [219]. Lysosomal [761]. Lyusi [1291].

M [989, 990]. M. [591]. MacDonald [989, 990]. mackerel [1312, 1229, 431, 809]. macloviana [895]. macro [731, 1265]. macro-epibionts [1265]. macro-invertebrate [731]. macroalgae [648, 725, 680, 1142, 2, 920, 88]. macroalgal [162, 609, 94, 11]. macrobenthic [672, 289, 1011, 208]. macrocephalus [788, 1289, 1314]. macroconsumers [1116]. Macrocystis [469]. macrodactylus [1119]. Macrodasyida [637, 207]. macrofauna [2, 228, 1243, 166]. macrofaunal [373]. macroinvertebrate [765, 1003, 846]. macroinvertebrates [1180, 880, 1325]. macronutrients [21]. macrophytes [376]. macroplankton [571]. macroplankton-pelagic [571]. macrosoma [40]. macrospina [1021]. macrothalli [1024]. Macrouridae [802]. Mactra [336, 1283]. mactroides [775]. maculatus [1140]. Madagascar [523, 526, 1128]. made [558]. Madeira [90]. maenas [1113, 820, 65, 903]. maeoticus [309]. Magdalena [1091]. Magellanic [1103, 971]. magellanica [866]. magellanicus [1166]. magnetic [1208]. magnets [1208]. magnitude [920]. Maharashtra [204]. main [551]. Maine [207]. mainensis [207]. mainland [1135]. maintaining [378]. Majidae [1120]. Majoidea [891]. major [171, 1195, 1297, 865]. Makaira [380]. makes [845]. making [745]. maklayi [662]. mako [1033]. Malay [778]. Malaysia [774, 951, 836, 1200]. Malaysian [878]. maldanids [1114]. Maldives [265, 1006]. male [1211, 325, 1113, 983, 252]. Mallotus [573]. Maltese [153]. Maluku [1148]. Mamanguape [1059]. mammals [873]. Man [937, 558]. man-made [558]. Management [317, 491, 200, 734, 721, 492, 272, 699, 1316, 336, 763, 108, 989, 990]. manatees [489]. manefishes [1121]. Mangrove [951, 1042, 672, 1321, 349, 1332, 785, 774, 281, 1339, 1242, 1243, 836, 1083]. mangroves [387]. Manipulation [71]. Mannar [387]. Manning [811, 1140]. manta [265, 265]. mantis [214]. mantle [953]. Manual [654]. Manus [30]. Mapping [190, 686]. MAR [138, 137, 127, 130, 548, 557]. MAR-ECO [138, 137, 127, 130]. marbled [855]. Marcello [1017, 1018]. Marcus [481]. MAREANO [686, 690]. Marenzeller [584]. margalefii [673]. margin [597, 757, 43, 1089, 796]. marginal [643]. margins [1286]. marhei [1006]. marina [1003, 2, 246, 678, 746, 707, 1024, 906, 1048, 655, 875]. marinas [760]. Marine [152, 179, 294, 427, 991, 100, 393, 792, 694, 1072, 810, 1216, 1363, 567, 769, 1163, 859, 390, 577, 916, 696, 391, 978, 635, 1214, 397, 894, 1045, 1361, 1142, 560, 744, 1144, 1190, 873, 201, 295, 535, 353, 148, 1369, 555, 949, 1342, 170, 568, 226, 10, 573, 207, 708, 15, 861, 1239, 283, 154, 80, 1124, 640, 737, 821, 309, 841, 123, 917, 7, 462, 526, 1117, 927, 1365, 1112, 847, 1303, 9, 827, 829, 1171, 198, 1276, 587, 807, 557, 657, 899, 776, 991, 1359, 1, 82, 100, 201, 295, 383, 745]. Marine [685, 885]. Marini [1074]. Marionina [117]. maritime [824]. mark [933]. mark-recapture [933]. marker [1162, 381]. markers [333, 785, 84, 1062, 665, 1031, 1213]. market [476]. marks [894, 144, 398]. marlin [1104, 380]. Marmara [1359]. marmoratus [855]. Marsupenaeus [327]. Martín [292, 580]. Martínez [656]. Martínez-Iglesias [656]. martinezii [911]. Maryland [1067]. Mass [744, 1071]. massive [966].

Massy [75]. matallanasi [703]. match [1196]. Mate [499]. material [1149]. Maternal [1139]. Matías [401]. mating [1102, 764, 1231]. matrix [1255, 344]. mats [11]. matter [929, 114, 1116, 1126, 159, 360]. Matthias [1017, 1018]. Maturation [511, 223, 404, 1176, 408, 124, 375, 327, 358]. mature [868]. maturing [594]. Maturity [1308, 325, 1278, 281, 625, 269, 773, 704, 937, 1066, 931, 870, 878, 1074]. Mauritania [1190]. Mauritius [1289]. may [22, 1042]. McEachran [413]. meadow [880, 812]. meadows [289, 246, 256]. measured [980, 497]. measurements [1212]. Measuring [742, 1212]. mechanical [1248, 1226]. mechanism [1364, 386, 378, 229]. mechanisms [432, 267]. mederi [1368]. media [1048, 732]. mediated [1133, 277]. Mediterranean [25, 966, 480, 1330, 1021, 424, 625, 930, 1290, 1008, 266, 732, 210, 298, 550, 557, 808, 562, 508, 671, 731, 1265, 264, 1025, 267, 1014, 872, 759, 855, 682, 310, 782, 275, 113, 1203, 890, 936, 365, 481, 889, 443, 967, 661, 749, 993, 205, 313, 1086, 849, 363, 168, 1269, 1374, 899, 1041]. mediterraneus [768]. Medusae [76, 973]. medusivorous [671]. Medusozoa [61]. Megabenthic [597, 1286]. megabenthos [1317]. Megafaunal [350, 796, 101]. Megalaspis [1204]. megalopa [1083]. megalopae [311, 455]. Meganyctiphanes [386, 445]. Megastygarctides [1019]. Meinert [342]. Meiobenthic [340, 1042, 779]. meiofaunal [908]. Melanitta [486]. melanochir [740]. melanophore [513]. melanopus [631]. melanostomus [1095, 314, 932]. melas [865]. melastomus [266]. meleagris [1199]. Melh [1346]. member [673]. Menor [557]. Menorca [1021]. mentella [177]. mercury [371]. meridionalis [237]. Merlangius [1352, 1268]. merlangus [1352, 1268]. Merluccius [531, 1074]. meroplankton [1198]. Mersin [480]. mesh [385, 663]. mesh-enclosed [385]. mesocosm [620]. mesocosms [197]. Mesodinium [42, 58, 229]. Mesograzer [826]. mesopelagic [634]. mesophotic [770]. Mesophyllaceae [1185, 1275]. Mesophyllum [1185]. Mesopodopsis [125]. Mesoscale [1092, 924]. mesoscales [628]. mesotrophic [528]. mesozooplankton [1335, 161, 921]. Messina [671]. mestrovi [150]. Metabolic [115]. metabolism [1090, 701, 1239, 795, 487, 832]. Metacarcinus [665]. Metacirolana [1259]. metagenomic [810]. Metal [200, 322]. metallothionein [322, 652]. metallothionein-like [322]. metals [1361, 986, 282]. metamorphic [660]. metamorphosis [974, 799, 1083]. metazooplankton [419, 941]. methane [952, 1240]. method [216, 827]. Methodical [798]. methods [137, 894, 1212, 692, 1210]. Metriatyla [1328]. Metridia [141, 578]. Mexican [501, 1091, 532, 1131, 1253, 292]. Mexico [1317, 1104, 1087, 765, 1063, 1266, 345, 209, 418, 1158, 934, 1167, 1131, 470, 1222, 1109]. meyeri [1148]. Mg [260]. MHC [865]. micans [787]. micro [1265]. micro- [1265]. microalgae [1303]. microarray [343]. microbes [360]. Microbial [35, 719, 1320, 781, 810, 1369, 495, 913, 111, 528, 1060, 529]. microbiology [393]. Microboring [863]. microcosm [285]. microenvironment [8]. Microhabitat [5, 184]. Micromesistius [431, 433]. Micromonas [418]. micronutrients [21]. microorganisms [385, 819]. Microphrys [369]. microplankton [1375]. microplastic [1318]. Microplastics [1280, 1246]. Micropogonias [784]. Microsatellite [381, 617, 785, 665]. microsatellites [876]. microscopy [944]. Microstomatidae [151]. microstructure [784, 1263]. microzooplankton [39, 836]. Mictyris [1261]. Mid [140, 1067, 542, 692, 487, 599, 138, 137, 1183, 131, 157, 134, 546, 127, 130, 537, 538, 539, 545, 23, 543, 540, 962, 128, 133, 548, 547, 544, 129, 541, 235, 1370]. Mid-Atlantic [140, 542, 692, 599, 138, 137, 1183, 131, 157, 134, 546, 127, 130, 537, 538, 539, 545, 23, 543, 540, 962, 128, 133, 548, 547, 544, 129, 541, 235, 1370]. mid-gut [487]. Mid-West [1067]. middle [479, 215]. Miers [1173]. might [54]. migrants [113]. Migration [841, 396, 432, 444, 1106, 678, 420]. migrations [594]. migratory [730, 1028]. million [54]. Milne [1368, 447, 622, 1160]. Milne-Edwards [447, 622, 1160]. mineralogy [883]. Minicoy [67, 1150]. minima [10]. minimum [1288]. minke [1264, 1000, 832, 50, 623, 600]. minor [1126, 564, 1043]. minuta [105]. minutus [1084, 182]. mirabilis [701, 126]. miraletus [424, 626, 521]. mironovi [134]. mirus [776]. Misidentification [1359]. Misool [1060]. Misophrioida [150]. Misund [179]. mite [526, 829]. Mithracidae [369, 1016, 1365]. Mithraculus [1365]. Mithrax [1016]. mitigate [1292]. mitigation [558]. Mito [1242]. Mito-nuclear [1242]. Mitochondrial [1098, 808, 759, 188, 483, 653, 487, 1242, 940, 553, 196]. mitogenomes [1329]. mixed [256]. mixed-species [256]. mixing [551]. Mixotrophy [1032]. Mn [217]. Mn-SOD [217]. mobile [731, 1372]. Mobula [1361]. mobular [1361]. mode [828]. Model [11, 692, 436, 953, 453, 437]. modeling [569]. Modelling [436, 1039, 438]. models [1122, 1058, 535]. moderates [1249]. Modiolinae [666]. Modiolus [603, 666]. modulated [1338]. modules [1138]. Mohn [244]. mojarra [411]. Molecular [264, 729, 84, 474, 24, 1140, 849, 652, 1125, 456, 1329, 51, 95, 1213, 647]. Molina [1366]. mollusc [822, 20]. Mollusca [22, 192, 1218, 336, 1071, 1283, 619, 69, 189, 940, 211, 994]. Molluscan [298, 550]. Molluscs [588, 824]. mollusk [149]. Monhysterida [757]. monitor [1292]. Monitoring [563, 814, 1027, 577, 1339, 108]. Monocorophium [835]. monodi [811]. monodon [876, 421]. Monokalliapseudes [1069]. monospecific [963]. monothalamid [869]. monothalamids [1099]. Monothalamous [28]. monsoon [1110, 187]. Monsoonal [836]. monstrosa [13]. Mooreonuphis [234]. moray [241, 661]. morhua [305, 197, 886, 328, 324, 530, 37, 422, 50, 176]. mormyrus [236]. Moroccan [1294, 1176, 1229, 946]. Morocco [1342]. Moronidae [410]. Morphological [852, 1021, 890, 1218, 1362, 264, 1279, 1302, 309, 1066, 752, 1213, 647, 275]. morphologically [1229, 716]. Morphology [713, 891, 1015, 888, 459, 712, 987, 666, 49, 905, 496, 1078, 1159, 1376, 32, 965, 831, 351, 1326, 391, 318, 421, 1237, 675]. Morphometric [1365, 1212, 1002]. morphometrics [308, 465]. Morphometry [1046]. morphotypes [1283]. mortalities [1036]. mortality [512, 701, 1023, 871, 236, 981, 613]. mortem [251]. Morton [391]. mosaicism [1108]. Motile [42, 12]. Moult [493]. Moulting [903, 1366, 494]. mounds [618]. Mousson [1332]. mouth [972]. Movement [985, 999, 177, 965, 756]. movements [1264]. mRNA [167, 217]. mtDNA [460, 1300]. mud [656, 1211, 180, 281, 630, 878, 358]. muddy [895]. mudflat [1298]. muelleri [156, 519]. Mugil [705]. Müller [778]. mullet [602, 705]. Mullidae [602, 1346, 108, 361, 647, 794, 738, 874]. Mullus [602, 1346]. Multi [415]. Multi-Level [415]. multibranchiata [1002]. multiparous [1262]. multiple [349, 397]. multiporate [1275]. Mumbai [286]. Munidopsis [385]. Munn [393]. Munnidae [856]. Muraena [661]. Muraenidae [241, 661]. murex [1290]. Muricidae [368, 160, 398]. murrayi [1315]. muscle [1245, 251, 467]. Musellifer [676]. museum [395]. Mushroom [767, 542]. mussel [1049, 158, 1012, 666, 1325, 365, 1304, 761, 302, 944, 591, 755]. Mussels [1240, 86, 1057, 722, 981, 453, 582, 1219]. Mussidae [691]. Mussismilia [691]. Mustelus [1191, 851, 1231]. muticum [1237]. Mya [1023]. Mycteroperca [404]. Myctophidae [1263]. Myrianida [771]. myriocystum [1337]. Mysid [1067, 649, 125]. Mysida [98]. Mysidacea [67, 649]. Mysids [204, 286, 331, 821]. Mytilicola [1057]. mytilid [603, 938]. Mytilidae [666, 302, 1240]. Mytilopsis [1325, 974]. Mytilus [1232, 909, 86, 1057, 155, 365, 761, 453, 582, 99, 96, 591]. Myxinidae [988].

n [1080, 74]. nagelfar [211]. namarensis [545]. name [23]. named [22]. naming [22]. Nan’ao [620]. Nannastacidae [140]. Nannochloropsis [1344]. nano [1375]. nano- [1375]. nanoflagellates [1362]. nanoparticles [1344]. nanoplankton [1197]. NaOH [223]. NaOH-alkalinized [223]. Narcomedusae [963]. narrownose [1191]. nassariid [954]. Nassariidae [950, 512]. nassariids [950]. National [1190, 1082, 88]. native [1063, 731, 1064, 1034, 1057, 746, 1168, 1237, 1004, 848]. Natural [916, 339, 1373, 1126, 720, 385, 309, 917]. naturally [827]. Nauplii [25]. Naushonia [656]. Navigation [432]. Naylor [1017, 1018]. NB [1156]. NE-Atlantic [242]. NEA [575]. neap [1272]. Near [1150, 242, 1025, 1334, 1370, 657]. near-bottom [242, 1025]. nearby [918]. nearshore [1009, 478]. needlefish [1348, 1179]. needs [1292]. Negaprion [270]. Neilonella [867, 895]. nekton [924]. nematifera [1082]. nematocysts [964]. Nematoda [706, 757, 581, 245]. Nematode [843, 308, 515, 556, 669]. nematodes [779]. Nemertea [1238, 301, 92]. nemertean [301]. Nemipterus [1300]. Nemopilema [882]. Nemopsis [912]. Neocallichirus [1147]. Neogastropoda [950, 831]. Neogobius [314, 932, 1095]. Neolepton [633]. Nephasoma [299]. Nephrops [216, 44, 251, 493, 750, 862]. Neptune [195]. Neptunea [144, 195]. Nereididae [1055]. nervous [149]. net [1306, 1364]. nets [663, 85]. neuroactive [974]. neuston [311]. Newfoundland [584, 468, 7]. newly [933, 767]. Nezumia [802]. Nha [1245]. Nibea [957]. niche [1178, 411]. niches [1257]. Nicippe [1184]. Nicobar [1154, 681]. Nicol [561]. Nicothoidae [237]. Nidalia [443]. Nidaliidae [443]. nidarosiensis [264]. nigellus [157]. Night [1244, 812]. nigra [486]. nigricans [380]. Niña [449, 259]. Nine [1236]. Niño [449]. Niño-La [449]. Niphates [62]. nitrate [651]. nitrogen [1321, 864, 467, 1195, 795, 1303]. nitrogen-fixing [1321]. Nivanteredo [1080]. No [721, 1365]. no-take [1365]. Noctiluca [402, 1248, 1287]. nodosa [1251, 1186]. Nodularia [1015]. nodularin [1015]. nodule [669, 843]. nodule-free [843]. noerrevangi [776]. Nogueira [976]. nom [23]. Nomeidae [671]. nomurai [882]. non [731, 680, 1034, 782, 746, 967, 1324, 1307, 1237, 617, 1004, 848]. non-destructive [617]. non-indigenous [782, 967, 1324]. non-native [731, 1034, 746, 1237, 1004, 848]. non-target [1307]. non-traditional [680]. Norileca [1258]. Normalizing [1209]. normani [75]. North [194, 860, 86, 1315, 1072, 681, 881, 299, 1059, 3, 244, 239, 494, 332, 731, 234, 984, 247, 10, 1285, 1205, 1370, 1235, 1020, 966, 999, 1064, 912, 771, 1076, 1172, 576, 980, 66, 575, 188, 528, 865, 135, 704, 946, 6, 464, 412, 728, 1171, 753, 34, 68, 931, 110, 721, 1036, 448, 136]. North-east [3, 234, 1076, 188, 1036]. North-eastern [1059, 1205]. north-Norwegian [247]. North-west [980]. north-western [731, 984, 966, 1171]. northeast [506, 900, 1112, 96, 75, 771, 176, 356, 175]. northeastern [689, 1121, 1202, 639, 205, 702, 231, 886]. Northern [604, 290, 211, 158, 1296, 825, 1183, 1312, 157, 590, 134, 546, 112, 718, 127, 130, 193, 1154, 934, 956, 1285, 371, 543, 1084, 368, 540, 982, 402, 132, 128, 133, 1160, 7, 6, 507, 824, 659, 690, 277, 777, 547, 474, 559, 1300, 34, 129, 541, 522, 1226, 235, 553, 971, 1287, 1041, 36, 649, 1251, 1191, 1274, 542]. northernmost [1074]. northward [1264]. northwest [653, 245, 69, 591, 468, 1043]. northwestern [39, 1373, 1100]. norvegica [386, 445]. norvegicus [216, 44, 251, 493, 750, 862]. Norway [1142, 609, 929, 216, 860, 1175, 112, 44, 607, 86, 2, 246, 251, 720, 534, 72, 162, 610, 371, 15, 416, 4, 1107, 1050, 1155, 690, 788, 728, 616, 613, 605, 606, 862]. Norwegian [1165, 179, 428, 1312, 339, 686, 1220, 1343, 264, 247, 444, 324, 170, 436, 10, 429, 434, 230, 431, 277, 435, 736, 437, 438, 430, 174]. nosed [764]. note [1172]. Notes [708, 136, 157, 771, 1078, 72, 1331, 1159, 1205, 1241, 259, 675, 278]. nourish [898]. nov [237, 656, 860, 1185, 1275, 472, 637, 131, 852, 91, 157, 703, 134, 374, 584, 681, 1238, 23, 1159, 150, 1096, 1006, 581, 854, 662, 709, 245, 412, 210, 589, 559, 68, 361, 1019, 693, 727]. Novaya [860]. novel [828, 1260, 665]. NPs [1344]. Nr [138, 179]. nuclear [556, 1242, 553]. nucleotide [216]. nudibranch [709]. Nudibranchia [22, 481]. Nuevo [512]. number [188]. nurseries [996, 664]. nursery [1206, 1312, 842, 1005, 877, 557]. nutrient [678, 528, 904, 363, 620, 1125]. nutrition [828, 257, 258]. Nutritional [21, 862, 1083]. nutritionally [862]. nutritive [1158]. NW [330, 228, 869, 860, 1008, 914, 363, 550].

O [1344, 130]. O. [83]. OAS [1110]. obliquimanus [248]. observation [651, 852, 326, 1279, 776]. observations [1342, 16, 889, 1255, 983, 569]. obtain [1071]. obtained [311]. occultus [92]. occupancy [972, 1281]. Occurrence [15, 327, 583, 1251, 264, 157, 534, 520, 580]. Ocean [1090, 100, 338, 1342, 556, 669, 911, 989, 990, 970, 572, 722, 230, 1249, 883, 632, 121, 952, 1223, 1296, 649, 973, 1075, 560, 797, 703, 838, 912, 782, 692, 681, 1154, 879, 881, 813, 1205, 265, 713, 1039, 468, 135, 142, 802, 1160, 993, 245, 943, 69, 421, 1092, 753, 76, 361, 1252, 870, 529, 291, 1100, 1327, 1235, 196]. oceanic [432, 1162, 881, 813, 1033, 1121, 1202]. oceanica [289, 73, 1249]. Oceanographic [1009, 1035, 531, 1317]. Oceans [885, 460]. octocoral [1022]. Octocorallia [423, 1254, 473, 443, 542]. octopod [75, 994]. octopus [999, 372, 79, 463, 79, 1043]. oculata [1344]. Ocypodidae [349, 972]. Odobenus [252]. odontocete [749]. Odontocymbiola [866]. Odontozona [533]. Oegopsida [1071]. off [952, 204, 320, 1204, 194, 332, 888, 686, 91, 185, 1110, 1176, 310, 834, 193, 312, 206, 936, 371, 813, 687, 717, 43, 1121, 1202, 503, 1035, 659, 690, 847, 395, 724, 1240, 632, 874, 380, 1289, 276]. off-shore [686]. officinalis [1271, 125]. offshore [1285, 747, 1150]. offspring [931]. Oikopleura [753, 33]. oil [1177]. Okinawa [563, 296]. old [54]. Ole [179]. Olencira [110]. Oligobrachia [675]. Oligobrachiidae [675]. oligoclora [1279]. oligotrophic [528]. olivacea [1211, 281, 878]. olivaceus [167, 513]. olive [167]. Oman [222, 1335]. ommastrephid [1071]. Ommastrephidae [704]. omnivorous [925]. Omura [813]. OMZ [1195]. on-offshore [1150]. onboard [1073]. Onchidium [658]. one [1081]. online [963]. only [1025]. onset [920, 937, 878]. ontogenesis [275, 670]. Ontogenetic [1130, 893, 619, 475, 13]. Ontogeny [410]. onuphid [1002, 668]. Onuphidae [234]. oocyte [223]. oogenesis [473]. open [71, 1144, 1117, 1100, 121]. open-water [71]. opercula [894]. opercularis [164]. operculata [716]. operculum [144, 398]. Ophiacanthida [1098]. Ophichtidae [1031]. ophidiid [698]. Ophidiidae [698]. Ophidiiformes [698]. Ophioderma [1225]. Ophiopholis [701]. Ophisurus [1031]. Ophiuroidea [934, 1225, 1098, 132, 1236]. Ophryotrocha [14, 119, 83]. opilio [325, 983]. opisthobranchs [619]. Oplegnathus [850]. opportunistic [1093]. opportunity [500, 283, 664]. opposite [1273]. optical [1334]. optically [1110]. optimal [754]. Optimization [1210]. orange [275, 323, 187, 878]. orange-spotted [187]. orangeclaw [1062]. oratoria [214]. Oratosquilla [214]. Orbiniidae [681]. Orbulina [8]. oregonensis [752]. organ [472]. organic [929, 114, 1361, 1136, 1116, 1126, 159, 360, 363]. organisms [289, 201, 353, 863]. organotin [650]. organs [630]. orientalis [1057]. orientation [1005]. origin [95, 1219, 456]. Orthoplana [1340]. Oslofjorden [4]. osmolality [1211]. osmoregulation [839, 262]. osmoregulatory [630]. osmotic [1026]. ossicles [1130]. ostracod [70]. Ostrea [1064]. Ostreidae [785]. ostreus [1140]. Otaria [711]. other [998, 954, 986, 943]. otolith [985, 784, 1135, 324, 1224, 1346, 1001, 465, 1263]. Otoliths [1094, 1148, 740, 636, 634, 175]. Otoplanidae [1340]. outer [1286]. outplanted [471]. oval [279]. Ovarian [281, 931, 327, 358]. Overlap [431, 1233, 809, 438, 430]. overstreeti [1222]. overwinter [435]. oviparous [877]. Oviposition [635, 777]. oweniana [388]. oxidase [1179]. oxidative [1232, 1344, 1318]. Oxycephalus [1139]. Oxygen [148, 17, 1288, 985, 10, 115, 1124]. oxyrinchus [1033]. oyster [1063, 1266, 1347, 694, 785, 1218, 115, 217, 624, 184]. oysters [1064, 377]. ozone [451].

P [1017, 1018]. P. [584, 513, 1257]. P4501A [167]. Pachycara [703]. Pachygrapsus [855, 1301]. Pacific [797, 703, 423, 881, 556, 421, 1100, 196, 1223, 1087, 352, 1061, 999, 234, 560, 1373, 984, 1358, 532, 729, 1347, 782, 1085, 1002, 1172, 385, 879, 1205, 1033, 1088, 30, 713, 953, 622, 698, 217, 1305, 245, 1091, 464, 292, 1240, 1171, 1080, 1314, 1081, 1043, 227]. Pagellus [808, 310]. Pagès [107]. Paguridae [1084]. Paguristes [935]. Pagurus [72, 1084, 450, 937]. PAHs [178]. pairing [893]. Pak [281]. Palaemon [305, 262, 520, 1119]. palaemonid [1331]. Palaemonidae [1119]. palatability [74]. Pallas [77, 314, 1095]. palmosa [1318]. palp [1172]. Palythoa [563]. panamense [934]. Pandalidae [7, 6, 931]. Pandalus [956, 1285, 7, 6, 307]. pandora [808]. panicea [1138, 898]. Panopea [592, 401]. pans [286]. pansies [519]. Pantopipetta [1172]. Panulirus [501, 499]. Papua [662]. Paracatenula [374]. Paraccacladiinae [1207]. Paraccacladium [1207]. Paracoryne [966]. Paradiopatra [1002]. Paradoxapseudes [1159]. parahaemolyticus [1356]. paralarvae [372]. Paralichthys [167, 513]. Paraliparis [157]. Paralithodes [49, 325]. parallel [946]. paramamosain [358]. parameters [881, 1221, 341, 1283, 1317, 870]. Paramuricea [1021]. Paranaguá [968]. Paranaitis [202]. parasite [506, 936, 59, 195, 1222]. Parasites [151, 729, 534, 633]. Parasitic [189, 237, 1188, 1173, 1204, 712, 1057, 1113, 837, 1258, 110]. parasitizing [1087, 1113, 1171]. Parastichopus [1175]. Pardaliscidae [1184]. parentage [617]. Parental [723, 258]. Parilia [1357]. Park [1190, 1082]. Parmales [1091]. paromolos [553]. part [855, 1084, 776, 540, 541]. partial [237, 80]. partially [525]. partitioning [1012, 446, 411, 1253, 450, 50]. partly [617]. Parupeneus [1374]. Pass [468]. passive [497]. pastinaca [448]. Patagonia [1274, 401, 590, 1166, 763, 954, 1035, 633, 777, 971]. Patagonian [1072, 986, 531, 847, 753, 634]. patagonicus [938]. patchiness [1164]. Patella [223, 899]. paternity [349, 665]. pathogenic [822]. Patos [926, 925, 928, 920, 919, 922, 916, 917, 921]. pattern [177, 1271, 1325, 1002, 670, 867, 748, 266, 139, 887, 401, 121]. Patterns [1129, 598, 1099, 765, 648, 1251, 506, 797, 1366, 923, 844, 615, 60, 812, 1203, 720, 373, 1316, 980, 230, 778, 1371, 1059, 118, 376, 531, 975, 34, 638, 414, 1042, 215, 1235, 36]. Paucitubulatina [676]. paucity [113]. Paul [873]. paulensis [922, 1257]. Paulo [977]. Pawson [713]. pCO [1319]. PCR [822]. pea [811]. pearl [1218, 702]. Pectinoidea [192]. peculiar [727]. pedunculate [645]. Pelagic [1163, 1164, 754, 478, 39, 571, 579, 139, 570, 168, 438, 203, 1353]. pelagic-benthic [139]. Peltidium [741]. pen [1315, 1046]. penaeid [426]. Penaeidae [977, 312]. penaeids [1257]. Penaeoidea [1013]. Penaeus [876, 977, 922, 789, 261, 1257]. penantis [51]. pencil [1202]. Penghu [1201]. Penguin [1103]. penguins [971]. Penilidia [584]. Peninsula [25, 474, 1367, 992, 778, 589, 724]. Peninsular [1200]. pennaki [708]. pennatulid [473]. Pennellidae [1204]. pens [1343]. peracarid [1069, 1168]. Peracarida [1174, 98, 997]. perch [446]. perches [830]. Perciformes [1109, 1234, 523, 1121, 1309, 1153, 1297, 1111]. Percnon [153]. Percoidei [830]. percoides [446]. Perdido [1317]. perennis [554]. performance [320, 383, 629, 239, 261]. performances [868]. Pericelis [1276]. perils [963]. Perinereis [864, 1055]. period [1375, 615, 313, 248, 176]. periods [378, 463, 677]. Peripheral [705]. Perithallis [1275]. Perkinsyllis [580]. perlevis [1072]. permanent [58, 1208]. Perna [158, 944, 755]. Persian [222, 465]. persistent [1194, 1361, 462]. perspective [1371, 577, 88]. perspectives [819]. perturbation [666]. pertusa [762]. Peru [1358, 688]. Peter [989, 990, 1017, 1018, 1084]. petronioi [1326]. Petrosia [719]. Petrovich [212]. Pettibonella [681]. pH [459, 947, 94, 123, 1293]. Phaeodactylum [1145]. Phaeophyceae [417, 524, 2, 1024, 1237]. Phaeophyta [679]. phaeton [1269, 1269]. phagocytes [1158]. Phakellia [91]. Phanang [281]. Phascolosoma [879, 59]. Phase [335, 1322, 1058]. phenological [1076]. phenotype [37]. Phenotypic [19, 1274, 258]. philippi [741]. Philippines [1341, 163, 471]. phlorotannin [74]. Phoca [45]. Phocoena [726, 1155, 457]. Phoenicopterus [908]. pholis [985]. phosphate [1321]. phospholipid [1143]. phospholipid-derived [1143]. photo [1289]. photo-identification [1289]. photosynthesis [795, 94]. photosynthetic [679, 947, 451, 868, 910, 766, 1313, 655, 875]. photosystem [1194]. Phoxocephalidae [1333]. Phragmope [1185]. phthalate [1327]. Phycogorgia [1254]. Phyllodoce [1311]. Phyllodocida [553]. Phyllodocidae [1311, 202]. Phyllophora [869]. PhyloCode [22]. Phylogenetic [744, 1178, 1329, 1062, 976, 1171, 1207, 553, 1362, 1331, 1098, 1006, 412, 770, 24, 1276]. phylogenetics [250]. Phylogeny [940, 119, 849, 1015, 456]. Phylogeographic [1234, 1062, 946]. Phylogeographical [321]. Phylogeography [808, 855, 536, 1072, 1354, 1140]. phylum [759, 301]. Physeter [788, 1289]. Physical [462, 1003, 479, 436, 1030]. physical/primary [436]. Physiological [80, 1368, 325, 1336, 217]. physiology [216, 167, 873, 853, 318]. Phytoplankton [387, 112, 1196, 1334, 1126, 400, 451, 837, 919, 516, 123, 1050, 898, 1100, 620, 1294]. pica [905]. pico [1375, 1197]. pico- [1375, 1197]. picocyanobacterial [80]. picoeukaryote [418]. picophytoplankton [418]. picoplanktonic [205]. piece [488]. piggybacking [756]. Pigment [919, 418, 655, 875]. Pigment-based [919]. pigmentation [1174]. pigmentatum [581]. pilchardus [1164, 1176, 1008]. pilot [865]. Pinctada [1218]. pink [922]. Pinnotheridae [811, 1140]. pipefish [764]. piscatorius [506]. Pisces [602, 687, 661, 118, 4, 110].

Piscicolidae [15]. Pisione [292, 24]. Pisionidae [292, 24]. pit [747]. Pitho [891]. Pituitary [594]. placement [1329, 676, 24]. Plagusia [1112]. plaice [886]. Plain [242]. plan [616]. planatus [870]. Planet [152]. planktivorous [437, 430]. plankton [1075, 242, 535, 1272, 811, 528, 285, 33]. plankton-collected [811]. planktonic [1345, 539, 8, 646]. plant [556]. plants [875]. planula [275]. planulae [1053]. plasticity [258]. plate [62]. platessa [886]. platform [1349]. Platichthys [594, 1001]. Platt [706]. Platyhelminthes [374, 1193, 914, 1276, 1089]. platypterus [345, 532, 1039]. Pleoticus [156]. Plesionika [931]. Pleuronectes [886]. plocamia [659]. plumage [1102]. plume [702]. plumes [185]. Plumularia [716]. Plumularioidea [1216]. pneumatophore [1298]. Pocillopora [797, 1088]. Pocilloporidae [1215, 1088]. Pocket [721]. poculostylis [637]. podas [280, 118]. Podocoryna [968]. Poecilasma [645]. Poecilosclerida [618]. Pogonophora [675]. Poland [486, 914]. Polar [537, 573, 1375, 538, 548]. Polarstern [1077]. policy [989, 990]. Polish [848, 776]. Pollachius [408]. pollutants [1317]. polluted [373, 220]. polyactis [1291, 1007]. Polyaromatic [178]. Polyarthra [770]. Polycera [481]. Polyceridae [481]. Polychaeta [243, 292, 14, 26, 1206, 234, 987, 190, 816, 864, 1055, 1002, 72, 442, 202, 639, 255, 367, 580]. Polychaete [833, 1162, 681, 373, 3, 300]. polychaetes [1123, 543, 668, 544]. polyclad [1193]. Polycladida [1193, 1276]. Polycystididae [1089]. polymerase [984, 1256]. Polymesoda [180]. polymetallic [669, 843]. polymorphic [84]. polymorphism [1102, 865, 1199, 300, 97]. Polynoid [544]. Polynoidae [1162, 255, 243]. polyp [762, 323]. Polyplacophora [189, 211]. polyplax [815]. Polyprion [1221]. Pomacentridae [1153, 1297]. Pomadasys [1330]. Ponganis [873]. Pontarachnidae [526, 829]. Pontella [1296]. Ponto [1034]. Ponto-Caspian [1034]. Pontogammarus [309]. pools [842, 1186]. poor [182, 97]. poor-cod [182]. poorly [1078, 539, 475, 673]. Population [876, 1135, 933, 1084, 1357, 967, 1348, 1013, 1066, 1155, 835, 786, 1119, 997, 1094, 158, 77, 985, 909, 1129, 426, 694, 193, 536, 711, 66, 813, 1224, 230, 889, 763, 1022, 714, 1339, 633, 1365, 1283, 665, 420, 772, 624, 983, 1092, 1300, 168, 1042, 870, 1297, 850, 957, 641, 1064]. Populations [428, 1045, 1328, 267, 1076, 310, 1103, 890, 1218, 429, 378, 1113, 188, 993, 1047, 827, 931, 1043, 1030]. Porcupine [242]. Porifera [1118, 1072, 104, 540, 129, 541]. Porphyra [84, 240]. porpoise [1155, 457]. porpoises [78, 467, 726]. Porsangerfjord [728]. Porter [1278]. ports [824]. Portugal [1135, 700, 236, 854, 1277, 145, 380]. Portuguese [388, 564, 87, 13, 40]. portunid [288]. Portunidae [783, 936]. Posidonia [289, 35, 73]. position [1006, 412, 891, 1207]. positioning [239]. positive [791]. Possible [751, 399, 85, 619]. Post [52, 1011, 1110, 251, 747, 1054, 1293, 1326, 981, 1228, 1073, 1026]. Post-bloom [52, 1026]. post-catch [1073]. post-dredging [747]. post-embryonic [1326]. post-ENSO [1054]. post-larvae [1293]. post-larval [1011]. post-mortem [251]. post-settlement [981]. pot [1354]. potassium [143]. potential [635, 500, 816, 864, 1230, 1369, 418, 386, 1195, 1088, 39, 1339, 527, 1317, 252, 657, 1268]. potentially [1174]. Pourtalesiid [128]. poutassou [431, 433]. power [556]. Practical [249]. practices [1073]. praegustator [110]. praeputialis [1054]. Prasinophyte [418]. prawn [262, 327]. prawns [426, 520]. Precision [736, 356]. Predation [956, 1095, 751, 897, 66, 820, 578, 1036]. Predator [932, 909, 444, 390, 164]. predator-prey [390]. Predators [6, 283, 1156, 7]. predatory [1047]. preference [37]. preferences [1064, 247, 661]. Preferential [1290]. Preliminary [624, 779, 28, 1243, 694]. Preparation [1337]. Presence [1246, 1186, 481, 1168, 1240]. preservation [1212]. preserved [1192]. preserving [899]. pressure [1284, 594, 80]. Prevalence [712, 1321, 1266, 859, 529]. previously [460, 1096, 1121, 522]. Prey [50, 78, 329, 602, 674, 444, 52, 1290, 390, 726, 1107, 932, 749, 253, 252, 1058]. preying [125]. Priacanthidae [1148]. prideaux [72]. primary [436, 402]. prime [378]. primiparous [1262]. Prince [903]. Printziana [1275]. Prionace [1033]. Prionospio [1154]. Pristigenys [1148]. probably [159]. probiotics [261]. Problem [415, 85]. process [711]. processes [251, 763, 531, 955]. Prochaetosoma [245]. production [459, 436, 610, 402, 405, 921, 1015]. production/individual [436]. production/individual-based [436]. productive [1334, 659]. productivity [387, 400, 403]. products [216]. profiling [1250, 1303]. profundorum [474]. programme [686, 442, 548]. progression [781]. Project [428, 537, 538, 429, 548]. projecting [576]. Projects [1292]. prokaryotic [1201]. prolifera [458, 181]. proliferation [1266, 240]. prolonged [1319, 677, 1242]. promotes [660]. propelling [229]. properties [1334, 951]. Prorocentrum [787, 1256, 800, 1125]. Proseriata [1340]. Protected [991, 1163, 859, 949, 640, 1292, 899]. protection [1273]. Protective [1356]. protein [452, 322, 493, 1125]. Protista [28, 539]. protists [657]. Protodrilidae [552]. Protodrilus [552]. protoporphyrin [1369]. Protoraphidaceae [1296]. Protoraphis [1296]. Protothaca [449]. provide [97]. providing [519]. province [1148, 1136, 1247]. proxy [1129]. Psammobatis [413]. Psammodrilidae [81]. Psednos [157]. Psenopsis [1302]. Pseudechinus [1166, 257, 258]. Pseudione [291]. Pseudocerotidae [1193]. Pseudomonas [822]. Pseudophycis [446]. Pseudorhombilidae [311]. Pseudothyone [1130]. Psychropotid [1081]. Pteriidae [1218]. Pteriomorphia [940]. Pterois [1087]. Pteromaja [662]. ptychodactiarian [287]. Ptychodactis [287]. publishing [807]. Puck [747, 779]. pudibundus [641]. puerilis [83]. puerulus [502]. pulchra [1320, 1194]. pulsed [1208]. pulses [502]. punctata [354]. punctatus [835]. punctulata [1158]. puncture [530]. puntazzo [890]. purple [1290]. Purposes [721]. purpurata [510]. pusilla [418]. pustulosum [466]. Putative [755, 1229, 1200]. putatively [464]. puzzle [488, 97]. Pycnogonida [1077, 1172, 1134, 690]. Pycnogonidae [122]. Pycnophyes [589]. pyrene [167]. pyrifera [469]. Pyropia [1108]. pyrosequencing [657]. Pyura [1054].

Qingdao [852]. quality [114, 335, 341]. quantitative [984]. Quatrefages [407]. Queen [164]. Queensland [829]. question [1075, 109]. quo [807].

R [1017, 1018]. R. [1296]. rabbitfishes [893]. radiata [1218]. radicans [417, 524, 389]. Radiolaria [61]. Radular [831]. RAFOS [435]. raggedtooth [396]. Rainbow [23]. Raja [424, 625, 626, 165, 521]. Rajidae [424, 626]. Randall [1206, 1300]. randalli [1300]. random [84]. Ranellidae [466]. range [560, 1338, 417, 1084, 831, 630, 1109, 253, 1179, 587]. ranges [1209]. ranging [268]. Rao [1350]. rapa [825, 1306, 1304, 799]. Rapana [825, 1304, 1307, 799]. Raphidophyceae [678]. Rapid [984, 1123, 386, 530, 643, 827, 1256]. rare [1193, 523, 1170, 278]. rate [1023, 582, 1086, 498, 1015]. rates [1224, 501, 1277, 445]. Rathke [305, 262, 520]. ratio [74, 898, 248, 870]. ratios [826]. Raunefjord [1050]. ray [625, 626, 265, 165, 521]. rayed [1218]. Rays [274, 263, 273, 1017, 1018]. razor [238]. rbc [852, 1279]. rDNA [84, 657, 719]. rDNA-DGGE [719]. re [1278, 584, 225, 243, 110]. re-description [584, 110]. re-establishment [243]. re-evaluating [1278]. re-vegetation [225]. reaction [984]. readings [174]. real [984, 822, 106]. real-time [984, 822]. reared [303, 197, 811, 1326, 494, 752]. Rearing [664]. rearrangement [1098]. rebound [1268]. recapture [933]. recently [1064, 507, 849]. receptors [755]. recognition [1275]. recognized [460]. Recolonization [1304]. recombinant [631]. Recombinase [1256]. recommendations [798, 1292, 752]. reconstruct [525]. reconstruction [970, 149]. Record [1350, 1333, 1148, 1193, 190, 1146, 523, 126, 481, 1239, 259, 802, 145, 245, 292, 1109, 1231, 817]. recorded [767]. records [231, 1183, 546, 1342, 1115, 1134, 135, 581, 1236, 395, 98, 587, 362]. Recovery [503, 143, 1194, 711, 1026]. Recruitment [86, 558, 469, 968, 574, 289, 156, 1053, 502, 118, 1156, 41, 497]. recruits [982]. recycling [1195]. Red [558, 304, 725, 426, 792, 961, 459, 156, 1175, 49, 602, 325, 715, 446, 1295, 868, 199, 993, 857, 1079, 113, 1203, 522, 849, 1374, 36]. Redescription [75, 1241, 1311, 988, 802, 83, 243, 342, 237, 131, 259, 406]. redfish [177]. Rediscovery [1078, 443]. reduce [820]. reduced [1293, 582]. reduces [1306]. reduction [1124, 1339]. Reef [1331, 1143, 1094, 558, 1101, 1064, 62, 563, 642, 1150, 1225, 1088, 717, 1134, 893, 664, 1060, 748, 643, 1322, 471, 1200, 717]. Reef-associated [1331]. Reefs [695, 558, 387, 1206, 1259, 1010, 1009, 295, 818, 306, 504, 900, 1070, 296, 770, 846, 31, 1210]. reference [113, 515, 478, 279]. reflect [832, 92]. refuges [998]. regeneration [510]. regime [765]. region [783, 369, 590, 483, 653, 858, 616, 1179]. Regional [610, 1099, 969]. regionalization [846]. regions [1321, 1375, 1233, 121]. regulate [1293]. regulation [755]. regulations [1273]. rehabilitation [664, 951]. rehydration [766]. reicherti [854]. Reinhardt [488]. Reinhardtius [371, 773, 103]. relate [37]. related [876, 1196, 1366, 247, 549, 101, 451, 916, 917, 927, 465, 409, 168]. relatedness [794]. relation [1264, 268, 444, 105, 1158, 640, 1067, 13, 341, 405]. relations [139]. Relationship [313, 1151, 1352, 574, 972, 509, 867, 1153, 1035, 1217, 1317, 1286, 878, 498, 20, 1281]. Relationships [1282, 1098, 640, 707]. Relative [897, 937, 258, 680, 308, 1085, 793, 845, 1066]. relatives [514]. released [677]. Relevance [971]. Relevant [721]. reliability [814]. relief [55]. remains [726, 525]. remarkable [662]. remarks [132, 1174, 770, 849, 361]. remediation [864]. remota [24]. removal [1337, 801, 906]. Renaudarctidae [942]. reniformis [519]. Renilla [519]. Renocila [1173]. repair [109]. repeats [188]. replacement [23]. replanting [951]. replicated [949]. report [1296, 759, 1295]. reported [1121]. reports [992, 587]. Reproduction [508, 156, 524, 473, 507, 1301, 367, 1308, 712, 388, 987, 532, 63, 105, 1166, 45, 447, 378, 424, 625, 236, 118, 815, 409, 1128, 254, 448, 1268]. Reproductive [320, 1332, 888, 592, 797, 1175, 1190, 775, 1278, 1166, 447, 1158, 214, 1167, 413, 1039, 518, 777, 1222, 420, 409, 1270, 401, 219, 364, 399, 825, 500, 268, 1262, 564, 1330, 524, 1315, 511, 509, 378, 687, 1084, 322, 1054, 195, 707, 257, 335, 313, 344, 1086, 187, 248, 160, 238]. Reptilia [329]. requirements [520]. Research [179, 274, 294, 491, 694, 391, 492, 177, 170, 1127, 263, 819, 807, 1359, 1, 82, 100, 201, 295, 383, 745, 685, 885]. Reserve [387, 501, 1365]. Residency [788]. Resilience [1003, 642, 285]. resistance [1347, 495]. resistant [1246]. resolution [328]. resolving [784]. resource [320, 973, 1253]. Resources [567, 568, 774, 577, 1210]. respiration [1239]. response [114, 1003, 705, 1319, 974, 576, 347, 125, 528, 1248, 1303, 164, 791, 261, 652, 1125]. Responses [572, 795, 1069, 1090, 1368, 77, 425, 679, 947, 167, 1336, 969, 217, 1208, 449, 630, 282, 1249, 901]. restoration [1088, 296, 1127]. Results [138, 428, 537, 538, 127, 429, 779]. Retention [62]. Retronectidae [374]. retrospective [699]. Return [1064]. reveal [1062, 794]. revealed [985, 880, 925, 1116, 483, 570, 456, 1040]. reveals [855, 1234, 1195, 946, 1252]. Review [146, 989, 990, 56, 1017, 1018, 830, 666, 682, 48, 38, 654, 698, 853, 191, 884, 1227, 47, 1350, 886, 1267, 481, 1292, 1037, 527, 741, 61, 110, 46, 276, 585]. Reviewer [169, 232, 293, 370, 490, 586, 683, 805, 896, 948]. revised [1259]. Revision [116, 91, 675, 907]. revisions [885]. revisited [950]. Reykjanes [60, 543]. Rhabdocoela [1089]. Rhabdoon [76]. Rhinobatidae [332]. Rhizaria [1099, 539]. Rhizocephala [936, 1113]. rhizophorae [785]. Rhizostomeae [409]. rhodaliids [962]. Rhodomonas [582]. Rhodophyta [860, 1097, 1185, 1149, 1275, 731, 852, 1279, 84, 910, 1096, 727]. rhombeus [411]. rhomboid [411]. ria [340, 228]. ribosomal [810, 1137]. Rica [1214]. Richness [640, 689, 1064, 2, 980, 1152, 36]. rides [1244]. Ridge [244, 138, 137, 1183, 131, 157, 140, 134, 546, 60, 127, 130, 537, 538, 539, 545, 23, 543, 540, 962, 128, 133, 542, 548, 547, 544, 129, 541, 235, 1370]. ridges [1007]. Ridgewayiidae [27]. ridley [1209]. Rimapenaeus [312, 1013]. Rincón [1191]. rings [894]. Rio [598]. Rise [446, 433, 1370]. Riseriellus [92]. risk [897]. risks [1327]. Risør [324]. River [615, 1059, 185, 1261, 1056, 341, 612, 457, 850, 702]. Robinione [1222]. Robsonella [372]. robustus [375]. Rocha [884]. Rocinela [1087]. rock [1027, 933, 502, 503, 12, 850]. Rockall [31]. rocky [369, 909, 1009, 222, 113, 1186, 1192, 1349, 1156, 318, 591, 846, 1281]. rodgersii [507]. Role [672, 895, 1156, 432, 909, 602, 39, 554, 932, 981, 360, 619, 253, 340, 557, 1026]. roles [1369]. Rondeletiola [564]. rooted [1321]. roots [35]. Rosa [330]. rosacea [404]. Roscoff [708]. rosefish [1255]. rosmarus [252, 252]. Ross [122]. Rossellidae [541]. Rossia [40]. Røttingen [179]. rotundicauda [1243]. roughneck [1013]. round [314, 932, 1095]. Rouville [667]. Royal [1093]. rubble [900, 1114]. rubens [1304]. rubicingulus [1078]. rubrum [42, 58, 229]. Rugiloricus [29]. Rugiloricus-species [29]. rupture [422]. Russell [1300]. Russian [179, 170, 198, 174]. Ryukyu [942].

S [989, 990, 853]. S. [186, 278]. s.l [562]. Sabatieria [667]. Sabellariid [1206]. Sabellidae [992, 987, 889, 367]. Saccharina [416, 318]. Sacculina [1113, 65]. Sacculinidae [936]. sacs [487]. Saduria [839, 143]. Sagami [1238]. sagittal [1263]. Sagmariasus [503]. Saharan [1229]. saida [573]. sailfish [345, 532, 1039]. sails [685]. saithe [408]. Sakai [1147]. salar [616, 617]. Saldanha [276]. Salento [25]. Salicornia [282]. salina [1356, 582, 1303]. salinities [1347, 868]. salinity [1211, 1368, 839, 1034, 1053, 701, 787, 520, 1261, 678, 821, 224, 630, 359, 652, 377]. Salish [752]. sallei [1325, 974]. Salmo [814, 616, 613, 617]. Salmon [605, 608, 615, 616, 613, 606, 617, 612]. salmonid [949, 614]. salmonids [617]. salmonis [814]. Salpida [1170]. salps [973]. salt [286]. same [711]. sampled [1370]. samples [1179, 31]. sampling [137, 1132, 902, 126, 548, 617, 12, 497]. San [1072, 401, 867, 895, 717, 292, 580, 291]. sand [523, 1137, 1293, 1208, 412, 1032]. sand-dwelling [412, 1032]. sand-eel [1137]. sandfish [1328]. sandhopper [1132]. sandskate [413]. sandy [53, 1244, 1132, 1116, 166, 779, 340, 1007]. sanguinea [787]. Santa [330]. Santana [991]. Santo [989, 990, 501]. santolla [1262, 1366, 291]. Sanya [504]. São [977, 441]. Sarawak [774]. Sarcocornia [554]. Sarcotragus [719]. Sarda [1106]. Sardina [1164, 1176, 1008]. sardine [1176, 1008]. Sardinella [1093, 168]. Sargassaceae [1223]. Sargassum [1223, 679, 1337, 1237, 1323]. sargus [321]. Saronikos [1361]. Sars [102, 130, 4]. sarsae [131]. sarsi [4]. Sasaki [1043]. Satellite [677]. saturation [582]. Saudi [1308]. Saurida [1213, 849]. Sauvage [1137]. savignyi [1157]. Saving [295]. saxatilis [1297]. Say [933, 1225, 1140]. scabbardfish [692]. scabra [1328]. scad [1204]. scale [814, 432, 348, 1012, 147, 1325, 793, 416, 255, 906, 34, 24, 88, 414, 646, 36]. scale-worms [255]. scaleless [24]. scales [773, 340]. scallop [670, 1274, 164]. Scalpellidae [1220]. Scaptrella [757]. scavenger [545]. scavenging [1364]. scenarios [1336]. Schander [440]. Schioedte [342]. Schizoporella [1086]. schmitti [977, 1191]. schubarti [1069]. sciaenid [730]. Sciaenidae [1224, 1111]. Science [200, 1071, 1170, 989, 990]. scientist [861]. scintillans [402, 1248, 1287]. Scleractinia [691, 1194, 1215, 275, 323, 1088]. scleractinian [316, 717, 1228]. Sclerodactylidae [1130]. sclerotic [472]. Scomber [1312, 431]. scombrus [1312, 431]. Scorpaenidae [1087]. Scorpaeniformes [1087, 157]. Scotland [190]. Scraping [1273]. screening [647]. Scyliorhinidae [559]. Scyliorhinus [269, 930, 484]. Scylla [1211, 281, 630, 878, 358]. scyphomedusa [912]. Scyphozoa [409, 1350, 209]. Scyphozoan [1217, 1371, 659]. SE [276]. Sea [1204, 1099, 25, 1061, 596, 1343, 122, 744, 480, 426, 697, 424, 625, 663, 556, 930, 1150, 1290, 410, 961, 713, 142, 225, 266, 494, 298, 550, 1269, 1374, 75, 814, 648, 725, 1265, 574, 102, 151, 346, 1175, 519, 703, 140, 1110, 1157, 1245, 287, 280, 1315, 1166, 1078, 1299, 515, 1158, 386, 446, 299, 722, 1357, 618, 30, 259, 677, 1121, 1202, 128, 542, 236, 468, 1046, 302, 815, 1120, 208, 507, 486, 257, 756, 898, 690, 801, 464, 313, 395, 360, 914, 189, 669, 843, 41, 461, 1246, 901, 791, 1080, 1270]. sea [612, 558, 304, 1335, 1288, 808, 869, 656, 428, 1079, 508, 1180, 1352, 671, 731, 966, 574, 1251, 999, 1010, 264, 1064, 1359, 1014, 703, 872, 178, 759, 1157, 771, 177, 1334, 742, 1330, 602, 644, 682, 674, 782, 325, 64, 29, 754, 1106, 1023, 173, 699, 1057, 113, 1203, 511, 1021, 567, 568, 1000, 373, 436, 573, 316, 504, 429, 434, 572, 757, 262, 520, 575, 126, 481, 171, 1084, 601, 747, 269, 150, 431, 889, 400, 451, 1195, 913, 402, 117, 485, 661, 577, 737, 714, 704, 937, 571, 1120, 749, 993, 403, 255]. Sea [518, 1008, 1253, 1217, 313, 732, 165, 210, 484, 521, 239, 1307, 983, 50, 768, 435, 1086, 752, 106, 1105, 375, 34, 1227, 819, 632, 522, 849, 743, 168, 779, 1019, 1226, 437, 438, 430, 1374, 557, 215, 1297, 529, 1095, 1007, 957, 569, 1028, 448, 719, 1268, 646, 1287, 1353, 899, 776, 121, 57, 1263, 702, 1041, 36, 146]. seabed [386]. seabird [621, 723, 1102]. seabirds [1093, 873, 527]. seabream [321, 890]. seagrass [304, 387, 765, 77, 1251, 1338, 880, 1136, 35, 812, 1186, 73, 906, 1004, 256]. Seagrasses [733]. seahorse [707]. seahorses [366]. seal [45]. Seals [85, 93, 1000, 726, 1107, 1165]. seamount [231, 276, 276]. seamounts [1121, 1202]. Seas [1189, 478, 254, 78, 1220]. seashore [392]. season [663]. Seasonal [419, 1335, 583, 968, 825, 1075, 1038, 532, 63, 590, 781, 644, 760, 621, 262, 575, 556, 924, 1050, 379, 550, 529, 862, 238, 658, 1287, 1321, 1352, 500, 742, 161, 447, 761, 277, 139, 248, 505, 569, 112]. Seasonality [378, 1069, 256]. seasonally [247]. seawater [223, 1144, 1201, 249, 285, 1200]. seaweed [1345, 1201, 417, 1295, 743]. seaweed-associated [1345]. Seaweeds [1250]. Sebastes [177, 1044]. Sebastião [441]. second [514, 513, 685]. second-generation [513]. Secondary [405, 436, 921]. Secor [841]. securis [1012]. sedentary [500]. sediment [929, 1144, 1201, 330, 1055, 308, 640, 554, 664, 1117, 756, 228, 318, 1317, 139, 843, 951, 1200]. sediment-impacted [664]. sedimentary [1198]. sedimented [996]. sediments [1321, 53, 744, 178, 148, 1369, 154, 1195, 43, 360, 779]. seedlings [679, 655, 875]. seep [1240]. seeps [952]. seilacheri [790, 1085]. Selected [212, 454, 1250, 34]. Selection [52, 5, 832, 309]. Selectivity [1055, 332, 280, 602, 607]. Selenidium [59]. Selenium [486]. Selvagens [689]. SEM [944]. semi [373, 1124, 531]. semi-enclosed [373, 1124, 531]. Senegal [636]. senegalensis [636]. sensitivity [650]. separation [827]. Sepia [1271, 125]. Sepietta [388]. Sepiola [379]. Sepiolidae [388, 564]. Sepioteuthis [279]. sequence [852, 1279, 84, 483, 865, 657]. sequences [1300, 553, 343]. sequencing [810, 653]. sequential [493]. Séret [1017, 1018]. sergeant [1297]. Seriatopora [1215]. series [177]. Seriolella [354]. serotonin [1245]. serotonin-containing [1245]. serpens [1031]. Serpula [190]. Serpulidae [190, 816]. Serra [1096]. Serranidae [687, 1109]. Serranus [871, 1109]. serrata [630]. Serratosagitta [1092]. serratus [2]. Sertularioidea [1363]. serve [1042]. sesarmid [1368]. Sesarmidae [1242]. setacea [716]. Setosabatieria [706]. Settlement [889, 680, 289, 974, 502, 981, 1004, 755]. Setúbal [700]. Sevastopol [1334]. seven [1214, 1201]. several [1169]. Sewell [1296]. Sex [333, 1102, 1031, 870, 848]. Sex-associated [333]. Sexual [269, 309, 248, 1128, 645, 1085, 625, 1159, 1357, 1102, 954, 937, 1066, 266, 409, 724, 405, 878, 254]. seychellensis [361]. Seychelles [361]. sezginii [1019]. Shabtay [694]. shading [41]. shallow [204, 588, 886, 452, 1225, 245, 98, 1128, 166, 36]. shallow-water [588, 452, 1128, 36]. shape [1135, 1224, 1274, 938, 1346, 465, 740, 1219]. shark [1308, 194, 828, 268, 1191, 881, 579, 1208, 1231]. Sharks [274, 273, 721, 272, 1329, 1033, 1141, 514, 263, 396, 270, 271, 482, 1327]. sharpsnout [890]. shelf [1099, 350, 872, 63, 781, 886, 357, 1078, 1021, 1316, 1035, 1286, 833, 103, 847]. shelf-break [357]. Shell [32, 1274, 1265, 894, 1129, 905, 1023, 938, 845, 450, 421, 1219]. shellfish [1064, 463]. shells [562, 539]. Shetland [386, 3, 20]. shift [1264, 416, 579, 1322, 168]. shifts [924, 893]. shimojiensis [1159]. ship [650]. shipworm [1080]. Shirayama [406]. Shoal [637]. shock [270, 1125, 1026]. shompens [681]. shore [995, 686, 909, 1113, 65, 1301]. shores [369, 383, 318, 591]. Short [310, 1318, 1011, 771, 1313]. Short-term [310, 1318, 1011, 1313]. shortfin [1033, 704]. shortnose [332]. show [1049, 689]. showing [34]. shrimp [876, 352, 533, 656, 156, 742, 447, 1085, 1147, 312, 1267, 956, 1285, 1331, 214, 1067, 922, 1013, 7, 6, 1066, 1326, 470, 341, 1222, 248, 931, 1119, 1073, 261, 652, 219]. shrimp-associated [352]. shrimps [21, 649, 1358, 1310, 1192]. shrubsolii [186]. Sian [501]. Sibling [372, 202]. Sicily [1014]. Siculo [855, 1164]. side [982, 470]. Siderastrea [335]. siderea [335]. Sidi [1294]. sieve [208]. Siganidae [893]. Siganus [187]. signal [970]. signata [1087]. signature [496]. signatures [1194, 1135]. significance [229]. sihama [522]. sikamea [377]. silent [146]. silico [1356]. silky [1327]. sill [10]. Sillaginidae [523]. Sillago [523, 522]. silveirai [1252]. silver [692, 354]. similarity [509]. Simonsen [1296]. simple [1049, 827]. simulate [435]. Simulating [318, 1313, 594]. simulations [11]. simultaneous [219]. sinensis [848]. Singapore [706, 1243]. single [58].

Sinonovacula [238]. sinuosa [1024]. siphon [510]. Siphonaria [1030]. Siphonophora [961, 962]. Siphonostomatoida [1204, 90]. Sipuncula [135]. sipunculid [59]. Sirembo [698]. sirenia [489]. site [1264, 635, 23, 126, 322, 1240, 1237, 1028]. site-tenacity [1264]. sites [162, 462, 643, 877]. situ [114, 651, 996, 1027, 664, 1255]. six [720, 66, 618]. Size [247, 164, 878, 702, 5, 1341, 999, 197, 1038, 718, 357, 323, 572, 281, 663, 1102, 400, 867, 111, 1022, 1346, 922, 773, 704, 937, 208, 411, 981, 1339, 801, 1228, 1152, 253, 848]. size-associated [5]. Size-based [702]. size-dependent [999]. size-fractioned [400]. size-specific [1341]. sized [320]. Skagerrak [78, 1343, 2, 324, 1165, 1143]. skate [264]. Skaters [478]. skates [1178]. skeleton [1358]. Skeletonema [800]. slabberi [125]. slender [1058]. slope [91, 584, 936, 13, 1005, 103]. slower [1065]. Small [906, 320, 689, 185, 178, 147, 1325, 793, 930, 372, 1137, 484, 168, 1007]. Small-scale [906, 1325, 793]. small-sized [320]. small-spotted [930, 484]. smelt [151]. smelts [1202]. Smirnov [675]. Smith [783]. smolts [814]. smooth [1191, 1141, 851, 943]. smooth-hound [1191]. smoothhound [1231]. snail [635, 954, 421]. snailfish [157]. snails [290, 19]. snake [1031]. snapping [1066, 1326]. snow [325, 983]. Snyder [978]. Social [1085, 271, 987, 893, 405]. Society [958, 959]. SOD [217]. Soft [1099, 10, 869, 1341, 798, 1023, 443, 726, 542, 1318, 843, 1322, 298, 550]. Soft-bottom [10]. soft-coral [1322]. soft-shell [1023]. Soft-walled [1099, 869]. Solander [180]. soldier [1261]. sole [375, 636]. Solea [636]. solely [898]. Solemdal [861]. Solenoceridae [156]. Solenogastres [149]. Solieriaceae [910]. solitaryensis [913]. Solomon [703]. solubilizing [1321]. solution [85, 97]. somatic [815, 257]. some [1079, 534, 774, 687, 15, 1250, 220, 221, 1305, 341, 266, 914, 1171]. Soomai [989, 990]. sorting [1073]. Sótano [401]. Sound [330, 405]. sounds [1237]. sources [880, 1136, 925, 73]. South [387, 968, 454, 91, 407, 1245, 856, 913, 503, 145, 1217, 695, 271, 482, 911, 743, 1026, 204, 25, 1272, 625, 236, 416, 1066, 266, 165, 1169, 380, 227, 1074, 1049, 1094, 237, 1185, 1275, 1352, 1061, 1075, 156, 1358, 703, 1157, 466, 1316, 385, 316, 504, 1234, 1250, 1033, 1141, 400, 451, 30, 924, 737, 554, 1305, 455, 19, 527, 927, 975, 1240, 396, 1219, 1252, 1007, 1263, 702, 641, 456]. south-central [625]. South-eastern [1217, 1066, 165, 1169, 1352, 1305, 1240]. South-west [1316]. South-western [911, 227, 1074, 1049, 1075, 30, 924, 1252]. Southeast [1332, 532, 622, 396, 526]. southeastern [783, 156, 1260, 650, 775, 816, 785, 312, 757, 154, 1192, 1120, 395, 420, 1122, 110, 421]. Southern [733, 237, 469, 1333, 1275, 765, 1063, 977, 1266, 784, 1262, 1366, 597, 926, 209, 280, 602, 460, 760, 418, 920, 117, 661, 988, 13, 1277, 1069, 858, 486, 1253, 914, 181, 741, 740, 779, 1314, 364, 1087, 1010, 928, 663, 517, 401]. southernmost [592]. Southwest [867, 895, 648, 667, 838, 162, 413, 1156, 1248, 256, 1367, 1134, 877, 1080, 529]. Southwestern [291, 991, 1147, 104, 284, 1309, 1283, 772, 833, 1276, 870, 888, 1178, 1224, 1130, 1013, 835, 1168, 1092, 997]. Sowinsky [309]. sp [71, 315, 237, 656, 472, 637, 311, 131, 852, 91, 157, 703, 134, 374, 584, 681, 1238, 1159, 1295, 150, 1096, 111, 1006, 767, 581, 854, 662, 11, 709, 245, 412, 210, 589, 559, 361, 1019, 1080, 693, 727, 719, 1026]. space [1039]. space-time [1039]. spacer [84]. spacer-1 [84]. Spain [228, 1246, 760, 517, 181, 406, 92]. Spanish [1283, 210]. Sparidae [310, 321, 890]. Spatial [1180, 506, 1011, 357, 980, 604, 220, 1371, 922, 411, 531, 228, 1247, 599, 740, 638, 1235, 765, 1265, 1012, 923, 1209, 742, 386, 572, 556, 867, 111, 1022, 341, 1283, 1105, 375, 34, 632, 340, 1058, 646]. spatially [1122]. Spatio [1288, 1251, 480, 1141, 1035, 1063]. Spatio-temporal [1288, 1251, 480, 1141, 1035, 1063]. Spatiotemporal [1056, 858, 1294, 872]. Spawning [368, 197, 840, 180, 444, 1271, 431, 473, 1039, 773, 1131, 462, 224, 1068, 736]. spec [68]. special [515, 478, 279, 1331, 589]. speciation [14]. Species [5, 2, 687, 1089, 215, 721, 952, 204, 26, 75, 352, 689, 1298, 1333, 1188, 1173, 1204, 1099, 1185, 1063, 1079, 1090, 90, 562, 1214, 1180, 671, 25, 67, 286, 331, 1361, 667, 425, 102, 1259, 691, 1077, 122, 852, 346, 454, 992, 1260, 706, 1064, 984, 480, 1359, 1201, 1038, 650, 407, 22, 63, 423, 926, 771, 134, 374, 1193, 1279, 287, 1306, 729, 856, 644, 1230, 233, 739, 89, 942, 29, 1044, 1072, 130, 681, 1154, 460, 1029, 720, 447, 790, 1147, 72, 573, 446, 545, 1172, 207, 162, 206, 478, 66, 23, 757, 1184]. species [793, 879, 956, 1331, 676, 1376, 1250, 299, 1205, 1241, 688, 1115, 1084, 27, 1056, 1290, 400, 778, 935, 104, 618, 30, 443, 1096, 1134, 1016, 1311, 186, 552, 132, 117, 259, 442, 1292, 468, 1046, 135, 142, 802, 163, 622, 1258, 581, 854, 698, 1037, 202, 1174, 288, 1120, 1160, 639, 526, 960, 1253, 464, 116, 943, 1168, 465, 243, 210, 1152, 69, 770, 724, 1307, 1242, 1169, 827, 829, 741, 278, 406, 1313, 301, 541, 591, 1140, 522, 198, 110, 342, 1276, 907, 361, 647, 794, 1252, 1080, 693, 557, 580, 276, 994]. species [256, 727, 1145, 81, 1081, 362, 561, 196, 36, 1257]. Species- [5]. species-specific [1090]. specific [1090, 1341, 938, 99]. specimens [811]. speed [390, 229]. speeds [328]. Speleophria [150]. sperm [499, 788, 1289]. Spermatogenesis [325]. Sphaerechinus [1299]. Sphaerodoridae [442]. Sphaerodoropsis [442]. Sphaeromatidae [1205, 1241]. Sphenoderes [278]. spheres [1071]. Sphyrna [943]. Sphyrnidae [1329]. spicule [619]. spicules [1118]. spider [163, 662, 1120, 870]. spiders [690]. spinax [194, 267, 268]. spinefoot [187]. spinosum [934]. spinosus [786, 786]. spiny [500, 511, 460, 897, 501, 499]. Spionidae [26, 1154, 186]. Spitsbergen [860, 1029]. splendid [830]. sponge [17, 111, 468, 1117, 1082, 1143]. Sponges [441, 560, 1144, 62, 1072, 540, 618, 1305, 1060, 129, 541, 211, 31, 276, 1200]. Spongosorites [900]. spontaneous [1239]. sporadic [502]. sporophytes [318]. spot [1267]. spotted [930, 499, 484, 187]. Spottobrotula [698]. spp [1093, 141, 1375, 35, 563, 1302, 822, 982, 709, 465, 628]. sprat [607]. Sprattus [607]. spread [1324]. spring [1196, 1176, 444, 431, 402, 892, 736, 505, 1151]. spring-spawning [444, 431, 736]. Springer [1191]. spumigena [1015]. Squalus [511]. squamatus [255]. squat [1278, 385]. squid [840, 388, 564, 279, 1167, 953, 87, 1071, 379, 40, 847, 1028, 1151]. squids [945, 704]. St [1046, 1026, 454, 407]. Stable [78, 185, 454, 985, 880, 644, 925, 1116, 621, 467, 826, 579, 847, 570, 1042]. stages [923, 1299, 1113, 1016, 125, 1326, 493, 307, 891, 892]. staghorn [1194]. star [815]. starch [1321]. Starfish [595, 1304]. stars [934, 722, 1236]. starvation [1303, 461]. state [977, 819, 1322]. States [395]. stations [918, 350]. Statolith [279, 1028]. Status [774, 656, 1063, 1254, 1334, 1176, 186, 486, 1207, 366, 862, 1268]. Stauromedusae [362]. Staurozoa [362, 362]. Stehmann [1017, 1018]. stenopodid [533]. steps [1309]. stigmatis [234]. Stimpson [312]. stimulating [594, 327]. stimuli [1248]. stingray [448]. stipulacea [1004]. Stock [1213, 1229, 173, 576, 572, 336, 953, 773, 740, 1268]. stocks [171, 483, 1274, 430]. Stolephorus [834]. stolonal [965]. stolons [965]. Stomach [623, 726, 749, 579, 847]. Stomatopoda [214]. Stomolophus [1199]. stonefish [1239]. stony [1322, 1041]. stony-coral [1322]. storage [1212]. Storm [318, 264]. Storm-generated [318]. story [1041]. stout [40]. strains [80]. Strait [671, 1014, 855, 1164, 51, 365, 1028, 1151]. Stramonita [425, 718, 368]. stranded [749]. strategies [1302, 763, 87, 1052, 724]. strategy [1075, 172, 1278, 1290, 257, 344, 983, 1086, 1314]. stratification [941]. stratospheric [451]. Streblospio [186]. strength [845]. Stress [901, 1232, 458, 1347, 1344, 452, 1336, 910, 1318, 1293, 449, 359, 652, 1083]. stresses [217, 1125]. stressors [251, 1030]. striae [398]. striatus [766]. stridens [1330]. Strike [140]. striped [1104, 751, 236]. strongly [776]. structural [519]. Structure [242, 700, 571, 1049, 1094, 876, 765, 77, 1328, 985, 882, 880, 742, 357, 1135, 598, 694, 810, 1023, 323, 1085, 66, 879, 1285, 900, 1224, 495, 283, 1062, 1022, 1199, 551, 1013, 1066, 1339, 1155, 297, 801, 1317, 181, 857, 624, 1092, 740, 638, 921, 779, 1119, 340, 1100, 850, 836, 641, 1040]. structured [953]. structures [1230, 353, 629, 55, 848]. struma [658]. student [392]. studeri [443]. Studies [418, 958, 285, 208, 959, 794, 848]. Study [1291, 1108, 1094, 1149, 77, 1359, 268, 902, 1034, 1136, 1230, 1044, 933, 530, 963, 322, 277, 614, 1122, 1322, 149, 1269, 355, 254, 646]. stultorum [1283]. Styela [334]. Stygarctidae [1019]. stylasterid [863]. Stylasteridae [964]. stylasterids [964]. stylet [637]. stylet-bearing [637]. Stylopus [1143]. Styraconyx [708]. Sub [537, 419, 479, 112, 538, 10, 463, 224, 728, 561, 548]. sub-Antarctic [419, 479, 561]. sub-arctic [10, 112, 224, 728]. sub-cellular [463]. Sub-Polar [537, 538, 548]. subfamily [1207]. sublittoral [860, 53, 162]. submarine [597, 1159]. Suboxic [106]. subsequent [904, 647]. subsidy [185]. substance [327]. substances [1110]. substrata [441, 816, 376]. substrate [1079, 1064, 155, 895, 1082, 1048, 254]. substrates [680, 289, 385]. subtidal [1011, 1186, 722, 1156, 245]. subtropical [369, 739, 715, 1272, 1116, 1316, 1261, 924, 821, 927, 1042]. success [583, 1203, 509, 1113, 517, 1297]. successful [1330, 694, 624]. successfully [1306]. succession [529, 800]. suction [497]. sudden [168]. sudhakari [1120]. Suez [426, 1203]. suffers [389]. sugar [416]. suggest [840]. suggesting [694, 624]. suggests [1242]. suitable [11]. Sulawesi [1332, 256]. sulculata [867, 895]. sulfide [143]. sulphide [487]. Summer [1314, 1375, 431, 400]. summers [430]. Sundarban [1110]. Sundarbans [1321]. Sunesonia [1275]. Supply [982, 1292]. Supply-side [982]. supported [1043]. supporting [1190]. supposedly [1162]. Suprabenthic [793]. surf [602]. Surface [120, 482, 1177, 1044, 451, 313, 398]. Survey [1123, 1214, 1142, 177, 28, 575, 34]. surveys [736]. Survival [1052, 1088, 199, 1005, 1228, 253, 461, 1226, 1073, 1048, 655, 875, 1083]. survivorship [982, 471]. susceptibility [1341]. suspended [86]. sustainability [902]. Sustainable [415, 734, 1190]. sustaining [1116]. Suzuette [989, 990]. Svalbard [952, 350, 28, 1184, 297, 252, 570]. SW [618, 973, 1133, 753, 1257]. swamps [281]. Sweden [638, 12, 166]. Swedish [588]. Swimbladder [1177, 422, 530]. swimming [783, 328, 410, 229, 313, 239]. swordfish [1104, 313]. swordtip [1028, 1151]. Syllidae [1079, 771, 639, 580]. Syllinae [1079]. Syllis [1079]. Symbiodiniaceae [1336]. Symbiodinium [563, 504]. symbiont [767]. symbionts [316]. symbiotic [691, 374, 504, 8, 1153, 645]. symmetricus [1206]. Sympagella [541]. Sympagic [570]. Sympagic-pelagic-benthic [570]. sympatric [808, 18, 1178, 1316, 446, 1253, 847, 553, 219, 1257]. Symphodus [1376, 768]. Symplocostoma [581]. Synanceia [1239]. Synchiropus [1269]. synchronized [1166]. Synchronous [473]. synchrony [903]. syndrome [1267]. Synechococcus [1375, 80, 205]. Syngnathidae [366]. Synodontidae [849]. synonymies [1020]. synonymized [243]. synonymy [1078, 1140]. Synthesis [439]. Syscenus [131]. System [1191, 320, 598, 1325, 730, 436, 334, 1159, 910, 1124, 1060, 1031, 753, 1231, 149, 437, 569, 645, 1200, 114]. Systematic [1037, 1340, 676]. systematics [1020]. systems [397, 902, 1257].

T [1017, 1018]. T. [553]. table [1005]. tabular [276]. Tachypleus [1187, 778]. tactile [578]. taeniopterus [874]. tagging [501]. Taiwan [1215, 1201]. takanoshimensis [712]. Takaomi [939]. take [1365]. tales [969]. Tamaulipas [1317]. Tanaidacea [1159, 1174, 1069]. tandem [188]. Tanzanian [1328, 1348]. Taranto [889]. Tardigrada [1019]. tardigrade [555]. tardigrades [1214, 776]. target [1307]. tasmanica [1092]. taurus [1208, 396]. taxa [1179]. Taxonomic [869, 70, 1254, 72, 907, 794, 1099, 656, 593, 834, 299, 132, 1174, 891, 647, 885]. Taxonomy [1079, 201, 104, 1161, 874]. Tc1 [343]. Tc1-like [343]. Teacuis [190]. technical [1127]. techniques [984, 1212, 951, 149]. technology [343]. Tehuantepec [345, 1167]. Tehuelche [1274]. Telemark [1165]. teleost [1164, 344]. Teleostei [1202, 508, 991, 1176, 1229, 1302, 410, 314]. Telothelepodidae [1037, 976]. Temnopleuridae [1166]. Temperate [733, 695, 973, 346, 1009, 466, 730, 378, 1301, 108, 55]. Temperature [1196, 999, 520, 1293, 224, 1284, 1266, 574, 839, 458, 1053, 701, 864, 754, 162, 572, 262, 37, 115, 910, 795, 1124, 1054, 811, 510, 904, 229, 313, 494, 983, 1151]. Temperature- [999]. Temperature-related [1196]. temperatures [679, 1347]. Temporal [348, 925, 1024, 341, 1288, 1063, 1180, 1251, 506, 1025, 480, 923, 310, 739, 386, 1141, 1371, 922, 1013, 1035, 1247, 599, 166, 438, 646, 888]. temporary [461]. tenacity [1264]. TEP [773]. Terebelliformia [976, 1037]. Terebratulida [1183]. Teredinidae [1080]. term [918, 966, 1011, 923, 177, 310, 1023, 928, 1318, 199, 916, 917, 927, 1313, 1327, 848]. termination [399]. Terns [1093]. Terrazoanthus [1252]. Terusan [836]. Testing [562, 680, 449]. testudinum [330]. Tetranchyroderma [472, 207]. Tetraodontiformes [786]. Tetraselmis [1052]. Tetrastemma [1238]. Texas [315]. thaca [449]. Thailand [281]. Thalassia [330]. thalassinidean [21]. Thaliacea [1170]. thalli [868]. Thallis [1275]. thallus [1237, 240]. Thaumastodermatidae [207]. thazard [483, 483]. Thecadinium [1032]. Thecostraca [25]. Their [70, 803, 710, 672, 78, 588, 998, 590, 1209, 1230, 720, 1088, 961, 1153, 1217, 955, 1286, 619, 514, 252, 240, 1327]. them [745]. Thematic [138, 179]. Themisto [525]. there [20]. thermal [941, 555, 279, 556, 449, 1125]. these [870]. third [82]. Thirty [958, 959]. Thomson [1092]. thoracic [358]. Thoracica [1220, 645]. thornback [625, 165]. those [843]. threadfin [1300]. threat [1280]. threatened [346, 1292]. Three [688, 483, 276, 918, 26, 1333, 671, 346, 1038, 1198, 1044, 173, 1362, 956, 1376, 1290, 186, 552, 132, 698, 1037, 847, 1313, 931, 1115]. thresher [579]. thrives [685]. throughout [1358, 6]. Thunnus [653, 1131]. Thyasiridae [952]. thynnus [1131]. Thyrsites [1234]. tibicen [1062]. tidal [708, 908, 554, 1001, 753, 184]. tide [1320, 458]. tides [1272, 685]. Tierra [909]. tiger [876, 828, 1208]. tile [241]. Tilesius [325]. time [984, 226, 221, 822, 1039, 1036, 1083]. timescales [1104]. timing [1196, 1262, 1166, 505]. tintinnid [535]. Tioman [1200]. Tisler [1143]. tissue [705, 109, 1040]. Tivela [775]. Tobago [676]. tobianus [1137]. together [936]. tolerance [555, 821, 123]. tolerances [983]. tonggol [653]. tonsa [888, 565]. Tony [146]. tool [1094, 902]. tools [899]. toothed [1354]. top [562]. top-shells [562]. Tore [179]. torpedo [1204]. tortugae [935, 1016]. Tourism [734]. Toxicity [1344]. Toxicology [200]. toxigenic [673]. toxin [463]. Toxopneustidae [1299]. trabeculata [840, 1305]. trace [1361, 1230]. tracer [53]. tracers [570]. Trachurus [1229]. Trachythyone [259]. tracking [328, 677]. trade [888]. trade-off [888]. traditional [680]. traffic [650]. traits [77, 715, 1315, 1070, 1054, 309, 409, 1297]. trammel [663]. Trang [1245]. trans [1214]. trans-isthmus [1214]. transcribed [84]. transcriptional [1125]. transect [20, 122]. Transfer [763]. Transplant [425, 996]. transplantation [1228]. transplants [629]. transport [590, 435]. transportation [901, 1051]. transversus [1301]. traps [814, 1027, 498]. Travin [177]. trawl [1306, 724, 1105]. trawled [44]. trawling [1014, 742, 1069]. trawls [470]. Treadwell [234]. treatment [270]. Tree [184]. Trematoda [1207]. Trematode [914]. trematodes [1171]. tremulus [1175]. trench [1343, 688]. trends [872, 886, 921]. Treptopale [553]. Tri [1162]. Tri-oceanic [1162]. Trichopeltariidae [622]. Trichopeltarion [622]. tricornutum [1145]. Trieste [1041]. triggers [1318]. trigonal [775]. Triple [140]. Tripodfish [326]. triquetra [904]. Trisopterus [182]. Trochidae [562]. Troglocarcinus [691]. Trondheimsfjord [720]. Trophic [1104, 1000, 579, 1253, 103, 1257, 1040, 185, 880, 1334, 900, 1033, 43, 571, 847, 181, 857, 1227]. Trophon [909]. Tropical [187, 1321, 218, 399, 1038, 880, 1136, 742, 739, 905, 1282, 447, 1085, 1002, 774, 1239, 1225, 284, 945, 1309, 124, 551, 1355, 411, 455, 748, 786, 883, 108, 1081, 796, 1111, 1088]. trossulus [591]. trout [814, 615, 616, 613, 612]. truncatus [1122, 700, 1122]. trunculus [1277, 1226, 160, 398]. trutta [814, 616, 613]. Tsawwassen [746]. Tsushima [1028, 1151]. tuberculata [291]. tuberculosa [563]. tubes [1260]. tubulariid [441]. Tubularioidea [76]. tumbil [1213]. tumefaciens [240]. Tumidotheres [1140]. tumour [859]. tuna [483, 1150, 653, 1131]. Tunicata [1324]. tunicate [1054]. Tunisia [941, 424, 626, 290, 322, 718, 368, 1290]. Tunisian [855, 1164, 310, 1218, 1346]. turbot [333]. turf [731]. turf-forming [731]. Turkey [663]. Turkish [165, 1105, 1019]. turnover [158, 739, 1195, 756]. Turritopsoides [1006]. Tursiops [700, 1122]. turtle [329, 1265]. Turtles [596, 346, 1209, 859, 677]. twenty [285]. twenty-first [285]. twinspot [1109]. Two [352, 90, 423, 676, 962, 442, 142, 829, 1276, 561, 952, 237, 1063, 350, 1045, 425, 1093, 650, 597, 112, 1076, 310, 1072, 460, 890, 1316, 897, 312, 446, 879, 900, 1056, 73, 889, 1114, 1311, 1062, 967, 1161, 716, 1192, 969, 854, 285, 1137, 639, 19, 1283, 629, 835, 589, 541, 505, 580, 1145, 81, 1058, 219, 196, 456]. two-phase [1058]. Tylosurus [1348]. type [1149, 1241, 228, 1237, 406, 342, 254]. types [864, 897]. Tyrrhenian [280, 602].

Ubatuba [783]. ucá [1339]. ucá’-crab [1339]. Uchida [362]. Ucides [349, 1339, 1083]. Ucididae [1083]. UK [386]. Ulmaridae [209]. Ultrastructure [59, 1182]. ultrastructures [1044]. Ulva [947, 458, 795]. Ulvales [947]. Umbrina [1224]. un-rooted [1321]. uncertainty [576]. uncommon [233]. uncontaminated [322]. unconventional [861]. underreported [1186]. undersized [347]. understanding [577]. understorey [857]. understudied [1070]. Underwater [1289, 1244]. undescribed [522]. uni [1242]. uni-directional [1242]. unicinctus [487]. uniform [1049]. Unique [695]. United [395]. units [70]. universa [8]. Unquantifiable [576]. Unusual [728]. unvegetated [246]. Unveiling [1070, 1309]. Upeneus [361, 647, 794, 738, 874]. Upogebiidae [1310]. upper [584, 936]. uptake [904]. Upwelling [114, 320, 1038, 193, 659, 702]. Uranoscopiformes [1137]. urbanized [758]. urchin [1157, 1166, 1078, 1299, 1158, 225, 507, 257, 801, 41, 1270]. urchins [648, 725, 1245, 713, 128]. Urechis [487]. Urodasys [637]. Uromunna [856]. Uroteuthis [1028, 1151]. Uruguay [332, 859]. USA [207, 793, 1067, 851].

Use [665, 21, 814, 397, 840, 784, 376, 904, 616, 405, 396, 1122]. used [617, 1219]. useful [1174]. Usefulness [899, 398]. Using [692, 1224, 898, 1071, 435, 1058, 1104, 223, 454, 984, 1135, 283, 822, 811, 1170, 944, 1068, 1213, 12, 414, 437, 438, 553, 799]. utility [720, 395]. utilization [1129, 565]. UV [451].

V [138]. vadis [807]. vaginalis [62]. vagrant [813]. Val [884]. Valenciennes [825, 1254]. valid [706, 790]. vanbrunti [1270]. vancouverensis [412]. vannamei [359, 789, 261, 652]. Vanuatu [738]. var. [673]. varia [124]. Variability [158, 981, 628, 176, 918, 419, 1063, 1011, 1025, 923, 742, 1076, 925, 1376, 400, 1054, 831, 1035, 1050, 858, 228, 836, 1294]. Variable [328, 320, 723, 356]. variables [71, 1286]. variance [1339]. variata [852]. Variation [1113, 188, 451, 822, 322, 124, 876, 5, 1075, 348, 1012, 1338, 496, 51, 621, 316, 1191, 575, 1141, 1357, 221, 778, 938, 1130, 716, 1346, 922, 1013, 527, 927, 421, 740, 248, 553, 184, 646]. Variations [1345, 1288, 1180, 1352, 888, 825, 112, 534, 556, 1248, 1287]. vary [905, 1237]. vegetated [246]. vegetation [225]. veined [825, 1306, 799]. velvet [194, 267, 268]. Veneridae [775, 510, 518]. Venezuela [1349, 98]. venosa [825, 1304, 799]. vent [140, 23]. ventalina [1022]. vents [30, 244, 227]. Venus [518]. Verde [859, 709]. Veretillum [473]. vermicularis [190]. verreauxi [503]. verrucos [518]. verrucosa [1239]. versus [855, 1116, 52, 390, 528, 258, 120, 898]. Vertical [141, 1110, 444, 161, 434, 239, 578, 242, 594, 590, 308, 549, 66, 678, 283, 1124, 737, 1197, 753, 638, 12]. Vetigastropoda [124]. via [1203]. viability [513]. Vibrio [1356, 822]. Victoria [122]. video [1027, 326]. Vietnam [1144]. view [360]. vigilance [390]. Vigo [228]. villosus [573]. violet [709]. Viral [43, 139]. virens [408]. Virginia [793]. virginica [1266]. virus [1267]. Viruses [819]. viscosity [229]. Visual [265]. vitelline [658]. vitellogenin [670]. Viti [1259]. vitro [1356]. vitulina [45]. vivax [59]. vivo [221, 1248]. volcanism [70]. volitans [1087]. volume [596]. volutator [77, 120]. Volutidae [831, 777]. vulgaris [463, 87]. vulgata [223]. vulnerability [375, 883, 578]. vulnerable [611, 1014].

W [1017, 1018, 1096]. Wadden [1057]. Walbaum [773]. Waldo [561]. Wales [856, 503, 92]. wall [12]. walled [869, 1099]. walruses [252]. Walther [1370]. Wang [1096]. warehou [354]. warm [1197]. warmer [575, 571]. warming [1090, 970, 722, 375, 632]. warmingii [1263]. warreni [441]. waste [762]. watching [734]. Water [744, 71, 533, 725, 1284, 611, 762, 1244, 268, 588, 1334, 1177, 452, 545, 900, 451, 132, 1160, 668, 462, 526, 1117, 341, 559, 239, 829, 34, 1128, 931, 1151, 36, 1143, 1040]. waterbodies [1280]. waters [204, 1104, 996, 419, 1308, 1188, 25, 1296, 1361, 1077, 346, 840, 926, 882, 886, 1334, 247, 466, 928, 93, 226, 408, 747, 230, 1302, 283, 1181, 381, 837, 183, 822, 1371, 528, 259, 1348, 1121, 1346, 851, 87, 622, 1131, 809, 1236, 1248, 453, 379, 772, 474, 559, 1247, 599, 623, 98, 570, 363, 600, 529, 878, 848, 362, 561].

Watersipora [782]. wave [905, 19, 591]. waves [685, 20]. weak [1025]. weakfish [751]. web [39, 297]. webs [1093, 573, 529]. wedge [1277, 1226]. weight [1352, 707, 878]. Wells [989, 990]. Weser [457]. West [1321, 1061, 1067, 929, 350, 588, 534, 980, 416, 1050, 470, 1213, 68, 693, 1093, 813, 698, 1242, 196, 1316]. Western [1021, 265, 605, 298, 550, 860, 731, 1175, 984, 703, 1198, 190, 1057, 28, 72, 601, 935, 1062, 1039, 622, 932, 749, 927, 266, 412, 724, 1307, 616, 606, 1171, 911, 1140, 12, 166, 227, 1074, 641, 352, 310, 881, 732, 361, 1095, 1081, 1268, 1049, 966, 1075, 30, 924, 1252]. wetland [554]. Whale [734, 1264, 588, 813, 1233, 283, 832, 865, 50]. whale-fall [588]. Whale-watching [734]. whales [1000, 788, 623, 1289, 600]. whelk [825, 1306, 1304, 195, 799]. where [502]. while [1090]. White [1017, 1018, 674, 601, 321, 354, 1267, 271, 482, 261, 652, 1023, 255]. White-beaked [674]. whitetip [881]. whiting [1352, 434, 523, 431, 433, 1268]. whitish [1283]. whole [1329]. Wide [560, 268, 280, 118, 1179]. wide-eyed [280, 118]. wide-ranging [268]. widely [1072]. widespread [1070]. wild [1266, 336, 1221, 396]. wild-caught [1266]. William [1017, 1018]. wind [920]. Winter [183, 451, 862]. wintering [486, 914]. Wirenia [149]. within [158, 1027, 1229, 730, 316, 676, 501, 117, 1131, 7, 1179, 612]. without [143]. wolffish [247]. wood [1373, 302, 1080]. wood-boring [1080]. woods [385]. Workshop [317]. world [476, 1241, 1017, 1018]. worldwide [276]. worm [987, 816, 487, 24, 367]. Worms [54, 59, 255]. Wrasse [614, 768]. wrasses [1047]. wreckfish [1221]. wurdemanni [219].

Xenobalanus [1169]. Xenostrobus [1012]. Xenoturbella [159]. Xiamen [381]. Xisha [913]. XX [1031]. XX/XY [1031]. XXIX/3 [1077]. XY [1031].

Y-Nauplii [25]. Y. [1096]. Yangtze [850]. year [82, 615, 933, 322, 1041]. Years [179, 958, 558, 54, 178, 170, 889, 1133, 959, 505, 174, 175]. yellow [1291, 1278, 1007, 957, 744, 646]. yezoensis [1108, 240]. yield [1277]. Yogyakarta [1295]. York [1233, 851]. young [784, 655, 875]. Yucatan [1286].

Zanclea [767, 960]. Zaops [1140]. Zapteryx [332]. Zealand [1275, 158, 446, 815, 507, 559]. Zemlya [860]. Zernov [869]. Zeuxo [1174]. Zhejiang [1247]. zhikong [670]. ZhongSha [1061]. Zidona [831, 777]. zinc [1337]. Zoantharia [748, 1252]. zoanthid [563]. Zoanthidea [227]. zoanthids [748]. Zoarcidae [4]. zoeae [752]. Zone [331, 151, 1131, 1288, 1038, 602, 798, 708, 551, 1301, 245, 106, 1323, 133, 548, 547]. zoobenthos [480]. Zoogeographic [1203]. zooplanktivorous [621]. Zooplankton [277, 57, 1051, 1196, 973, 1025, 549, 823, 1124, 1139, 284, 737, 505, 529, 1235, 702]. Zooplankton-mediated [277]. zooxanthellae [563]. Zosime [1115]. Zosimeidae [1115]. Zostera [1003, 2, 246, 746, 707, 906, 1048, 655, 875]. Zosteraceae [655]. zygaena [943]. Zyzzyzus [441].

References

Anonymous:2005:MBR

[1] Anonymous. Marine Biology Research — a new journal. Marine Biology Research, 1(1):1, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510018980.

Fredriksen:2005:SRM

[2] Stein Fredriksen, Hartvig Christie, and Bjarte Andre Sæthre. Species richness in macroalgae and macrofauna assemblages on Fucus serratus L. (Phaeophyceae) and Zostera marina L. (Angiospermae) in Skagerrak, Norway. Marine Biology Research, 1(1):2–19, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000510018953.

Narayanaswamy:2005:EBP

[3] Bhavani E. Narayanaswamy, Brian J. Bett, and John D. Gage. Ecology of bathyal polychaete fauna at an Arctic-Atlantic boundary (Faroe– Shetland Channel, North-east Atlantic). Marine Biology Research, 1 (1):20–32, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510018971.

Nash:2005:AGB

[4] Richard D. M. Nash and Audrey J. Geffen. Aspects of the general biology of Sars’ eelpout, Lycenchelys sarsi (Collett, 1871) (Pisces, Zoarcidae), in Oslofjorden, Norway. Marine Biology Research, 1(1):33–38, 2005. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000510018926.

Arakaki:2005:MSI

[5] S. Arakaki and M. Tokeshi. Microhabitat selection in intertidal gobiid fishes: Species- and size-associated variation. Marine Biology Research, 1(1):39–47, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510019024.

Parsons:2005:PNS

[6] D. G. Parsons. Predators of northern shrimp, Pandalus borealis (Pandalidae), throughout the North Atlantic. Marine Biology Research, 1 (1):48–58, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510018944.

Parsons:2005:IBN

[7] D. G. Parsons. Interactions between northern shrimp, Pandalus borealis (Pandalidae), and its key predators within the eastern Newfoundland and Labrador marine ecosystem. Marine Biology Research, 1 (1):59–67, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510019060.

Kohler-Rink:2005:CMS

[8] Stephanie Köhler-Rink and Michael Kühl. The chemical microenvironment of the symbiotic planktonic foraminifer Orbulina universa. Marine Biology Research, 1(1):68–78, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000510019015.

Schander:2005:WCB

[9] Christoffer Schander and Endre Willassen. What can biological barcoding do for marine biology? Marine Biology Research, 1(1): 79–83, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510018962.

Holte:2005:SBF

[10] Børge Holte, Eivind Oug, and Salve Dahle. Soft-bottom fauna and oxygen minima in sub-arctic north Norwegian marine sill basins. Marine Biology Research, 1(2):85–96, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000510019033.

Oberg:2005:MSC

[11] Jörgen Öberg. Model simulations of conditions suitable for the establishment of Enteromorpha sp. (Chlorophyta) macroalgal mats. Marine Biology Research, 1(2):97–106, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000510019042.

Thomasson:2005:CVD

[12] Maria A. Thomasson and Björn G. Tunberg. Composition and vertical distribution of the motile epifauna on a vertical rock wall in Gullmarsfjorden, western Sweden, using an alternative sampling approach. Marine Biology Research, 1(2):107–117, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000510019051.

Moura:2005:FHC

[13] Teresa Moura, Ivone Figueiredo, Pedro Bordalo-Machado, and Leonel Serrano Gordo. Feeding habits of Chimaera monstrosa L. (Chimaeridae) in relation to its ontogenetic development on the southern Portuguese continental slope. Marine Biology Research, 1(2):118–126, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000510019079.

Aakesson:2005:BIS

[14] Bertil Åkesson and Hannelore Paxton. Biogeography and incipient speciation in Ophryotrocha labronica (Polychaeta, Dorvilleidae). Marine Biology Research, 1(2):127–139, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000510019088.

Karlsbakk:2005:OLH

[15] Egil Karlsbakk. Occurrence of leeches (Hirudinea, Piscicolidae) on some marine fishes in Norway. Marine Biology Research, 1(2):140– 148, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510019105.

Heinz:2005:LOB

[16] Petra Heinz, Emmanuelle Geslin, and Christoph Hemleben. Laboratory observations of benthic foraminiferal cysts. Marine Biology Research, 1 (2):149–159, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510019114.

Hoffmann:2005:ODC

[17] Friederike Hoffmann, Ole Larsen, Hans Tore Rapp, and Ronald Osinga. Oxygen dynamics in choanosomal sponge explants. Marine Biology Research, 1(2):160–163, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000510019006.

Bearzi:2005:DSE

[18] Maddalena Bearzi. Dolphin sympatric ecology. Marine Biology Research, 1(3):165–175, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510019132.

Prowse:2005:PHT

[19] T. A. A. Prowse and A. J. Pile. Phenotypic homogeneity of two intertidal snails across a wave exposure gradient in South Australia. Marine Biology Research, 1(3):176–185, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000510018999.

Witbaard:2005:MFA

[20] Rob Witbaard, Rogier Daan, Maarten Mulder, and Marc Lavaleye. The mollusc fauna along a depth transect in the Faroe Shetland Channel: Is there a relationship with internal waves? Marine Biology Research, 1 (3):186–201, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000510018935.

Abed-Navandi:2005:NET

[21] Daniel Abed-Navandi, Helmut Koller, and Peter C. Dworschak. Nutritional ecology of thalassinidean shrimps constructing burrows with debris chambers: the distribution and use of macronutrients and micronutrients. Marine Biology Research, 1(3):202–215, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000510019123.

Dayrat:2005:ANS

[22] Benoît Dayrat. Advantages of naming species under the PhyloCode: An example of how a new species of Discodorididae (Mollusca, Gastropoda, Euthyneura, Nudibranchia, Doridina) may be named. Marine Biology Research, 1(3):216–232, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000510019141.

Ivanenko:2005:NSA

[23] Viatcheslav N. Ivanenko, Danielle Defaye, and Rony Huys. A new species of Ambilimbus nom. nov., a replacement name for Amphicrossus Huys, 1991 (Copepoda, Cyclopoida, Erebonasteridae), from the Mid-Atlantic Ridge at 36^{o}N (hydrothermal vent site Rainbow). Marine Biology Research, 1(3):233–241, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000510019097.

Struck:2005:SSW

[24] Torsten H. Struck, Günter Purschke, and Kenneth M. Halanych. A scaleless scale worm: Molecular evidence for the phylogenetic placement of Pisione remota (Pisionidae, Annelida). Marine Biology Research, 1 (4):243–253, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500261951.

Belmonte:2005:NCT

[25] Genuario Belmonte. Y-Nauplii (Crustacea, Thecostraca, Facetotecta) from coastal waters of the Salento Peninsula (south eastern Italy, Mediterranean Sea) with descriptions of four new species. Marine Biology Research, 1(4):254–266, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000500202518.

Aguirrezabalaga:2005:SAP

[26] Florencio Aguirrezabalaga and Argiloa Ceberio. Spionidae (Annelida: Polychaeta) from the Capbreton Canyon (Bay of Biscay, NE Atlantic) with descriptions of a new genus and three new species. Marine Biology Research, 1(4):267–280, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000500262066.

Krsinic:2005:BJN

[27] Frano Krsinic. Badijella jalzici — a new genus and species of calanoid copepod (Calanoida, Ridgewayiidae) from an anchialine cave on the Croatian Adriatic coast. Marine Biology Research, 1(4):281– 289, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500262025.

Gooday:2005:MFG

[28] Andrew J. Gooday, Samuel S. Bowser, Tomas Cedhagen, Nils Cornelius, Morten Hald, Sergei Korsun, and Jan Pawlowski. Monothalamous foraminiferans and gromiids (Protista) from western Svalbard: a preliminary survey. Marine Biology Research, 1(4):290–312, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000510019150.

Gad:2005:FDA

[29] Gunnar Gad and Pedro Martínez Arbizu. First description of an Arctic Loricifera — a new Rugiloricus-species from the Laptev Sea. Marine Biology Research, 1(5):313–325, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000500327455.

Lopez-Gonzalez:2005:NSS

[30] Pablo J. López-González, Estefanía Rodríguez, and Michel Segonzac. A new species of sea anemone (Cnidaria: Anthozoa: Actiniaria) from Manus Basin hydrothermal vents, South-western Pacific. Marine Biology Research, 1(5):326–337, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000500380306.

VanSoest:2005:DAS

[31] Rob W. M. Van Soest and Marc S. S. Lavaleye. Diversity and abundance of sponges in bathyal coral reefs of Rockall Bank, NE Atlantic, from boxcore samples. Marine Biology Research, 1(5):338–349, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000500380322.

Malchus:2005:SHM

[32] Nikolaus Malchus and Anders Warén. Shell and hinge morphology of juvenile Limopsis (Bivalvia: Arcoida) — implications for limopsid evolution. Marine Biology Research, 1(5):350–364, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000500384100.

Tonnesson:2005:GIO

[33] Kajsa Tönnesson, Marie Maar, Cristian Vargas, Eva Friis Møller, Suree Satapoomin, Sultana Zervoudaki, Epaminondas Christou, Antonia Giannakourou, Anne Sell, Jens Kjerulf Petersen, Torkel Gissel Nielsen, and Peter Tiselius. Grazing impact of Oikopleura dioica and copepods on an autumn plankton community. Marine Biology Research, 1 (5):365–373, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500396435.

Stelzenmuller:2005:ESS

[34] Vanessa Stelzenmüller, Siegfried Ehrich, and Gerd-Peter Zauke. Effects of survey scale and water depth on the assessment of spatial distribution patterns of selected fish in the northern North Sea showing different levels of aggregation. Marine Biology Research, 1(6):375–387, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000500361009.

Garcia-Martinez:2005:MCS

[35] Micaela García-Martínez, John Kuo, Kieryn Kilminster, Di Walker, Ramón Rosselló-Mora, and Carlos M. Duarte. Microbial colonization in the seagrass Posidonia spp. roots. Marine Biology Research, 1(6): 388–395, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500443419.

Zuschin:2005:DPB

[36] Martin Zuschin and P. Graham Oliver. Diversity patterns of bivalves in a coral dominated shallow-water bay in the northern Red Sea — high species richness on a local scale. Marine Biology Research, 1 (6):396–410, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500456262.

Jordan:2005:DTP

[37] Anders D. Jordan, Maria F. Steinhausen, Neill A. Herbert, Barbara Grisdale-Helland, Ståle J. Helland, and John F. Steffensen. Does temperature preference relate to the anaerobic capacity of Atlantic cod (Gadus morhua L.) with different haemoglobin phenotype? Marine Biology Research, 1(6):411–416, 2005. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000500459795.

Nielsen:2005:BR

[38] Torkel Gissel Nielsen. Book review. Marine Biology Research, 1 (6):417, 2005. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500420615.

Moritz:2006:PRM

[39] C. M. Moritz, D. Montagnes, J. H. Carleton, D. Wilson, and A. D. McKinnon. The potential role of microzooplankton in a northwestern Australian pelagic food web. Marine Biology Research, 2(1):1–13, 2006. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000500497159.

Rosa:2006:DAB

[40] R. Rosa, J. Pereira, P. R. Costa, and M. L. Nunes. Distribution, abundance, biology and biochemistry of the stout bobtail squid Rossia macrosoma from the Portuguese coast. Marine Biology Research, 2 (1):14–23, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500519846.

Sjotun:2006:CEC

[41] Kjersti Sjøtun, Hartvig Christie, and Jan Helge Fosså. The combined effect of canopy shading and sea urchin grazing on recruitment in kelp forest (Laminaria hyperborea). Marine Biology Research, 2 (1):24–32, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500537418.

Fenchel:2006:MBB

[42] Tom Fenchel and Per Juel Hansen. Motile behaviour of the bloomforming ciliate Mesodinium rubrum. Marine Biology Research, 2(1): 33–40, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600571044.

Middelboe:2006:VAA

[43] Mathias Middelboe and Ronnie N. Glud. Viral activity along a trophic gradient in continental margin sediments off central Chile. Marine Biology Research, 2(1):41–51, 2006. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000600620650.

Eriksson:2006:DCN

[44] Susanne P. Eriksson. Differences in the condition of Norway lobsters (Nephrops norvegicus (L.)) from trawled and creeled fishing areas. Marine Biology Research, 2(1):52–58, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000600623803.

Hauksson:2006:GRI

[45] Erlingur Hauksson. Growth and reproduction in the Icelandic common seal (Phoca vitulina l., 1758). Marine Biology Research, 2(1): 59–73, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600650038.

Uiblein:2006:BR

[46] Franz Uiblein. Book review. Marine Biology Research, 2(1):74– 75, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500495880.

Sundby:2006:BR

[47] Svein Sundby. Book review. Marine Biology Research, 2(1):75– 76, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500495369.

Hansen:2006:BR

[48] Per Juel Hansen. Book review. Marine Biology Research, 2(1): 76, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600649931.

Epelbaum:2006:FBF

[49] A. Epelbaum and R. Borisov. Feeding behaviour and functional morphology of the feeding appendages of red king crab Paralithodes camtschaticus larvae. Marine Biology Research, 2(2):77–88, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000600672529.

Sivertsen:2006:PPB

[50] Steffen P. Sivertsen, Torstein Pedersen, Ulf Lindstrøm, and Tore Haug. Prey partitioning between cod (Gadus morhua) and minke whale (Balaenoptera acutorostrata) in the Barents Sea. Marine Biology Research, 2(2):89–99, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600670952.

Guerra-Garcia:2006:CPL

[51] José M. Guerra-García, Susana Redondo-Gómez, Águeda G. Espina, Jesús M. Castillo, Teresa Luque, J. Carlos García-Gómez, and M. Enrique Figueroa. Caprella penantis Leach, 1814 and Caprella dilatata Kroyer, 1843 (Crustacea: Amphipoda) from the Strait of Gibraltar: a molecular approach to explore intra- and interspecific variation. Marine Biology Research, 2(2):100–108, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000600672511.

Koski:2006:PBF

[52] Marja Koski and Christian Wexels Riser. Post-bloom feeding of Calanus finmarchicus copepodites: Selection for autotrophic versus heterotrophic prey. Marine Biology Research, 2(2):109–119, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000600684367.

Buhring:2006:EBA

[53] S. I. Bühring, S. Ehrenhauss, A. Kamp, L. Moodley, and U. Witte. Enhanced benthic activity in sandy sublittoral sediments: Evidence from ^{13}C tracer experiments. Marine Biology Research, 2(2):120– 129, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600678773.

Boggemann:2006:WMM

[54] Markus Böggemann. Worms that might be 300 million years old. Marine Biology Research, 2(2):130–135, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000600696304.

Wilhelmsson:2006:EHR

[55] Dan Wilhelmsson, Saleh A. S. Yahya, and Marcus C. Öhman. Effects of high-relief structures on cold temperate fish assemblages: a field experiment. Marine Biology Research, 2(2):136–147, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000600684359.

Boggemann:2006:BR

[56] Markus Böggemann. Book review. Marine Biology Research, 2 (2):148, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000500396393.

Zervoudaki:2006:ZDD

[57] S. Zervoudaki, T. G. Nielsen, E. D. Christou, and I. Siokou-Frangou. Zooplankton distribution and diversity in a frontal area of the Aegean Sea. Marine Biology Research, 2(3):149–168, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000600702037.

Hansen:2006:BFC

[58] Per Juel Hansen and Tom Fenchel. The bloom-forming ciliate Mesodinium rubrum harbours a single permanent endosymbiont. Marine Biology Research, 2(3):169–177, 2006. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000600719577.

Leander:2006:UAS

[59] Brian S. Leander. Ultrastructure of the archigregarine Selenidium vivax (Apicomplexa) — a dynamic parasite of sipunculid worms (host: Phascolosoma agassizii). Marine Biology Research, 2(3):178–190, 2006. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000600724395.

Fock:2006:FLP

[60] Heino Fock and Hans-Christian John. Fish larval patterns across the Reykjanes Ridge. Marine Biology Research, 2(3):191–199, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000600784159.

Stepanjants:2006:RBC

[61] S. D. Stepanjants, G. Cortese, S. B. Kruglikova, and K. R. Bjørklund. A review of bipolarity concepts: History and examples from Radiolaria and Medusozoa (Cnidaria). Marine Biology Research, 2(3): 200–241, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600781767.

Duckworth:2006:REC

[62] A. R. Duckworth, W. M. Brück, K. E. Janda, T. P. Pitts, and P. J. McCarthy. Retention efficiencies of the coral reef sponges Aplysina lacunosa, Callyspongia vaginalis and Niphates digitalis determined by Coulter counter and plate culture analysis. Marine Biology Research, 2 (4):243–248, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600781775.

Debes:2006:SAR

[63] Høgni Hammershaimb Debes and Kirstin Eliasen. Seasonal abundance, reproduction and development of four key copepod species on the Faroe shelf. Marine Biology Research, 2(4):249–259, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000600798787.

Fossheim:2006:FAB

[64] Maria Fossheim, Einar M. Nilssen, and Michaela Aschan. Fish assemblages in the Barents Sea. Marine Biology Research, 2(4):260– 269, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600815698.

Mouritsen:2006:ESC

[65] Kim N. Mouritsen and Tomas Jensen. The effect of Sacculina carcini infections on the fouling, burying behaviour and condition of the shore crab, Carcinus maenas. Marine Biology Research, 2(4):270– 275, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600874752.

Irigoien:2006:CPS

[66] Xabier Irigoien and Roger P. Harris. Comparative population structure, abundance and vertical distribution of six copepod species in the North Atlantic: Evidence for intraguild predation? Marine Biology Research, 2(4):276–290, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600865321.

Biju:2006:MCM

[67] A. Biju, S. U. Panampunnayil, and M. P. Prabhakaran. Mysidacea (Crustacea) from the Minicoy lagoon (Lakshadweep, India) with description of a new species of Anisomysis. Marine Biology Research, 2 (4):291–295, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600717431.

Sterrer:2006:CDG

[68] Wolfgang Sterrer and Martin V. Sørensen. Chirognathia dracula gen. et spec. nov. (Gnathostomulida) from the west coast of North America. Marine Biology Research, 2(4):296–302, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000600895013.

Schander:2006:FHN

[69] Christoffer Schander, Amélie H. Scheltema, and Dimitry L. Ivanov. Falcidens halanychi, a new species of Chaetodermomorpha (=Caudofoveata) (Mollusca) from the northwest Atlantic Ocean. Marine Biology Research, 2(5):303–315, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600798795.

Bodergat:2006:TUC

[70] Anne-Marie Bodergat, Kimihiko Oki, Michel Rio, and Michel Tabarant. Taxonomic units, civilization, volcanism: Their influence on the chemical composition of ostracod carapaces (Kagoshima Bay, Kyushu Island, Japan). Marine Biology Research, 2(5):316–325, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000600801300.

Abdo:2006:MEV

[71] D. A. Abdo, C. N. Battershill, and E. S. Harvey. Manipulation of environmental variables and the effect on the growth of Haliclona sp.: Implications for open-water aquaculture. Marine Biology Research, 2 (5):326–332, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600820946.

Hoisaeter:2006:TBN

[72] Tore Høisæter and Tor J. Samuelsen. Taxonomic and biological notes on a species of Iphitime (Polychaeta, Eunicida) associated with Pagurus prideaux from western Norway. Marine Biology Research, 2(5): 333–354, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600867277.

Lepoint:2006:FST

[73] G. Lepoint, A.-S. Cox, P. Dauby, M. Poulicek, and S. Gobert. Food sources of two detritivore amphipods associated with the seagrass Posidonia oceanica leaf litter. Marine Biology Research, 2(5): 355–365, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600962797.

Norderhaug:2006:IPC

[74] K. M. Norderhaug, K. Nygaard, and S. Fredriksen. Importance of phlorotannin content and C : N ratio of Laminaria hyperborea in determining its palatability as food for consumers. Marine Biology Research, 2(6):367–371, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000600962789.

Allcock:2006:RDS

[75] A. L. Allcock, J. M. Strugnell, H. Ruggiero, and M. A. Collins. Redescription of the deep-sea octopod Benthoctopus normani (Massy 1907) and a description of a new species from the Northeast Atlantic. Marine Biology Research, 2(6):372–387, 2006. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000600973315.

Stepanjants:2006:MGR

[76] S. D. Stepanjants and K. N. Kosobokova. Medusae of the genus Rhabdoon (Hydrozoa: Anthomedusae: Tubularioidea) in the Arctic Ocean. Marine Biology Research, 2(6):388–397, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000601012402.

Bostrom:2006:IRS

[77] Christoffer Boström, Maria Lastuniemi, and Erik Bonsdorff. Infaunal responses to seagrass habitat structure: a study of life-history traits and population dynamics of Corophium volutator (Pallas). Marine Biology Research, 2(6):398–410, 2006. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000601021692.

Angerbjorn:2006:SIA

[78] Anders Angerbjörn, Patrik Börjesson, and Katrin Brandberg. Stable isotope analysis of harbour porpoises and their prey from the Baltic and Kattegat/Skagerrak Seas. Marine Biology Research, 2(6): 411–419, 2006. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000601023896.

Pascual:2006:GLH

[79] S. Pascual, F. Rocha, and A. Guerra. Gross lesions in the hubb octopus Octopus hubbsorum. Marine Biology Research, 2(6):420–423, 2006. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000601050451.

Lu:2006:PCM

[80] Zhiqiang Lu, Nianzhi Jiao, and Huiyong Zhang. Physiological changes in marine picocyanobacterial Synechococcus strains exposed to elevated CO_{2} partial pressure. Marine Biology Research, 2(6):424–430, 2006. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000601055419.

Worsaae:2006:DTN

[81] Katrine Worsaae and Wolfgang Sterrer. Description of two new interstitial species of Psammodrilidae (Annelida) from Bermuda. Marine Biology Research, 2(6):431–445, 2006. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000600973000.

Fenchel:2007:MBRa

[82] Tom Fenchel and Franz Uiblein. Marine Biology Research — the third year. Marine Biology Research, 3(1):1–2, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000701225896.

Paxton:2007:ROP

[83] Hannelore Paxton and Bertil Åkesson. Redescription of Ophryotrocha puerilis and O. labronica (Annelida, Dorvilleidae). Marine Biology Research, 3(1):3–19, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000601024373.

Hu:2007:MIC

[84] Zimin Hu, Yingjun He, Peng Xia, and Delin Duan. Molecular identification of Chinese cultivated Porphyra (Bangiaceae, Rhodophyta) based on the rDNA internal transcribed spacer-1 sequence and random amplified polymorphic DNA markers. Marine Biology Research, 3(1):20–28, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000601070863.

Konigson:2007:SFN

[85] Sara Königson, Malin Hemmingsson, Sven-Gunnar Lunneryd, and Karl Lundström. Seals and fyke nets: An investigation of the problem and its possible solution. Marine Biology Research, 3(1):29–36, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000601072596.

Frantzen:2007:RBM

[86] Sylvia Frantzen. Recruitment of blue mussels, Mytilus edulis L., on suspended collectors in Finnmark, North Norway (70–71^{o}N). Marine Biology Research, 3(1):37–48, 2007. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000601182627.

Moreno:2007:GSS

[87] Ana Moreno, Manuela Azevedo, João Pereira, and Graham J. Pierce. Growth strategies in the squid Loligo vulgaris from Portuguese waters. Marine Biology Research, 3(1):49–59, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000601129115.

Thomsen:2007:AMD

[88] Mads S. Thomsen, Thomas Wernberg, Peter Stææhr, Dorte Krause-Jensen, Nils Risgaard-Petersen, and Brian R. Silliman. Alien macroalgae in Denmark — a broad-scale national perspective. Marine Biology Research, 3(2):61–72, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000701213413.

Fosshagen:2007:NSE

[89] Audun Fosshagen and Thomas M. Iliffe. New species of epacteriscids (Copepoda, Calanoida) from anchialine caves in the Caicos Islands and the Bahamas. Marine Biology Research, 3(2):73–92, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000701274571.

Bandera:2007:TNA

[90] M. E. Bandera, M. Conradi, and P. J. López-González. Two new asterocherid species (Siphonostomatoida: Asterocheridae) from Madeira and the Canary Islands (eastern Atlantic). Marine Biology Research, 3(2):93–108, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701274589.

Carvalho:2007:PSN

[91] Mariana S. Carvalho, Ruth Desqueyroux-Faúndez, and Eduardo Hajdu. Phakellia sur sp. nov. (Demospongiae, Halichondrida, Axinellidae) from the lower slope off Cape Horn (South America), with a revision of the genus. Marine Biology Research, 3(2):109–116, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000701276063.

Sundberg:2007:GDR

[92] Per Sundberg and Malin Strand. Genetics do not reflect habitat differences in Riseriellus occultus (Heteronemertea, Nemertea) from Spain and Wales. Marine Biology Research, 3(2):117–122, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000601182619.

Haug:2007:DHS

[93] Tore Haug, Kjell Tormod Nilssen, Lotta Lindblom, and Ulf Lindstrøm. Diets of hooded seals (Cystophora cristata) in coastal waters and drift ice waters along the east coast of Greenland. Marine Biology Research, 3(3):123–133, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701358531.

Middelboe:2007:DEP

[94] Anne Lise Middelboe and Per Juel Hansen. Direct effects of pH and inorganic carbon on macroalgal photosynthesis and growth. Marine Biology Research, 3(3):134–144, 2007. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000701320556.

Hu:2007:DCG

[95] Zimin Hu, Alan T. Critchley, Tianxiang Gao, Xiaoqi Zeng, Steve L. Morrell, and Delin Duan. Delineation of Chondrus (Gigartinales, Florideophyceae) in China and the origin of C. crispus inferred from molecular data. Marine Biology Research, 3(3):145–154, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000701335679.

Sukhotin:2007:GLM

[96] A. A. Sukhotin, P. P. Strelkov, N. V. Maximovich, and H. Hummel. Growth and longevity of Mytilus edulis (L.) from northeast Europe. Marine Biology Research, 3(3):155–167, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000701364869.

Wares:2007:AFL

[97] J. P. Wares and A. M. H. Blakeslee. Amplified fragment length polymorphism data provide a poor solution to the Littorina littorea puzzle. Marine Biology Research, 3(3):168–174, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000701320572.

Sorbe:2007:NRM

[98] Jean Claude Sorbe, Alberto Martin, and Yusbelly Diaz. New records of Mysida (Crustacea: Peracarida) from shallow waters of the Caribbean coast of Venezuela. Marine Biology Research, 3(3):175–181, 2007. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000701397513.

Springer:2007:ESH

[99] Stevan A. Springer and Daniel D. Heath. Environment-specific heterozygote deficiency and developmental instability in hybrid Mytilus. Marine Biology Research, 3(3):182–187, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701320564.

Fenchel:2007:MBRb

[100] Tom Fenchel and Franz Uiblein. Marine Biology Research — ocean exploration and marine ecosystems. Marine Biology Research, 3(4): 189–190, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701496422.

Jones:2007:DRC

[101] Daniel O. B. Jones, Brian J. Bett, and Paul A. Tyler. Depthrelated changes in the Arctic epibenthic megafaunal assemblages of Kangerdlugssuaq, East Greenland. Marine Biology Research, 3(4): 191–204, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701455287.

Brix:2007:FNS

[102] Saskia Brix. Four new species of Desmosomatidae Sars, 1897 (Crustacea: Isopoda) from the deep sea of the Angola Basin. Marine Biology Research, 3(4):205–230, 2007. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000701490987.

Solmundsson:2007:TEG

[103] Jon Solmundsson. Trophic ecology of Greenland halibut (Reinhardtius hippoglossoides) on the Icelandic continental shelf and slope. Marine Biology Research, 3(4):231–242, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701477513.

Lopes:2007:TEP

[104] Daniela A. Lopes, Eduardo Hajdu, and Henry M. Reiswig. Taxonomy of Euretidae (Porifera, Hexactinellida, Hexactinosida) of Campos Basin, southwestern Atlantic, with a description of a new species. Marine Biology Research, 3(4):243–255, 2007. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000701468777.

Gil:2007:RBF

[105] Damian G. Gil and Héctor E. Zaixso. The relation between feeding and reproduction in Anasterias minuta (Asteroidea: Forcipulata). Marine Biology Research, 3(4):256–264, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701472035.

Sorokin:2007:SZB

[106] Yuri I. Sorokin. Suboxic zone in the Black Sea: real fact or analytical artefact. Marine Biology Research, 3(4):265–271, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000701358549.

Youngbluth:2007:BFP

[107] Marsh J. Youngbluth. Bibliography of Francesc Pagès (1986–2007). Marine Biology Research, 3(4):272–274, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000701496430.

Uiblein:2007:GMI

[108] Franz Uiblein. Goatfishes (Mullidae) as indicators in tropical and temperate coastal habitat monitoring and management. Marine Biology Research, 3(5):275–288, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000701687129.

Skage:2007:DFF

[109] Morten Skage and Christoffer Schander. DNA from formalin-fixed tissue: extraction or repair? That is the question. Marine Biology Research, 3 (5):289–295, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701473942.

Trilles:2007:OPC

[110] Jean-Paul Trilles. Olencira praegustator (Crustacea, Isopoda, Cymothoidae) parasitic on Brevoortia species (Pisces, Clupeidae) from the southeastern coasts of North America: review and re-description. Marine Biology Research, 3(5):296–311, 2007. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000701543439.

Luter:2007:CAM

[111] Heidi M. Luter, Alan R. Duckworth, and Craig Syms. Cytotoxic and antimicrobial activity of the sponge Iotrochota sp. as a function of size and spatial competitors. Marine Biology Research, 3(5):312–318, 2007. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000701635102.

Eilertsen:2007:PTS

[112] Hans C. Eilertsen and Sylvia Frantzen. Phytoplankton from two sub-Arctic fjords in northern Norway 2002–2004: I. Seasonal variations in chlorophyll a and bloom dynamics. Marine Biology Research, 3(5): 319–332, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701632877.

Golani:2007:IRC

[113] Daniel Golani, Ruth Reef-Motro, Shiri Ekshtein, Albert Baranes, and Ariel Diamant. Ichthyofauna of the rocky coastal littoral of the Israeli Mediterranean, with reference to the paucity of Red Sea (Lessepsian) migrants in this habitat. Marine Biology Research, 3(5):333–341, 2007. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000701635110.

Aspetsberger:2007:IBR

[114] Fanni Aspetsberger, Matthias Zabel, Timothy Ferdelman, Ulrich Struck, Andreas Mackensen, Astrid Ahke, and Ursula Witte. Instantaneous benthic response to different organic matter quality: In situ experiments in the Benguela Upwelling System. Marine Biology Research, 3 (5):342–356, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701632885.

LeMoullac:2007:MAO

[115] Gilles Le Moullac, Isabelle Quéau, Pierrick Le Souchu, Stéphane Pouvreau, Jeanne Moal, Jean René Le Coz, and Jean François Samain. Metabolic adjustments in the oyster Crassostrea gigas according to oxygen level and temperature. Marine Biology Research, 3(5):357– 366, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701635128.

Rogacheva:2007:RAG

[116] Antonina V. Rogacheva. Revision of the Arctic group of species of the family Elpidiidae (Elasipodida, Holothuroidea). Marine Biology Research, 3(6):367–396, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000701781880.

Matamoros:2007:NSW

[117] Lisa Matamoros, Seray Yildiz, and Christer Erséus. A new species within the genus Marionina (Enchytraeidae: Annelida: Clitellata) from the southern Black Sea. Marine Biology Research, 3(6):397–402, 2007. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000701694844.

Morato:2007:GRR

[118] T. Morato, P. Afonso, N. Carvalho, P. Lourinho, R. S. Santos, H. M. Krug, and R. D. M. Nash. Growth, reproduction and recruitment patterns of the wide-eyed flounder, Bothus podas Delaroche (Pisces: Bothidae), from the Azores. Marine Biology Research, 3(6): 403–411, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701712331.

Heggoy:2007:PAG

[119] Kristin Kvalø Heggøy, Christoffer Schander, and Bertil Åkesson. The phylogeny of the annelid genus Ophryotrocha (Dorvilleidae). Marine Biology Research, 3(6):412–420, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701695361.

Riisgaard:2007:SDF

[120] Hans Ulrik Riisgård and Peer Schotge. Surface deposit feeding versus filter feeding in the amphipod Corophium volutator. Marine Biology Research, 3(6):421–427, 2007. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000701696302.

Zeng:2007:CDP

[121] Yonghui Zeng, Xihan Chen, and Nianzhi Jiao. Contrasting diversity pattern of Cytophaga–Flavobacteria in the estuarine and open ocean regions of the East China Sea. Marine Biology Research, 3(6):428–437, 2007. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000701696278.

Cano:2007:AAP

[122] Esperanza Cano and Pablo J. López-González. Ammotheidae (Arthropoda: Pycnogonidae) collected during the Victoria Land Transect cruise 2004 (Ross Sea, Antarctica), with a description of a new species of Ammothea. Marine Biology Research, 3(6):438–445, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000701696294.

Nielsen:2007:DII

[123] Lasse Tor Nielsen, Nina Lundholm, and Per Juel Hansen. Does irradiance influence the tolerance of marine phytoplankton to high pH? Marine Biology Research, 3(6):446–453, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701711820.

Najmudeen:2007:VBC

[124] T. M. Najmudeen. Variation in biochemical composition during gonad maturation of the tropical abalone Haliotis varia Linnaeus 1758 (Vetigastropoda: Haliotidae). Marine Biology Research, 3(6):454–461, 2007. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000701696252.

Marquez:2007:FRE

[125] Lorenzo Márquez, Mónica Caballos, and Pedro Domingues. Functional response of early stages of the cuttlefish Sepia officinalis preying on the mysid Mesopodopsis slabberi. Marine Biology Research, 3 (6):462–467, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701751479.

Katsanevakis:2007:FRA

[126] Stelios Katsanevakis and Maria Thessalou-Legaki. First record of Alicia mirabilis (Anthozoa: Actiniaria) from the Aegean Sea and density assessment with distance sampling in a site of high abundance. Marine Biology Research, 3(6):468–472, 2007. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701695379.

Gebruk:2008:BFN

[127] Andrey V. Gebruk, Tom Fenchel, and Franz Uiblein. Benthic fauna of the northern Mid-Atlantic Ridge: results of the MAR-ECO expedition. Marine Biology Research, 4(1–2):1–2, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000701870881.

Mironov:2008:PSU

[128] Alexander N. Mironov. Pourtalesiid sea urchins (Echinodermata: Echinoidea) of the northern Mid-Atlantic Ridge. Marine Biology Research, 4(1–2):3–24, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701847863.

Tabachnick:2008:GSP

[129] Konstantin R. Tabachnick and Allen G. Collins. Glass sponges (Porifera, Hexactinellida) of the northern Mid-Atlantic Ridge. Marine Biology Research, 4(1–2):25–47, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000701847848.

Gebruk:2008:HHE

[130] Andrey V. Gebruk. Holothurians (Holothuroidea, Echinodermata) of the northern Mid-Atlantic Ridge collected by the G. O. Sars MAR-ECO expedition with descriptions of four new species. Marine Biology Research, 4(1–2):48–60, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701842898.

Brandt:2008:DAS

[131] Angelika Brandt and Hans Georg Andres. Description of Aega sarsae sp. nov. and redescription of Syscenus atlanticus Kononenko, 1988 (Crustacea, Isopoda, Aegidae) from the Mid-Atlantic Ridge. Marine Biology Research, 4(1–2):61–75, 2008. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000701842906.

Martynov:2008:DWO

[132] Alexander V. Martynov and Nina M. Litvinova. Deep-water Ophiuroidea of the northern Atlantic with descriptions of three new species and taxonomic remarks on certain genera and species. Marine Biology Research, 4(1–2):76–111, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701840066.

Molodtsova:2008:ANM

[133] Tina N. Molodtsova, Nadezhda P. Sanamyan, and Natalya B. Keller. Anthozoa from the northern Mid-Atlantic Ridge and Charlie-Gibbs Fracture Zone. Marine Biology Research, 4(1–2):112–130, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000701821744.

Dilman:2008:AFN

[134] Anna B. Dilman. Asteroid fauna of the northern Mid-Atlantic Ridge with description of a new species Hymenasterides mironovi sp. nov. Marine Biology Research, 4(1–2):131–151, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701821736.

Murina:2008:NRE

[135] Vantsetti V. Murina. New records of Echiura and Sipuncula in the North Atlantic Ocean, with the description of a new species of Jacobia. Marine Biology Research, 4(1–2):152–156, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701864553.

Zezina:2008:NBN

[136] Olga N. Zezina. Notes on brachiopods from the North Atlantic. Marine Biology Research, 4(1–2):157–159, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000701819011.

Bergstad:2008:AMS

[137] Odd Aksel Bergstad and Andrey V. Gebruk. Approach and methods for sampling of benthic fauna on the 2004 MAR-ECO expedition to the Mid-Atlantic Ridge. Marine Biology Research, 4(1–2):160–163, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000701851477.

Anonymous:2008:TIN

[138] Anonymous. Thematic issue nr. 1: Benthic fauna of the Mid-Atlantic Ridge: Results of the MAR-ECO Expedition, Andrey V. Gebruk (Guest Editor). Marine Biology Research, 4(1–2):??, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802009439.

Siem-Jorgensen:2008:VDC

[139] Maria Siem-Jørgensen, Ronnie N. Glud, and Mathias Middelboe. Viral dynamics in a coastal sediment: seasonal pattern, controlling factors and relations to the pelagic-benthic coupling. Marine Biology Research, 4(3):165–179, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000801888718.

Corbera:2008:NDS

[140] Jordi Corbera, Michel Segonzac, and Marina R. Cunha. A new deep-sea genus of Nannastacidae (Crustacea, Cumacea) from the Lucky Strike hydrothermal vent field (Azores Triple Junction, Mid-Atlantic Ridge).

Marine Biology Research, 4(3):180–192, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000801898576.

Daase:2008:VDC

[141] Malin Daase, Ketil Eiane, Dag L. Aksnes, and Daniel Vogedes. Vertical distribution of Calanus spp. and Metridia longa at four Arctic locations. Marine Biology Research, 4(3):193–207, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000801907948.

Nair:2008:TNS

[142] Vijayalakshmi R. Nair, Saramma U. Panampunnayil, Honey U. K. Pillai, and Raghavan Gireesh. Two new species of chaetognatha from the Andaman Sea, Indian Ocean. Marine Biology Research, 4(3): 208–214, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000701696260.

Vismann:2008:RHS

[143] Bent Vismann and Lars Hagerman. Recovery from hypoxia with and without sulfide in Saduria entomon: potassium, ATP and behavior. Marine Biology Research, 4(3):215–223, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000701877282.

Miranda:2008:AGN

[144] Richard M. Miranda, Katsuaki Fujinaga, and Shigeru Nakao. Age and growth of Neptunea arthritica estimated from growth marks in the operculum. Marine Biology Research, 4(3):224–235, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000701881706.

Pereira:2008:FRA

[145] Alexandre M. Pereira, Marta M. Rufino, and Miguel B. Gaspar. First record of Albunea carabus (Decapoda, Anomura, Hippidea) from the Algarve coast, South Portugal. Marine Biology Research, 4(3): 236–238, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000801907955.

Anonymous:2008:BRS

[146] Anonymous. Book review: The silent deep: the discovery, ecology and conservation of the deep sea, by Tony Koslow. Marine Biology Research, 4(3):239, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000801891670.

Fenchel:2008:HSS

[147] Tom Fenchel and Franz Uiblein. Heterogeneity on a small scale. Marine Biology Research, 4(4):241–242, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000802123164.

Glud:2008:ODM

[148] Ronnie N. Glud. Oxygen dynamics of marine sediments. Marine Biology Research, 4(4):243–289, 2008. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000801888726.

Todt:2008:NSB

[149] Christiane Todt, Thomas Büchinger, and Andreas Wanninger. The nervous system of the basal mollusk Wirenia argentea (Solenogastres): a study employing immunocytochemical and 3D reconstruction techniques. Marine Biology Research, 4(4):290–303, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000801922830.

Krsinic:2008:DSM

[150] Frano Krsinić. Description of Speleophria mestrovi sp. nov., a new copepod (Misophrioida) from an anchialine cave in the Adriatic Sea. Marine Biology Research, 4(4):304–312, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000801930072.

Busch:2008:PDS

[151] Markus Wilhelm Busch, Sven Klimpel, Tracey Sutton, and Uwe Piatkowski. Parasites of the deep-sea smelt Bathylagus euryops (Argentiniformes: Microstomatidae) from the Charlie-Gibbs Fracture Zone (CGFZ). Marine Biology Research, 4(4):313–317, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000801907963.

Anonymous:2008:AMG

[152] Anonymous. Arc marine — GIS for a blue planet. Marine Biology Research, 4(4):318–319, 2008. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000802126712.

Sciberras:2008:BII

[153] Marija Sciberras and Patrick J. Schembri. Biology and interspecific interactions of the alien crab Percnon gibbesi in the Maltese Islands. Marine Biology Research, 4(5):321–332, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000801964923.

Lee:2008:FLH

[154] Won Je Lee. Free-living heterotrophic euglenids from marine sediments of the Gippsland Basin, southeastern Australia. Marine Biology Research, 4(5):333–349, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802040137.

Joschko:2008:CAH

[155] Tanja J. Joschko, Bela H. Buck, Lars Gutow, and Alexander Schröder. Colonization of an artificial hard substrate by Mytilus edulis in the German Bight. Marine Biology Research, 4(5):350–360, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000801947043.

Castilho:2008:RRS

[156] Antonio Leão Castilho, Rogério Caetano Da Costa, Adilson Fransozo, and Maria Lucia Negreiros-Fransozo. Reproduction and recruitment of the South American red shrimp, Pleoticus muelleri (Crustacea: Solenoceridae), from the southeastern coast of Brazil. Marine Biology Research, 4(5):361–368, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802029536.

Chernova:2008:NSP

[157] Natalia V. Chernova and Peter R. Møller. A new snailfish, Paraliparis nigellus sp. nov. (Scorpaeniformes, Liparidae), from the northern Mid-Atlantic Ridge — with notes on occurrence of Psednos in the area. Marine Biology Research, 4(5):369–375, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000802017507.

Alfaro:2008:VGH

[158] Andrea C. Alfaro, Steve C. Webb, and Claire Barnaby. Variability of growth, health, and population turnover within mussel beds of Perna canaliculus in northern New Zealand. Marine Biology Research, 4(5): 376–383, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802022879.

Israelsson:2008:XDP

[159] Olle Israelsson. Xenoturbella (Deuterostomia) probably feeds on dissolved organic matter. Marine Biology Research, 4(5):384–391, 2008. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000802036085.

Vasconcelos:2008:CIA

[160] Paulo Vasconcelos, Belisandra Lopes, Margarida Castro, and Miguel B. Gaspar. Comparison of indices for the assessment of reproductive activity in Hexaplex trunculus (Gastropoda: Muricidae). Marine Biology Research, 4(5):392–399, 2008. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000802020436.

Gislason:2008:VDS

[161] Astthor Gislason. Vertical distribution and seasonal dynamics of mesozooplankton in the Iceland Basin. Marine Biology Research, 4(6): 401–413, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802232882.

Husa:2008:CMB

[162] Vivian Husa, Kjersti Sjøtun, Narve Brattenborg, and Tor Eiliv Lein. Changes of macroalgal biodiversity in sublittoral sites in southwest Norway: impact of an introduced species or higher temperature? Marine Biology Research, 4(6):414–428, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000802232874.

Naruse:2008:DFN

[163] Tohru Naruse, Jose Christopher E. Mendoza, and Peter K. L. Ng. Descriptions of five new species of false spider crabs (Decapoda: Brachyura: Hymenosomatidae) from the Philippines. Marine Biology Research, 4 (6):429–441, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802232866.

Schmidt:2008:SAD

[164] Maike Schmidt, Eva E. R. Philipp, and Doris Abele. Size and age-dependent changes of escape response to predator attack in the Queen scallop Aequipecten opercularis. Marine Biology Research, 4 (6):442–450, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802270346.

Saglam:2008:FET

[165] Hacer Saglam and Nimet Selda Bascinar. Feeding ecology of thornback ray (Raja clavata Linnaeus, 1758) on the Turkish coast of the southeastern Black Sea. Marine Biology Research, 4(6):451–457, 2008. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000802233393.

Tunberg:2008:CTC

[166] Björn G. Tunberg and Daniel Krång. Composition and temporal change of the macrofauna in a shallow, sandy bottom in the Lysekil archipelago, western Sweden. Marine Biology Research, 4(6):458–469, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802208973.

Choi:2008:CCP

[167] Cheol Young Choi, Kwang Wook An, Hyun Suk Shin, Myung In An, and Pil Gue Jo. Changes of cytochrome P4501A mRNA expression and physiology responses in the olive flounder, Paralichthys olivaceus, exposed to benzo[a]pyrene. Marine Biology Research, 4(6):470–476, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802232890.

Tsikliras:2008:CRG

[168] Athanassios C. Tsikliras. Climate-related geographic shift and sudden population increase of a small pelagic fish (Sardinella aurita) in the eastern Mediterranean Sea. Marine Biology Research, 4(6): 477–481, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802291292.

Anonymous:2008:RA

[169] Anonymous. Reviewer acknowledgement 2007. Marine Biology Research, 4(6):482, 2008. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802610897.

Haug:2009:FYN

[170] Tore Haug, Ingolf Røttingen, Harald Gjøsæter, Ole Arve Misund, Tom Fenchel, and Franz Uiblein. Fifty years of Norwegian–Russian collaboration in marine research. Marine Biology Research, 5(1):1–3, 2009. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000802512747.

Klyashtorin:2009:CCC

[171] Leonid B. Klyashtorin, Vladimir Borisov, and Alexey Lyubushin. Cyclic changes of climate and major commercial stocks of the Barents Sea. Marine Biology Research, 5(1):4–17, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000802512283.

Falk-Petersen:2009:LLS

[172] Stig Falk-Petersen, Patrick Mayzaud, Gerhard Kattner, and John R. Sargent. Lipids and life strategy of Arctic Calanus. Marine Biology Research, 5(1):18–39, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000802512267.

Gjosaeter:2009:EET

[173] Harald Gjøsæter, Bjarte Bogstad, and Sigurd Tjelmeland. Ecosystem effects of the three capelin stock collapses in the Barents Sea. Marine Biology Research, 5(1):40–53, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000802454866.

Yaragina:2009:FYA

[174] Natalia A. Yaragina, Kjell H. Nedreaas, Valentina P. Koloskova, Hildegunn Mjanger, Harald Senneset, Natalya V. Zuykova, and Perå Gotnes. Fifteen years of annual Norwegian–Russian cod comparative age readings. Marine Biology Research, 5(1):54–65, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802454882.

Zuykova:2009:ADN

[175] Natalya V. Zuykova, Valentina P. Koloskova, Hildegunn Mjanger, Kjell H. Nedreaas, Harald Senneset, Natalia A. Yaragina, Per Aagotnes, and Sondre Aanes. Age determination of Northeast Arctic cod otoliths through 50 years of history. Marine Biology Research, 5(1): 66–74, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802454874.

Yaragina:2009:VCN

[176] Natalia A. Yaragina, Bjarte Bogstad, and Yurt A. Kovalev. Variability in cannibalism in Northeast Arctic cod (Gadus morhua) during the period 1947–2006. Marine Biology Research, 5(1):75–85, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802512739.

Drevetnyak:2009:HMP

[177] Konstantin Drevetnyak and Kjell H. Nedreaas. Historical movement pattern of juvenile beaked redfish (Sebastes mentella Travin) in the Barents Sea as inferred from long-term research survey series. Marine Biology Research, 5(1):86–100, 2009. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000802534865.

Dahle:2009:PHP

[178] Salve Dahle, Vladimir Savinov, Jarle Klungsøyr, Stepan Boitsov, Natalia Plotitsyna, Andrey Zhilin, Tatiana Savinova, and Vera Petrova. Polyaromatic hydrocarbons (PAHs) in the Barents Sea sediments: small changes over the recent 10 years. Marine Biology Research, 5(1): 101–108, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802512259.

Anonymous:2009:TIN

[179] Anonymous. Thematic issue nr. 2: Fifty years of Norwegian–Russian collaboration in marine research, Tore Haug, Ingolf Røttingen, Harald Gjøsæter and Ole Arve Misund (Guest Editors). Marine Biology Research, 5(1):??, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802512754.

Clemente:2009:GDS

[180] Sandhya Clemente and Baban Ingole. Gametogenic development and spawning of the mud clam, Polymesoda erosa (Solander, 1876) at Chorao Island, Goa. Marine Biology Research, 5(2):109–121, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802317709.

Sanchez-Moyano:2009:DTS

[181] Juan Emilio Sánchez-Moyano and Isabel García-Asencio. Distribution and trophic structure of annelid assemblages in a Caulerpa prolifera bed from southern Spain. Marine Biology Research, 5(2):122–132, 2009. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000802317691.

Magnussen:2009:EPC

[182] Eydfinn Magnussen and Marita D. Magnussen. Ecology of poor-cod (Trisopterus minutus) on the Faroe Bank. Marine Biology Research, 5 (2):133–142, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802345841.

Lilliendahl:2009:WDA

[183] Kristjan Lilliendahl. Winter diets of auks in Icelandic coastal waters. Marine Biology Research, 5(2):143–154, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000802279636.

Wilk:2009:EVF

[184] John Wilk and Rüdiger Bieler. Ecophenotypic variation in the Flat Tree Oyster, Isognomon alatus (Bivalvia: Isognomonidae), across a tidal microhabitat gradient. Marine Biology Research, 5(2):155–163, 2009. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000802279644.

Connolly:2009:SIE

[185] Rod M. Connolly, Thomas A. Schlacher, and Troy F. Gaston. Stable isotope evidence for trophic subsidy of coastal benthic fisheries by river discharge plumes off small estuaries. Marine Biology Research, 5 (2):164–171, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802266625.

Mahon:2009:DGB

[186] Andrew R. Mahon, Heidi K. Mahon, Daniel M. Dauer, and Kenneth M. Halanych. Discrete genetic boundaries of three Streblospio (Spionidae, Annelida) species and the status of S. shrubsolii. Marine Biology Research, 5(2):172–178, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000802317683.

Susilo:2009:TME

[187] Endang Sri Susilo, Lilik Harnadi, and Akihiro Takemura. Tropical monsoon environments and the reproductive cycle of the orangespotted spinefoot Siganus guttatus. Marine Biology Research, 5(2): 179–185, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802266633.

Kijewska:2009:VCN

[188] Agnieszka Kijewska, Artur Burzynski, and Roman Wenne. Variation in the copy number of tandem repeats of mitochondrial DNA in the North-East Atlantic cod populations. Marine Biology Research, 5(2): 186–192, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802277408.

Sigwart:2009:PFD

[189] Julia D. Sigwart. Parasitic foraminifers on a deep-sea chiton (Mollusca, Polyplacophora, Leptochitonidae) from Iceland. Marine Biology Research, 5(2):193–199, 2009. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000802266641.

Dodd:2009:MSV

[190] Jane Dodd, Laura Baxter, and David J. Hughes. Mapping Serpula vermicularis (Polychaeta: Serpulidae) aggregations in Loch Teacuis, western Scotland, a new record. Marine Biology Research, 5(2): 200–205, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802345858.

Sjotun:2009:BR

[191] Kjersti Sjøtun. Book review. Marine Biology Research, 5(2): 206, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902740859.

Dijkstra:2009:PMB

[192] Henk H. Dijkstra, Anders Warén, and Gudmundur Gudmundsson. Pectinoidea (Mollusca: Bivalvia) from Iceland. Marine Biology Research, 5 (3):207–243, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802425643.

Giraldo:2009:CUC

[193] Alan Giraldo, Ruben Escribano, Victor Marín, and Pamela Hidalgo. Coastal upwelling circulation and its influence on the population dynamics of Calanus chilensis (Brodski, 1959) off northern Chile (23^{o}S). Marine Biology Research, 5(3):244–256, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000802425635.

Aranha:2009:BAV

[194] Ana Aranha, Gui Menezes, and Mário Rui Pinho. Biological aspects of the velvet belly lantern shark, Etmopterus spinax (Linnaeus, 1758) off the Azores, North East Atlantic. Marine Biology Research, 5(3): 257–267, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802433175.

Miranda:2009:EIP

[195] Richard M. Miranda, Katsuaki Fujinaga, Anthony S. Ilano, and Shigeru Nakao. Effects of imposex and parasite infection on the reproductive features of the Neptune whelk Neptunea arthritica. Marine Biology Research, 5(3):268–277, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000802419422.

Zhong:2009:CMG

[196] Jing Zhong, Qiujin Zhang, Qunshan Xu, Michael Schubert, Vincent Laudet, and Yiquan Wang. Complete mitochondrial genomes defining two distinct lancelet species in the West Pacific Ocean. Marine Biology Research, 5(3):278–285, 2009. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000802430817.

Busch:2009:EFS

[197] Kjersti Eline Tønnessen Busch, Arild Folkvord, Håkon Otterå, William F. Hutchinson, and Terje Svåsand. Effects of female spawning experience and larval size on feeding and growth of cod larvae Gadus morhua l. reared in mesocosms. Marine Biology Research, 5 (3):286–296, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802441319.

Todaro:2009:NSA

[198] M. Antonio Todaro, Matteo Dal Zotto, Anastassya S. Maiorova, and Andrey V. Adrianov. A new species of Aspidiophorus (Gastrotricha, Chaetonotida) from the Russian Far East with a key to marine species of the genus. Marine Biology Research, 5(3):297–303, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802419430.

Nyberg:2009:LTS

[199] Cecilia D. Nyberg and Inger Wallentinus. Long-term survival of an introduced red alga in adverse conditions. Marine Biology Research, 5 (3):304–308, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802428159.

Hylland:2009:TFM

[200] Ketil Hylland. The toxicology of fishes; metal contamination in aquatic environments — science and lateral management. Marine Biology Research, 5(3):309–312, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000902859089.

Fenchel:2009:MBR

[201] Tom Fenchel and Franz Uiblein. Marine Biology Research: Taxonomy of marine organisms. Marine Biology Research, 5(4):313–314, 2009. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000902984689.

Nygren:2009:ABS

[202] Arne Nygren, Jenny Eklöf, and Fredrik Pleijel. Arctic-boreal sibling species of Paranaitis (Polychaeta, Phyllodocidae). Marine Biology Research, 5(4):315–327, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000802441301.

Wienerroither:2009:CDD

[203] Rupert Wienerroither, Franz Uiblein, Fernando Bordes, and Teresa Moreno. Composition, distribution, and diversity of pelagic fishes around the Canary Islands, Eastern Central Atlantic. Marine Biology Research, 5(4):328–344, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802478055.

Abraham:2009:MCS

[204] Biju Abraham and Saramma U. Panampunnayil. Mysids (Crustacea) from the shallow waters off Maharashtra and south Gujarat, India, with description of a new species. Marine Biology Research, 5(4): 345–362, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802454890.

Polat:2009:ABP

[205] Sevim Polat and Zahit Uysal. Abundance and biomass of picoplanktonic Synechococcus (Cyanobacteria) in a coastal ecosystem of the northeastern Mediterranean, the Bay of Iskenderun. Marine Biology Research, 5 (4):363–373, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802512275.

Ibanez:2009:DDC

[206] Christian M. Ibáñez, Patricio A. Camus, and Francisco J. Rocha. Diversity and distribution of cephalopod species off the coast of Chile. Marine Biology Research, 5(4):374–384, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000802534873.

Hummon:2009:NSM

[207] William D. Hummon and Christopher J. Guadiz. A new species of marine Gastrotricha from Maine, USA: Tetranchyroderma mainensis (Macrodasyida, Thaumastodermatidae). Marine Biology Research, 5 (4):385–390, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802441327.

Pavithran:2009:ISS

[208] Sini Pavithran, Baban Ingole, Mandar Nanajkar, and Reshma Goltekar. Importance of sieve size in deep-sea macrobenthic studies. Marine Biology Research, 5(4):391–398, 2009. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000802441285.

Diupotex-Chong:2009:CAA

[209] Maria Esther Diupotex-Chong, Alberto Ocaña-Luna, and Marina Sánchez-Ramírez. Chromosome analysis of Aurelia aurita Linné, 1758 (Scyphozoa: Ulmaridae), southern Gulf of Mexico. Marine Biology Research, 5(4):399–403, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000802534907.

Sanchez-Tocino:2009:CKG

[210] Luís Sánchez-Tocino and J. Manuel Tierno de Figueroa. Contribution to the knowledge of the genus Loxosomella (Entoprocta) from the Spanish Mediterranean Sea, with the description of a new species, Loxosomella ameliae sp. nov. Marine Biology Research, 5(4):404–407, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802534899.

Todt:2009:CHN

[211] Christiane Todt, Paco Cárdenas, and Hans Tore Rapp. The chiton Hanleya nagelfar (Polyplacophora, Mollusca) and its association with sponges in the European Northern Atlantic. Marine Biology Research, 5(4):408–411, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802572394.

Chernova:2009:SBA

[212] Natalia V. Chernova. Selected bibliography of Anatoly Petrovich Andriashev (1910–2009). Marine Biology Research, 5(4):412–413, 2009. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000902984671.

Moller:2009:BF

[213] Peter Rask Møller. Biology of fishes. Marine Biology Research, 5 (4):414, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902889821.

Kodama:2009:RBJ

[214] Keita Kodama, Hiroaki Shiraishi, Masatoshi Morita, and Toshihiro Horiguchi. Reproductive biology of the Japanese mantis shrimp Oratosquilla oratoria (Crustacea Stomatopoda): Annual cycle of gonadal development and copulation. Marine Biology Research, 5(5): 415–426, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802644714.

Vidjak:2009:SCD

[215] Olja Vidjak and Natalia Bojanić. Species composition and distribution patterns of the family Corycaeidae Dana, 1852 (Copepoda: Cyclopoida) in the middle Adriatic Sea. Marine Biology Research, 5(5): 427–440, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802603629.

Albalat:2009:ECM

[216] Amaya Albalat, Sebastian G. Gornik, Robert J. A. Atkinson, Graham H. Coombs, and Douglas M. Neil. Effect of capture method on the physiology and nucleotide breakdown products in the Norway lobster (Nephrops norvegicus). Marine Biology Research, 5(5):441–450, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802603637.

Park:2009:CME

[217] Mi Seon Park, Pil Gue Jo, Yong Ki Choi, Kwang Wook An, and Cheol Young Choi. Characterization and mRNA expression of Mn-SOD and physiological responses to stresses in the Pacific oyster Crassostrea gigas. Marine Biology Research, 5(5):451–461, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802626554.

Bapary:2009:ECG

[218] Mohammad Abu Jafor Bapary, Pouvave Fainuulelei, and Akihiro Takemura. Environmental control of gonadal development in the tropical damselfish Chrysiptera cyanea. Marine Biology Research, 5(5): 462–469, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802644722.

Zhang:2009:RIB

[219] Dong Zhang, Junda Lin, Jörg D. Hardege, and Andrew L. Rhyne. Reproductive isolation between two sympatric simultaneous hermaphroditic shrimp, Lysmata wurdemanni and L. boggessi. Marine Biology Research, 5(5):470–477, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802626547.

Ladhar-Chaabouni:2009:SDC

[220] Rim Ladhar-Chaabouni, Monia Machreki-Ajmi, and Amel Hamza-Chaffai. Spatial distribution of cadmium and some biomarkers in Cerastoderma glaucum living in a polluted area. Marine Biology Research, 5 (5):478–486, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802683985.

Ladhar-Chaabouni:2009:VVS

[221] Rim Ladhar-Chaabouni, Wafa Smaoui-Damak, and Amel Hamza-Chaffai. In vivo variation of some biomarkers with time and cadmium concentration in the cockle Cerastoderma glaucum. Marine Biology Research, 5(5):487–495, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000802534881.

Ghanbarifardi:2009:DDA

[222] Mehdi Ghanbarifardi and Masoumeh Malek. Distribution, diversity, and abundance of rocky intertidal fishes in the Persian Gulf and Gulf of Oman, Iran. Marine Biology Research, 5(5):496–502, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000802441293.

AquinoDeSouza:2009:AOM

[223] Rosana Aquino De Souza, Paul Tyler, and Stephen John Hawkins. Artificial oocyte maturation in Patella depressa and Patella vulgata using NaOH-alkalinized seawater. Marine Biology Research, 5(5): 503–510, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000802603645.

Praebel:2009:TSC

[224] Kim Præbel, Jørgen S. Christiansen, and Svein-Erik Fevolden. Temperature and salinity conditions in a sub-Arctic intertidal spawning habitat for capelin. Marine Biology Research, 5(5):511–514, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000902729670.

Norderhaug:2009:SUG

[225] Kjell Magnus Norderhaug and Hartvig C. Christie. Sea urchin grazing and kelp re-vegetation in the NE Atlantic. Marine Biology Research, 5 (6):515–528, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902932985.

Heiner:2009:FTD

[226] Iben Heiner, Tom M. Boesgaard, and Reinhardt M. Kristensen. First time discovery of loricifera from Australian waters and marine caves. Marine Biology Research, 5(6):529–546, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000902933009.

Zelnio:2009:HAA

[227] Kevin A. Zelnio, Estefania Rodriguez, and Marymegan Daly. Hexacorals (Anthozoa: Actiniaria, Zoanthidea) from hydrothermal vents in the south-western Pacific. Marine Biology Research, 5(6):547–571, 2009. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000902729662.

Rodil:2009:SVB

[228] Iván F. Rodil, Mariano Lastra, and Jesús López. Spatial variability of benthic macrofauna in the Ria of Vigo (NW Spain): Effect of sediment type and food availability. Marine Biology Research, 5(6): 572–584, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902729688.

Riisgaard:2009:CPM

[229] Hans Ulrik Riisgård and Poul S. Larsen. Ciliary-propelling mechanism, effect of temperature and viscosity on swimming speed, and adaptive significance of ‘jumping’ in the ciliate Mesodinium rubrum. Marine Biology Research, 5(6):585–595, 2009. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000902729704.

Kristoffersen:2009:DGP

[230] Jon Bent Kristoffersen and Anne Gro Vea Salvanes. Distribution, growth, and population genetics of the glacier lanternfish (Benthosema glaciale) in Norwegian waters: Contrasting patterns in fjords and the ocean. Marine Biology Research, 5(6):596–604, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000903042479.

Abecasis:2009:NRI

[231] David Abecasis, Frederico Cardigos, Frederico Almada, and Jorge Manuel Santos Gonçalves. New records on the ichthyofauna of the Gorringe seamount (Northeastern Atlantic). Marine Biology Research, 5 (6):605–611, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902729696.

Anonymous:2009:RA

[232] Anonymous. Reviewer acknowledgement 2008. Marine Biology Research, 5(6):612, 2009. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903239521.

Fenchel:2010:BCC

[233] Tom Fenchel and Franz Uiblein. Biodiversity in concert: Common, uncommon, and new species. Marine Biology Research, 6(1):1–5, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000903468856.

Budaeva:2010:LDM

[234] Nataliya Budaeva and Kristian Fauchald. Larval development of Mooreonuphis stigmatis (Treadwell, 1922) (Polychaeta: Onuphidae) from the north-east Pacific. Marine Biology Research, 6(1):6–24, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000902932977.

Vecchione:2010:CNM

[235] Michael Vecchione, Richard E. Young, and Uwe Piatkowski. Cephalopods of the northern Mid-Atlantic Ridge. Marine Biology Research, 6(1): 25–52, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902810751.

Monteiro:2010:AGM

[236] Pedro Monteiro, Luis Bentes, Rui Coelho, Carla Correia, Karim Erzini, Pedro G. Lino, Joaquim Ribeiro, and Jorge M. S. Gonçalves. Age and growth, mortality and reproduction of the striped sea bream, Lithognathus mormyrus Linnaeus 1758, from the south coast of Portugal (Algarve). Marine Biology Research, 6(1):53–65, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000903039731.

Alberico:2010:HMS

[237] Natalia A. Alberico and Daniel Roccatagliata. Homoeoscelis meridionalis sp. nov. (Copepoda, Nicothoidae) parasitic on two cumaceans from southern South America with comments on its biology, and partial redescription of H. frigida Hansen, 1923. Marine Biology Research, 6(1):66–77, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902810777.

Yan:2010:SCR

[238] Hongwei Yan, Qi Li, Wenguang Liu, Ruihai Yu, and Lingfeng Kong. Seasonal changes in reproductive activity and biochemical composition of the razor clam Sinonovacula constricta (Lamarck 1818). Marine Biology Research, 6(1):78–88, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000903039756.

Schmalenbach:2010:VPS

[239] Isabel Schmalenbach and Friedrich Buchholz. Vertical positioning and swimming performance of lobster larvae (Homarus gammarus) in an artificial water column at Helgoland, North Sea. Marine Biology Research, 6(1):89–99, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902810769.

Wang:2010:ECK

[240] Ping Wang, Gang Wang, Yajuan Teng, Xinshu Li, Jing Ji, Xinghong Xu, and Yongzhong Li. Effects of cefotaxime and kanamycin on thallus proliferation and differentiation in Porphyra yezoensis and their inhibition on Agrobacterium tumefaciens. Marine Biology Research, 6 (1):100–105, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000902933017.

Coluccia:2010:CCM

[241] Elisabetta Coluccia, Anna Maria Deiana, Angelo Libertini, and Susanna Salvadori. Cytogenetic characterization of the moray eel Gymnothorax tile and chromosomal banding comparison in Muraenidae (Anguilliformes). Marine Biology Research, 6(1):106–111, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000902932993.

Christiansen:2010:NBP

[242] Bernd Christiansen, Solveig I. Bühring, Olaf Pfannkuche, and Horst Weikert. The near-bottom plankton community at the Porcupine Abyssal Plain, NE-Atlantic: Structure and vertical distribution. Marine Biology Research, 6(2):113–124, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000903150363.

Salazar-Silva:2010:RHC

[243] Patricia Salazar-Silva. Redescription of Harmothoe crucis (Annelida, Polychaeta, Polynoidae), and re-establishment of synonymized species from the Grand Caribbean with descriptions of four new species. Marine Biology Research, 6(2):125–154, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000903037248.

Schander:2010:FHV

[244] Christoffer Schander, Hans Tore Rapp, Jon Anders Kongsrud, Torkild Bakken, Jørgen Berge, Sabine Cochrane, Eivind Oug, Ingvar Byrkjedal, Christiane Todt, Tomas Cedhagen, Audun Fosshagen, Andrey Gebruk, Kim Larsen, Lisa Levin, Matthias Obst, Fredrik Pleijel, Sabine Stöhr, Anders Warén, Nina Therese Mikkelsen, Silje Hadler-Jacobsen, Rozemarijn Keuning, Kristin Heggøy Petersen, Ingunn H. Thorseth, and Rolf B. Pedersen. The fauna of hydrothermal vents on the Mohn Ridge (North Atlantic). Marine Biology Research, 6(2):155–171, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000903147450.

Rho:2010:PDS

[245] Hyun Soo Rho, Won Gi Min, Wilfrida Decraemer, and Dong Sung Kim. Prochaetosoma dokdoense sp. nov. (Nematoda: Draconematidae) from Dokdo, Korea: First record of the genus Prochaetosoma from a shallow subtidal zone in the northwest Pacific Ocean. Marine Biology Research, 6(2):172–188, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903078622.

Fredriksen:2010:IZM

[246] Stein Fredriksen, Annelies De Backer, Christoffer Boström, and Hartvig Christie. Infauna from Zostera marina L. meadows in Norway. Differences in vegetated and unvegetated areas. Marine Biology Research, 6 (2):189–200, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903042461.

Falk-Petersen:2010:SLS

[247] Inger-Britt Falk-Petersen, Premasany Kanapathippilai, Raul Primicerio, and Tove K. Hansen. Size, locality and seasonally related feeding preferences of common wolffish (Anarhichas lupus L.) from north-Norwegian waters. Marine Biology Research, 6(2):201–212, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000903078614.

Terossi:2010:SRR

[248] Mariana Terossi and Fernando L. Mantelatto. Sexual ratio, reproductive period and seasonal variation of the gonochoric shrimp Hippolyte obliquimanus (Caridea: Hippolytidae). Marine Biology Research, 6 (2):213–219, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903078630.

Markager:2010:PGA

[249] Stiig Markager. Practical guidelines for the analysis of seawater. Marine Biology Research, 6(2):220, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000903514220.

Berggren:2010:DCP

[250] Matz S. Berggren. Decapod crustacean phylogenetics. Marine Biology Research, 6(2):221–222, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451001003604796.

Gornik:2010:IDA

[251] Sebastian G. Gornik, Amaya Albalat, Robert J. A. Atkinson, Graham H. Coombs, and Douglas M. Neil. The influence of defined ante-mortem stressors on the early post-mortem biochemical processes in the abdominal muscle of the Norway lobster, Nephrops norvegicus (Linnaeus, 1758). Marine Biology Research, 6(3):223–238, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000903147468.

Skoglund:2010:FAC

[252] Eli G. Skoglund, Christian Lydersen, Otto Grahl-Nielsen, Tore Haug, and Kit M. Kovacs. Fatty acid composition of the blubber and dermis of adult male Atlantic walruses (Odobenus rosmarus rosmarus) in Svalbard, and their potential prey. Marine Biology Research, 6(3):239–250, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000903233755.

Seljeset:2010:RPC

[253] Oddbjørn Seljeset, Knut Wiik Vollset, Arild Folkvord, and Audrey J. Geffen. The role of prey concentration and size range in the growth and survival of larval cod. Marine Biology Research, 6(3):251–262, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000903150355.

Yagunova:2010:IST

[254] Ekaterina B. Yagunova and Andrew N. Ostrovsky. The influence of substrate type on sexual reproduction of the bryozoan Cribrilina annulata (Gymnolaemata, Cheilostomata): a case study from Arctic seas. Marine Biology Research, 6(3):263–270, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000903147443.

Plyuscheva:2010:DAS

[255] Maria Plyuscheva, Daniel Martin, and Temir Britayev. Diet analyses of the scale-worms Lepidonotus squamatus and Harmothoe imbricata (Polychaeta, Polynoidae) in the White Sea. Marine Biology Research, 6(3):271–281, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903334694.

Vonk:2010:AEE

[256] Jan Arie Vonk, Marjolijn J. A. Christianen, and Johan Stapel. Abundance, edge effect, and seasonality of fauna in mixed-species seagrass meadows in southwest Sulawesi, Indonesia. Marine Biology Research, 6 (3):282–291, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903233789.

Poorbagher:2010:ENS

[257] Hadi Poorbagher, Miles D. Lamare, and Mike F. Barker. Effects of nutrition on somatic growth and reproductive strategy of the sea urchin Pseudechinus huttoni. Marine Biology Research, 6(3):292–301, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000903233797.

Poorbagher:2010:RIP

[258] Hadi Poorbagher, Miles D. Lamare, Mike F. Barker, and Will Rayment. Relative importance of parental diet versus larval nutrition on development and phenotypic plasticity of Pseudechinus huttoni larvae (Echinodermata: Echinoidea). Marine Biology Research, 6(3): 302–314, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903300877.

Mercier:2010:FRS

[259] Annie Mercier, David L. Pawson, Doris J. Pawson, and Jean-François Hamel. First record of the sea cucumber Trachythyone nina (Echinodermata: Holothuroidea) in Canadian waters, with a redescription of the species and notes on its distribution and biology. Marine Biology Research, 6(3):315–320, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000903233805.

Guerra-Garcia:2010:KCM

[260] José M. Guerra-García, M. Pilar Cabezas, Elena Baeza-Rojano, and J. Carlos García-Gómez. Na, K, Ca and Mg of intertidal caprellids (Crustacea: Amphipoda). Marine Biology Research, 6(3):321–326, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000903300901.

Wang:2010:EPW

[261] Yanbo Wang and Qing Gu. Effect of probiotics on white shrimp (Penaeus vannamei) growth performance and immune response. Marine Biology Research, 6(3):327–332, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000903300893.

Janas:2010:STE

[262] Urszula Janas and John I. Spicer. Seasonal and temperature effects on osmoregulation by the invasive prawn Palaemon elegans Rathke, 1837 in the Baltic Sea. Marine Biology Research, 6(3):333–337, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451001003670086.

Seret:2010:ERF

[263] Bernard Séret, Tom Fenchel, and Franz Uiblein. European research focus on sharks and rays. Marine Biology Research, 6(4):339–340, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451001003657729.

Cannas:2010:MME

[264] Rita Cannas, Maria Cristina Follesa, Serenella Cabiddu, Cristina Porcu, Susanna Salvadori, Samuel Paco Iglésias, Anna Maria Deiana, and Angelo Cau. Molecular and morphological evidence of the occurrence of the Norwegian skate Dipturus nidarosiensis (Storm, 1881) in the Mediterranean Sea. Marine Biology Research, 6(4):341–350, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000903428496.

Kitchen-Wheeler:2010:VII

[265] Anne-Marie Kitchen-Wheeler. Visual identification of individual manta ray (Manta alfredi) in the Maldives Islands, Western Indian Ocean. Marine Biology Research, 6(4):351–363, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000903233763.

Rey:2010:DPG

[266] Javier Rey, Rui Coelho, Domingo Lloris, Bernard Séret, and Luis Gil de Sola. Distribution pattern of Galeus atlanticus in the Alborán Sea (south western Mediterranean) and some sexual character comparison with Galeus melastomus. Marine Biology Research, 6(4): 364–372, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903042487.

Coelho:2010:CAM

[267] Rui Coelho, Javier Rey, Luis Gil de Sola, Joana Fernandez de Carvalho, and Karim Erzini. Comparing Atlantic and Mediterranean populations of the velvet belly lanternshark, Etmopterus spinax, with comments on the efficiency of density-dependent compensatory mechanisms. Marine Biology Research, 6(4):373–380, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000903300885.

Coelho:2010:DDV

[268] Rui Coelho and Karim Erzini. Depth distribution of the velvet belly, Etmopterus spinax, in relation to growth and reproductive cycle: the case study of a deep-water lantern shark with a wide-ranging critical habitat. Marine Biology Research, 6(4):381–389, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000802644706.

Kousteni:2010:SMF

[269] Vasiliki Kousteni, Mary Kontopoulou, and Persefoni Megalofonou. Sexual maturity and fecundity of Scyliorhinus canicula (Linnaeus, 1758) in the Aegean Sea. Marine Biology Research, 6(4):390–398, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000903233771.

Spaet:2010:LHA

[270] Julia L. Y. Spaet, Steven T. Kessel, and Samuel H. Gruber. Learned hook avoidance of lemon sharks (Negaprion brevirostris) based on electroreception and shock treatment. Marine Biology Research, 6(4): 399–407, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903039749.

Sperone:2010:SIA

[271] Emilio Sperone, Primo Micarelli, Sara Andreotti, Sara Spinetti, Andrea Andreani, Fabrizio Serena, Elvira Brunelli, and Sandro Tripepi. Social interactions among bait-attracted white sharks at Dyer Island (South Africa). Marine Biology Research, 6(4):408–414, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000903078648.

Coelho:2010:BMD

[272] Rui Coelho. Biology and management of dogfish sharks. Marine Biology Research, 6(4):415–416, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451001003629314.

Seret:2010:SRA

[273] Bernard Séret. Sharks and rays of Australia. Marine Biology Research, 6(4):417–420, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903505624.

Anonymous:2010:ERF

[274] Anonymous. European research focus on sharks and rays. Marine Biology Research, 6(4):??, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451001003657711.

Goffredo:2010:GPB

[275] Stefano Goffredo, Gabriella Gasparini, Giulia Marconi, Maria Teresa Putignano, Claudia Pazzini, and Francesco Zaccanti. Gonochorism and planula brooding in the Mediterranean endemic orange coral Astroides calycularis (Scleractinia: Dendrophylliidae). Morphological aspects of gametogenesis and ontogenesis. Marine Biology Research, 6 (5):421–436, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903428488.

Vieira:2010:TNE

[276] Wellington F. Vieira, Bruno Cosme, and Eduardo Hajdu. Three new Erylus (Demospongiae, Astrophorida, Geodiidae) from the Almirante Saldanha Seamount (off SE Brazil), with further data for a tabular review of worldwide species and comments on Brazilian seamount sponges. Marine Biology Research, 6(5):437–460, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000903334728.

Riser:2010:ZMC

[277] Christian Wexels Riser, Marit Reigstad, and Paul Wassmann. Zooplanktonmediated carbon export: a seasonal study in a northern Norwegian fjord. Marine Biology Research, 6(5):461–471, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000903437067.

Sorensen:2010:NSR

[278] Martin V. Sørensen, Hyun Soo Rho, and Dongsung Kim. A new species of the rare genus Sphenoderes (Cyclorhagida, Kinorhyncha), with differential notes on S. indicus Higgins, 1969. Marine Biology Research, 6(5):472–484, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903334702.

Ikeda:2010:SGJ

[279] Yuzuru Ikeda and Masato Kobayashi. Statolith growth of juvenile oval squid Sepioteuthis lessoniana (Cephalopoda: Loliginidae) with special reference to ambient thermal condition. Marine Biology Research, 6 (5):485–495, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903334710.

Esposito:2010:FHS

[280] Valentina Esposito, Luca Castriota, Pierpaolo Consoli, Teresa Romeo, Manuela Falautano, and Franco Andaloro. Feeding habits and selectivity of the wide-eyed flounder, Bothus podas (Delaroche, 1809) (Bothidae) from the southern Tyrrhenian sea. Marine Biology Research, 6 (5):496–502, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903438495.

Islam:2010:ODS

[281] Md. Sherazul Islam, Keita Kodama, and Hisashi Kurokura. Ovarian development and size at maturity of the mud crab Scylla olivacea in Pak Phanang mangrove swamps, Thailand. Marine Biology Research, 6(5):503–510, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903335113.

Sharma:2010:AHM

[282] Anubha Sharma, Iti Gontia, Pradeep K. Agarwal, and Bhavanath Jha. Accumulation of heavy metals and its biochemical responses in Salicornia brachiata, an extreme halophyte. Marine Biology Research, 6 (5):511–518, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903434064.

Laidre:2010:WOE

[283] Kristin L. Laidre, Mads Peter Heide-Jørgensen, Miles L. Logsdon, Leon Delwiche, and Torkel Gissel Nielsen. A whale of an opportunity: Examining the vertical structure of chlorophyll- a in high Arctic waters using instrumented marine predators. Marine Biology Research, 6(6): 519–529, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903505616.

Melo:2010:DZC

[284] Pedro A. M. C. Melo, Tâmara A. Silva, Sigrid Neumann-Leitão, Ralf Schwamborn, Lucia M. O. Gusmão, and Fernando Porto Neto. Demersal zooplankton communities from tropical habitats in the southwestern Atlantic. Marine Biology Research, 6(6):530–541, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000903426557.

Nielsen:2010:HRT

[285] Lasse Tor Nielsen, Hans Henrik Jakobsen, and Per Juel Hansen. High resilience of two coastal plankton communities to twenty-first century seawater acidification: Evidence from microcosm studies. Marine Biology Research, 6(6):542–555, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000903476941.

Biju:2010:MCS

[286] Abraham Biju and Saramma Usha Panampunnayil. Mysids (Crustacea) from the salt pans of Mumbai, India, with a description of a new species. Marine Biology Research, 6(6):556–569, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000903527701.

Eash-Loucks:2010:PAN

[287] Wendy E. Eash-Loucks, Stephen C. Jewett, Daphne G. Fautin, Max K. Hoberg, and Héloïse Chenelot. Ptychodactis aleutiensis, a new species of ptychodactiarian sea anemone (Cnidaria: Anthozoa: Actiniaria) from the Aleutian Islands, Alaska. Marine Biology Research, 6(6): 570–578, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903476933.

Padate:2010:NSP

[288] Vinay P. Padate, Chandrashekher U. Rivonker, A. Chandrashekar Anil, Subhash S. Sawant, and Venkat Krishnamurthy. A new species of portunid crab of the genus Charybdis (De Haan, 1833) (Crustacea: Decapoda: Brachyura) from Goa, India. Marine Biology Research, 6(6): 579–590, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903505608.

Cabanellas-Reboredo:2010:IDS

[289] Miguel Cabanellas-Reboredo, Salud Deudero, Josep Alós, and Iris Hendriks. Initial data on settlement and recruitment of macrobenthic organisms on artificial substrates located over Posidonia oceanica meadows. Marine Biology Research, 6(6):591–599, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000903524690.

Lahbib:2010:IBC

[290] Youssef Lahbib, Sami Abidli, Jean-Francois Chiffoleau, Bernard Averty, and Najoua Trigui El Menif. Imposex and butyltin concentrations in snails from the lagoon of Bizerta (Northern Tunisia). Marine Biology Research, 6(6):600–607, 2010. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000903437075.

Vinuesa:2010:ILS

[291] Julio H. Vinuesa and Pamela Balzi. Infestation of Lithodes santolla (Decapoda: Lithodidae) by Pseudione tuberculata (Isopoda: Bopyridae) in San Jorge Gulf, Southwestern Atlantic Ocean. Marine Biology Research, 6(6):608–612, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000903478327.

Salcedo-Oropeza:2010:FRD

[292] Diana Leticia Salcedo-Oropeza, Pablo Hernández-Alcántara, and Vivianne Solís-Weiss. First record and distribution of Pisione longispinulata Aguado and San Martín (Annelida: Polychaeta: Pisionidae) in the Mexican Pacific coasts. Marine Biology Research, 6(6):613–617, 2010. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000903524708.

Anonymous:2010:RA

[293] Anonymous. Reviewer acknowledgement 2009. Marine Biology Research, 6(6):618, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.515999.

Anonymous:2010:MBR

[294] Anonymous. Marine biology research. Marine Biology Research, 6 (6):??, 2010. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.526819.

Fenchel:2011:SCR

[295] Tom Fenchel and Franz Uiblein. Saving coral reefs — and applied marine biology in Marine Biology Research. Marine Biology Research, 7(1):1–2, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.496854.

Omori:2011:DRC

[296] Makoto Omori. Degradation and restoration of coral reefs: Experience in Okinawa, Japan. Marine Biology Research, 7(1):3–12, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451001003642317.

Renaud:2011:BFW

[297] Paul E. Renaud, Michael Tessmann, Anita Evenset, and Guttorm N. Christensen. Benthic food-web structure of an Arctic fjord (Kongsfjorden, Svalbard). Marine Biology Research, 7(1):13–26, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451001003671597.

Urra:2011:MAL

[298] Javier Urra, Serge Gofas, José L. Rueda, and Pablo Marina. Molluscan assemblages in littoral soft bottoms of the Alboran Sea (Western Mediterranean Sea). Marine Biology Research, 7(1):27–42, 2011. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451001003660301.

Kedra:2011:DSN

[299] Monika Kedra and Mark A. Shields. On the deep-sea Nephasoma species in the Eastern North Atlantic — a taxonomic guide. Marine Biology Research, 7(1):43–53, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451001003660293.

Nygren:2011:CPP

[300] Arne Nygren, Erika Norlinder, Marina Panova, and Fredrik Pleijel. Colour polymorphism in the polychaete Harmothoe imbricata (Linnaeus, 1767). Marine Biology Research, 7(1):54–62, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451001003713555.

Strand:2011:DBD

[301] Malin Strand and Per Sundberg. A DNA-based description of a new nemertean (phylum Nemertea) species. Marine Biology Research, 7 (1):63–70, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451001003713563.

Ockelmann:2011:LWC

[302] Kurt W. Ockelmann and Grete E. Dinesen. Life on wood — the carnivorous deep-sea mussel Idas argenteus (Bathymodiolinae, Mytilidae, Bivalvia). Marine Biology Research, 7(1):71–84, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451001003714504.

Baeza-Rojano:2011:LHC

[303] Elena Baeza-Rojano, José M. Guerra-García, M. Pilar Cabezas, and Isabel Pacios. Life history of Caprella grandimana (Crustacea: Amphipoda) reared under laboratory conditions. Marine Biology Research, 7(1):85–92, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451001003713571.

Al-Rousan:2011:ASC

[304] Saber Al-Rousan, Fuad Al-Horani, Ehab Eid, and Maroof Khalaf. Assessment of seagrass communities along the Jordanian coast of the Gulf of Aqaba, Red Sea. Marine Biology Research, 7(1):93–99, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451001003660319.

Gruszka:2011:PER

[305] Piotr Gruszka and Beata Wiecaszek. Palaemon elegans Rathke, 1837 in the food of Baltic cod (Gadus morhua callarias L., 1758) from the Gulf of Gdańsk. Marine Biology Research, 7(1):100–105, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451001003660285.

Hoeksema:2011:BCR

[306] Bert W. Hoeksema. The biology of coral reefs. Marine Biology Research, 7(1):106–107, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.510890.

Rasmussen:2011:LSP

[307] Tone Rasmussen and Michaela Aschan. Larval stages of Pandalus borealis. Marine Biology Research, 7(2):109–121, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451001003764004.

Fleeger:2011:WRI

[308] John W. Fleeger, Mark A. Grippo, and Stephen Todd Pastorick. What is the relative importance of sediment granulometry and vertical gradients to nematode morphometrics? Marine Biology Research, 7(2): 122–134, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.492221.

Nahavandi:2011:SNS

[309] Nahid Nahavandi, Martin Plath, Ralph Tiedemann, and Ali R. Mirzajani. Sexual and natural selection on morphological traits in a marine amphipod, Pontogammarus maeoticus (Sowinsky, 1894). Marine Biology Research, 7(2):135–146, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451001003713589.

Fassatoui:2011:STT

[310] Chiheb Fassatoui, Esma Mdelgi, and Mohamed S. Romdhane. Shortterm temporal genetic investigation of two populations of Pagellus erythrinus (Linnaeus, 1758, Sparidae) from Western and Eastern Mediterranean basins off the Tunisian coast. Marine Biology Research, 7(2): 147–158, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451001003762792.

BertaciniDeMoraes:2011:JDC

[311] Juliana Cristina Bertacini De Moraes, Maria Lucia Negreiros Fransozo, and Gustavo Augusto Schmidt De Melo. Juvenile development of the crab Bathyrhombila sp. (Crustacea, Decapoda, Pseudorhombilidae) from megalopae obtained in the neuston. Marine Biology Research, 7(2): 159–175, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.489754.

Hiroki:2011:BDS

[312] Katia Aparecida Nunes Hiroki, Adilson Fransozo, Rogério Caetano Costa, Antonio Leão Castilho, Roberto Munehisa Shimizu, Ariádine Cristine Almeida, and Michele Furlan. Bathymetric distribution of the shrimp Rimapenaeus constrictus (Stimpson, 1874) (Decapoda, Penaeidae) in two locations off the southeastern Brazilian coast. Marine Biology Research, 7(2):176–185, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.489614.

Romeo:2011:RBS

[313] Teresa Romeo, Patrizia Perzia, Valentina Esposito, Danilo Malara, Pietro Battaglia, Pierpaolo Consoli, Simone Canese, and Franco Andaloro. Relationship between swordfish swimming behaviour and sea surface temperature in the central Mediterranean Sea during the reproductive period. Marine Biology Research, 7(2):186–194, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2010.489615.

Ocalewicz:2011:CCR

[314] Konrad Ocalewicz and Mariusz Sapota. Cytogenetic characteristics of the round goby Neogobius melanostomus (Pallas, 1814) (Teleostei: Gobiidae: Benthophilinae). Marine Biology Research, 7(2):195–201, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2010.489613.

Adkins:2011:DCS

[315] Merritt Adkins and Anja Schulze. Development of Capitella sp. G from Galveston Bay, Texas. Marine Biology Research, 7(2):202–207, 2011. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2010.489612.

Huang:2011:LVA

[316] Hui Huang, Zhijun Dong, Liangmin Huang, Jianhui Yang, Baoping Di, Yuanchao Li, Guowei Zhou, and Chenglong Zhang. Latitudinal variation in algal symbionts within the scleractinian coral Galaxea fascicularis in the South China Sea. Marine Biology Research, 7(2):208–211, 2011. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2010.489616.

Anonymous:2011:ICW

[317] Anonymous. The 9th international conference and workshop on lobster biology and management. Marine Biology Research, 7(2):212, 2011. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/17451000.2011.562383.

Roleda:2011:SGS

[318] Michael Y. Roleda and Dirk Dethleff. Storm-generated sediment deposition on rocky shores: Simulating burial effects on the physiology and morphology of Saccharina latissima sporophytes. Marine Biology Research, 7(3):213–223, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.497189.

Eilertsen:2011:DAF

[319] Mette Eilertsen, Kjell Magnus Norderhaug, and Kjersti Sjøtun. Does the amphipod fauna associated with epiphytes on kelp (Laminaria hyperborea) change with depth? Marine Biology Research, 7(3): 224–234, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.492592.

Aguilera:2011:RPS

[320] Victor M. Aguilera, Katty Donoso, and Rubén Escribano. Reproductive performance of small-sized dominant copepods with a highly variable food resource in the coastal upwelling system off the Chilean Humboldt Current. Marine Biology Research, 7(3):235–249, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2010.499437.

Gonzalez-Wanguemert:2011:PHW

[321] Mercedes González-Wangüemert, Elsa Froufe, Angel Pérez-Ruzafa, and Paulo Alexandrino. Phylogeographical history of the white seabream Diplodus sargus (Sparidae): Implications for insularity. Marine Biology Research, 7(3):250–260, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.499438.

Machreki-Ajmi:2011:VML

[322] Monia Machreki-Ajmi, Tarek Rebai, and Amel Hamza-Chaffai. Variation of metallothionein-like protein and metal concentrations during the reproductive cycle of the cockle Cerastoderma glaucum from an uncontaminated site: a 1-year study in the Gulf of Gabès area (Tunisia). Marine Biology Research, 7(3):261–271, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.497187.

Goffredo:2011:CPB

[323] Stefano Goffredo, Erik Caroselli, Gabriella Gasparini, Giulia Marconi, Maria Teresa Putignano, Claudia Pazzini, and Francesco Zaccanti. Colony and polyp biometry and size structure in the orange coral Astroides calycularis (Scleractinia: Dendrophylliidae). Marine Biology Research, 7(3):272–280, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2010.492222.

Gjosaeter:2011:AGO

[324] Jakob Gjøsæter and Didrik S. Danielssen. Age, growth and otolith annulus formation of cod (Gadus morhua) in the Risør area on the Norwegian Skagerrak coast during 1986–1996. Marine Biology Research, 7 (3):281–288, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.492225.

Filina:2011:SPM

[325] Elena A. Filina. Spermatogenesis and physiological maturity of male red king crab (Paralithodes camtschaticus Tilesius, 1815) and snow crab (Chionoecetes opilio Fabricius, 1788) in the Barents Sea. Marine Biology Research, 7(3):289–296, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.497188.

Davis:2011:TAB

[326] Matthew P. Davis and Prosanta Chakrabarty. Tripodfish (Aulopiformes: Bathypterois) locomotion and landing behaviour from video observation at bathypelagic depths in the Campos Basin of Brazil. Marine Biology Research, 7(3):297–303, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.515231.

Ye:2011:OFS

[327] Haihui Ye, Jun Ma, Qiongwu Lin, and Guizhong Wang. Occurrence of follicle-stimulating hormone-like substance in the kuruma prawn, Marsupenaeus japonicus, during ovarian maturation. Marine Biology Research, 7(3):304–309, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.515230.

Ferno:2011:VSS

[328] Anders Fernö, Terje Jørgensen, Svein Løkkeborg, and Paul D. Winger. Variable swimming speeds in individual Atlantic cod (Gadus morhua L.) determined by high-resolution acoustic tracking. Marine Biology Research, 7(3):310–313, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2010.492223.

Bello:2011:HBC

[329] Giambattista Bello, Andrea Travaglini, and Flegra Bentivegna. Histioteuthis bonnellii (Cephalopoda: Histioteuthidae): a new prey item of the leatherback turtle Dermochelys coriacea (Reptilia: Dermochelidae). Marine Biology Research, 7(3):314–316, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2010.492224.

Devereux:2011:ITT

[330] Richard Devereux, Diane F. Yates, Jessica Aukamp, Robert L. Quarles, Stephen J. Jordan, Roman S. Stanley, and Peter M. Eldridge. Interactions of Thalassia testudinum and sediment biogeochemistry in Santa Rosa Sound, NW Florida. Marine Biology Research, 7(4): 317–331, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.515227.

Biju:2011:MCE

[331] Abraham Biju and Saramma Usha Panampunnayil. Mysids (Crustacea) from the Exclusive Economic Zone of India with description of a new species. Marine Biology Research, 7(4):332–364, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2010.515228.

Barbini:2011:FHH

[332] Santiago A. Barbini, Luis O. Lucifora, and Natalia M. Hozbor. Feeding habits and habitat selectivity of the shortnose guitarfish, Zapteryx brevirostris (Chondrichthyes, Rhinobatidae), off north Argentina and Uruguay. Marine Biology Research, 7(4):365–377, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2010.515229.

Casas:2011:SAD

[333] Laura Casas, Laura Sánchez, and László Orbán. Sex-associated DNA markers from turbot. Marine Biology Research, 7(4):378–387, 2011. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2010.515226.

Jiang:2011:ECA

[334] Aili Jiang, Zhen Yu, and Changhai Wang. Effect of cadmium on the antioxidant system of the ascidian, Styela clava. Marine Biology Research, 7(4):388–397, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.515685.

Ramos:2011:PIB

[335] Ruth Ramos, Carolina Bastidas, Denise Debrot, and Elia García. Phase I and II biotransformation and antioxidant enzymes in the coral Siderastrea siderea act as biomarkers for reproductive condition and habitat quality. Marine Biology Research, 7(4):398–406, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2010.515684.

Li:2011:GCB

[336] Qi Li, Lin Yang, Qiaozhen Ke, and Lingfeng Kong. Gametogenic cycle and biochemical composition of the clam Mactra chinensis (Mollusca: Bivalvia): Implications for aquaculture and wild stock management. Marine Biology Research, 7(4):407–415, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2010.515686.

Whitfield:2011:CLC

[337] Alan K. Whitfield. Coastal lagoons — critical habitats of environmental change. Marine Biology Research, 7(4):416–417, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2010.538064.

Fenchel:2011:OA

[338] Tom Fenchel. Ocean acidification. Marine Biology Research, 7(4): 418–419, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.550051.

Brattegard:2011:NFN

[339] Torleiv Brattegard, Tore Høisæter, Kjersti Sjøtun, Tom Fenchel, and Franz Uiblein. Norwegian fjords: From natural history to ecosystem ecology and beyond. Marine Biology Research, 7(5):421–424, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.561355.

Veiga:2011:MCS

[340] Puri Veiga, Celia Besteiro, and Marcos Rubal. Meiobenthic communities structure at different spatial scales on sandy beaches from a ria environment: the role of abiotic factors. Marine Biology Research, 7 (5):425–435, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.527987.

Rashed-Un-Nabi:2011:TSD

[341] Md. Rashed-Un-Nabi, Md. Abdulla Al-Mamun, Md. Hadayet Ullah, and M. Golam Mustafa. Temporal and spatial distribution of fish and shrimp assemblage in the Bakkhali river estuary of Bangladesh in relation to some water quality parameters. Marine Biology Research, 7 (5):436–452, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.527988.

Trilles:2011:REE

[342] Jean-Paul Trilles and John E. Randall. Redescription of Elthusa emarginata (Bleeker, 1857) (Crustacea, Isopoda, Cymothoidae), type species of the genus Elthusa Schioedte and Meinert, 1884. Marine Biology Research, 7(5):453–465, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.528770.

Wenne:2011:AMT

[343] Roman Wenne, Luiza Handschuh, Anita Pocwierz-Kotus, Radosław Urbaniak, Piotr Formanowicz, Joanna Całkiewicz, Katarzyna Brzozowska, Marek Figlerowicz, Grzegorz Wegrzyn, and Borys Wróbel. The application of microarray technology to the identification of tc1-like element sequences in fish genomes. Marine Biology Research, 7(5): 466–477, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.527990.

Sequeira:2011:GMT

[344] Vera Sequeira, Silvia Vila, Ana Neves, Pedro Rifes, Maria Leonor Nunes, Ana Rita Vieira, Rafaela Barros Paiva, Marta Muñoz, and Leonel Serrano Gordo. The gelatinous matrix of the teleost Helicolenus dactylopterus dactylopterus (Delaroche, 1809) in the context of its reproductive strategy. Marine Biology Research, 7(5):478–487, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2010.528769.

Cerdenares-LadronDeGuevara:2011:AGS

[345] Genoveva Cerdenares-Ladrón De Guevara, Enrique Morales-Bojórquez, and Rubén Rodríguez-Sánchez. Age and growth of the sailfish Istiophorus platypterus (Istiophoridae) in the Gulf of Tehuantepec, Mexico. Marine Biology Research, 7(5):488–499, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.528201.

Carman:2011:ACW

[346] Victoria González Carman, Karina C. Álvarez, Laura Prosdocimi, María C. Inchaurraga, Rubén F. Dellacasa, Adrián Faiella, Cintia Echenique, Raúl González, Julián Andrejuk, Hermes W. Mianzan, Claudio Campagna, and Diego A. Albareda. Argentinian coastal waters: a temperate habitat for three species of threatened sea turtles. Marine Biology Research, 7(5):500–508, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.528772.

Leitao:2011:CBR

[347] Francisco Miguel Leitão and Miguel Baptista Gaspar. Comparison of the burrowing response of undersized cockles (Cerastoderma edule) after fishing disturbance caused by hand dredge and harvesting knife. Marine Biology Research, 7(5):509–514, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.527989.

Cagnon:2011:TGV

[348] Christine Cagnon, Beatrice Lauga, Solange Karama, and Claude Mouches. Temporal genetic variation in European eel Anguilla anguilla (Linnaeus, 1748): a fine scale investigation in the adour estuary. Marine Biology Research, 7(5):515–519, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.535832.

Baggio:2011:GEM

[349] Rafael A. Baggio, Maria W. Pil, Walter A. Boeger, Luciana A. Patella, Antonio Ostrensky, and Marcio R. Pie. Genetic evidence for multiple paternity in the mangrove land crab Ucides cordatus (Decapoda: Ocypodidae). Marine Biology Research, 7(5):520–524, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2010.528771.

Bergmann:2011:MAT

[350] Melanie Bergmann, Nina Langwald, Jörg Ontrup, Thomas Soltwedel, Ingo Schewe, Michael Klages, and Tim W. Nattkemper. Megafaunal assemblages from two shelf stations west of Svalbard. Marine Biology Research, 7(6):525–539, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.535834.

Morton:2011:BFM

[351] Brian Morton. The biology and functional morphology of Arctica islandica (Bivalvia: Arcticidae) — a gerontophilic living fossil. Marine Biology Research, 7(6):540–553, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.535833. See comments [391].

Allen:2011:TNS

[352] Gerald R. Allen and John E. Randall. Two new species of shrimpassociated gobies (Gobiidae: Cryptocentrus) from the Western Pacific. Marine Biology Research, 7(6):554–564, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2010.530545.

Findlay:2011:CIH

[353] Helen S. Findlay, Hannah L. Wood, Michael A. Kendall, John I. Spicer, Richard J. Twitchett, and Stephen Widdicombe. Comparing the impact of high CO_{2} on calcium carbonate structures in different marine organisms. Marine Biology Research, 7(6):565–575, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2010.547200.

Horn:2011:CDS

[354] Peter L. Horn, Tim Burrell, Amelia Connell, and Matthew R. Dunn. A comparison of the diets of silver (Seriolella punctata) and white (Seriolella caerulea) warehou. Marine Biology Research, 7(6):576– 591, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.528776.

Xu:2011:ASC

[355] Zhi-Huan Xu, Gui-Zhong Wang, Qian Mu, Li-Sheng Wu, and Shao-Jing Li. An approach to the study of copepod egg banks based on efficient DNA extraction from individual copepod eggs. Marine Biology Research, 7(6):592–598, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.528773.

Yaragina:2011:PNA

[356] Natalia A. Yaragina, Hans Høie, Valentina P. Koloskova, Hildegunn Mjanger, Kjell H. Nedreaas, Harald Senneset, Natalya V. Zuykova, and Per Ågotnes. Precision of the Northeast Arctic cod age determination under variable environmental and information conditions. Marine Biology Research, 7(6):599–607, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.538062.

Escolar:2011:SDB

[357] Mariana Escolar, Daniel R. Hernández, and Claudia Bremec. Spatial distribution, biomass and size structure of Ctenodiscus australis (Echinodermata: Asteroidea) in shelf-break areas, Argentine. Marine Biology Research, 7(6):608–616, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.528774.

Zeng:2011:IDE

[358] Hui Zeng, Jingru Huang, Wenxing Li, Huiyang Huang, and Haihui Ye. Identification of differentially expressed genes in the thoracic ganglion of the mud crab, Scylla paramamosain during ovarian maturation. Marine Biology Research, 7(6):617–622, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2010.528775.

Sun:2011:EKA

[359] Hongyan Sun, Lvping Zhang, Chunhua Ren, Chang Chen, Sigang Fan, Jian Jun Xia, Hongjun Lin, and Chaoqun Hu. The expression of Na, K-ATPase in Litopenaeus vannamei under salinity stress. Marine Biology Research, 7(6):623–628, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2010.538063.

Rowe:2011:AVR

[360] Gilbert T. Rowe and Jody W. Deming. An alternative view of the role of heterotrophic microbes in the cycling of organic matter in deepsea sediments. Marine Biology Research, 7(7):629–636, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.560269.

Uiblein:2011:NGU

[361] Franz Uiblein and Phillip C. Heemstra. A new goatfish, Upeneus seychellensis sp. nov. (Mullidae), from the Seychelles Bank, with remarks on Upeneus guttatus and a key to Western Indian Ocean Upeneus species. Marine Biology Research, 7(7):637–650, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2010.547202.

Zagal:2011:NRS

[362] Carolina J. Zagal, Yayoi M. Hirano, Claudia E. Mills, Graham J. Edgar, and Neville S. Barrett. New records of Staurozoa from Australian coastal waters, with a description of a new species of Lucernariopsis Uchida, 1929 (Cnidaria, Staurozoa, Stauromedusae) and a key to Australian Stauromedusae. Marine Biology Research, 7(7):651–666, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.558097.

Trabelsi:2011:INC

[363] Asma Trabelsi and Fereidoun Rassoulzadegan. Inorganic nutrient control of dissolved organic carbon (DOC) dynamics in NW Mediterranean waters: An experimental approach. Marine Biology Research, 7(7): 667–676, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.554557.

Zuniga-Flores:2011:RDC

[364] Marcela S. Zúñiga-Flores, Sofía Ortega-García, Ma. Del Carmen Rodríguez-Jaramillo, and Juana López-Martínez. Reproductive dynamics of the common dolphinfish Coryphaena hippurus in the southern Gulf of California. Marine Biology Research, 7(7):677–689, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.554558.

Kalkan:2011:BSB

[365] Evrim Kalkan, Asli Kurtulus, Öncü Maraci, and Rasit Bilgin. Is the Bosphorus Strait a barrier to gene flow for the Mediterranean mussel, Mytilus galloprovincialis (Lamarck, 1819)? Marine Biology Research, 7 (7):690–700, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.558098.

Szabo:2011:SHS

[366] Zoltán Szabó, Bethany K. Kimokeo, Robert J. Toonen, and John E. Randall. On the status of the Hawaiian seahorses Hippocampus hilonis, H. histrix and H. fisheri (Syngnathidae). Marine Biology Research, 7 (7):701–709, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.558096.

Tovar-Hernandez:2011:RIF

[367] MarÍa Ana Tovar-Hernández, Beatriz Yáñez-Rivera, and José Luis Bortolini-Rosales. Reproduction of the invasive fan worm Branchiomma bairdi (Polychaeta: Sabellidae). Marine Biology Research, 7(7):710– 718, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2010.547201.

Lahbib:2011:SID

[368] Youssef Lahbib, Sami Abidli, and Najoua Trigui El Menif. Spawning and intracapsular development of Stramonita haemastoma haemastoma (Gastropoda: Muricidae) collected in northern Tunisia. Marine Biology Research, 7(7):719–726, 2011. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.558099.

Blanco:2011:HID

[369] Camila G. Blanco, João B. L. Gusmão-Junior, Ronaldo A. Christofoletti, and Tãnia M. Costa. Hydrodynamism and its influence on the density of the decorator crab Microphrys bicornutus (Mithracidae) on intertidal rocky shores from a subtropical region. Marine Biology Research, 7(7):727–731, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.558100.

Anonymous:2011:RA

[370] Anonymous. Reviewer acknowledgement 2010. Marine Biology Research, 7(7):732, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.617374.

Julshamn:2011:CMA

[371] Kaare Julshamn, Sylvia Frantzen, Stig Valdersnes, Bente Nilsen, Amund Maage, and Kjell Nedreaas. Concentrations of mercury, arsenic, cadmium and lead in Greenland halibut (Reinhardtius hippoglossoides) caught off the coast of northern Norway. Marine Biology Research, 7 (8):733–745, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.594893.

Miranda:2011:SCS

[372] Richard M. Miranda, Viviana Espinoza, Jessica Dörner, Ana FarÍas, and Iker Uriarte. Sibling cannibalism on the small octopus Robsonella fontaniana (d’Orbigny, 1834) paralarvae. Marine Biology Research, 7 (8):746–756, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.596543.

Helguera:2011:DPM

[373] Yusmila Helguera, Lisbet Díaz-Asencio, Raúl Fernández-Garcés, Miguel Gómez-Batista, Aniel Guillén, Misael Díaz-Asencio, and Maickel Armenteros. Distribution patterns of macrofaunal polychaete assemblages in a polluted semi-enclosed bay: Cienfuegos, Caribbean Sea. Marine Biology Research, 7(8):757–768, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2011.569552.

Dirks:2011:NSS

[374] Ulrich Dirks, Harald R. Gruber-Vodicka, Nikolaus Leisch, Wolfgang Sterrer, and Jörg A. Ott. A new species of symbiotic flatworms, Paracatenula galateia sp. nov. (Platyhelminthes: Catenulida: Retronectidae) from Belize (Central America). Marine Biology Research, 7(8): 769–777, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.574880.

Stark:2011:CMG

[375] James W. Stark. Contrasting the maturation, growth, spatial distribution and vulnerability to environmental warming of Hippoglossoides robustus (Bering flounder) with H. elassodon (flathead sole) in the eastern Bering Sea. Marine Biology Research, 7(8):778–785, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.569554.

MullerPatraoDeOliveira:2011:GLP

[376] Otto Müller Patrão De Oliveira and Antonio Carlos Marques. Global and local patterns in the use of macrophytes as substrata by hydroids (Hydrozoa: Anthoathecata and Leptothecata). Marine Biology Research, 7(8):786–795, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.578647.

Xu:2011:ESL

[377] Fei Xu, Ximing Guo, Li Li, and Guofan Zhang. Effects of salinity on larvae of the oysters Crassostrea ariakensis, C. sikamea and the hybrid cross. Marine Biology Research, 7(8):796–803, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2011.569555.

Jaubet:2011:SRP

[378] MarÍa L. Jaubet and Gabriel N. Genzano. Seasonality and reproductive periods of the hydroid Clytia gracilis in temperate littoral ecosystems. Is asexual reproduction the prime mechanism in maintaining populations? Marine Biology Research, 7(8):804–811, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.578650.

Rodrigues:2011:SAA

[379] Marcelo Rodrigues, Manuel E. Garci, Jesús S. Troncoso, and Ángel Guerra. Seasonal abundance of the Atlantic bobtail squid Sepiola atlantica in Galician waters (NE Atlantic). Marine Biology Research, 7 (8):812–819, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.578649.

Veiga:2011:DBM

[380] Pedro Veiga, José C. Xavier, Carlos A. Assis, and Karim Erzini. Diet of the blue marlin, Makaira nigricans, off the south coast of Portugal. Marine Biology Research, 7(8):820–825, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.578648.

Li:2011:MDM

[381] Wei-Ye Li, Jing Zhong, Wei Xu, and Yi-Quan Wang. Microsatellite DNA marker development and genetic diversity of Branchiostoma belcheri in Xiamen waters. Marine Biology Research, 7(8):826–831, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.569553.

Anonymous:2011:EB

[382] Anonymous. Editorial board. Marine Biology Research, 7(8): ??, 2011. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
17451000.2011.633400.

Fenchel:2012:MBR

[383] Tom Fenchel and Franz Uiblein. Marine Biology Research — with good performance towards new shores! Marine Biology Research, 8 (1):1–2, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.623704.

Lee:2012:FLB

[384] Won Je Lee. Free-living benthic heterotrophic euglenids from Botany Bay, Australia. Marine Biology Research, 8(1):3–27, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.596545.

Hoyoux:2012:DGM

[385] Caroline Hoyoux, Magali Zbinden, Sarah Samadi, Françoise Gaill, and Philippe Compère. Diet and gut microorganisms of Munidopsis squat lobsters associated with natural woods and mesh-enclosed substrates in the deep South Pacific. Marine Biology Research, 8(1):28–47, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2011.605144.

Hirai:2012:TSD

[386] Junya Hirai and Daniel O. B. Jones. The temporal and spatial distribution of krill (Meganyctiphanes norvegica) at the deep seabed of the Faroe– Shetland Channel, UK: a potential mechanism for rapid carbon flux to deep sea communities. Marine Biology Research, 8(1):48–60, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2011.594891.

Alikunhi:2012:PPI

[387] Nabeel M. Alikunhi and Kandasamy Kathiresan. Phytoplankton productivity in interlinked mangroves, seagrass and coral reefs and its ecotones in Gulf of Mannar Biosphere Reserve South east India. Marine Biology Research, 8(1):61–73, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.596544.

Czudaj:2012:DAR

[388] Stephanie Czudaj, João Pereira, Ana Moreno, Ana Maria Costa, Ulrich Saint-Paul, and Rui Rosa. Distribution, abundance, reproduction and ageing of the common bobtail squid Sepietta oweniana (Sepiolidae, Cephalopoda) from the Portuguese Coast. Marine Biology Research, 8(1):74–86, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.594892.

Gunnarsson:2012:BAF

[389] Karl Gunnarsson and Anders Berglund. The brown alga Fucus radicans suffers heavy grazing by the isopod Idotea baltica. Marine Biology Research, 8(1):87–89, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.594890.

Martin:2012:MPP

[390] R. Aidan Martin and Neil Hammerschlag. Marine predator-prey contests: Ambush and speed versus vigilance and agility. Marine Biology Research, 8(1):90–94, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.614255.

Ridgway:2012:CBF

[391] Iain D. Ridgway. Comments on ‘The biology and functional morphology of Arctica islandica ’ by Brian Morton, Marine Biology Research, 2011. Marine Biology Research, 8(1):95–97, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.618189. See [351].

Sneli:2012:SGS

[392] Jon-Arne Sneli. A student’s guide to the seashore. Marine Biology Research, 8(1):98–99, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2011.616208.

Fenchel:2012:MME

[393] Tom Fenchel. Marine microbiology: ecology and applications Colin munn (with foreword by farooq azam). Marine Biology Research, 8 (1):100, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.626054.

Anonymous:2012:IA

[394] Anonymous. Instructions for authors. Marine Biology Research, 8 (1):??, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
17451000.2012.637399.

Ross:2012:UMR

[395] Steve W. Ross, Michael C. T. Carlson, and Andrea M. Quattrini. The utility of museum records for documenting distributions of deep-sea corals off the southeastern United States. Marine Biology Research, 8 (2):101–114, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.628680.

Smale:2012:MHU

[396] Malcolm J. Smale, Anthony J. Booth, Michael R. Farquhar, Michael R. Meÿer, and Lesley Rochat. Migration and habitat use of formerly captive and wild raggedtooth sharks (Carcharias taurus) on the southeast coast of South Africa. Marine Biology Research, 8(2):115–128, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.617756.

Bedini:2012:ECU

[397] Roberto Bedini and Luigi Piazzi. Evaluation of the concurrent use of multiple descriptors to detect anthropogenic impacts in marine coastal systems. Marine Biology Research, 8(2):129–140, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.615324.

Vasconcelos:2012:AUO

[398] Paulo Vasconcelos, Ines H. Gharsallah, Paula Moura, Nedra Zamouri-Langar, Adel Gaamour, Hechmi Missaoui, Othman Jarboui, and Miguel B. Gaspar. Appraisal of the usefulness of operculum growth marks for ageing Hexaplex trunculus (Gastropoda: Muricidae): Comparison between surface striae and adventitious layers. Marine Biology Research, 8(2):141–153, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.616896.

Bapary:2012:FAP

[399] Mohammad Abu Jafor Bapary, Md. Nurul Amin, and Akihiro Takemura. Food availability as a possible determinant for initiation and termination of reproductive activity in the tropical damselfish Chrysiptera cyanea. Marine Biology Research, 8(2):154–162, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.605146.

Li:2012:LVE

[400] Gang Li, Liangmin Huang, Huaxue Liu, Zhixin Ke, Qiang Lin, Guangyan Ni, Jianqiang Yin, Kaizhi Li, Xingyu Song, Pingping Shen, and Yehui Tan. Latitudinal variability (6^{o}S–20^{o}N) of early summer phytoplankton species compositions and size-fractioned productivity from Java Sea to South China Sea. Marine Biology Research, 8(2):163–171, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.615323.

Zaidman:2012:RPS

[401] Paula C. Zaidman, Marina A. Kroeck, Erica M. Oehrens Kissner, and Enrique M. Morsan. Reproductive pattern of Southern geoduck, Panopea abbreviata, at El Sótano (San Matías Gulf, Patagonia, Argentina). Marine Biology Research, 8(2):172–181, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.605145.

Madhu:2012:ECP

[402] Nikathithara V. Madhu, Retnamma Jyothibabu, Padinjaratte A. Maheswaran, Kadeparambil A. Jayaraj, and Chittur T. Achuthankutty. Enhanced chlorophyll a and primary production in the northern Arabian Sea during the spring intermonsoon due to green Noctiluca scintillans bloom. Marine Biology Research, 8(2):182–188, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.605143.

Piontkovski:2012:ICA

[403] Sergey A. Piontkovski and Michel R. Claereboudt. Interannual changes of the Arabian Sea productivity. Marine Biology Research, 8(2): 189–194, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.615325.

Burgos-Aceves:2012:LDM

[404] Mario A. Burgos-Aceves, Rafael Campos-Ramos, and Danitzia A. Guerrero-Tortolero. Leukophilia during maturation of the leopard grouper Mycteroperca rosacea. Marine Biology Research, 8(2):195– 200, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.596546.

Skjaeraasen:2012:SSC

[405] Jon Egil Skjæraasen, Justin J. Meager, and Mikko Heino. Secondary sexual characteristics in codfishes (Gadidae) in relation to sound production, habitat use and social behaviour. Marine Biology Research, 8 (3):201–209, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.637562.

Sorensen:2012:GDH

[406] Martin V. Sørensen, María Herranz, Hyun Soo Rho, Won-Gi Min, Hiroshi Yamasaki, Nuria Sánchez, and Fernando Pardos. On the genus Dracoderes Higgins and Shirayama, 1990 (Kinorhyncha: Cyclorhagida) with a redescription of its type species, D. abei, and a description of a new species from Spain. Marine Biology Research, 8(3):210–232, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2011.615328.

Daly:2012:DEN

[407] Marymegan Daly, Renzo Perissinotto, Megan Laird, David Dyer, and Antonio Todaro. Description and ecology of a new species of Edwardsia de Quatrefages, 1842 (Anthozoa, Actiniaria) from the St Lucia Estuary, South Africa. Marine Biology Research, 8(3):233–245, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.617757.

IHomrum:2012:GMD

[408] Eydna Í Homrum, Bogi Hansen, Petur Steingrund, and Hjálmar Hátún. Growth, maturation, diet and distribution of saithe (Pollachius virens) in Faroese waters (NE Atlantic). Marine Biology Research, 8(3): 246–254, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.627921.

Schiariti:2012:RBL

[409] AgustÍn Schiariti, Ernesto Christiansen, Andre Carrara Morandini, Fábio Lang da Silveira, Diego Agustin Giberto, and Hermes Walter Mianzan. Reproductive biology of Lychnorhiza lucerna (Cnidaria: Scyphozoa: Rhizostomeae): Individual traits related to sexual reproduction. Marine Biology Research, 8(3):255–264, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.616897.

Leis:2012:OSA

[410] Jeffrey M. Leis, Philippe Balma, Rémi Ricoux, and René Galzin. Ontogeny of swimming ability in the European Sea Bass, Dicentrarchus labrax (L.) (Teleostei: Moronidae). Marine Biology Research, 8(3): 265–272, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.616898.

Pessanha:2012:SSF

[411] André Luiz Machado Pessanha and Francisco Gerson Araújo. Spatial and size feeding niche partitioning of the rhomboid mojarra Diapterus rhombeus (Cuvier, 1829) in a tropical Brazilian Bay. Marine Biology Research, 8(3):273–283, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.615326.

Rundell:2012:DPP

[412] Rebecca J. Rundell and Brian S. Leander. Description and phylogenetic position of the first sand-dwelling entoproct from the western coast of North America: Loxosomella vancouverensis sp. nov. Marine Biology Research, 8(3):284–291, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.619545.

Mabragana:2012:RBA

[413] Ezequiel Mabragaña, Pedro Ibáñez, and María B. Cousseau. Reproductive biology and abundance of the freckled sandskate Psammobatis lentiginosa McEachran,1983 in the southwest Atlantic. Marine Biology Research, 8(3):292–299, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.617043.

Thornycroft:2012:CAI

[414] Rosanne E. Thornycroft and Anthony J. Booth. Computer-aided identification of coelacanths, Latimeria chalumnae, using scale patterns. Marine Biology Research, 8(3):300–306, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2011.628679.

McBride:2012:SFM

[415] Margaret Mary McBride. Sustainable fisheries: Multi-level approaches to a global problem. Marine Biology Research, 8(3):307–308, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.647920.

Moy:2012:LSS

[416] Frithjof E. Moy and Hartvig Christie. Large-scale shift from sugar kelp (Saccharina latissima) to ephemeral algae along the south and west coast of Norway. Marine Biology Research, 8(4):309–321, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.637561.

Forslund:2012:GGR

[417] Helena Forslund, Ove Eriksson, and Lena Kautsky. Grazing and geographic range of the Baltic seaweed Fucus radicans (Phaeophyceae). Marine Biology Research, 8(4):322–330, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.637565.

Hernandez-Becerril:2012:SPS

[418] David U. Hernández-Becerril, Aldo Aquino-Cruz, David A. Salas-De-León, Martha Signoret-Poillon, and M. Adela Monreal-Gómez. Studies on picophytoplankton in the southern Gulf of Mexico: pigment analysis and potential importance of the picoeukaryote prasinophyte Micromonas pusilla. Marine Biology Research, 8(4):331–340, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.638639.

Aguirre:2012:SVM

[419] Gastón E. Aguirre, Fabiana L. Capitanio, Gustavo A. Lovrich, and Graciela B. Esnal. Seasonal variability of metazooplankton in coastal sub-Antarctic waters (Beagle Channel). Marine Biology Research, 8 (4):341–353, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.627922.

SampaioSantAnna:2012:RMP

[420] Bruno Sampaio Sant’Anna, Alexander Turra, and Fernando José Zara. Reproductive migration and population dynamics of the blue crab Callinectes danae in an estuary in southeastern Brazil. Marine Biology Research, 8(4):354–362, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2011.637563.

Sepulveda:2012:CVS

[421] Roger D. Sepúlveda and Christian M. Ibáñez. Clinal variation in the shell morphology of intertidal snail Acanthina monodon in the Southeastern Pacific Ocean. Marine Biology Research, 8(4):363–372, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.634813.

Midling:2012:SHA

[422] Kjell Øivind Midling, Christian Koren, Odd-Børre Humborstad, and Bjørn-Steinar Sæther. Swimbladder healing in Atlantic cod (Gadus morhua), after decompression and rupture in capture-based aquaculture. Marine Biology Research, 8(4):373–379, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.638640.

DelMarSoler-Hurtado:2012:TNG

[423] Maria Del Mar Soler-Hurtado and Pablo J. López-González. Two new gorgonian species (Anthozoa: Octocorallia: Gorgoniidae) from Ecuador (Eastern Pacific). Marine Biology Research, 8(4):380–387, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.634814.

Kadri:2012:AGR

[424] Hasna Kadri, Sondes Marouani, Bechir Saïdi, Mohamed Nejmeddine Bradai, Mohamed Ghorbel, Abderrahmen Bouaïn, and Eric Morize. Age, growth and reproduction of Raja miraletus (Linnaeus, 1758) (Chondrichthyes: Rajidae) of the Gulf of Gabès (Tunisia, Central Mediterranean Sea). Marine Biology Research, 8(4):388–396, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.619546.

BragaCastro:2012:TBI

[425] Ítalo Braga Castro, Cristina De Almeida Rocha-Barreira, Marcos Antônio Fernandez, and Gregório Bigatti. Transplant bioassay induces different imposex responses in two species of the genus Stramonita. Marine Biology Research, 8(4):397–404, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.627923.

El-Ganainy:2012:PBP

[426] Azza A. El-Ganainy and Mohamed H. Yassien. The population biology of penaeid prawns in the Gulf of Suez, Red Sea, Egypt. Marine Biology Research, 8(4):405–411, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.637564.

Anonymous:2012:MF

[427] Anonymous. Marine fish 2012. Marine Biology Research, 8(4): 412, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
17451000.2012.669605.

Anonymous:2012:EIB

[428] Anonymous. Effects of interactions between fish populations on ecosystem dynamics in the Norwegian Sea — results of the INFERNO project. Marine Biology Research, 8(5–6):413–414, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2012.660165.

Huse:2012:EIB

[429] Geir Huse, Jens Christian Holst, Kjell Utne, Leif Nøttestad, Webjørn Melle, Aril Slotte, Geir Ottersen, Tom Fenchel, and Franz Uiblein. Effects of interactions between fish populations on ecosystem dynamics in the Norwegian Sea — results of the INFERNO project. Marine Biology Research, 8(5–6):415–419, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.653372.

Utne:2012:HDO

[430] Kjell Rong Utne, Geir Huse, Geir Ottersen, Jens Christian Holst, Vladimir Zabavnikov, Jan Arge Jacobsen, Gudmundur J. Óskarsson, and Leif Nøttestad. Horizontal distribution and overlap of planktivorous fish stocks in the Norwegian Sea during summers 1995–2006. Marine Biology Research, 8(5–6):420–441, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.640937.

Langoy:2012:ODD

[431] Herdis Langøy, Leif Nøttestad, Georg Skaret, Cecilie Broms, and Anders Fernö. Overlap in distribution and diets of Atlantic mackerel (Scomber scombrus), Norwegian spring-spawning herring (Clupea harengus) and blue whiting (Micromesistius poutassou) in the Norwegian Sea during late summer. Marine Biology Research, 8(5–6):442–460, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.642803.

Broms:2012:NMH

[432] Cecilie Broms, Webjørn Melle, and John K. Horne. Navigation mechanisms of herring during feeding migration: the role of ecological gradients on an oceanic scale. Marine Biology Research, 8(5–6):461–474, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2011.640689.

Payne:2012:RFA

[433] Mark R. Payne, Afra Egan, Sascha M. M. Fässler, Hjálmar Hátún, Jens Christian Holst, Jan Arge Jacobsen, Aril Slotte, and Harald Loeng. The rise and fall of the NE Atlantic blue whiting (Micromesistius poutassou). Marine Biology Research, 8(5–6):475–487, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2011.639778.

Huse:2012:VDH

[434] Geir Huse, Kjell Rong Utne, and Anders Fernö. Vertical distribution of herring and blue whiting in the Norwegian Sea. Marine Biology Research, 8(5–6):488–501, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.639779.

Soiland:2012:URF

[435] Henrik Søiland and Geir Huse. Using RAFOS floats to simulate overwinter transport of Calanus finmarchicus in the Norwegian Sea. Marine Biology Research, 8(5–6):502–507, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2011.639780.

Hjollo:2012:MSP

[436] Solfrid Sætre Hjøllo, Geir Huse, Morten D. Skogen, and Webjørn Melle. Modelling secondary production in the Norwegian Sea with a fully coupled physical/primary production/individual-based Calanus finmarchicus model system. Marine Biology Research, 8(5–6):508–526, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2011.642805.

Utne:2012:ECC

[437] Kjell Rong Utne, Solfrid Sætre Hjøllo, Geir Huse, and Morten Skogen. Estimating the consumption of Calanus finmarchicus by planktivorous fish in the Norwegian Sea using a fully coupled 3D model system. Marine Biology Research, 8(5–6):527–547, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2011.642804.

Utne:2012:EHT

[438] Kjell Rong Utne and Geir Huse. Estimating the horizontal and temporal overlap of pelagic fish distribution in the Norwegian Sea using individualbased modelling. Marine Biology Research, 8(5–6):548–567, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2011.639781.

Huse:2012:EEN

[439] Geir Huse. Ecosystem ecology: a new synthesis. Marine Biology Research, 8(5–6):568–569, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.639782.

Anonymous:2012:CS

[440] Anonymous. Christoffer Schander (1960–2012). Marine Biology Research, 8(5–6):570–572, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.685559.

DeCampos:2012:SSE

[441] Carlos José Alexandre De Campos, Alvaro Esteves Migotto, Ulisses Pinheiro, and Antonio Carlos Marques. Sponges as substrata and early life history of the tubulariid Zyzzyzus warreni (Cnidaria: Hydrozoa) in the São Sebastião Channel, Brazil. Marine Biology Research, 8(7): 573–583, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.638641.

Moreira:2012:TNS

[442] Juan Moreira and Julio Parapar. Two new species of Sphaerodoropsis Hartman and Fauchald, 1971 (Polychaeta: Sphaerodoridae) from Iceland (BIOICE programme). Marine Biology Research, 8(7):584– 593, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.638929.

Lopez-Gonzalez:2012:RCS

[443] Pablo J. López-González, Jordi Grinyó, and Josep-Maria Gili. Rediscovery of Cereopsis studeri Koch, 1891, a forgotten Mediterranean soft coral species, and its inclusion in the genus Nidalia gray, 1835 (Octocorallia, Alcyonacea, Nidaliidae). Marine Biology Research, 8(7):

594–604, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2011.650178.

Ferreira:2012:VMN

[444] A. S. de A. Ferreira, Erling Kåre Stenevik, Knut Wiik Vollset, Rolf Korneliussen, and Arild Folkvord. Vertical migration of Norwegian springspawning herring larvae in relation to predator and prey distribution. Marine Biology Research, 8(7):605–614, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2011.653370.

Pond:2012:DGR

[445] David W. Pond, Geraint A. Tarling, Katrin Schmidt, and Inigo Everson. Diet and growth rates of Meganyctiphanes norvegica in autumn. Marine Biology Research, 8(7):615–623, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2011.653366.

Horn:2012:DPT

[446] Peter L. Horn, Jeff S. Forman, and Matthew R. Dunn. Dietary partitioning by two sympatric fish species, red cod (Pseudophycis bachus) and sea perch (Helicolenus percoides), on Chatham Rise, New Zealand. Marine Biology Research, 8(7):624–634, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2011.653543.

Hernaez:2012:RBG

[447] Patricio Hernáez, Edgar Villegas-Jiménez, Fresia Villalobos-Rojas, and Ingo S. Wehrtmann. Reproductive biology of the ghost shrimp Lepidophthalmus bocourti (A. Milne-Edwards, 1870) (Decapoda: Axiidea: Callianassidae): a tropical species with a seasonal reproduction. Marine Biology Research, 8(7):635–643, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2011.653369.

Yigin:2012:AGR

[448] C. Cigdem Yigin and Ali Ismen. Age, growth and reproduction of the common stingray, Dasyatis pastinaca from the North Aegean Sea. Marine Biology Research, 8(7):644–653, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.659667.

Riascos:2012:TSR

[449] Jose M. Riascos, Cecilia M. Avalos, Aldo S. Pacheco, and Olaf Heilmayer. Testing stress responses of the bivalve Protothaca thaca to El niño-la niña thermal conditions. Marine Biology Research, 8(7):654–661, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2011.653367.

SampaioSantAnna:2012:SPB

[450] Bruno Sampaio Sant’Anna, Laura Cristina Da Cruz Dominciano, Sabrina Fernandes Buozi, and Alexander Turra. Is shell partitioning between the hermit crabs Pagurus brevidactylus and Pagurus criniticornis explained by interference and/or exploitation competition? Marine Biology Research, 8(7):662–669, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.653371.

Li:2012:VUI

[451] Gang Li and Kunshan Gao. Variation in UV irradiance related to stratospheric ozone levels affects photosynthetic carbon fixation of winter phytoplankton assemblages from surface coastal water of the South China Sea. Marine Biology Research, 8(7):670–676, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2011.653368.

Heinz:2012:KSP

[452] Petra Heinz, Roman A. Marten, Valiyaparambil N. Linshy, Timo Haap, Emmanuelle Geslin, and Heinz-R. Köhler. 70 kD stress protein (Hsp70) analysis in living shallow-water benthic foraminifera. Marine Biology Research, 8(7):677–681, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2011.650177.

Riisgaard:2012:FDG

[453] Hans Ulrik Riisgård, Kim Lundgreen, and Poul S. Larsen. Field data and growth model for mussels Mytilus edulis in Danish waters. Marine Biology Research, 8(8):683–700, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.678857.

Carrasco:2012:DSF

[454] Nicola K. Carrasco, Renzo Perissinotto, and Holly A. Nel. Diet of selected fish species in the freshwater-deprived St Lucia Estuary, South Africa, assessed using stable isotopes. Marine Biology Research, 8(8):

701–714, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.678855.

Pralon:2012:HDM

[455] Bruno Gabriel Nunes Pralon, Gustavo Luis Hirose, and Maria Lucia Negreiros-Fransozo. Horizontal distribution of megalopae of Brachyura (Crustacea, Decapoda) in a South American tropical estuary. Marine Biology Research, 8(8):715–726, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.676183.

derMerwe:2012:MPS

[456] Aletta Elizabeth Bester van der Merwe, Maria Eugenia D’Amato, Belinda Louisa Swart, and Rouvay Roodt-Wilding. Molecular phylogeny of South African abalone, its origin and evolution as revealed by two genes. Marine Biology Research, 8(8):727–736, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.678856.

Wenger:2012:HPP

[457] Denise Wenger and Sven Koschinski. Harbour porpoise (Phocoena phocoena Linnaeus, 1758) entering the Weser river after decades of absence. Marine Biology Research, 8(8):737–745, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.676184.

Dong:2012:IEA

[458] Meitao Dong, Xiaowen Zhang, Jian Zou, Naihao Ye, Dong Xu, Shanli Mou, Chengwei Liang, and Wenqi Wang. Identification and expression analysis of the gene lhcSR associated with adaptation to light and low temperature stress in the green tide forming alga Ulva prolifera. Marine Biology Research, 8(8):746–755, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.676188.

Burdett:2012:ECA

[459] Heidi L. Burdett, Elena Aloisio, Piero Calosi, Helen S. Findlay, Stephen Widdicombe, Angela D. Hatton, and Nicholas A. Kamenos. The effect of chronic and acute low pH on the intracellular DMSP production and epithelial cell morphology of red coralline algae. Marine Biology Research, 8(8):756–763, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.676189.

Groeneveld:2012:HCL

[460] Johan C. Groeneveld, Sophie Von der Heyden, and Conrad A. Matthee. High connectivity and lack of mtDNA differentiation among two previously recognized spiny lobster species in the southern Atlantic and Indian Oceans. Marine Biology Research, 8(8):764–770, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.676185.

Sun:2012:ETS

[461] Xiujun Sun and Qi Li. Effects of temporary starvation on larval growth, survival and development of the sea cucumber Apostichopus japonicus. Marine Biology Research, 8(8):771–777, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.676186.

Penton:2012:PCP

[462] Paulette M. Penton and Gail K. Davoren. Physical characteristics of persistent deep-water spawning sites of capelin: Importance for delimiting critical marine habitats. Marine Biology Research, 8(8): 778–783, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.678858.

Lage:2012:DSC

[463] Sandra Lage, Joana Raimundo, Vanda Brotas, and Pedro Reis Costa. Detection and sub-cellular distribution of the amnesic shellfish toxin, domoic acid, in the digestive gland of Octopus vulgaris during periods of toxin absence. Marine Biology Research, 8(8):784–789, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.659668.

Rodriguez:2012:FES

[464] Estefanía Rodríguez. First endomyarian sea anemone (Cnidaria, Actiniaria) putatively from chemosynthetic environments: a new deep-sea genus and species from the North Pacific. Marine Biology Research, 8 (9):791–801, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.692162.

Sadighzadeh:2012:CDO

[465] Zahra Sadighzadeh, Victor M. Tuset, Tooraj Valinassab, Mohammad R. Dadpour, and Antoni Lombarte. Comparison of different otolith shape descriptors and morphometrics for the identification of closely related species of Lutjanus spp. from the Persian Gulf. Marine Biology Research, 8(9):802–814, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.692163.

Gallardo:2012:ELB

[466] Carlos S. Gallardo, Daniela Haro, Carolina Wagner, Orlando Garrido, and Juan I. Cañete. Egg-laying behaviour and intracapsular development of Argobuccinum pustulosum (Gastropoda: Ranellidae) in temperate waters at the South coast of Chile. Marine Biology Research, 8 (9):815–828, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.693615.

Jansen:2012:FEH

[467] Okka E. Jansen, Geert M. Aarts, Krishna Das, Gilles Lepoint, Loïc Michel, and Peter J. H. Reijnders. Feeding ecology of harbour porpoises: stable isotope analysis of carbon and nitrogen in muscle and bone. Marine Biology Research, 8(9):829–841, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.692164.

Murillo:2012:DSS

[468] Francisco Javier Murillo, Pablo Durán Muñoz, Javier Cristobo, Pilar Ríos, Concepción González, Ellen Kenchington, and Alberto Serrano. Deep-sea sponge grounds of the Flemish Cap, Flemish Pass and the Grand Banks of Newfoundland (Northwest Atlantic Ocean): Distribution and species composition. Marine Biology Research, 8(9): 842–854, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.682583.

Almanza:2012:CGK

[469] Viviana Almanza, Alejandro H. Buschmann, María C. Hernández-González, and Luís A. Henríquez. Can giant kelp (Macrocystis pyrifera) forests enhance invertebrate recruitment in southern Chile? Marine Biology Research, 8(9):855–864, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.692159.

Rabago-Quiroz:2012:FAB

[470] Carlos H. Rábago-Quiroz, Juana López-Martínez, José E. Valdez-Holguín, Manuel O. Nevárez-Martínez, and Alejandro Acevedo-Cervantes. Fish assemblages in the bycatch of bottom shrimp trawls on the west side of the Gulf of California, Mexico. Marine Biology Research, 8(9):865–876, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.692161.

Villanueva:2012:GSJ

[471] Ronald D. Villanueva, Maria Vanessa B. Baria, and Dexter W. dela Cruz. Growth and survivorship of juvenile corals outplanted to degraded reef areas in Bolinao-Anda Reef Complex, Philippines. Marine Biology Research, 8(9):877–884, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.682582.

Atherton:2012:TBS

[472] Sarah Atherton and Rick Hochberg. Tetranchyroderma bronchostylus sp. nov., the first known gastrotrich (Gastrotricha) with a sclerotic canal in the caudal organ. Marine Biology Research, 8(9):885–892, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.676187.

Lopes:2012:ROS

[473] Vanessa M. Lopes, Miguel Baptista, Marta S. Pimentel, Tiago Repolho, Luis Narciso, and Rui Rosa. Reproduction in Octocorallia: Synchronous spawning and asynchronous oogenesis in the pennatulid Veretillum cynomorium. Marine Biology Research, 8(9):893–900, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.693616.

Sanjuan:2012:MIA

[474] Andrés Sanjuán, Alejandro De Carlos, Cristina Rodríguez-Cabello, Rafael Bañón, Francisco Sánchez, and Alberto Serrano. Molecular identification of the arrowhead dogfish Deania profundorum (Centrophoridae) from the northern waters of the Iberian peninsula. Marine Biology Research, 8(9):901–905, 2012. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.692160.

LeaoFeitosa:2012:BOC

[475] João Lucas Leão Feitosa, Laís De Carvalho Teixeira Chaves, Pedro Henrique Cipresso Pereira, Rodrigo Lima Guerra Moraes, and Beatrice Padovani Ferreira. Behavioral and ontogenetic colour changes of a poorly known lutjanid. Marine Biology Research, 8(9):906–911, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.702914.

Jawad:2012:GWF

[476] Laith A. Jawad. Groupers of the world: a field and market guide. Marine Biology Research, 8(9):912–913, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.703781.

Anonymous:2012:C

[477] Anonymous. Corrigendum. Marine Biology Research, 8(9):914, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.703818.

Ikawa:2012:SSC

[478] Terumi Ikawa, Hidehiko Okabe, and Lanna Cheng. Skaters of the seas — comparative ecology of nearshore and pelagic Halobates species (Hemiptera: Gerridae), with special reference to Japanese species. Marine Biology Research, 8(10):915–936, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.705848.

Calcagno:2012:EPD

[479] Javier Angel Calcagno, Jessica Natalia Curelovich, Vanessa Maribel Fernandez, Sven Thatje, and Gustavo Alejandro Lovrich. Effects of physical disturbance on a sub-Antarctic middle intertidal bivalve assemblage. Marine Biology Research, 8(10):937–953, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.702911.

Cinar:2012:STD

[480] Melih Ertan Çinar, Tuncer Katagan, Bilal Öztürk, Ertan Dagli, Sermin Açik, Banu Bitlis, Kerem Bakir, and Alper Dogan. Spatio-temporal distributions of zoobenthos in Mersin Bay (Levantine Sea, eastern Mediterranean) and the importance of alien species in benthic communities. Marine Biology Research, 8(10):954–968, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.706305.

Keppel:2012:NGR

[481] Erica Keppel, Marco Sigovini, and Davide Tagliapietra. A new geographical record of Polycera hedgpethi er. Marcus, 1964 (Nudibranchia: Polyceridae) and evidence of its established presence in the Mediterranean Sea, with a review of its geographical distribution. Marine Biology Research, 8(10):969–981, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.706306.

Sperone:2012:SBB

[482] Emilio Sperone, Primo Micarelli, Sara Andreotti, Pietro Brandmayr, Ilaria Bernabò, Elvira Brunelli, and Sandro Tripepi. Surface behaviour of bait-attracted white sharks at Dyer Island (South Africa). Marine Biology Research, 8(10):982–991, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.708043.

Kumar:2012:TGS

[483] Girish Kumar, Swaraj Priyaranjan Kunal, Maria Rosalia Menezes, and Ram Murti Meena. Three genetic stocks of frigate tuna Auxis thazard thazard (Lacepede, 1800) along the Indian coast revealed from sequence analyses of mitochondrial DNA D-loop region. Marine Biology Research, 8(10):992–1002, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.702913.

Santic:2012:FHS

[484] Mate Santić, Biljana Rada, and Armin Pallaoro. Feeding habits of small-spotted catshark (Scyliorhinus canicula Linnaeus, 1758) from the eastern central Adriatic Sea. Marine Biology Research, 8(10): 1003–1011, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.702912.

Matic-Skoko:2012:AGF

[485] Sanja Matić-Skoko, Josipa Ferri, Pero Tutman, Daria Skaramuca, Domagoj Dikić, Duje Lisicić, Zdenko Franić, and Bosko Skaramuca. The age, growth and feeding habits of the European conger eel, Conger conger (L.) in the Adriatic Sea. Marine Biology Research, 8(10): 1012–1018, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.706307.

Pilarczyk:2012:SSS

[486] Bogumiła Pilarczyk, Agnieszka Tomza-Marciniak, Renata Pilarczyk, Katarzyna Kavetska, Izabella Rzad, Diana Hendzel, and Andrzej Marciniak. Selenium status in sea ducks (Melanitta fusca, Melanitta nigra and Clangula hyemalis) wintering on the southern Baltic coast, Poland. Marine Biology Research, 8(10):1019–1025, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2012.706304.

Ma:2012:FAS

[487] Yu-Bin Ma, Zhi-Feng Zhang, Ming-Yu Shao, Kyoung-Ho Kang, Li-Tao Zhang, Xiao-Li Shi, and Ying-Ping Dong. Function of the anal sacs and mid-gut in mitochondrial sulphide metabolism in the echiuran worm Urechis unicinctus. Marine Biology Research, 8(10):1026–1031, 2012. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2012.707320.

Ghigliotti:2012:BAF

[488] Laura Ghigliotti, Jørgen Schou Christiansen, Svein-Erik Fevolden, and Eva Pisano. Biodiversity of Arctic fishes: first karyological information on Gaidropsarus argentatus (Reinhardt, 1837), a new piece to the puzzle. Marine Biology Research, 8(10):1032–1035, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2012.708419.

Dill:2012:ECS

[489] Lawrence M. Dill. Ecology and conservation of the sirenia: dugongs and manatees. Marine Biology Research, 8(10):1036–1037, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.727231.

Anonymous:2012:RA

[490] Anonymous. Reviewer acknowledgement 2011. Marine Biology Research, 8(10):1038, 2012. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.718788.

Buchholz:2013:LRIa

[491] Fred Buchholz, Adrian Linnane, and Gro van der Meeren. Lobster research integrated: From biology to management. Marine Biology Research, 9(1):1–2, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.727439.

Buchholz:2013:LRIb

[492] Fred Buchholz, Adrian Linnane, Gro van der Meeren, Tom Fenchel, and Franz Uiblein. Lobster research integrated: From biology to management. Marine Biology Research, 9(1):3–6, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.727438.

Philp:2013:MCD

[493] Heather Philp and Gudrun Marteinsdottír. Moult cycle of the decapod Nephrops norvegicus (Linnaeus, 1758): Description of stages and sequential changes in haemolymph protein and haemocyanin. Marine Biology Research, 9(1):7–18, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2012.727435.

Schmalenbach:2013:ETM

[494] Isabel Schmalenbach and Friedrich Buchholz. Effects of temperature on the moulting and locomotory activity of hatchery-reared juvenile lobsters (Homarus gammarus) at Helgoland (North Sea). Marine Biology Research, 9(1):19–26, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2012.727433.

Kunkel:2013:CAF

[495] Joseph G. Kunkel and Michael J. Jercinovic. Carbonate apatite formulation in cuticle structure adds resistance to microbial attack for American lobster. Marine Biology Research, 9(1):27–34, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.727427.

Gherardi:2013:LIS

[496] Francesca Gherardi, Giuliana Parisi, and Laura Aquiloni. Looking for ‘identity signatures’ in the American lobster (Homarus americanus): Interindividual variation in body colour and in facial and chelar morphology. Marine Biology Research, 9(1):35–41, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.727426.

Wahle:2013:GBA

[497] Richard A. Wahle, Charlene Bergeron, John Tremblay, Carl Wilson, Victoria Burdett-Coutts, Michel Comeau, Rémy Rochette, Peter Lawton, Robert Glenn, and Mark Gibson. The geography and bathymetry of American lobster benthic recruitment as measured by diver-based suction sampling and passive collectors. Marine Biology Research, 9 (1):42–58, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.727428.

Watson:2013:RBA

[498] Win Watson and Steven H. Jury. The relationship between American lobster catch, entry rate into traps and density. Marine Biology Research, 9(1):59–68, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.727430.

Robertson:2013:MCS

[499] Denice N. Robertson and Mark J. Butler Iv. Mate choice and sperm limitation in the spotted spiny lobster, Panulirus guttatus. Marine Biology Research, 9(1):69–76, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.727429.

Briones-Fourzan:2013:IRO

[500] Patricia Briones-Fourzán, Erika Magallón-Gayón, and Enrique Lozano-Álvarez. Increased reproductive opportunity: a potential benefit of seasonal aggregation for a little-gregarious and highly sedentary spiny lobster. Marine Biology Research, 9(1):77–87, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.727431.

Ley-Cooper:2013:EER

[501] Kim Ley-Cooper, Simon De Lestang, Bruce F. Phillips, and Enrique Lozano-Álvarez. Estimates of exploitation rates of the spiny lobster fishery for Panulirus argus from tagging within the Bahía Espíritu Santo ‘Sian Ka’an’ Biosphere Reserve, Mexican Caribbean. Marine Biology Research, 9(1):88–96, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2012.727434.

Linnane:2013:ICR

[502] Adrian Linnane, Richard McGarvey, Matthew Hoare, and Peter Hawthorne. The importance of conserving recruitment pulses in rock lobster (Jasus edwardsii) fisheries where puerulus settlement is low or highly sporadic. Marine Biology Research, 9(1):97–103, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.727432.

Montgomery:2013:RER

[503] Steven S. Montgomery and Geoffery W. Liggins. Recovery of the eastern rock lobster Sagmariasus verreauxi off New South Wales, Australia.

Marine Biology Research, 9(1):104–115, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.727436.

Huang:2013:DFL

[504] Hui Huang, Guowei Zhou, Jianhui Yang, Sheng Liu, Feng You, and Xinming Lei. Diversity of free-living and symbiotic Symbiodinium in the coral reefs of sanya, South China Sea. Marine Biology Research, 9 (2):117–128, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.708045.

Tommasi:2013:SDZ

[505] Désirée A. G. Tommasi, Richard D. Routledge, Brian P. V. Hunt, and Evgeny A. Pakhomov. The seasonal development of the zooplankton community in a British Columbia (Canada) fjord during two years with different spring bloom timing. Marine Biology Research, 9(2): 129–144, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.708044.

Canas:2013:STB

[506] Lucía Cañás, Paz Sampedro, A. Celso Fariña, and Jorge Landa. Spatial, temporal and bathymetric distribution patterns of the parasite Chondracanthus lophii of anglerfish, Lophius piscatorius, in the northeast Atlantic. Marine Biology Research, 9(2):145–156, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.731691.

Pecorino:2013:RDS

[507] Danilo Pecorino, Miles D. Lamare, and Mike F. Barker. Reproduction of the Diadematidae sea urchin Centrostephanus rodgersii in a recently colonized area of northern New Zealand. Marine Biology Research, 9 (2):157–168, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.708046.

Bariche:2013:RIB

[508] Michel Bariche, Antranik Kajajian, and Ernesto Azzurro. Reproduction of the invasive bluespotted cornetfish Fistularia commersonii (Teleostei, Fistulariidae) in the Mediterranean Sea. Marine Biology Research, 9 (2):169–180, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.707322.

Isomura:2013:RBG

[509] Naoko Isomura, Yuichiro Baba, Shunsuke Nagata, Masanori Nonaka, and Hiromi Hannah Yamamoto. The relationship between genetic similarity and reproductive success in the branching coral Acropora intermedia. Marine Biology Research, 9(2):181–188, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.707321.

Nunez:2013:ETI

[510] Jesús DarÍo Nuñez, Emiliano Hernán Ocampo, Nicolás Mariano Chiaradia, Enrique Morsan, and Maximiliano Cledón. The effect of temperature on the inhalant siphon regeneration of Amiantis purpurata (Lamarck, 1818) (Bivalvia; Veneridae). Marine Biology Research, 9 (2):189–197, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.707323.

Gracan:2013:MFR

[511] Romana Gracan, Bojan Lazar, Ivan Posavec, Gordana Gregorović, and Gordana Lacković. Maturation, fecundity and reproductive cycle of spiny dogfish, Squalus acanthias, in the Adriatic Sea. Marine Biology Research, 9(2):198–207, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.729956.

Avaca:2013:AGM

[512] María Soledad Avaca, Maite Narvarte, and Pablo Martín. Age, growth and mortality in Buccinanops globulosus (Gastropoda: Nassariidae) from Golfo Nuevo (Argentina). Marine Biology Research, 9 (2):208–219, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.708420.

Sui:2013:VME

[513] Juan Sui, Qinghua Liu, Zhizhong Xiao, Daoyuan Ma, Shihong Xu, Tao He, Yifan Liu, Yongshuang Xiao, and Jun Li. The viability, melanophore and embryo genesis of first- and second-generation hybrids between Paralichthys olivaceus and P. dentatus. Marine Biology Research, 9(2): 220–226, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.729843.

Sato:2013:BST

[514] Keiichi Sato. Biology of sharks and their relatives, second edition. Marine Biology Research, 9(2):227, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.745005.

Hasemann:2013:EDI

[515] Christiane Hasemann, Melanie Bergmann, Corinna Kanzog, Normen Lochthofen, Eberhard Sauter, Ingo Schewe, and Thomas Soltwedel. Effects of dropstone-induced habitat heterogeneity on Arctic deep-sea benthos with special reference to nematode communities. Marine Biology Research, 9(3):229–245, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.739694.

Mutshinda:2013:ECD

[516] Crispin M. Mutshinda, Luis Troccoli-Ghinaglia, Zoe V. Finkel, Frank E. Müller-Karger, and Andrew J. Irwin. Environmental control of the dominant phytoplankton in the Cariaco basin: a hierarchical Bayesian approach. Marine Biology Research, 9(3):246–260, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.731693.

Ros:2013:IFC

[517] Macarena Ros, José M. Guerra-García, Manuel González-Macías, Ángela Saavedra, and Carlos M. López-Fe. Influence of fouling communities on the establishment success of alien caprellids (Crustacea: Amphipoda) in Southern Spain. Marine Biology Research, 9(3):261–273, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.739695.

Popovic:2013:RCG

[518] Zvjezdana Popović, Ivona Mladineo, Daria Ezgeta-Balić, Zeljka Trumbić, Nedo Vrgoc, and Melita Peharda. Reproductive cycle and gonad development of Venus verrucos a L. (Bivalvia: Veneridae) in Kastela Bay, Adriatic Sea. Marine Biology Research, 9(3):274–284, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.731690.

Clavico:2013:ICS

[519] Etiene E. G. Clavico, Bernardo A. P. Da Gama, Angélica R. Soares, Keila Mara Cassiano, and Renato C. Pereira. Interaction of chemical and structural components providing defences to sea pansies Renilla reniformis and Renilla muelleri. Marine Biology Research, 9(3): 285–292, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.739693.

Janas:2013:TSR

[520] Urszula Janas, Małgorzata Piłka, and Dagmara Lipińska. Temperature and salinity requirements of Palaemon adspersus Rathke, 1837 and Palaemon elegans Rathke, 1837. Do they explain the occurrence and expansion of prawns in the Baltic Sea? Marine Biology Research, 9 (3):293–300, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.739699.

Santic:2013:FHB

[521] Mate Santić, Biljana Rada, and Armin Pallaoro. Feeding habits of brown ray (Raja miraletus Linnaeus, 1758) from the eastern central Adriatic Sea. Marine Biology Research, 9(3):301–308, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.739698.

Tikochinski:2013:GEU

[522] Yaron Tikochinski, Idan Shainin, Yossi Hyams, Uzi Motro, and Daniel Golani. Genetic evidence for an undescribed species previously considered as Sillago sihama from the northern Red Sea. Marine Biology Research, 9(3):309–315, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2012.731692.

Kaga:2013:FRR

[523] Tatsuya Kaga and Elaine Heemstra. First record of a rare sand whiting, Sillago caudicula (Perciformes: Sillaginidae), from Madagascar. Marine Biology Research, 9(3):316–320, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.742547.

Forslund:2013:RRI

[524] Helena Forslund and Lena Kautsky. Reproduction and reproductive isolation in Fucus radicans (Phaeophyceae). Marine Biology Research, 9 (3):321–326, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.731694.

Temperoni:2013:EFD

[525] Brenda Temperoni, María Delia Viñas, and Daniel Hernández. Enhancing fish diet analysis: equations to reconstruct Themisto gaudichaudii and Euphausia lucens length from partially digested remains. Marine Biology Research, 9(3):327–332, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.739696.

Pesic:2013:NMW

[526] Vladimir Pesić. A new marine water mite species (Acari, Hydrachnidia, Pontarachnidae) from a coastal lake in Southeast Madagascar. Marine Biology Research, 9(3):333–336, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.739697.

Quillfeldt:2013:ICV

[527] Petra Quillfeldt and Juan F. Masello. Impacts of climate variation and potential effects of climate change on South American seabirds — a review. Marine Biology Research, 9(4):337–357, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.756982.

Martinez-Garcia:2013:DRM

[528] Sandra Martínez-García, Emilio Fernández, Alejandra Calvo-Díaz, Pedro Cermeño, Emilio Marañón, Xose Anxelu G. Morán, and Eva Teira. Differential response of microbial plankton to nutrient inputs in oligotrophic versus mesotrophic waters of the North Atlantic. Marine Biology Research, 9(4):358–370, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.745002.

Vinas:2013:SSZ

[529] María Delia Viñas, Rubén Mario Negri, Georgina Daniela Cepeda, Daniel Hernández, Ricardo Silva, María Cristina Daponte, and Fabiana Lía Capitanio. Seasonal succession of zooplankton in coastal waters of the Argentine Sea (Southwest Atlantic Ocean): prevalence of classical or microbial food webs. Marine Biology Research, 9(4):371–382, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.745003.

Humborstad:2013:BAA

[530] Odd-Børre Humborstad and Anders Mangor-Jensen. Buoyancy adjustment after swimbladder puncture in cod Gadus morhua: An experimental study on the effect of rapid decompression in capture-based aquaculture. Marine Biology Research, 9(4):383–393, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.742546.

Reinaldo:2013:SPA

[531] Matías Ocampo Reinaldo, Raúl González, Gabriela Williams, Lorena Pía Storero, María Alejandra Romero, Maite Narvarte, and Domingo Antonio Gagliardini. Spatial patterns of the Argentine hake Merluccius hubbsi and oceanographic processes in a semi-enclosed Patagonian ecosystem. Marine Biology Research, 9(4):394–406, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.739700.

DeGuevara:2013:SRS

[532] Genoveva Cerdenares-Ladrón De Guevara, Enrique Morales-Bojórquez, Carmen Rodríguez-Jaramillo, Agustín Hernández-Herrera, and Andrés Abitia-Cárdenas. Seasonal reproduction of sailfish Istiophorus platypterus from the southeast Mexican Pacific. Marine Biology Research, 9(4): 407–420, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.742545.

Anker:2013:DND

[533] Arthur Anker and Marcos Tavares. Description of a new deep-water stenopodid shrimp of the genus Odontozona Holthuis, 1946 (Crustacea, Decapoda) from Brazil. Marine Biology Research, 9(4):421– 430, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.745004.

Hemmingsen:2013:LVO

[534] Willy Hemmingsen and Ken Mackenzie. Latitudinal variations in the occurrence of some cod parasites along the west coast of Norway. Marine Biology Research, 9(4):431–436, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.742548.

Fenchel:2013:BET

[535] Tom Fenchel. The biology and ecology of tintinnid ciliates. Models for marine plankton. Marine Biology Research, 9(4):437, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.760738.

Gouws:2013:CIP

[536] Gavin Gouws. Crustacean issues 19: Phylogeography and population genetics in crustacea. Marine Biology Research, 9(4):437–439, 2013. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2012.756983.

Gebruk:2013:BSPa

[537] Andrey V. Gebruk and Imants G. Priede. Benthos of the sub-polar front area on the Mid-Atlantic Ridge: Results of the ECOMAR project. Marine Biology Research, 9(5–6):441–442, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.750000.

Gebruk:2013:BSPb

[538] Andrey V. Gebruk, Imants G. Priede, Tom Fenchel, and Franz Uiblein. Benthos of the sub-polar front area on the Mid-Atlantic Ridge: Results of the ECOMAR project. Marine Biology Research, 9(5–6): 443–446, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.749999.

Gooday:2013:NPK

[539] Andrew J. Gooday, Nina Rothe, and Richard B. Pearce. New and poorly known benthic foraminifera (Protista, Rhizaria) inhabiting the shells of planktonic foraminifera on the bathyal Mid-Atlantic Ridge. Marine Biology Research, 9(5–6):447–461, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.750365.

Lopes:2013:NDG

[540] Daniela A. Lopes and Konstantin R. Tabachnick. New data on glass sponges (Porifera, Hexactinellida) of the northern Mid-Atlantic Ridge. Part 1. Farreidae. Marine Biology Research, 9(5–6):462–468, 2013. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2012.749995.

Tabachnick:2013:NDG

[541] Konstantin R. Tabachnick and Larisa L. Menshenina. New data on glass sponges (Porifera, Hexactinellida) of the northern Mid-Atlantic Ridge. Part 2. Aphrocallistidae, Euretidae, Euplectellidae and Rossellidae (with descriptions of two new species of Sympagella). Marine Biology Research, 9(5–6):469–487, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.749996.

Molodtsova:2013:DSM

[542] Tina N. Molodtsova. Deep-sea mushroom soft corals (Octocorallia: Alcyonacea: Alcyoniidae) of the Northern Mid-Atlantic Ridge. Marine Biology Research, 9(5–6):488–515, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2012.750427.

Kongsrud:2013:BPN

[543] Jon A. Kongsrud, Nataliya Budaeva, Ruth Barnich, Eivind Oug, and Torkild Bakken. Benthic polychaetes from the northern Mid-Atlantic Ridge between the Azores and the Reykjanes Ridge. Marine Biology Research, 9(5–6):516–546, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2012.749997.

Shields:2013:PPM

[544] Mark A. Shields, Adrian G. Glover, and Helena Wiklund. Polynoid polychaetes of the Mid-Atlantic Ridge and a new holothurian association. Marine Biology Research, 9(5–6):547–553, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2012.749992.

Horton:2013:HNC

[545] Tammy Horton and Michael Thurston. Hirondellea namarensis (Crustacea: Amphipoda: Lysianassoidea: Hirondelleidae), a new deep-water scavenger species from the Mid-Atlantic Ridge. Marine Biology Research, 9(5–6):554–562, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2012.749994.

Dilman:2013:AFN

[546] Anna B. Dilman. Asteroid fauna of the northern Mid-Atlantic Ridge: additional records. Marine Biology Research, 9(5–6):563–586, 2013. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2012.749993.

Rogacheva:2013:HCG

[547] Antonina Rogacheva, Andrey Gebruk, and Claudia H. S. Alt. Holothuroidea of the Charlie Gibbs Fracture Zone area, northern Mid-Atlantic Ridge. Marine Biology Research, 9(5–6):587–623, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.750428.

Priede:2013:EME

[548] Imants G. Priede, David S. M. Billett, Andrew S. Brierley, A. Rus Hoelzel, Mark Inall, and Peter I. Miller. The ECO-MAR (ecosystem of the Mid-Atlantic Ridge at the Sub-Polar Front and Charlie Gibbs Fracture Zone) project: description of the benthic sampling programme 2007–2010. Marine Biology Research, 9(5–6):624–628, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.749998.

Hoffle:2013:DVH

[549] Hannes Höffle, Richard D. M. Nash, Tone Falkenhaug, and Peter Munk. Differences in vertical and horizontal distribution of fish larvae and zooplankton, related to hydrography. Marine Biology Research, 9(7): 629–644, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.765576.

Urra:2013:SDM

[550] Javier Urra, Pablo Marina, Carmen Salas, Serge Gofas, and José L. Rueda. Seasonal dynamics of molluscan assemblages associated with littoral soft bottoms of the NW Alboran Sea (Western Mediterranean Sea). Marine Biology Research, 9(7):645–660, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.765573.

Neves:2013:FCA

[551] Leonardo Mitrano Neves, Tatiana Pires Teixeira, Taynara Pontes Franco, Hamilton Hissa Pereira, and Francisco Gerson Araújo. Fish composition and assemblage structure in the estuarine mixing zone of a tropical estuary: comparisons between the main channel and an adjacent lagoon. Marine Biology Research, 9(7):661–675, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.765575.

Martinez:2013:DTN

[552] Alejandro Martínez, Maikon Di Domenico, Katharina Jörger, Jon Norenburg, and Katrine Worsaae. Description of three new species of Protodrilus (Annelida, Protodrilidae) from Central America. Marine Biology Research, 9(7):676–691, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.765574.

Wei:2013:PGV

[553] Nu-Wei V. Wei, Charlotte Watson, and Karen S. Gibb. Phylogenetic and geographic variation of northern Australian sympatric lineages of Treptopale homalos and T. paromolos (Annelida: Phyllodocida: Chrysopetalidae) using mitochondrial and nuclear sequences. Marine Biology Research, 9(7):692–702, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2013.765578.

Negrin:2013:RSP

[554] Vanesa L. Negrin, Carla V. Spetter, Valeria A. Guinder, Gerardo M. E. Perillo, and Jorge E. Marcovecchio. The role of Sarcocornia perennis and tidal flooding on sediment biogeochemistry in a South American wetland. Marine Biology Research, 9(7):703–715, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.765579.

Halberg:2013:ETT

[555] Kenneth A. Halberg, Dennis K. Persson, Aslak Jørgensen, Reinhardt M. Kristensen, and Nadja Møbjerg. Ecology and thermal tolerance of the marine tardigrade Halobiotus crispae (Eutardigrada: Isohypsibiidae). Marine Biology Research, 9(7):716–724, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.765577.

Kim:2013:SSV

[556] Hyeong gi Kim, Hyun Soo Rho, and Chul-Woong Oh. Seasonal and spatial variations in nematode assemblages affected by thermal influence of nuclear power plant in Korea (East Sea, Pacific Ocean). Marine Biology Research, 9(8):725–738, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.766346.

Verdiell-Cubedo:2013:ANR

[557] David Verdiell-Cubedo, Francisco J. Oliva-Paterna, Ana Ruiz-Navarro, and Mar Torralva. Assessing the nursery role for marine fish species in a hypersaline coastal lagoon (Mar Menor, Mediterranean Sea). Marine Biology Research, 9(8):739–748, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.765580.

Al-Horani:2013:DAR

[558] Fuad A. Al-Horani and Maroof A. Khalaf. Developing artificial reefs for the mitigation of man-made coral reef damages in the Gulf of Aqaba, Red Sea: coral recruitment after 3.5 years of deployment. Marine Biology Research, 9(8):749–757, 2013. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.765582.

Sato:2013:AGS

[559] Keiichi Sato, Andrew L. Stewart, and Kazuhiro Nakaya. Apristurus garricki sp. nov., a new deep-water catshark from the northern New Zealand waters (Carcharhiniformes: Scyliorhinidae). Marine Biology Research, 9 (8):758–767, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.765586.

Carballo:2013:WDR

[560] José Luis Carballo, Jose María Aguilar-Camacho, Ingrid S. Knapp, and James J. Bell. Wide distributional range of marine sponges along the Pacific Ocean. Marine Biology Research, 9(8):768–775, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.765581.

Zelaya:2013:TNS

[561] Diego G. Zelaya and Cristián Ituarte. Two new species of Waldo Nicol, 1966 from sub-Antarctic waters (Bivalvia: Galeommatoidea). Marine Biology Research, 9(8):776–784, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.765584.

Barco:2013:TAD

[562] Andrea Barco, Julian Evans, Patrick J. Schembri, Marco Taviani, and Marco Oliverio. Testing the applicability of DNA barcoding for Mediterranean species of top-shells (Gastropoda, Trochidae, Gibbula s.l.). Marine Biology Research, 9(8):785–793, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.765585.

Hibino:2013:MCC

[563] Yuya Hibino, Peter Todd, Carey D. Ashworth, Masami Obuchi, and James Davis Reimer. Monitoring colony colour and zooxanthellae (Symbiodinium spp.) condition in the reef zoanthid Palythoa tuberculosa in Okinawa, Japan. Marine Biology Research, 9(8):794–801, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.766344.

Czudaj:2013:DRB

[564] Stephanie Czudaj, João Pereira, Ana Moreno, Ulrich Saint-Paul, and Rui Rosa. Distribution and reproductive biology of the lentil bobtail squid, Rondeletiola minor (Cephalopoda: Sepiolidae) from the Portuguese Atlantic Coast. Marine Biology Research, 9(8):802–808, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.765583.

Dong:2013:DEA

[565] Yuan Dong, Gui-Peng Yang, and Kam W. Tang. Dietary effects on abundance and carbon utilization ability of DMSP-consuming bacteria associated with the copepod Acartia tonsa Dana. Marine Biology Research, 9(8):809–814, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.765587.

Whitfield:2013:EE

[566] Alan K. Whitfield. Estuarine ecology. Marine Biology Research, 9 (8):815–816, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.789913.

Haug:2013:CEBa

[567] Tore Haug, Andrey Dolgov, Ingolf Røttingen, Knut Sunnanå, and Oleg Titov. Climate effects on the Barents Sea marine living resources. Marine Biology Research, 9(9):817–818, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.775461.

Haug:2013:CEBb

[568] Tore Haug, Andrey Dolgov, Ingolf Røttingen, Knut Sunnanå, Oleg Titov, Tom Fenchel, and Franz Uiblein. Climate effects on the Barents Sea marine living resources. Marine Biology Research, 9(9):819–821, 2013. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.775460.

Yakushev:2013:SCC

[569] Evgeniy V. Yakushev and Kai Sørensen. On seasonal changes of the carbonate system in the Barents Sea: observations and modeling. Marine Biology Research, 9(9):822–830, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.775454.

Soreide:2013:SPB

[570] Janne E. Søreide, Michael L. Carroll, Haakon Hop, William G. Ambrose, Jr., Else N. Hegseth, and Stig Falk-Petersen. Sympagic-pelagic-benthic coupling in Arctic and Atlantic waters around Svalbard revealed by stable isotopic and fatty acid tracers. Marine Biology Research, 9 (9):831–850, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.775457.

Orlova:2013:SMP

[571] Emma L. Orlova, Andrey V. Dolgov, Paul E. Renaud, Vladimir D. Boitsov, Irina P. Prokopchuk, and Mira V. Zashihina. Structure of the macroplankton-pelagic fish-cod trophic complex in a warmer Barents Sea. Marine Biology Research, 9(9):851–866, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.775453.

Ingvaldsen:2013:RSD

[572] Randi B. Ingvaldsen and Harald Gjøsæter. Responses in spatial distribution of Barents Sea capelin to changes in stock size, ocean temperature and ice cover. Marine Biology Research, 9(9):867–877, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.775450.

Hop:2013:PCB

[573] Haakon Hop and Harald Gjøsæter. Polar cod (Boreogadus saida) and capelin (Mallotus villosus) as key species in marine food webs of the Arctic and the Barents Sea. Marine Biology Research, 9(9): 878–894, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.775458.

Bogstad:2013:CRB

[574] Bjarte Bogstad, Gjert Endre Dingsør, Randi B. Ingvaldsen, and Harald Gjøsæter. Changes in the relationship between sea temperature and recruitment of cod, haddock and herring in the Barents Sea. Marine Biology Research, 9(9):895–907, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.775451.

Johansen:2013:SVG

[575] Geir Odd Johansen, Edda Johannesen, Kathrine Michalsen, Asgeir Aglen, and Åge Fotland. Seasonal variation in geographic distribution of North East Arctic (NEA) cod — survey coverage in a warmer Barents Sea. Marine Biology Research, 9(9):908–919, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.775456.

Howell:2013:UUP

[576] Daniel Howell, Anatoly A. Filin, Bjarte Bogstad, and Jan Erik Stiansen. Unquantifiable uncertainty in projecting stock response to climate change: Example from North East Arctic cod. Marine Biology Research, 9(9):920–931, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.775452.

Michalsen:2013:MLR

[577] Kathrine Michalsen, Padmini Dalpadado, Elena Eriksen, Harald Gjøsæter, Randi B. Ingvaldsen, Edda Johannesen, Lis Lindal Jørgensen, Tor Knutsen, Dmitry Prozorkevich, and Mette Skern-Mauritzen. Marine living resources of the Barents Sea — ecosystem understanding and monitoring in a climate change perspective. Marine Biology Research, 9(9):932–947, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.775459.

Vestheim:2013:VDF

[578] Hege Vestheim, Sandra Brucet, and Stein Kaartvedt. Vertical distribution, feeding and vulnerability to tactile predation in Metridia longa (Copepoda, Calanoida). Marine Biology Research, 9(10):949– 957, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.793806.

Polo-Silva:2013:TSD

[579] Carlos Polo-Silva, Seth D. Newsome, Felipe Galván-Magaña, Marcela Grijalba-Bendeck, and Adolfo Sanjuan-Muñoz. Trophic shift in the diet of the pelagic thresher shark based on stomach contents and stable isotope analyses. Marine Biology Research, 9(10):958–971, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.793802.

VeronesiFukuda:2013:FOP

[580] Marcelo Veronesi Fukuda and João Miguel De Matos Nogueira. First occurrence of Perkinsyllis San Martín, López and Aguado, 2009 (Polychaeta: Syllidae) in Brazil, with descriptions of two new species. Marine Biology Research, 9(10):972–989, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.793807.

Neres:2013:NRS

[581] PatrÍcia Fernandes Neres, Verônica G. Fonsêca-Genevois, and André Morgado Esteves. New records and species of Enchelidiidae (Nematoda): Symplocostoma brasiliensis sp. nov. and a first description of the female of Calyptronema pigmentatum Gerlach, 1957. Marine Biology Research, 9(10):990–1004, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.793803.

Riisgaard:2013:GMM

[582] Hans Ulrik Riisgård, Daniel Pleissner, Kim Lundgreen, and Poul S. Larsen. Growth of mussels Mytilus edulis at algal (Rhodomonas salina) concentrations below and above saturation levels for reduced filtration rate. Marine Biology Research, 9(10):1005–1017, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2012.742549.

Berasategui:2013:SOH

[583] Anabela A. Berasategui, M. Sofia Dutto, Javier Chazarreta, and Monica S. Hoffmeyer. Seasonal occurrence and hatching success of benthic eggs of calanoid copepods in the bahía blanca estuary, Argentina. Marine Biology Research, 9(10):1018–1028, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.793801.

Gebruk:2013:PDS

[584] Andrey V. Gebruk, Antonina V. Rogacheva, David L. Pawson, Jean-François Hamel, Kevin G. Macisaac, and Annie Mercier. Penilidia desbarresi sp. nov. (Echinodermata: Holothuroidea: Elasipodida) from the upper slope of Newfoundland and re-description of P. ludwigi (von Marenzeller, 1893). Marine Biology Research, 9(10):1029– 1036, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.793810.

White:2013:BRA

[585] William T. White. Book review anthiines. Marine Biology Research, 9 (10):1037–1038, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.797589.

Anonymous:2013:RA

[586] Anonymous. Reviewer acknowledgement 2012. Marine Biology Research, 9(10):1039–1040, 2013. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.810763.

Uiblein:2014:NRR

[587] Franz Uiblein. New records and range extension reports towards elucidating marine ecosystems. Marine Biology Research, 10(1):1–2, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.840732.

Danise:2014:MSW

[588] Silvia Danise, Stefano Dominici, Adrian G. Glover, and Thomas G. Dahlgren. Molluscs from a shallow-water whale-fall and their affinities with adjacent benthic communities on the Swedish west coast. Marine Biology Research, 10(1):3–16, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.793811.

Sanchez:2014:PAS

[589] Nuria Sánchez, María Herranz, Jesús Benito, and Fernando Pardos. Pycnophyes almansae sp. nov. and Pycnophyes lageria sp. nov., two new homalorhagid kinorhynchs (Kinorhyncha, Homalorhagida) from the Iberian Peninsula, with special focus on introvert features. Marine Biology Research, 10(1):17–36, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.793804.

Dellatorre:2014:SAV

[590] Fernando Gaspar Dellatorre, Gustavo Lovrich, Laura Rojas, Viviana Milano, Nerina Delourdes Figueroa, and Pedro José Barón. Seasonal abundance and vertical distribution of crab larvae from northern Patagonia (Argentina): implications for their transport in a biogeographic boundary region. Marine Biology Research, 10(1):37–50, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.793805.

Tam:2014:DCM

[591] Jamie C. Tam and Ricardo A. Scrosati. Distribution of cryptic mussel species (Mytilus edulis and M. trossulus) along wave exposure gradients on northwest Atlantic rocky shores. Marine Biology Research, 10(1):51–60, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.793809.

Calderon-Aguilera:2014:RCG

[592] Luis E. Calderon-Aguilera, Eugenio A. Aragón-Noriega, Enrique Morales-Bojórquez, Edgar Alcántara-Razo, and Jorge Chávez-Villalba. Reproductive cycle of the geoduck clam Panopea generosa at its southernmost distribution limit. Marine Biology Research, 10(1): 61–72, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.793808.

Culverhouse:2014:EAC

[593] Philip F. Culverhouse, Norman Macleod, Robert Williams, Mark C. Benfield, Rubens M. Lopes, and Marc Picheral. An empirical assessment of the consistency of taxonomic identifications. Marine Biology Research, 10(1):73–84, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.810762.

Damasceno-Oliveira:2014:PFS

[594] Alfredo Damasceno-Oliveira, Berta Levavi-Sivan, Joseph Aizen, José Gonçalves, Begoña Fernández-Durán, and João Coimbra. Pituitary follicle-stimulating hormone (FSH) and luteinizing hormone (LH) levels in maturing female flounder Platichthys flesus under hydrostatic pressure simulating vertical migrations. Marine Biology Research, 10 (1):85–92, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.793813.

Villier:2014:SBE

[595] Loïc Villier. Starfish — biology and ecology of the Asteroidea. Marine Biology Research, 10(1):93–94, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.820323.

Bugoni:2014:BST

[596] Leandro Bugoni. The biology of sea turtles, volume III. Marine Biology Research, 10(1):94–95, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.826369.

Currie:2014:MBT

[597] David R. Currie and Shirley J. Sorokin. Megabenthic biodiversity in two contrasting submarine canyons on Australia’s southern continental margin. Marine Biology Research, 10(2):97–110, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2013.797586.

Fortes:2014:PFC

[598] Wagner L. S. Fortes, Pedro H. Almeida-Silva, Luana Prestrelo, and Cassiano Monteiro-Neto. Patterns of fish and crustacean community structure in a coastal lagoon system, Rio de Janeiro, Brazil. Marine Biology Research, 10(2):111–122, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.797645.

Silva:2014:STD

[599] Mónica A. Silva, Rui Prieto, Irma Cascão, Maria Inês Seabra, Miguel Machete, Mark F. Baumgartner, and Ricardo S. Santos. Spatial and temporal distribution of cetaceans in the mid-Atlantic waters around the Azores. Marine Biology Research, 10(2):123–137, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.793814.

Vikingsson:2014:RCD

[600] Gísli A. Víkingsson, Bjarki Tór Elvarsson, Droplaug Ólafsdóttir, Jóhann Sigurjónsson, Valerie Chosson, and Anton Galan. Recent changes in the diet composition of common minke whales (Balaenoptera acutorostrata) in Icelandic waters. A consequence of climate change? Marine Biology Research, 10(2):138–152, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.793812.

Korsun:2014:IFF

[601] Sergei Korsun, Morten Hald, Elena Golikova, Anna Yudina, Ivan Kuznetsov, Dmitry Mikhailov, and Olga Knyazeva. Intertidal foraminiferal fauna and the distribution of Elphidiidae at Chupa Inlet, western White Sea. Marine Biology Research, 10(2):153–166, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.814786.

Esposito:2014:DPS

[602] Valentina Esposito, Franco Andaloro, Daniela Bianca, Antonino Natalotto, Teresa Romeo, Gianfranco Scotti, and Luca Castriota. Diet and prey selectivity of the red mullet, Mullus barbatus (Pisces: Mullidae), from the southern Tyrrhenian Sea: the role of the surf zone as a feeding ground. Marine Biology Research, 10(2):167–178, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.797585.

Curin:2014:IDE

[603] Mira Ćurin, Melita Peharda, Barbara Calcinai, and Stjepko Golubić. Incidence of damaging endolith infestation of the edible mytilid bivalve Modiolus barbatus. Marine Biology Research, 10(2):179–189, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.814793.

Jonsdottir:2014:SDN

[604] Ingibjörg G. Jónsdóttir. Spatial distribution of northern ambereye (Hymenodora glacialis) around Iceland. Marine Biology Research, 10(2): 190–196, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.797588.

Skaala:2014:IBSa

[605] Øystein Skaala, Kjersti Sjøtun, Einar Dahl, Vivian Husa, and Arne Bjørge. Interactions between salmon farming and the ecosystem: Lessons from the Hardangerfjord, Western Norway. Marine Biology Research, 10 (3):197–198, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.840731.

Skaala:2014:IBSb

[606] Øystein Skaala, Kjersti Sjøtun, Einar Dahl, Vivian Husa, Arne Bjørge, and Franz Uiblein. Interactions between salmon farming and the ecosystem: Lessons from the Hardangerfjord, western Norway. Marine Biology Research, 10(3):199–202, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.840730.

Falkenhaug:2014:DCF

[607] Tone Falkenhaug and Padmini Dalpadado. Diet composition and food selectivity of sprat (Sprattus sprattus) in Hardangerfjord, Norway. Marine Biology Research, 10(3):203–215, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.810752.

Asplin:2014:DSL

[608] Lars Asplin, Ingrid A. Johnsen, Anne D. Sandvik, Jon Albretsen, Vibeke Sundfjord, Jan Aure, and Karin K. Boxaspen. Dispersion of salmon lice in the Hardangerfjord. Marine Biology Research, 10(3):216–225, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.810755.

Husa:2014:HCM

[609] Vivian Husa, Henning Steen, and Kjersti Sjøtun. Historical changes in macroalgal communities in Hardangerfjord (Norway). Marine Biology Research, 10(3):226–240, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.810751.

Husa:2014:RIF

[610] Vivian Husa, Tina Kutti, Arne Ervik, Kjersti Sjøtun, Pia Kupka Hansen, and Jan Aure. Regional impact from fin-fish farming in an intensive production area (Hardangerfjord, Norway). Marine Biology Research, 10 (3):241–252, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.810754.

Buhl-Mortensen:2014:DVD

[611] Pål Buhl-Mortensen and Lene Buhl-Mortensen. Diverse and vulnerable deep-water biotopes in the Hardangerfjord. Marine Biology Research, 10(3):253–267, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.810759.

Vollset:2014:CRL

[612] Knut W. Vollset, Helge Skoglund, Bjørn T. Barlaup, Ulrich Pulg, Sven-Erik Gabrielsen, Tore Wiers, Bjørnar Skår, and Gunnar B. Lehmann. Can the river location within a fjord explain the density of Atlantic salmon and sea trout? Marine Biology Research, 10(3): 268–278, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.810761.

Skaala:2014:ESL

[613] Øystein Skaala, Steinar Kålås, and Reidar Borgstrøm. Evidence of salmon lice-induced mortality of anadromous brown trout (Salmo trutta) in the Hardangerfjord, Norway. Marine Biology Research, 10 (3):279–288, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.810756.

Skiftesvik:2014:WLC

[614] Anne Berit Skiftesvik, Geir Blom, Ann-Lisbeth Agnalt, Caroline M. F. Durif, Howard I. Browman, Reidun M. Bjelland, Lisbeth S. Harkestad, Eva Farestveit, Ole Ingar Paulsen, Merete Fauske, Trond Havelin, Knut Johnsen, and Stein Mortensen. Wrasse (Labridae) as cleaner fish in salmonid aquaculture — the Hardangerfjord as a case study. Marine Biology Research, 10(3):289–300, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.810760.

Fjortoft:2014:DCG

[615] Helene Børretzen Fjørtoft, Reidar Borgstrøm, and Øystein Skaala. Differential changes in growth patterns of anadromous brown trout and Atlantic salmon from the river etneelva over a 25-year period. Marine Biology Research, 10(3):301–307, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.810753.

Skaala:2014:CPA

[616] Øystein Skaala, Geir Helge Johnsen, Håvard Lo, Reidar Borgstrøm, Vidar Wennevik, Michael Møller Hansen, Joseph E. Merz, Kevin A. Glover, and Bjørn T. Barlaup. A conservation plan for Atlantic salmon (Salmo salar) and anadromous brown trout (Salmo trutta) in a region with intensive industrial use of aquatic habitats, the Hardangerfjord, western Norway. Marine Biology Research, 10(3):308–322, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.810758.

Skaala:2014:MDU

[617] Øystein Skaala, Kevin A. Glover, Bjørn T. Barlaup, and Reidar Borgstrøm. Microsatellite DNA used for parentage identification of partly digested Atlantic salmon (Salmo salar) juveniles through nondestructive diet sampling in salmonids. Marine Biology Research, 10 (3):323–328, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.810757.

Lopes:2014:CSD

[618] Daniela A. Lopes and Eduardo Hajdu. Carnivorous sponges from deepsea coral mounds in the Campos Basin (SW Atlantic), with the description of six new species (Cladorhizidae, Poecilosclerida, Demospongiae). Marine Biology Research, 10(4):329–356, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.797587.

Sanchez-Tocino:2014:OCS

[619] Luis Sánchez-Tocino, J. Manuel Tierno de Figueroa, and Juan Lucas Cervera. Ontogenetic changes in the spicule formation and their possible role in chromodorid opisthobranchs (Mollusca, Chromodorididae). Marine Biology Research, 10(4):357–373, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.815782.

Xu:2014:INE

[620] Zhi-Huan Xu, Zhen-Ren Guo, Xu Xu, Dao-Jian Huang, Xiao-Xin Sun, Li-Ming Jiang, and Ji-Fang Yang. The impact of nutrient enrichment on the phytoplankton and bacterioplankton community during a mesocosm experiment in Nan’ao of Daya Bay. Marine Biology Research, 10 (4):374–382, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.814792.

Hipfner:2014:SVF

[621] J. Mark Hipfner, Laura Mcfarlane-Tranquilla, Brianne Addison, and Keith A. Hobson. Seasonal variation in the foraging ecology of a zooplanktivorous seabird assessed with stable isotope analysis. Marine Biology Research, 10(4):383–390, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.814791.

Naruse:2014:DNS

[622] Tohru Naruse and Jun Hashimoto. Description of a new species of the genus Trichopeltarion A. Milne-Edwards, 1880 (Decapoda: Brachyura: Trichopeltariidae) from western Pacific and southeast Asian waters. Marine Biology Research, 10(4):391–399, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.814789.

Song:2014:SCB

[623] Kyung-Jun Song and Chang Ik Zhang. Stomach contents of bycaught minke whales (Balaenoptera acutorostrata) in Korean waters. Marine Biology Research, 10(4):400–406, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.814787.

Shabtay:2014:PDG

[624] Ateret Shabtay, Yaron Tikochinski, Yehuda Benayahu, and Gil Rilov. Preliminary data on the genetic structure of a highly successful invading population of oyster suggesting its establishment dynamics in the Levant. Marine Biology Research, 10(4):407–415, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.814790. See comments [694].

Kadri:2014:AGS

[625] Hasna Kadri, Sondes Marouani, Bechir Saïdi, Mohamed Nejmeddine Bradai, Abderrahmen Bouaïn, and Eric Morize. Age, growth, sexual maturity and reproduction of the thornback ray, Raja clavata (L.), of the Gulf of Gabès (south-central Mediterranean Sea). Marine Biology Research, 10(4):416–425, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.797584.

Kadri:2014:FHB

[626] Hasna Kadri, Sondes Marouani, Mohamed Nejmeddine Bradai, and Abderrahmen Bouaïn. Food habits of the brown ray Raja miraletus (Chondrichthyes: Rajidae) from the Gulf of Gabès (Tunisia). Marine Biology Research, 10(4):426–434, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.797583.

Anonymous:2014:CA

[627] Anonymous. Conference announcement. Marine Biology Research, 10 (4):435, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.865305.

Vogedes:2014:VCS

[628] Daniel Vogedes, Ketil Eiane, Anna S. Båtnes, and Jørgen Berge. Variability in Calanus spp. abundance on fine- to mesoscales in an Arctic fjord: implications for little auk feeding. Marine Biology Research, 10 (5):437–448, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.815781.

Romatzki:2014:IEF

[629] Sascha Bernd Carsten Romatzki. Influence of electrical fields on the performance of Acropora coral transplants on two different designs of structures. Marine Biology Research, 10(5):449–459, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2013.814794.

Romano:2014:GOR

[630] Nicholas Romano, Xugan Wu, Chaoshu Zeng, Jerome Genodepa, and Jennifer Elliman. Growth, osmoregulatory responses and changes to the lipid and fatty acid composition of organs from the mud crab, Scylla serrata, over a broad salinity range. Marine Biology Research, 10(5): 460–471, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.819981.

Kim:2014:ERG

[631] Na Na Kim, Hyun Suk Shin, Jehee Lee, and Cheol Young Choi. Effects of recombinant growth hormone on growth factor and immune component levels in the cinnamon clownfish, Amphiprion melanopus. Marine Biology Research, 10(5):472–481, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.819982.

Tiedemann:2014:SAL

[632] Maik Tiedemann, Matthias Kloppmann, Jens Ulleweit, Joachim Paul Gröger, and Wilhelm Hagen. A spatial analysis of larval fish assemblages in the Celtic Sea off Great Britain (47^{o} to 51^{o}N): implications of bathymetry and ocean warming. Marine Biology Research, 10(5): 482–493, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.819983.

Presta:2014:FBP

[633] MarÍa Laura Presta, Florencia Cremonte, and Cristián Ituarte. The fit between parasites and intermediate host population dynamics: larval digeneans affecting the bivalve Neolepton cobbi (Galeommatoidea) from Patagonia. Marine Biology Research, 10(5):494–503, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2013.831175.

Zenteno:2014:LGC

[634] José I. Zenteno, Claudia A. Bustos, and Mauricio F. Landaeta. Larval growth, condition and fluctuating asymmetry in the otoliths of a mesopelagic fish in an area influenced by a large Patagonian glacier. Marine Biology Research, 10(5):504–514, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.831176.

Allen:2014:OSI

[635] Richard M. Allen. Oviposition site influences dispersal potential in a marine bubble snail. Marine Biology Research, 10(5):515–522, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.831177.

Vinagre:2014:AOJ

[636] Catarina Vinagre, Anabela Maia, Rachid Amara, and Henrique N. Cabral. Anomalous otoliths in juveniles of common sole, Solea solea, and Senegal sole, Solea senegalensis. Marine Biology Research, 10(5): 523–529, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.831178.

Atherton:2014:UPS

[637] Sarah Atherton. Urodasys poculostylis sp. nov., a new stylet-bearing gastrotrich (Macrodasyida) from Capron Shoal, Florida. Marine Biology Research, 10(5):530–536, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.831179.

Tarling:2014:SPV

[638] Geraint A. Tarling, Finlo R. Cottier, and Inigo Everson. Spatial patterns in the vertical structure of euphausiids in gullmarsfjord, Sweden: Identifying influences on bilayer formation and distribution. Marine Biology Research, 10(6):537–553, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.831180.

Paresque:2014:GHL

[639] Karla Paresque and João Miguel de Matos Nogueira. The genus Haplosyllis Langerhans, 1879 (Polychaeta: Syllidae) from northeastern Brazil, with descriptions of two new species. Marine Biology Research, 10 (6):554–576, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.841941.

Lutz-Collins:2014:ASR

[640] Vanessa Lutz-Collins and Pedro A. Quijón. Animal-sediment relationships in an Atlantic Canada marine protected area: Richness, composition and abundance in relation to sediment food indicators. Marine Biology Research, 10(6):577–588, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.833337.

deLima:2014:DPS

[641] Paloma Aparecida de Lima, Vivian Fransozo, Luciana Segura de Andrade, Ariádine Cristine Almeida, Michele Furlan, and Adilson Fransozo. Distribution and population structure of the flecked box crab Hepatus pudibundus (Decapoda, Brachyura) in the western South Atlantic. Marine Biology Research, 10(6):589–600, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.841940.

Huang:2014:BCC

[642] Hui Huang, Yi Yang, Xiubao Li, Jianhui Yang, Jiansheng Lian, Xinming Lei, Daoru Wang, and Jing Zhang. Benthic community changes following the 2010 Hainan flood: Implications for reef resilience. Marine Biology Research, 10(6):601–611, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.841942.

Rogers:2014:CHR

[643] Ricardo Rogers, Gabriel de Oliveira Correal, Thiago Cunha de Oliveira, Leonardo Lara de Carvalho, Patrícia Mazurek, Juliana Eymara Fernandes Barbosa, Luciana Chequer, Thaisa Francielle Souza Domingos, Kelly de Andrade Jandre, Luciana Sanches Dourado Leão, Laura de Andrade Moura, Gisele Exel Occhioni, Viviane Martins de Oliveira, Elisabetta Santos Silva, Alan Motta Cardoso, Ana de Castro e Costa, and Carlos Eduardo Leite Ferreira. Coral health rapid assessment in marginal reef sites. Marine Biology Research, 10(6):612–624, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2013.841944.

Ezgeta-Balic:2014:SDS

[644] Daria Ezgeta-Balić, Sonja Lojen, Tadej Dolenec, Petra Zvab Rozic, Matej Dolenec, Mirjana Najdek, and Melita Peharda. Seasonal differences of stable isotope composition and lipid content in four bivalve species from the Adriatic Sea. Marine Biology Research, 10(6):625–634, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.833338.

Yamaguchi:2014:SSS

[645] Sachi Yamaguchi, Sachi Yoshida, Atsushi Kaneko, Kota Sawada, Keiko Yasuda, and Yoichi Yusa. Sexual system of a symbiotic pedunculate barnacle Poecilasma kaempferi (Cirripedia: Thoracica). Marine Biology Research, 10(6):635–640, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.841943.

Yu:2014:BSV

[646] Ying Yu, Wuchang Zhang, Cuixia Zhang, Feng Zhou, Nan Zhao, and Tian Xiao. Basin-scale variation in planktonic ciliate distribution: a detailed temporal and spatial study of the Yellow Sea. Marine Biology Research, 10(7):641–654, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.852683.

Uiblein:2014:NGS

[647] Franz Uiblein and Gavin Gouws. A new goatfish species of the genus Upeneus (Mullidae) based on molecular and morphological screening and subsequent taxonomic analysis. Marine Biology Research, 10(7): 655–681, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.850515.

Azzarello:2014:LHC

[648] Jacob J. Azzarello, Dan A. Smale, Timothy J. Langlois, and Eckart Håkansson. Linking habitat characteristics to abundance patterns of canopy-forming macroalgae and sea urchins in southwest Australia. Marine Biology Research, 10(7):682–693, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.841945.

Biju:2014:DAM

[649] Abraham Biju, Loganathan Jagadeesan, and Saramma U. Panampunnayil. Distribution and abundance of mysid shrimps (Crustacea: Mysidacea) in the Northern Indian Ocean. Marine Biology Research, 10 (7):694–711, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.841946.

Costa:2014:DOS

[650] Mércia B. Costa, Gabriela C. Zamprogno, Felipe C. Pedruzzi, Letícia de Morais, Mônica M. P. Tognella, Ana Flavia L. Godoi, Dayana M. Santos, Mary Rosa R. de Marchi, and Marcos A. Fernandez. Differential organotin sensitivity in two Leucozonia species from a ship traffic area in southeastern Brazil. Marine Biology Research, 10(7):712–724, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.850512.

Abdullah:2014:ENK

[651] Mohamed I. Abdullah and Stein Fredriksen. The exudation of nitrate by the kelp Laminaria hyperborea, an observation during in situ incubation experiments. Marine Biology Research, 10(7):725–730, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.850514.

Wang:2014:MCM

[652] Yanhong Wang, Zhe Zhao, Luping Zhang, Chaoqu Hu, Chunhua Ren, and Lihong Yuan. Molecular characterization of metallothionein from white shrimp, Litopenaeus vannamei and its expression response to salinity stress. Marine Biology Research, 10(7):731–737, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2013.850513.

Kunal:2014:GHL

[653] Swaraj Priyaranjan Kunal, Girish Kumar, Maria Rosalia Menezes, and Ram Murti Meena. Genetic homogeneity in longtail tuna Thunnus tonggol (Bleeker, 1851) from the northwest coast of India inferred from direct sequencing analysis of the mitochondrial DNA D-loop region. Marine Biology Research, 10(7):738–743, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.852682.

Nielsen:2014:BRH

[654] Claus Nielsen. Book review: The Helgoland Manual of Animal Development. Marine Biology Research, 10(7):744, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.879989.

Zhang:2014:ELI

[655] Pei-Dong Zhang, Ling-Yu Zhang, Shu-Na Niu, Yan Sun, and Lu Tian. Effect of light intensity on survival, growth, and photosynthetic pigment of young seedlings of eelgrass Zostera marina Linnaeus, 1753 (Alismatales:

Zosteraceae). Marine Biology Research, 10(8):745–754, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.854917.

Anker:2014:NDS

[656] Arthur Anker. Naushonia draconis sp. nov., a heavily armoured mud shrimp from the Caribbean Sea, and taxonomic status of Espeleonaushonia Juarrero and Martínez-Iglesias, 1997 (Decapoda: Gebiidea: Laomediidae). Marine Biology Research, 10(8):755–770, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.852684.

Wolf:2014:EPR

[657] Christian Wolf, Estelle Silvia Kilias, and Katja Metfies. Evaluating the potential of 18S rDNA clone libraries to complement pyrosequencing data of marine protists with near full-length sequence information. Marine Biology Research, 10(8):771–780, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.852685.

Ying:2014:SCF

[658] Xue-Ping Ying, Hans-Uwe Dahms, Xiao-Ying Ni, Cong-Cong Hou, Yu-Jie Zhang, and Wan-Xi Yang. Seasonal changes of the fatty acid composition in the hepatopancreas and vitelline gland of the gastropod Onchidium struma. Marine Biology Research, 10(8):781–790, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.853125.

Riascos:2014:DCL

[659] Jose M. Riascos, Viviana Villegas, and Aldo S. Pacheco. Diet composition of the large scyphozoan jellyfish Chrysaora plocamia in a highly productive upwelling centre off northern Chile. Marine Biology Research, 10(8):791–798, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.863353.

Jin:2014:HPL

[660] Cuili Jin, Jingjing Qiu, Siyu Yu, Li Miao, and Xiaojian Zhou. Histamine promotes the larval metamorphic competence of barnacle Amphibalanus amphitrite. Marine Biology Research, 10(8):799–806, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.863352.

Matic-Skoko:2014:FPM

[661] Sanja Matić-Skoko, Pero Tutman, Dubravka Bojanić Varezić, Daria Skaramuca, Domagoj Dikić, Duje Lisicić, and Bosko Skaramuca. Food preferences of the Mediterranean moray eel, Muraena helena (Pisces: Muraenidae), in the southern Adriatic Sea. Marine Biology Research, 10(8):807–815, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.863351.

Ng:2014:PMG

[662] Peter K. L. Ng and Arthur Anker. Pteromaja maklayi gen. et sp. nov., a remarkable new spider crab (Crustacea: Brachyura: Epialtidae) from Papua New Guinea. Marine Biology Research, 10(8):816–823, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.853127.

Kalayci:2014:ISD

[663] Ferhat Kalayci and Tuncay Yesilçiçek. Influence of season, depth and mesh size on the trammel nets catch composition and discard in the Southern Black Sea, Turkey. Marine Biology Research, 10(8): 824–832, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.853126.

Ng:2014:RJC

[664] Chin Soon Lionel Ng and Loke Ming Chou. Rearing juvenile ‘corals of opportunity’ in in situ nurseries — a reef rehabilitation approach for sediment-impacted environments. Marine Biology Research, 10(8): 833–838, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.853124.

Rojas-Hernandez:2014:UNM

[665] Noemi Rojas-Hernandez, David Véliz, and Luis Miguel Pardo. Use of novel microsatellite markers for population and paternity analysis in the commercially important crab Metacarcinus edwardsii (Brachyura: Cancridae). Marine Biology Research, 10(8):839–844, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.863350.

Dinesen:2014:RFM

[666] Grete E. Dinesen and Brian Morton. Review of the functional morphology, biology and perturbation impacts on the boreal, habitat-forming horse mussel Modiolus modiolus (Bivalvia: Mytilidae: Modiolinae). Marine Biology Research, 10(9):845–870, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.866250.

Botelho:2014:KNS

[667] Alessandra Prates Botelho, André Morgado Esteves, and Verônica da Fonsêca-Genevois. Known and new species of Sabatieria Rouville, 1903 (Araeolaimida: Comesomatidae) from the southwest Atlantic (Campos Basin, Brazil). Marine Biology Research, 10(9):871– 891, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.866249.

Paxton:2014:BDW

[668] Hannelore Paxton and Andrés Arias. Brooding deep-water onuphid polychaetes (Annelida) from the Bay of Biscay. Marine Biology Research, 10 (9):892–905, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.863354.

Singh:2014:DSN

[669] Ravail Singh, Dmitry M. Miljutin, Maria A. Miljutina, Pedro Martinez Arbizu, and Baban S. Ingole. Deep-sea nematode assemblages from a commercially important polymetallic nodule area in the Central Indian Ocean Basin. Marine Biology Research, 10(9):906–916, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.866251.

Li:2014:EPG

[670] Yun Li, Dapeng Sun, Zhenkui Qin, and Zhifeng Zhang. Expression pattern of the vitellogenin gene in the zhikong scallop, Chlamys farreri, during ontogenesis. Marine Biology Research, 10(9):917–926, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.863356.

Battaglia:2014:FHJ

[671] Pietro Battaglia, Simona Musolino, Valentina Esposito, Giovanni Ammendolia, Pierpaolo Consoli, Franco Andaloro, and Teresa Romeo. Feeding habits of juvenile fishes belonging to three medusivorous species (Centrolophidae and Nomeidae) from the Strait of Messina (central Mediterranean Sea). Marine Biology Research, 10(9):927–933, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.863357.

An:2014:RMF

[672] Tianchen An, Junchen Lyu, Wanxian Jia, Manchuriga Wang, Shuangshuang Wei, and Yingxia Zhang. Role of macrobenthic fauna in mangrove carbon fluxes indicated by their cellulase and hemicellulase activities. Marine Biology Research, 10(9):934–940, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2013.863355.

Tillmann:2014:ACV

[673] Urban Tillmann, Bernd Krock, and Bettina B. Taylor. Azadinium caudatum var. margalefii, a poorly known member of the toxigenic genus Azadinium (Dinophyceae). Marine Biology Research, 10(10): 941–956, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.866252.

Fall:2014:WBD

[674] Johanna Fall and Mette Skern-Mauritzen. White-beaked dolphin distribution and association with prey in the Barents Sea. Marine Biology Research, 10(10):957–971, 2014. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2013.872796.

Smirnov:2014:ROA

[675] Roman V. Smirnov. A revision of the Oligobrachiidae (Annelida: Pogonophora), with notes on the morphology and distribution of Oligobrachia haakonmosbiensis Smirnov. Marine Biology Research, 10(10): 972–982, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.872799.

Kaanneby:2014:TNS

[676] Tobias Kånneby, Sarah Atherton, and Rick Hochberg. Two new species of Musellifer (Gastrotricha: Chaetonotida) from Florida and Tobago and the systematic placement of the genus within Paucitubulatina. Marine Biology Research, 10(10):983–995, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2013.872797.

Mestre:2014:STS

[677] Frederico Mestre, Marco P. Bragança, Antonieta Nunes, and Manuel E. dos Santos. Satellite tracking of sea turtles released after prolonged captivity periods. Marine Biology Research, 10(10):996–1006, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2013.872801.

Katano:2014:ENC

[678] Toshiya Katano, Makoto Yoshida, Souichi Yamaguchi, Kenji Yoshino, Takaharu Hamada, Masumi Koriyama, and Yuichi Hayami. Effect of nutrient concentration and salinity on diel vertical migration of Chattonella marina (Raphidophyceae). Marine Biology Research, 10(10): 1007–1018, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.879987.

Chen:2014:GPA

[679] Binbin Chen and Dinghui Zou. Growth and photosynthetic activity of Sargassum henslowianum (Fucales, Phaeophyta) seedlings in responses to different light intensities, temperatures and CO_{2} levels under laboratory conditions. Marine Biology Research, 10(10):1019–1026, 2014. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2013.872798.

Brant:2014:TRE

[680] Kristopher M. Brant and Matthew R. Gilg. Testing the relative effectiveness of traditional and non-traditional antifouling substrates on barnacle and macroalgae settlement. Marine Biology Research, 10(10): 1027–1032, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.872800.

Gopal:2014:NSP

[681] Aiswarya Gopal, Abdul Jaleel Koovapurath Useph, Saramma Aikkarakunnath Varghese, and Sanjeevan Veloorkirakathil Narayana. A new species of polychaete, Pettibonella shompens sp. nov. (Orbiniidae), from the Nicobar Islands, North Indian Ocean. Marine Biology Research, 10(10): 1033–1037, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2013.879988.

Falkenhaug:2014:RJB

[682] Tone Falkenhaug. Review of jellyfish blooms in the Mediterranean and Black Sea. Marine Biology Research, 10(10):1038–1039, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.880790.

Anonymous:2014:RA

[683] Anonymous. Reviewer acknowledgement 2013. Marine Biology Research, 10(10):1040, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.914227.

Anonymous:2014:EB

[684] Anonymous. Editorial board. Marine Biology Research, 10(10): ??, 2014. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
17451000.2014.924764.

Uiblein:2015:WTB

[685] Franz Uiblein. What thrives between the waves and beneath the tides? — Marine Biology Research sails into its second decade. Marine Biology Research, 11(1):1–3, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2014.969737.

Buhl-Mortensen:2015:MPF

[686] Lene Buhl-Mortensen, Pål Buhl-Mortensen, Margaret F. J. Dolan, and Børge Holte. The MAREANO programme — a full coverage mapping of the Norwegian off-shore benthic environment and fauna. Marine Biology Research, 11(1):4–17, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2014.952312.

Kandula:2015:SDS

[687] Sujatha Kandula, Kantimahanti V. L. Shrikanya, and V. A. Iswarya Deepti. Species diversity and some aspects of reproductive biology and life history of groupers (Pisces: Serranidae: Epinephelinae) off the central eastern coast of India. Marine Biology Research, 11(1):18–33, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.949271.

Kilgallen:2015:TNS

[688] Niamh M. Kilgallen. Three new species of Hirondellea (Crustacea, Amphipoda, Hirondelleidae) from hadal depths of the Peru–Chile Trench. Marine Biology Research, 11(1):34–48, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.889309.

Almada:2015:ISI

[689] Frederico Almada, David Abecasis, David Villegas-Ríos, Sofia Henriques, Miguel P. Pais, Marisa Batista, Bárbara Horta e Costa, Joana Martins, Inês Tojeira, Nuno Vasco Rodrigues, Ricardo Araújo, Miguel Souto, Hany Alonso, Jesús M. Falcón, Filipe Henriques, Paulo Catry, Henrique Cabral, Manuel Biscoito, and Vitor C. Almada. Ichthyofauna of the Selvagens Islands. Do small coastal areas show high species richness in the northeastern Atlantic? Marine Biology Research, 11(1):49–61, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.889306.

Ringvold:2015:DSS

[690] Halldis Ringvold, Arne Hassel, Roger N. Bamber, and Lene Buhl-Mortensen. Distribution of sea spiders (Pycnogonida, Arthropoda) off northern Norway, collected by MAREANO. Marine Biology Research, 11(1):62–75, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.889308.

Canario:2015:NST

[691] Roberta Canário, Marcelle F. S. Badaró, Rodrigo Johnsson, and Elizabeth G. Neves. A new species of Troglocarcinus (Decapoda: Brachyura: Cryptochiridae) symbiotic with the Brazilian endemic coral Mussismilia (Anthozoa: Scleractinia: Mussidae). Marine Biology Research, 11(1):76–85, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.894243.

Figueiredo:2015:UIB

[692] Cátia Figueiredo, Hugo Diogo, João G. Pereira, and Ruth M. Higgins. Using information-based methods to model age and growth of the silver scabbardfish, Lepidopus caudatus, from the mid-Atlantic Ocean. Marine Biology Research, 11(1):86–96, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.889307.

Velip:2015:HBS

[693] Dinesh T. Velip and Chandrashekher U. Rivonker. Hexapus bidentatus sp. nov. (Crustacea: Decapoda: Brachyura: Hexapodidae), a new species from Goa, west coast of India. Marine Biology Research, 11 (1):97–105, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.889305.

Galil:2015:CPD

[694] Bella S. Galil. Comments on ‘Preliminary data on the genetic structure of a highly successful invading population of oyster suggesting its establishment dynamics in the Levant’ by Ateret Shabtay et al., Marine Biology Research, 2014. Marine Biology Research, 11(1): 106–108, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.949776. See [624].

Smale:2015:EAT

[695] Dan Smale. Ecology of Australian temperate reefs: the unique south. Marine Biology Research, 11(1):109–110, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.890288.

Pavlov:2015:CHI

[696] Dimitri A. Pavlov. Condition and health indicators of exploited marine fishes. Marine Biology Research, 11(1):110–112, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.904886.

Fenchel:2015:ELS

[697] Tom Fenchel. The extreme life of the sea. Marine Biology Research, 11(1):112, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.914226.

Nielsen:2015:RIW

[698] Jørgen G. Nielsen, Werner Schwarzhans, and Franz Uiblein. Review of the Indo-West Pacific ophidiid genera Sirembo and Spottobrotula (Ophidiiformes, Ophidiidae), with description of three new species. Marine Biology Research, 11(2):113–134, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.904885.

Gjosaeter:2015:REB

[699] Harald Gjøsæter, Bjarte Bogstad, Sigurd Tjelmeland, and Samuel Subbey. A retrospective evaluation of the Barents Sea capelin management advice. Marine Biology Research, 11(2):135–143, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.928414.

Martinho:2015:SAB

[700] Francisco Martinho, Andreia Pereira, Cristina Brito, Raquel Gaspar, and Inês Carvalho. Structure and abundance of bottlenose dolphins (Tursiops truncatus) in coastal Setúbal Bay, Portugal. Marine Biology Research, 11(2):144–156, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.894244.

Fang:2015:ETS

[701] Jinghui Fang, Jihong Zhang, Yi Liu, Zengjie Jiang, Yuze Mao, and Jianguang Fang. Effects of temperature and salinity on mortality and metabolism of Ophiopholis mirabilis. Marine Biology Research, 11(2): 157–167, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.904884.

Zhou:2015:SBA

[702] Linbin Zhou, Liangmin Huang, Yehui Tan, Xiping Lian, and Kaizhi Li. Size-based analysis of a zooplankton community under the influence of the Pearl River plume and coastal upwelling in the northeastern South China Sea. Marine Biology Research, 11(2):168–179, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.904882.

Corbella:2015:DND

[703] Cecília Corbella and Peter R. Møller. Description of a new deepsea eelpout Pachycara matallanasi sp. nov. from the Solomon Sea (western South Pacific Ocean). Marine Biology Research, 11(2):180– 187, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.894245.

Oesterwind:2015:DSM

[704] Daniel Oesterwind, Uwe Piatkowski, and Heinz Brendelberger. On distribution, size and maturity of shortfin squids (Cephalopoda, Ommastrephidae) in the North Sea. Marine Biology Research, 11(2): 188–196, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.894246.

Fazio:2015:PBH

[705] Francesco Fazio, Giuseppe Piccione, Francesca Arfuso, and Caterina Faggio. Peripheral blood and head kidney haematopoietic tissue response to experimental blood loss in mullet (Mugil cephalus). Marine Biology Research, 11(2):197–202, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.898850.

Chen:2015:NSS

[706] Cheng Ann Chen and Shabdin Mohd Long. A new species of Setosabatieria Platt, 1985 (Nematoda: Comesomatidae) from Chek Jawa, Singapore, with a key to valid species of the genus. Marine Biology Research, 11(2):203–208, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.898849.

Park:2015:LWR

[707] Joo Myun Park and Seok Nam Kwak. Length–weight relationships and reproductive characteristics of the crowned seahorse (Hippocampus coronatus) in eelgrass beds (Zostera marina) of Dongdae Bay, Korea. Marine Biology Research, 11(2):209–213, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.902536.

Jorgensen:2015:NCC

[708] Aslak Jørgensen and Nadja Møbjerg. Notes on the cryptobiotic capability of the marine arthrotardigrades Styraconyx haploceros (Halechiniscidae) and Batillipes pennaki (Batillipedidae) from the tidal zone in Roscoff, France. Marine Biology Research, 11(2):214–217, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.904883.

Pola:2015:NNF

[709] Marta Pola, Leila Carmona, Gonçalo Calado, and Juan Lucas Cervera. A new nudibranch, Flabellina albomaculata sp. nov. (Flabellinidae), from the Cape Verde Archipelago with comparisons among all eastern Atlantic violet Flabellina spp. Marine Biology Research, 11(2):218–222, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.923102.

Edgar:2015:FTA

[710] Kirsty Edgar. Foraminifera and their applications. Marine Biology Research, 11(2):223–224, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.946424.

Grandi:2015:RPP

[711] M. Florencia Grandi, Silvana L. Dans, and Enrique A. Crespo. The recovery process of a population is not always the same: the case of Otaria flavescens. Marine Biology Research, 11(3):225–235, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.932912.

Cericola:2015:PRM

[712] Michael James Cericola and Jason D. Williams. Prevalence, reproduction and morphology of the parasitic isopod Athelges takanoshimensis Ishii, 1914 (Isopoda: Bopyridae) from Hong Kong hermit crabs. Marine Biology Research, 11(3):236–252, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.928415.

Minin:2015:SUG

[713] Kirill V. Minin, Nikolay B. Petrov, and Irina P. Vladychenskaya. Sea urchins of the genus Gracilechinus Fell and Pawson, 1966 from the Pacific Ocean: Morphology and evolutionary history. Marine Biology Research, 11(3):253–268, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.928413.

Nair:2015:DCP

[714] Vijayalakshmi R. Nair, Karati Komal Kusum, Raghavan Gireesh, and Maheswari Nair. The distribution of the chaetognath population and its interaction with environmental characteristics in the Bay of Bengal and the Arabian Sea. Marine Biology Research, 11(3):269–282, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.914224.

Frameschi:2015:LHT

[715] Israel Fernandes Frameschi, Luciana Segura de Andrade, Carlos Eduardo Rocha Duarte Alencar, Gustavo Monteiro Teixeira, Vivian Fransozo, Lissandra Corrêa Fernandes-Góes, and Adilson Fransozo. Life-history traits of the red brocade hermit crab Dardanus insignis on the subtropical Brazilian coast. Marine Biology Research, 11(3):283–293, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.923100.

Meretta:2015:ECV

[716] Pablo Ezequiel Meretta and Gabriel Néstor Genzano. Epibiont community variation on two morphologically different hydroid colonies: Amphisbetia operculata and Plumularia setacea (Cnidaria, Hydrozoa). Marine Biology Research, 11(3):294–303, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.923101.

Lozano-Cortes:2015:ACJ

[717] Diego F. Lozano-Cortés and Fernando A. Zapata. Abundance and composition of juvenile scleractinian corals on a fringing reef (Little Reef) off San Andres Island, Colombian Caribbean. Marine Biology Research, 11(3):304–312, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.914222.

ElAyari:2015:ESE

[718] Tahani El Ayari, Sami Abidli, Youssef Lahbib, Pablo Rodríguez González, José Ignacio García Alonso, and Najoua Trigui-El Menif. The effect of size and epibiotic barnacles on imposex in Stramonita haemastoma collected from the northern coast of Tunisia. Marine Biology Research, 11(3):313–320, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.914223.

Yigitturk:2015:MCD

[719] Gürkan Yigittürk and Ataç Uzel. Microbial community diversity associated with Sarcotragus sp. and Petrosia ficiformis from the Aegean Sea based on 16S rDNA-DGGE fingerprinting. Marine Biology Research, 11 (3):321–329, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.914225.

He:2015:CLI

[720] Song He and Jarle Mork. Characteristic LDH isozyme electrophoretic patterns in six flatfish species in the Trondheimsfjord, Norway and their utility for the detection of natural species hybrids. Marine Biology Research, 11(3):330–334, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.940976.

Williams:2015:SBC

[721] Tom Williams. Sharks, batoids and chimaeras of the North Atlantic — FAO species catalogue for fisheries purposes no. 7 / North Atlantic sharks relevant to fisheries management: a pocket guide / North Atlantic batoids and chimaeras relevant to fisheries management: a pocket guide.

Marine Biology Research, 11(3):335–336, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.950588.

Keppel:2015:IEO

[722] Elise A. Keppel, Ricardo A. Scrosati, and Simon C. Courtenay. Interactive effects of ocean acidification and warming on subtidal mussels and sea stars from Atlantic Canada. Marine Biology Research, 11 (4):337–348, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.932914.

Kidawa:2015:PEA

[723] Dorota Kidawa, Dariusz Jakubas, Katarzyna Wojczulanis-Jakubas, Lech Stempniewicz, Emilia Trudnowska, Rafał Boehnke, Liliana Keslinka-Nawrot, and Katarzyna Błachowiak-Samołyk. Parental efforts of an Arctic seabird, the little auk Alle alle, under variable foraging conditions. Marine Biology Research, 11(4):349–360, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.940974.

Sequeira:2015:DSS

[724] Vera Sequeira, Ana Neves, Rafaela Barros Paiva, Ana Rita Vieira, Adelino Canário, and Leonel Serrano Gordo. Diversity of sexual strategies of fish species caught by bottom trawl off the western Iberian Peninsula. Marine Biology Research, 11(4):361–374, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2014.952308.

Bekkby:2015:RSU

[725] Trine Bekkby, Gro Angeltveit, Hege Gundersen, Lise Tveiten, and Kjell Magnus Norderhaug. Red sea urchins (Echinus esculentus) and water flow influence epiphytic macroalgae density. Marine Biology Research, 11(4):375–384, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.943239.

Meheust:2015:DBI

[726] Eléonore Méheust, Eric Alfonsi, Patrick Le Ménec, Sami Hassani, and Jean-Luc Jung. DNA barcoding for the identification of soft remains of prey in the stomach contents of grey seals (Halichoerus grypus) and harbour porpoises (Phocoena phocoena). Marine Biology Research, 11 (4):385–395, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.943240.

Wang:2015:GHS

[727] Hong Wei Wang, Yun Guan, Feng Qin Zhao, and Dan Zhao. Grateloupia huanghaiensis sp. nov. (Halymeniaceae, Rhodophyta), a peculiar new species from China. Marine Biology Research, 11(4):396–404, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.932913.

Sivertsen:2015:BAU

[728] Knut Sivertsen and Arne Bjørge. On the brink of the Arctic: Unusual intertidal sub-Arctic kelp associations in the porsangerfjord, North Norway. Marine Biology Research, 11(4):405–413, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.940975.

Eshragh:2015:MCS

[729] Roya Eshragh and Brian S. Leander. Molecular contributions to species boundaries in dicyemid parasites from eastern Pacific cephalopods. Marine Biology Research, 11(4):414–422, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.943241.

Henriques:2015:IGI

[730] Romina Henriques, Warren M. Potts, Warwick H. H. Sauer, and Paul W. Shaw. Incipient genetic isolation of a temperate migratory coastal sciaenid fish (Argyrosomus inodorus) within the Benguela Cold Current system. Marine Biology Research, 11(4):423–429, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.952309.

Bedini:2015:ENN

[731] Roberto Bedini, Marco Bedini, Lisa Bonechi, and Luigi Piazzi. Effects of non-native turf-forming Rhodophyta on mobile macro-invertebrate assemblages in the north-western Mediterranean Sea. Marine Biology Research, 11(4):430–437, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.952310.

Rosas-Luis:2015:FFH

[732] Rigoberto Rosas-Luis and Pilar Sánchez. Food and feeding habits of Alloteuthis media in the Western Mediterranean Sea. Marine Biology Research, 11(4):438–442, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.962543.

Hessing-Lewis:2015:GST

[733] Margot Hessing-Lewis. A guide to southern temperate seagrasses. Marine Biology Research, 11(4):443–444, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.963825.

Reeves:2015:WWS

[734] Randall Reeves. Whale-watching: Sustainable tourism and ecological management. Marine Biology Research, 11(4):444–446, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.983525.

Diefenderfer:2015:CC

[735] Heida L. Diefenderfer. Coastal conservation. Marine Biology Research, 11(4):446–448, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.999095.

Stenevik:2015:PED

[736] Erling Kåre Stenevik, Jon Helge Vølstad, Åge Høines, Sondre Aanes, Gudmundur J. Óskarsson, Jan Arge Jacobsen, and Øyvind Tangen. Precision in estimates of density and biomass of Norwegian spring-spawning herring based on acoustic surveys. Marine Biology Research, 11(5): 449–461, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.995672.

Miloslavic:2015:IVH

[737] Marijana Miloslavić, Davor Lucić, Marko Zarić, Barbara Gangai, and Ivona Onofri. The importance of vertical habitat gradients on zooplankton distribution in an enclosed marine environment (South Adriatic Sea). Marine Biology Research, 11(5):462–474, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.955802.

Uiblein:2015:NGG

[738] Franz Uiblein and Daniel C. Gledhill. A new goatfish of the genus Upeneus (Mullidae) from Australia and Vanuatu, with inter- and intraspecific comparisons. Marine Biology Research, 11(5):475–491, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.958088.

Fernandez:2015:CTT

[739] Marina O. Fernandez, Sergio A. Navarrete, and Antonio C. Marques. A comparison of temporal turnover of species from benthic cnidarian assemblages in tropical and subtropical harbours. Marine Biology Research, 11(5):492–503, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.955804.

Steer:2015:SVS

[740] Michael A. Steer and Anthony J. Fowler. Spatial variation in shape of otoliths for southern garfish Hyporhamphus melanochir — contribution to stock structure. Marine Biology Research, 11(5):504–515, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.952313.

Song:2015:DNS

[741] Sung Joon Song, Hans-Uwe Dahms, Pawana Kangtia, Supawadee Chullasorn, Jongseong Ryu, and Jong Seong Khim. Description of a new species of the genus Peltidium philippi, 1839 from southern Korea (Copepoda, Harpacticoida) — with a review of the genus. Marine Biology Research, 11(5):516–527, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.955801.

Escobar-Toledo:2015:MSS

[742] Fabián Escobar-Toledo, Manuel J. Zetina-Rejón, and Luis Orlando Duarte. Measuring the spatial and seasonal variability of community structure and diversity of fish by-catch from tropical shrimp trawling in the Colombian Caribbean Sea. Marine Biology Research, 11 (5):528–539, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.952311.

Titlyanov:2015:RSF

[743] Eduard A. Titlyanov, Tamara V. Titlyanova, Xiubao Li, Tatyana L. Kalita, and Hui Huang. Recent (2008–2012) seaweed flora of Hainan Island, South China Sea. Marine Biology Research, 11(5):540– 550, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.955803.

Chang:2015:PDB

[744] Xianbo Chang, Wenzheng Liu, Qi Yin, Hong Wang, and Xiaohua Zhang. Phylogenetic diversity and biological activities of marine actinomycetes isolated from sediments of the Yellow Sea Cold Water Mass, China. Marine Biology Research, 11(5):551–560, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2014.962540.

Uiblein:2015:CAM

[745] Franz Uiblein. Comparative approaches in Marine Biology Research — and making them comparable. Marine Biology Research, 11(6): 561–563, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1041533.

Knight:2015:CEI

[746] Nicole S. Knight, Carolyn Prentice, Michelle Tseng, and Mary I. O’Connor. A comparison of epifaunal invertebrate communities in native eelgrass Zostera marina and non-native Zostera japonica at tsawwassen, BC. Marine Biology Research, 11(6):564–571, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.985230.

Kotwicki:2015:DEC

[747] Lech Kotwicki, Maria Szymelfenig, Frank Fiers, and Bozena Graca. Diversity and environmental control of benthic harpacticoids of an offshore post-dredging pit in coastal waters of Puck Bay, Baltic Sea. Marine Biology Research, 11(6):572–583, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.962541.

Rabelo:2015:DPZ

[748] Emanuelle F. Rabelo, Marcelo de O. Soares, Luis Ernesto A. Bezerra, and Helena Matthews-Cascon. Distribution pattern of zoanthids (Cnidaria: Zoantharia) on a tropical reef. Marine Biology Research, 11 (6):584–592, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.962542.

Peda:2015:CPS

[749] Cristina Pedà, Pietro Battaglia, Alessia Scuderi, Alessandro Voliani, Cecilia Mancusi, Franco Andaloro, and Teresa Romeo. Cephalopod prey in the stomach contents of odontocete cetaceans stranded in the western Mediterranean Sea. Marine Biology Research, 11(6):593–602, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.966724.

Philp:2015:LHN

[750] Heather Philp, Amaya Albalat, and Gudrún Marteinsdóttir. Live holding of Nephrops norvegicus (Linnaeus, 1758) in land-based facilities: Health and condition effects. Marine Biology Research, 11(6): 603–612, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.971814.

Blasina:2015:PPS

[751] Gabriela E. Blasina, Andrea C. Lopez Cazorla, and Juan M. Díaz de Astarloa. Possible predation by the striped weakfish Cynoscion guatucupa on estuary-associated fishes in an Argentinian coastal lagoon. Marine Biology Research, 11(6):613–623, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.973417.

Sorochan:2015:DMC

[752] Kevin A. Sorochan, William D. P. Duguid, and Pedro A. Quijón. Diagnostic morphological characteristics of laboratory-reared Cancer oregonensis (Brachyura: Cancridae) with recommendations for identifying cancrid zoeae in the Salish Sea. Marine Biology Research, 11(6): 624–632, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.978337.

Spinelli:2015:DVD

[753] Mariela Spinelli, Carla Derisio, Patricia Martos, Marcelo Pájaro, Graciela Esnal, Hermes Mianzán, and Fabiana Capitanio. Diel vertical distribution of the larvacean Oikopleura dioica in a North Patagonian tidal frontal system (42^{o}–45^{o}S) of the SW Atlantic Ocean. Marine Biology Research, 11(6):633–643, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.978338.

Gao:2015:EOT

[754] Qian Gao, Zhao-Li Xu, Jia-Jie Chen, and Jing Zhu. Estimating the optimal temperature for pelagic Copepoda in the East China Sea. Marine Biology Research, 11(6):644–654, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2014.978339.

Young:2015:PIA

[755] Tim Young, Andrea C. Alfaro, Clara Sánchez-Lazo, and John Robertson. Putative involvement of adrenergic receptors in regulation of mussel (Perna canaliculus) larval settlement. Marine Biology Research, 11 (6):655–665, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.979833.

Purcell:2015:EPS

[756] Steven W. Purcell and Hampus Eriksson. Echinoderms piggybacking on sea cucumbers: Benign effects on sediment turnover and movement of hosts. Marine Biology Research, 11(6):666–670, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2014.962544.

Jacob:2015:NSS

[757] Jini Jacob, Koovapurath Useph Abdul Jaleel, Rosamma Philip, and Rayaroth Damodaran. A new species of Scaptrella (Nematoda: Monhysterida) from the continental margin of the southeastern Arabian Sea. Marine Biology Research, 11(6):671–676, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2015.1009468.

Demirel:2015:IDH

[758] Nazli Demirel. Ichthyoplankton dynamics in a highly urbanized estuary. Marine Biology Research, 11(7):677–688, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2015.1007873.

DalZotto:2015:ACN

[759] Matteo Dal Zotto. Antygomonas caeciliae, a new kinorhynch from the Mediterranean Sea, with report of mitochondrial genetic data for the phylum. Marine Biology Research, 11(7):689–702, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1007872.

Guerra-Garcia:2015:SFD

[760] José M. Guerra-García, Macarena Ros, and Elena Baeza-Rojano. Seasonal fluctuations and dietary analysis of fouling caprellids (Crustacea: Amphipoda) from marinas of southern Spain. Marine Biology Research, 11(7):703–715, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.985231.

Kovacic:2015:LDI

[761] Ines Kovacić, Maja Fafandel, and Nevenka Bihari. Lysosomal deoxyribonuclease II from the mussel Mytilus galloprovincialis: Characterization and seasonal activity. Marine Biology Research, 11(7): 716–724, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2014.985232.

Buhl-Mortensen:2015:EWF

[762] Pål Buhl-Mortensen, Eirik Tenningen, and Anne Britt Skar Tysseland. Effects of water flow and drilling waste exposure on polyp behaviour in Lophelia pertusa. Marine Biology Research, 11(7):725–737, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/17451000.2014.993651.

Lisnizer:2015:TPD

[763] Nora Lisnizer, Pablo García-Borboroglu, Miguel Pascual, and Pablo Yorio. Transfer processes drive population dynamics of kelp gull colonies in Patagonia: Implications for management strategies. Marine Biology Research, 11(7):738–746, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2014.993652.

Sundin:2015:HDM

[764] Josefin Sundin, Gunilla Rosenqvist, and Anders Berglund. Hypoxia delays mating in the broad-nosed pipefish. Marine Biology Research, 11 (7):747–754, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1007874.

Avila:2015:IHS

[765] Enrique Ávila, Benjamín Yáñez, and Laura Elena Vazquez-Maldonado. Influence of habitat structure and environmental regime on spatial distribution patterns of macroinvertebrate assemblages associated with seagrass beds in a southern Gulf of Mexico coastal lagoon. Marine Biology Research, 11(7):755–764, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1007875.

Pang:2015:DPB

[766] Tong Pang, Litao Zhang, Jianguo Liu, Hu Li, and Junpeng Li. Differences in photosynthetic behaviour of Kappaphycus alvarezii and Kappaphycus striatus during dehydration and rehydration. Marine Biology Research, 11(7):765–772, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1007876.

Montano:2015:MCN

[767] Simone Montano, Davide Seveso, Paolo Galli, Stefania Puce, and Bert W. Hoeksema. Mushroom corals as newly recorded hosts of the hydrozoan symbiont Zanclea sp. Marine Biology Research, 11(7): 773–779, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1009467.

Skeljo:2015:AGA

[768] Frane Skeljo, Jure Brcić, Vedran Vuletin, and Josipa Ferri. Age and growth of the axillary wrasse, Symphodus mediterraneus (L.) from the eastern Adriatic Sea. Marine Biology Research, 11(7):780–784, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2015.1016963.

Hauge:2015:GMF

[769] Kjellrun Hiis Hauge. Governance of marine fisheries and biodiversity conservation. Marine Biology Research, 11(7):785–787, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1025078.

Schizas:2015:NSL

[770] Nikolaos V. Schizas, Hans-Uwe Dahms, Pawana Kangtia, Paulo H. C. Corgosinho, and Alexandra M. Galindo Estronza. A new species of Longipedia Claus, 1863 (Copepoda: Harpacticoida: Longipediidae) from Caribbean mesophotic reefs with remarks on the phylogenetic affinities of Polyarthra. Marine Biology Research, 11(8):789–803, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1013556.

Dietrich:2015:NSM

[771] Anna Dietrich, Tatjana Hager, Regine Bönsch, Charlotte Winkelmann, Andreas Schmidt, and Arne Nygren. A new species of Myrianida (Autolytinae, Syllidae, Annelida) from the North Sea, with short notes on the distribution and habitat of Northeast Atlantic autolytines. Marine Biology Research, 11(8):804–813, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1016971.

SampaioSantAnna:2015:DPB

[772] Bruno Sampaio Sant’Anna, Joaquim Olinto Branco, Michel Mattar de Oliveira, Harry Boos, and Alexander Turra. Diet and population biology of the invasive crab Charybdis hellerii in southwestern Atlantic waters. Marine Biology Research, 11(8):814–823, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1024134.

Nunez:2015:DMS

[773] Lara Agulló Núñez, Elvar H. Hallfredsson, and Inger-Britt Falk-Petersen. Different maturity scales affect estimations of fecundity, TEP and spawning stock size of Greenland halibut, Reinhardtius hippoglossoides (Walbaum, 1792). Marine Biology Research, 11(8):824–833, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1031797.

Hoque:2015:SSF

[774] Mohammed Muzammel Hoque, Abu Hena Mustafa Kamal, Mohd Hanafi Idris, Osumanu Haruna Ahmed, A. S. M. Saifullah, and Md. Masum Billah. Status of some fishery resources in a tropical mangrove estuary of Sarawak, Malaysia. Marine Biology Research, 11(8):834–846, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2015.1016970.

Denadai:2015:RCT

[775] Marcia R. Denadai, Luciana Le Sueur-Maluf, Carla G. Marques, Antonia C. Z. Amaral, Isadora Adamo, Leonardo Q. Yokoyama, and Alexander Turra. Reproductive cycle of the trigonal clam Tivela mactroides (Bivalvia, Veneridae) in Caraguatatuba Bay, southeastern Brazil. Marine Biology Research, 11(8):847–858, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1024130.

Zawierucha:2015:FOM

[776] Krzysztof Zawierucha, Katarzyna Grzelak, Lech Kotwicki, Lukasz Kaczmarek, and Reinhardt Møbjerg Kristensen. First observation of the marine tardigrades Batillipes mirus and Batillipes noerrevangi (Arthrotardigrada, Batillipedidae) from a strongly brackish part of the Polish Baltic Sea coast. Marine Biology Research, 11(8):859–868, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1024133.

Roche:2015:RCO

[777] Andrea Roche, Juliana Giménez, and Maite A. Narvarte. Reproductive cycle and oviposition in Zidona dufresnei (Caenogastropoda: Volutidae) from northern Patagonia, Argentina. Marine Biology Research, 11(8):869–878, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1024132.

Liew:2015:LPG

[778] Pui Ling Liew, Wei Lun Ng, and Soon Guan Tan. Levels and patterns of genetic variation in an Asian horseshoe crab species, Tachypleus gigas Müller, from the Malay Peninsula. Marine Biology Research, 11 (8):879–886, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1024135.

Urban-Malinga:2015:BMF

[779] Barbara Urban-Malinga. Biodiversity of meiobenthic free-living nematodes of the southern Baltic Sea — preliminary results of the community composition and structure in sandy sediments of Puck Bay. Marine Biology Research, 11(8):887–892, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1041530.

Anonymous:2015:C

[780] Anonymous. Corrigendum. Marine Biology Research, 11(8):893, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1049837.

Djurhuus:2015:SPM

[781] Anni Djurhuus, Jóhanna Jørgensen, Hjálmar Hátún, Høgni Hammershaimb Debes, and Debes Hammershaimb Christiansen. Seasonal progression of microbial communities on the Faroe shelf. Marine Biology Research, 11(9):895–908, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1041532.

Ferrario:2015:POM

[782] Jasmine Ferrario, Jean-Loup d’Hondt, Agnese Marchini, and Anna Occhipinti-Ambrogi. From the Pacific Ocean to the Mediterranean Sea: Watersipora arcuata, a new non-indigenous bryozoan in Europe. Marine Biology Research, 11(9):909–919, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1041531.

Antunes:2015:ADS

[783] Mariana Antunes, Vivian Fransozo, Giovana Bertini, Gustavo Monteiro Teixeira, and Adilson Fransozo. Abundance and distribution of the swimming crab Callinectes danae Smith, 1869 (Crustacea, Decapoda, Portunidae) in the Ubatuba region, southeastern Brazil. Marine Biology Research, 11(9):920–932, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1044998.

Cavole:2015:UOM

[784] Letícia Maria Cavole and Manuel Haimovici. The use of otolith microstructure in resolving issues of ageing and growth of young Micropogonias furnieri from southern Brazil. Marine Biology Research, 11 (9):933–943, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1031799.

Galvao:2015:AGD

[785] Márcia Santos Nunes Galvão and Alexandre Wagner Silva Hilsdorf. Assessing the genetic diversity of the mangrove oyster Crassostrea rhizophorae (Bivalvia, Ostreidae) by microsatellite markers in southeastern Brazil. Marine Biology Research, 11(9):944–954, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1047381.

Santos:2015:PEB

[786] Sérgio Ricardo Santos, Amanda Carvalho de Andrade, José Roberto Verani, and Marcelo Vianna. Population explosion of the burrfish Chilomycterus spinosus spinosus (Diodontidae, Tetraodontiformes) in a eutrophic tropical estuary. Marine Biology Research, 11(9):955–964, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1045000.

Islabao:2015:ISG

[787] Carolina Antuarte Islabão and Clarisse Odebrecht. Influence of salinity on the growth of Akashiwo sanguinea and Prorocentrum micans (Dinophyta) under acclimated conditions and abrupt changes. Marine Biology Research, 11(9):965–973, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1031798.

Rodland:2015:RAS

[788] Elisabeth Strandbråten Rødland and Arne Bjørge. Residency and abundance of sperm whales (Physeter macrocephalus) in the bleik canyon, Norway. Marine Biology Research, 11(9):974–982, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1031800.

Sun:2015:CDE

[789] Hongyan Sun, Junde Dong, Chunhua Ren, Lvping Zhang, Xueming Dan, Chang Chen, Yuanzhou Zhang, and Chaoqun Hu. Cloning and differential expression of Na,K-ATPase in Penaeus vannamei. Marine Biology Research, 11(9):983–989, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1024131.

Hernaez:2015:CGV

[790] Patricio Hernáez, Alexandra Gamboa-González, and Sammy De Grave. Callichirus garthi is a valid species, distinct from C. seilacheri (Decapoda: Axiidea: Callianassidae). Marine Biology Research, 11(9): 990–997, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1044999.

Tsutsui:2015:DSA

[791] Kenta Tsutsui, Eriko Shimada, and Yusuke Tsuruwaka. Deep-sea anemone (Cnidaria: Actiniaria) exhibits a positive behavioural response to blue light. Marine Biology Research, 11(9):998–1003, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1047382.

Fricke:2015:ERL

[792] Ronald Fricke. European red list of marine fishes. Marine Biology Research, 11(9):1004–1007, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1064535.

Johnson:2015:SSD

[793] William S. Johnson. Suprabenthic species distributions relative to smallscale bathymetric features along the Virginia coast, USA. Marine Biology Research, 11(10):1009–1020, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1055277.

Uiblein:2015:DRT

[794] Franz Uiblein and Gavin Gouws. Distinction and relatedness — taxonomic and genetic studies reveal a new species group of goatfishes (Upeneus; Mullidae). Marine Biology Research, 11(10):1021–1042, 2015. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2015.1064963.

Liu:2015:REC

[795] Chunxiang Liu and Dinghui Zou. Responses of elevated CO_{2} on photosynthesis and nitrogen metabolism in Ulva lactuca (Chlorophyta) at different temperature levels. Marine Biology Research, 11(10):1043– 1052, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1062520.

daCosta:2015:MCA

[796] Paulo Alberto Silva da Costa, Michael Maia Mincarone, Adriana da Costa Braga, Agnaldo Silva Martins, Helena Passeri Lavrado, Manuel Haimovici, and Ana Paula da Costa Falcão. Megafaunal communities along a depth gradient on the tropical Brazilian continental margin. Marine Biology Research, 11(10):1053–1064, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2015.1062521.

Castrillon-Cifuentes:2015:RPC

[797] Ana Lucia Castrillón-Cifuentes, Carlos G. Muñoz, and Fernando A. Zapata. Reproductive patterns of the coral Pocillopora damicornis at Gorgona Island, Colombian Pacific Ocean. Marine Biology Research, 11(10): 1065–1075, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1056798.

Filippova:2015:MRD

[798] Nadezhda A. Filippova, Alexandra V. Gerasimova, and Nikolay V. Maximovich. Methodical recommendations for the description of soft bottom communities in the littoral zone. Marine Biology Research, 11(10): 1076–1084, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1056799.

Yang:2015:IML

[799] Zhipeng Yang, Hong Yu, Ruihai Yu, and Qi Li. Induced metamorphosis in larvae of the veined rapa whelk Rapana venosa using chemical cues. Marine Biology Research, 11(10):1085–1092, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1062518.

Zhang:2015:EAC

[800] Yiwen Zhang, Jiangtao Wang, Liju Tan, Jing Cao, and Hui Li. Effect of allelopathy on the competition and succession of Skeletonema costatum and Prorocentrum donghaiense. Marine Biology Research, 11(10): 1093–1099, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1062519.

Rodriguez-Barreras:2015:EFR

[801] Ruber Rodríguez-Barreras, Alain Durán, Julián Lopéz-Morell, and Alberto M. Sabat. Effect of fish removal on the abundance and size structure of the sea urchin Diadema antillarum: a field experiment. Marine Biology Research, 11(10):1100–1107, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2015.1064140.

Nakayama:2015:RNI

[802] Naohide Nakayama and Hiromitsu Endo. Redescription of Nezumia infranudis (Gilbert and Hubbs, 1920), with the first record of the species from the Eastern Indian Ocean (Actinopterygii: Gadiformes: Macrouridae). Marine Biology Research, 11(10):1108–1115, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1064962.

DiDario:2015:FGT

[803] Fabio Di Dario. Fishes: a guide to their diversity. Marine Biology Research, 11(10):1116–1117, 2015. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1071851.

Boenigk:2015:CEB

[804] Jens Boenigk. The choanoflagellates: Evolution, biology and ecology. Marine Biology Research, 11(10):1118–1119, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1080846.

Anonymous:2015:RA

[805] Anonymous. Reviewer acknowledgement 2014. Marine Biology Research, 11(10):1120, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1083360.

Anonymous:2015:EB

[806] Anonymous. Editorial board. Marine Biology Research, 11(10): ??, 2015. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1114293.

Uiblein:2016:DAP

[807] Franz Uiblein. Developmental assistance in publishing marine biology research — quo vadis? Marine Biology Research, 12(1):1–3, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1115644.

Angiulli:2016:PCP

[808] Elisa Angiulli, Luciana Sola, Giandomenico Ardizzone, Chiheb Fassatoui, and Anna Rita Rossi. Phylogeography of the common pandora Pagellus erythrinus in the central Mediterranean Sea: sympatric mitochondrial lineages and genetic homogeneity. Marine Biology Research, 12 (1):4–15, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1069355.

Oskarsson:2016:FEM

[809] Gudmundur J. Óskarsson, Asta Gudmundsdottir, Sveinn Sveinbjörnsson, and Torsteinn Sigurdsson. Feeding ecology of mackerel and dietary overlap with herring in Icelandic waters. Marine Biology Research, 12 (1):16–29, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1073327.

Genitsaris:2016:MMC

[810] Savvas Genitsaris, Sebastien Monchy, Jeremie Denonfoux, Stephanie Ferreira, Konstantinos Ar. Kormas, Telesphore Sime-Ngando, Eric Viscogliosi, and Urania Christaki. Marine microbial community structure assessed from combined metagenomic analysis and ribosomal amplicon deep-sequencing. Marine Biology Research, 12(1):30–42, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1084425.

Marco-Herrero:2016:LDP

[811] Elena Marco-Herrero, Pilar Drake, J. Ignacio González-Gordillo, and Jose A. Cuesta. Larval development of the pea crab Afropinnotheres monodi Manning, 1993 (Decapoda, Pinnotheridae) using planktoncollected and laboratory-reared specimens: effects of temperature. Marine Biology Research, 12(1):43–55, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1080369.

Garcia-Sanz:2016:CPA

[812] Sara García-Sanz, Pablo G. Navarro, Lydia Png-Gonzalez, and Fernando Tuya. Contrasting patterns of amphipod dispersion in a seagrass meadow between day and night: consistency through a lunar cycle. Marine Biology Research, 12(1):56–65, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1069357.

Jung:2016:OWW

[813] Jean-Luc Jung, Wim C. Mullié, Koen Van Waerebeek, Moulaye Mohamed Wagne, Abdellahi Samba Ould Bilal, Zein El Abidine Ould Sidaty, Lola Toomey, Eléonore Méheust, and Frédéric Marret. Omura’s whale off West Africa: autochthonous population or inter-oceanic vagrant in the Atlantic Ocean? Marine Biology Research, 12(1):66–75, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1084424.

Arechavala-Lopez:2016:LSU

[814] Pablo Arechavala-Lopez, Ingebrigt Uglem, Marius Berg, Pål Arne Bjørn, and Bengt Finstad. Large-scale use of fish traps for monitoring sea trout (Salmo trutta) smolts and sea lice (Lepeophtheirus salmonis) infestations: efficiency and reliability. Marine Biology Research, 12(1): 76–84, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1069356.

Overdyk:2016:ARS

[815] Lauren M. Overdyk, Robert E. Scheibling, and Michael F. Barker. Asexual reproduction and somatic growth of the fissiparous sea star Allostichaster polyplax in New Zealand. Marine Biology Research, 12 (1):85–95, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1084423.

DosSantosSchwan:2016:IWH

[816] Isabela Dos Santos Schwan, Ana Claudia Dos Santos Brasil, David Neves, and Gustavo M. Dias. The invasive worm Hydroides elegans (Polychaeta — Serpulidae) in southeastern Brazil and its potential to dominate hard substrata. Marine Biology Research, 12(1):96–103, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1080370.

Vijaylaxmi:2016:FDR

[817] Jayaram Vijaylaxmi, Vinay P. Padate, and Chandrashekher U. Rivonker. First distributional record of Carupella banlaensis from India. Marine Biology Research, 12(1):104–111, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1073328.

Fulton:2016:EFC

[818] Christopher J. Fulton. Ecology of fishes on coral reefs. Marine Biology Research, 12(1):112–113, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1112018.

Sulcius:2016:VMB

[819] Sigitas Sulcius and Karin Holmfeldt. Viruses of microorganisms in the Baltic Sea: current state of research and perspectives. Marine Biology Research, 12(2):115–124, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1118514.

Mendonca:2016:CAC

[820] Vanessa Mendonça, Catarina Vinagre, Diana Boaventura, Henrique Cabral, and Ana C. F. Silva. Chitons’ apparent camouflage does not reduce predation by green crabs Carcinus maenas. Marine Biology Research, 12(2):125–132, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1088951.

Miyashita:2016:DST

[821] Leonardo K. Miyashita and Danilo Calliari. Distribution and salinity tolerance of marine mysids from a subtropical estuary, Brazil. Marine Biology Research, 12(2):133–145, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1099678.

Liu:2016:VAP

[822] Rui Liu, Limei Qiu, Xin Zhao, Huan Zhang, Lingling Wang, Zhanhui Hou, Dahai Gao, and Linsheng Song. Variation analysis of pathogenic Vibrio spp. and Pseudomonas spp. in Changhai mollusc farming waters using real-time PCR assay during 2011–2014. Marine Biology Research, 12(2):146–157, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1099679.

Jones:2016:IFE

[823] Salome Jones, Renzo Perissinotto, Nicola Kim Carrasco, and Andre Vosloo. Impact of a flood event on the zooplankton of an estuarine lake. Marine Biology Research, 12(2):158–167, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2015.1104417.

Pejovic:2016:DBA

[824] Ivana Pejovic, Alba Ardura, Laura Miralles, Andres Arias, Yaisel J. Borrell, and Eva Garcia-Vazquez. DNA barcoding for assessment of exotic molluscs associated with maritime ports in northern Iberia. Marine Biology Research, 12(2):168–176, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1112016.

Bi:2016:SVB

[825] Jinhong Bi, Qi Li, Hong Yu, Zhixin Zhang, Yan Lian, Renjie Wang, and Tingyu Wang. Seasonal variations in biochemical composition during the reproductive cycle of the veined rapa whelk Rapana venosa (Valenciennes, 1846) from the northern coast of China. Marine Biology Research, 12(2):177–185, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2015.1125002.

Nordstrom:2016:MIH

[826] Marie C. Nordström, Katri Aarnio, and Erik Bonsdorff. Mesograzer identity, not host algae, determines consumer stable isotope ratios. Marine Biology Research, 12(2):186–192, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1112017.

Sidari:2016:SRM

[827] Rossana Sidari. A simple and rapid method for separation and isolation of marine algal species from naturally evolved populations. Marine Biology Research, 12(2):193–199, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1125003.

Castro:2016:NME

[828] José I. Castro, Keiichi Sato, and Ashby B. Bodine. A novel mode of embryonic nutrition in the tiger shark, Galeocerdo cuvier. Marine Biology Research, 12(2):200–205, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1099677.

Smit:2016:TNS

[829] Harry Smit. Two new species of the marine water mite family Pontarachnidae from Queensland, Australia (Acari: Hydrachnidia). Marine Biology Research, 12(2):206–209, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2015.1118515.

Craig:2016:RSP

[830] Matthew T. Craig. Review of the splendid perches, Callanthias (Percoidei: Callanthiidae). Marine Biology Research, 12(2):210, 2016. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2016.1141430.

Medina:2016:RMZ

[831] Alonso I. Medina, M. Alejandra Romero, and Maite A. Narvarte. Radular morphology of Zidona dufresnei (Neogastropoda: Volutidae) and an analysis of its variability along the distributional range. Marine Biology Research, 12(3):211–220, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1131297.

Meier:2016:FAC

[832] Sonnich Meier, Stig Falk-Petersen, Lars Aage Gade-Sørensen, Michael Greenacre, Tore Haug, and Ulf Lindstrøm. Fatty acids in common minke whale (Balaenoptera acutorostrata) blubber reflect the feeding area and food selection, but also high endogenous metabolism. Marine Biology Research, 12(3):221–238, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1118513.

Shimabukuro:2016:PDS

[833] Maurício Shimabukuro, Sandra Bromberg, and Ana M. S. Pires-Vanin. Polychaete distribution on the southwestern Atlantic continental shelf. Marine Biology Research, 12(3):239–254, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2015.1131299.

Gao:2016:DEL

[834] Dongkui Gao, Ruijing Wan, Qian Ma, Xiumei Zhang, and Xiaodong Bian. Development of eggs and larvae of Stolephorus commersonnii and taxonomic key to fish eggs of the Clupeidae and Engraulidae off China. Marine Biology Research, 12(3):255–267, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2015.1125004.

Rumbold:2016:PDT

[835] Carlos E. Rumbold, Trinidad Ruíz Barlett, María A. Gavio, and Sandra M. Obenat. Population dynamics of two invasive amphipods in the Southwestern Atlantic: Monocorophium acherusicum and Ericthonius punctatus (Crustacea). Marine Biology Research, 12(3): 268–277, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1142091.

Yong:2016:MLV

[836] Yu Lin Yong, Li Lee Chew, Choon Weng Lee, and Ving Ching Chong. Monsoonal and lunar variability in microzooplankton abundance and community structure in the Terusan mangrove creek (Malaysia). Marine Biology Research, 12(3):278–293, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1143107.

Li:2016:DPF

[837] Qian Li, Xin Wang, Xianhua Liu, Nianzhi Jiao, and Guangyi Wang. Diversity of parasitic fungi associated with phytoplankton in Hawaiian waters. Marine Biology Research, 12(3):294–303, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1088950.

Costa:2016:FFI

[838] Eudriano F. S. Costa, June F. Dias, and Hilario Murua. Fecundity of fishes inhabiting coastal and estuarine environments in the southwest Atlantic Ocean. Marine Biology Research, 12(3):304–315, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1136065.

Borecka:2016:CET

[839] Anna Borecka, Urszula Janas, and Halina Kendzierska. The combined effect of temperature and salinity changes on osmoregulation and haemocyanin concentration in Saduria entomon (Linnaeus, 1758). Marine Biology Research, 12(3):316–322, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1142092.

Carrasco:2016:ISG

[840] Sergio A. Carrasco and Alejandro Pérez-Matus. Inshore spawning grounds of the squid Doryteuthis gahi suggest the consistent use of defoliated kelp Lessonia trabeculata in Central Chilean waters. Marine Biology Research, 12(3):323–328, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2015.1136064.

Nazari:2016:MEM

[841] Hassan Nazari. Migration ecology of marine fishes, by David hallock Secor. Marine Biology Research, 12(3):329–330, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1140899.

Dias:2016:IPA

[842] Marta Dias, Joana Roma, Catarina Fonseca, Maria Pinto, Henrique N. Cabral, Ana Silva, and Catarina Vinagre. Intertidal pools as alternative nursery habitats for coastal fishes. Marine Biology Research, 12 (4):331–344, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1143106.

Singh:2016:NCI

[843] Ravail Singh, Dmitry M. Miljutin, Ann Vanreusel, Teresa Radziejewska, Maria M. Miljutina, Alexei Tchesunov, Christian Bussau, Valentina Galtsova, and Pedro Martinez Arbizu. Nematode communities inhabiting the soft deep-sea sediment in polymetallic nodule fields: do they differ from those in the nodule-free abyssal areas? Marine Biology Research, 12(4):345–359, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1148822.

Denisenko:2016:DBP

[844] Nina V. Denisenko, Peter J. Hayward, Ole S. Tendal, and Jan Sørensen. Diversity and biogeographic patterns of the bryozoan fauna of the Faroe Islands. Marine Biology Research, 12(4):360–378, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1148817.

NagataRagagnin:2016:WMG

[845] Marilia Nagata Ragagnin, Bruno Sampaio Sant’Anna, Daniel Gorman, Claudio Campi De Castro, André Paulo Tschiptschin, and Alexander Turra. What makes a good home for hermits? Assessing gastropod shell density and relative strength. Marine Biology Research, 12(4): 379–388, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1148818.

Ulate:2016:LRE

[846] Karol Ulate, Carlos Sánchez, Alexandra Sánchez-Rodríguez, David Alonso, Octavio Aburto-Oropeza, and Leonardo Huato-Soberanis. Latitudinal regionalization of epibenthic macroinvertebrate communities on rocky reefs in the Gulf of California. Marine Biology Research, 12 (4):389–401, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1143105.

Rosas-Luis:2016:ATE

[847] Rigoberto Rosas-Luis, Joan Navarro, Pilar Sánchez, and José L. Del Río. Assessing the trophic ecology of three sympatric squid in the marine ecosystem off the Patagonian Shelf by combining stomach content and stable isotopic analyses. Marine Biology Research, 12(4): 402–411, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1142094.

Wojcik-Fudalewska:2016:LTS

[848] Dagmara Wójcik-Fudalewska and Monika Normant-Saremba. Longterm studies on sex and size structures of the non-native crab Eriocheir sinensis from Polish coastal waters. Marine Biology Research, 12(4):412–418, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1148820.

Tikochinski:2016:MAR

[849] Yaron Tikochinski, Barry Russell, Yossi Hyams, Uzi Motro, and Daniel Golani. Molecular analysis of the recently described lizardfish Saurida lessepsianus (Synodontidae) from the Red Sea and the Mediterranean, with remarks on its phylogeny and genetic bottleneck effect. Marine Biology Research, 12(4):419–425, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1148821.

Xiao:2016:IYR

[850] Yongshuang Xiao, Daoyuan Ma, Ming Dai, Qinghua Liu, Zhizhong Xiao, Jun Li, Jing Liu, Shihong Xu, and Qiqun Cheng. The impact of Yangtze River discharge on the genetic structure of a population of the rock bream, Oplegnathus fasciatus. Marine Biology Research, 12 (4):426–434, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1154576.

Montemarano:2016:DCS

[851] Justin J. Montemarano, Jason Havelin, and Matthew Draud. Diet composition of the smooth dogfish (Mustelus canis) in the waters of Long Island, New York, USA. Marine Biology Research, 12(4): 435–442, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1148819.

Cao:2016:MOL

[852] Cui Cui Cao, Ya Zhuo Li, and Hong Wei Wang. Morphological observation, life history and rbc L gene sequence analysis of a new species, Grateloupia variata sp. nov. (Halymeniaceae, Rhodophyta) from Qingdao, China. Marine Biology Research, 12(4):443–453, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2015.1104416.

Nilsson:2016:BRP

[853] Göran Nilsson. Book review: The physiology of fishes, fourth edition, edited by D. H. Evans, J. B. Claiborne and S. Currie. Marine Biology Research, 12(4):454, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2016.1169299.

Neves:2016:FAK

[854] Ricardo C. Neves, Martin V. Sørensen, and Maria Herranz. First account on kinorhynchs from Portugal, with the description of two new species: Echinoderes lusitanicus sp. nov. and E. reicherti sp. nov. Marine Biology Research, 12(5):455–470, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1154973.

Deli:2016:PMC

[855] Temim Deli, Sara Fratini, Lapo Ragionieri, Khaled Said, Noureddine Chatti, and Christoph D. Schubart. Phylogeography of the marbled crab Pachygrapsus marmoratus (Decapoda, Grapsidae) along part of the African Mediterranean coast reveals genetic homogeneity across the Siculo–Tunisian Strait versus heterogeneity across the Gibraltar Strait. Marine Biology Research, 12(5):471–487, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1154972.

Esquete:2016:GUC

[856] Patricia Esquete and George D. F. Wilson. The genus Uromunna (Crustacea: Isopoda: Munnidae) in New South Wales, Australia, with a key for all known species. Marine Biology Research, 12(5):488–512, 2016. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2015.1131298.

Schaal:2016:BTS

[857] Gauthier Schaal, Jean-Charles Leclerc, Gabin Droual, Cedric Leroux, and Pascal Riera. Biodiversity and trophic structure of invertebrate assemblages associated with understorey red algae in a Laminaria digitata bed. Marine Biology Research, 12(5):513–523, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2016.1164318.

Peiro-Alcantar:2016:SVD

[858] M. Teresa Peiro-Alcantar, René Funes-Rodríguez, Rogelio González-Armas, Reginaldo Durazo, and Pablo Del Monte Luna. Spatiotemporal variability of demersal fish larvae assemblages in the southern region of the California Current. Marine Biology Research, 12(5): 524–540, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1164319.

Lopez-Mendilaharsu:2016:DTP

[859] Milagros López-Mendilaharsu, Gabriela M. Vélez-Rubio, Cecilia Lezama, Anita Aisenberg, Antonia Bauzá, Laura Berrondo, Victoria Calvo, Noel Caraccio, Andrés Estrades, Martín Hernández, Martín Laporta, Gustavo Martínez-Souza, Melisa Morales, Verónica Quirici, Mariana Ríos, and Alejandro Fallabrino. Demographic and tumour prevalence data for juvenile green turtles at the coastal-marine protected area of cerro verde, Uruguay. Marine Biology Research, 12(5):541–550, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1169294.

Athanasiadis:2016:LFC

[860] Athanasios Athanasiadis. Leptophytum flavescens comb. nov. (Corallinales, Rhodophyta), an Arctic endemic from the sublittoral of NW Spitsbergen, North Norway, and western Novaya Zemlya, with epitypification of L. laeve. Marine Biology Research, 12(5):551–558, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1164320.

Kjesbu:2016:PSU

[861] Olav Sigurd Kjesbu and Svein Sundby. Per Solemdal (1941–2016) — an unconventional marine scientist. Marine Biology Research, 12(5): 559–562, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1183793.

Watts:2016:SNS

[862] Andrew J. R. Watts, Amaya Albalat, Ian P. Smith, Robert J. A. Atkinson, and Douglas M. Neil. Seasonal nutritional status in Norway lobsters, Nephrops norvegicus (L.): are females nutritionally compromised over the winter? Marine Biology Research, 12(6):563–572, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1174337.

Pica:2016:MOL

[863] Daniela Pica, Aline Tribollet, Stjepko Golubic, Marzia Bo, Cristina Gioia Di Camillo, Giorgio Bavestrello, and Stefania Puce. Microboring organisms in living stylasterid corals (Cnidaria, Hydrozoa). Marine Biology Research, 12(6):573–582, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1169298.

Fang:2016:ERP

[864] Jinghui Fang, Jihong Zhang, Zengjie Jiang, Meirong Du, Yi Liu, Yuze Mao, Yaping Gao, and Jianguang Fang. Environmental remediation potential of Perinereis aibuhitensis (Polychaeta) based on the effects of temperature and feed types on its carbon and nitrogen budgets. Marine Biology Research, 12(6):583–594, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1177653.

Monteiro:2016:MHC

[865] Sílvia S. Monteiro, José V. Vingada, Alfredo López, Graham J. Pierce, Marisa Ferreira, Andrew Brownlow, Bjarni Mikkelsen, Misty Niemeyer, Robert J. Deaville, Catarina Eira, and Stuart Piertney. Major Histocompatibility Complex (MHC) class II sequence polymorphism in long-finned pilot whale (Globicephala melas) from the North Atlantic. Marine Biology Research, 12(6):595–607, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1174266.

DelBrio:2016:DBB

[866] Federico Del Brio, Marta Commendatore, Italo B. Castro, Patricia Gomes Costa, Gilberto Fillmann, and Gregorio Bigatti. Distribution and bioaccumulation of butyltins in the edible gastropod Odontocymbiola magellanica. Marine Biology Research, 12(6):608–620, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1169296.

Liuzzi:2016:CNS

[867] María G. Liuzzi, Juan López-Gappa, and Vanesa Olivieri. The clam Neilonella sulculata in San Jorge Gulf (Argentina, Southwest Atlantic): spatial distribution pattern and inverse relationship between size and density. Marine Biology Research, 12(6):621–630, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1174336.

Li:2016:CPP

[868] Yongfu Li, Jianguo Liu, Litao Zhang, and Tong Pang. Changes of photosynthetic performances in mature thalli of the red alga Gelidium amansii (Gelidiaceae) exposed to different salinities. Marine Biology Research, 12(6):631–639, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1177654.

Anikeeva:2016:TCD

[869] Oksana V. Anikeeva and Andrew J. Gooday. Taxonomic composition and distribution of soft-walled monothalamid foraminifera in the area of Zernov’s Phyllophora field (NW Black Sea). Marine Biology Research, 12(6):640–646, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1174267.

Varisco:2016:GMS

[870] Martín Varisco, Julia Colombo, Tomás Isola, and Julio Vinuesa. Growth and maturity of the spider crab Halicarcinus planatus (Brachyura: Hymenosomatidae) females in the southwestern Atlantic Ocean. Can these parameters be influenced by the population sex ratio? Marine Biology Research, 12(6):647–655, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1183792.

Gordo:2016:AGM

[871] Leonel Serrano Gordo, Ana Neves, Ana Rita Vieira, Rafaela Barros Paiva, and Vera Sequeira. Age, growth and mortality of the comber Serranus cabrilla (Linnaeus, 1758) in the Eastern Atlantic. Marine Biology Research, 12(6):656–662, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1169295.

Corbera:2016:CAL

[872] Jordi Corbera and Bella S. Galil. Cumacean assemblages on the Levantine shelf (Mediterranean Sea) — spatiotemporal trends between 2005 and 2012. Marine Biology Research, 12(6):663–672, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2016.1169297.

Dunphy:2016:DPM

[873] Brendon Dunphy. Diving physiology of marine mammals and seabirds, by Paul ponganis. Marine Biology Research, 12(6):673–674, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1185667.

Uiblein:2016:JBI

[874] Franz Uiblein, Gavin Gouws, Daniel C. Gledhill, and Karen Stone. Just off the beach: intrageneric distinctiveness of the bandtail goatfish Upeneus taeniopterus (Mullidae) based on a comprehensive alphataxonomy and barcoding approach. Marine Biology Research, 12(7): 675–694, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1190458.

Zhao:2016:ECE

[875] Jian-Sheng Zhao, Qian Zhang, Jie Liu, Pei-Dong Zhang, and Wen-Tao Li. Effects of copper enrichment on survival, growth and photosynthetic pigment of seedlings and young plants of the eelgrass Zostera marina. Marine Biology Research, 12(7):695–705, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1189083.

Alam:2016:PSG

[876] M. M. Mahbub Alam and Snæbjörn Pálsson. Population structure of the giant tiger shrimp Penaeus monodon in Bangladesh based on variation in microsatellites and immune-related genes. Marine Biology Research, 12(7):706–714, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1196820.

Vazquez:2016:ENS

[877] Diego M. Vazquez, Ezequiel Mabragaña, Valeria Gabbanelli, and Juan M. Díaz de Astarloa. Exploring nursery sites for oviparous chondrichthyans in the Southwest Atlantic (36^{o}S–41^{o}S). Marine Biology Research, 12(7):715–725, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1203948.

Waiho:2016:SDL

[878] Khor Waiho, Hanafiah Fazhan, and Mhd Ikhwanuddin. Size distribution, length–weight relationship and size at the onset of sexual maturity of the orange mud crab, Scylla olivacea, in Malaysian waters. Marine Biology Research, 12(7):726–738, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1200726.

Johnson:2016:GST

[879] Nathan D. Johnson and Anja Schulze. Genetic structure in two Phascolosoma species in the Pacific Ocean. Marine Biology Research, 12 (7):739–747, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1196819.

Du:2016:FST

[880] Jianguo Du, Xinqing Zheng, Teguh Peristiwady, Jianji Liao, Petrus Ch. Makatipu, Xijie Yin, Wenjia Hu, Wulan Koagouw, and Bin Chen. Food sources and trophic structure of fishes and benthic macroinvertebrates in a tropical seagrass meadow revealed by stable isotope analysis. Marine Biology Research, 12(7):748–757, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1183791.

Joung:2016:ELH

[881] Shoou-Jeng Joung, Nien-Fu Chen, Hua-Hsun Hsu, and Kwang-Ming Liu. Estimates of life history parameters of the oceanic whitetip shark, Carcharhinus longimanus, in the Western North Pacific Ocean. Marine Biology Research, 12(7):758–768, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1203947.

Dong:2016:LGD

[882] Zhijun Dong, Zhongyuan Liu, Dongyan Liu, Qingqing Liu, and Tingting Sun. Low genetic diversity and lack of genetic structure in the giant jellyfish Nemopilema nomurai in Chinese coastal waters. Marine Biology Research, 12(7):769–775, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1196818.

Taylor:2016:CMT

[883] Paul D. Taylor, Aileen Tan Shau-Hwai, Anatoliy B. Kudryavstev, and J. William Schopf. Carbonate mineralogy of a tropical bryozoan biota and its vulnerability to ocean acidification. Marine Biology Research, 12(7):776–780, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1203951.

Sobocinski:2016:BRF

[884] Kathryn L. Sobocinski. Book review: A Field Guide to Coastal Fishes: From Alaska to California, by Val Kells, Luiz A. Rocha and Larry G. Allen. Marine Biology Research, 12(7):781, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1208826.

Uiblein:2016:OHC

[885] Franz Uiblein. Oceans of hypotheses — comprehensive alpha-taxonomic accounts and integrative taxonomic revisions in Marine Biology Research. Marine Biology Research, 12(8):783–784, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1228979.

Dutz:2016:RTA

[886] Jörg Dutz, Josianne G. Støttrup, Claus Stenberg, and Peter Munk. Recent trends in the abundance of plaice Pleuronectes platessa and cod Gadus morhua in shallow coastal waters of the Northeastern Atlantic continental shelf — a review. Marine Biology Research, 12(8): 785–796, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1210806.

Sivadas:2016:BBP

[887] Sanitha K. Sivadas and Baban S. Ingole. Biodiversity and biogeography pattern of benthic communities in the coastal basins of India. Marine Biology Research, 12(8):797–816, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1203949.

Berasategui:2016:RTC

[888] Anabela A. Berasategui, Melisa D. Fernandez-Severini, M. Clara Menendez, Florencia Biancalana, M. Sofia Dutto, Valeria Guinder, M. Celeste Lopez-Abbate, Javier Chazarreta, and Monica S. Hoffmeyer. Reproductive trade-off of the copepod Acartia tonsa in a hypersaline estuary of the Southwestern Atlantic. Temporal variations in the morphology of eggs. Marine Biology Research, 12(8):817–829, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1203950.

Lezzi:2016:SPD

[889] Marco Lezzi, Michela Del Pasqua, Cataldo Pierri, and Adriana Giangrande. Settlement and population dynamics of the alien invasive Branchiomma bairdi (Annelida: Sabellidae) in the Mediterranean Sea: two years of observations in the Gulf of Taranto (Italy). Marine Biology Research, 12(8):830–841, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1206940.

Hammami:2016:MGC

[890] Ibtissem Hammami, Oum Kalthoum Ben Hassine, Myriam Kaouèche, and Lilia Bahri-Sfar. Morphological and genetic characterization of the sharpsnout seabream populations (Diplodus puntazzo, Sparidae) along a boundary area between the two Mediterranean basins. Marine Biology Research, 12(8):842–853, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1189080.

Santana:2016:MLS

[891] William Santana, Jessica Colavite, Juan Bolaños, Jesús E. Hernández, and Miguel Canepa. Morphology of the larval stages of Pitho aculeata (Gibbes, 1850) (Crustacea, Brachyura, Majoidea) and its implications on the taxonomic position of the genus. Marine Biology Research, 12 (8):854–863, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1209524.

Silva:2016:ADE

[892] Teresa Silva, Astthor Gislason, Olafur S. Astthorsson, and Gudrun Marteinsdottir. Abundance and distribution of early life stages of krill around Iceland during spring. Marine Biology Research, 12(8): 864–873, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1210808.

Mirbach:2016:OSS

[893] Christopher E. Mirbach and Simon J. Brandl. Ontogenetic shifts in the social behaviour of pairing coral reef rabbitfishes (Siganidae). Marine Biology Research, 12(8):874–880, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1203952.

Bokenhans:2016:AEM

[894] Verena Bökenhans, Gregorio Bigatti, and Andres Averbuj. Age estimation methods in the marine gastropod Buccinanops globulosus comparing shell marks and opercula growth rings. Marine Biology Research, 12(8):881–887, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1209526.

Lopez-Gappa:2016:RCN

[895] Juan López-Gappa, María G. Liuzzi, and Gabriel Genzano. Role of the clam Neilonella sulculata as substrate for the hydroid Bougainvillia macloviana on muddy bottoms of San Jorge Gulf (Argentina, Southwest Atlantic). Marine Biology Research, 12(8):888–893, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1210809.

Anonymous:2016:RA

[896] Anonymous. Reviewer acknowledgement 2015. Marine Biology Research, 12(8):894, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1229960.

Hesse:2016:RPR

[897] Jan Hesse, Jenni A. Stanley, and Andrew G. Jeffs. Relative predation risk in two types of habitat for juvenile Australasian spiny lobsters, Jasus edwardsii. Marine Biology Research, 12(9):895–906, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1236200.

Riisgaard:2016:URE

[898] Hans Ulrik Riisgård, Lars Kumala, and Katerina Charitonidou. Using the F/R-ratio for an evaluation of the ability of the demosponge Halichondria panicea to nourish solely on phytoplankton versus free-living bacteria in the sea. Marine Biology Research, 12(9):907–916, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1206941.

Zarrouk:2016:UMP

[899] Anis Zarrouk, Mohamed Salah Romdhane, and Free Espinosa. Usefulness of marine protected areas as tools for preserving the highly endangered limpet, Patella ferruginea, in the Mediterranean Sea. Marine Biology Research, 12(9):917–931, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1213396.

Kazanidis:2016:TSS

[900] Georgios Kazanidis and Ursula F. M. Witte. The trophic structure of Spongosorites coralliophaga-coral rubble communities at two northeast Atlantic cold water coral reefs. Marine Biology Research, 12(9): 932–947, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1216569.

Tonn:2016:SRS

[901] Nina Tonn, Sara C. Novais, Cátia S. E. Silva, Hugo A. Morais, João P. S. Correia, and Marco F. L. Lemos. Stress responses of the sea cucumber Holothuria forskali during aquaculture handling and transportation. Marine Biology Research, 12(9):948–957, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1218030.

DeBiasi:2016:ASD

[902] Anna Maria De Biasi, Lorenzo Pacciardi, and Luigi Piazzi. An asymmetrical sampling design as a tool for sustainability assessment of human activities in coastal systems: a fish farming study case. Marine Biology Research, 12(9):958–968, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1225958.

Poirier:2016:MSG

[903] Luke A. Poirier, Joshua Mohan, Rachael Speare, Jeff Davidson, Pedro A. Quijón, and Sophie St-Hilaire. Moulting synchrony in green crabs (Carcinus maenas) from Prince Edward Island, Canada. Marine Biology Research, 12(9):969–977, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1210810.

Ojamae:2016:DNU

[904] Karin Ojamäe, Anneliis Peterson, and Inga Lips. Dark nutrient uptake at low temperature and subsequent light use efficiency by the dinoflagellate Heterocapsa triquetra. Marine Biology Research, 12(9): 978–985, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1216570.

Forrester:2016:DFS

[905] Graham E. Forrester, Reuben J. A. Macfarlan, Allison J. Holevoet, and Sarah Merolla. Dislodgement force and shell morphology vary according to wave exposure in a tropical gastropod (Cittarium pica). Marine Biology Research, 12(9):986–992, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1225956.

Silberberger:2016:SSR

[906] Marc Jürgen Silberberger, Jonas Thormar, and Stein Fredriksen. Smallscale removal of seagrass (Zostera marina L.): effects on the infaunal community. Marine Biology Research, 12(9):993–1002, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1225959.

Tu:2016:TRC

[907] Tzu-Hsuan Tu, Chang-Feng Dai, and Ming-Shiou Jeng. Taxonomic revision of Coralliidae with descriptions of new species from New Caledonia and the Hawaiian Archipelago. Marine Biology Research, 12(10): 1003–1038, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1241411.

Lee:2016:IFC

[908] Matthew R. Lee and Jaen Mayorga-Dussarrat. The impact of feeding by Chilean flamingos (Phoenicopterus chilensis) on the meiofaunal assemblage of a tidal flat. Marine Biology Research, 12(10): 1039–1052, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1228975.

Curelovich:2016:RPT

[909] Jessica Curelovich, Gustavo A. Lovrich, and Javier A. Calcagno. The role of the predator Trophon geversianus in an intertidal population of Mytilus chilensis in a rocky shore of the Beagle Channel, Tierra del Fuego, Argentina. Marine Biology Research, 12(10): 1053–1063, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1228976.

Li:2016:ELT

[910] Hu Li, Jianguo Liu, Litao Zhang, and Tong Pang. Effects of low temperature stress on the antioxidant system and photosynthetic apparatus of Kappaphycus alvarezii (Rhodophyta, Solieriaceae). Marine Biology Research, 12(10):1064–1077, 2016. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1232827.

Tapia:2016:ABH

[911] Francisco Alejandro Puente Tapia, Luciana Mabel Díaz Briz, Agustín Schiariti, Rebeca Gasca, and Gabriel Genzano. Association between Hyperoche martinezii (Amphipoda: Hyperiidae) and ctenophores from the Buenos Aires coast, Argentina (South-western Atlantic Ocean). Marine Biology Research, 12(10):1078–1087, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1232829.

Daley:2016:BAH

[912] Meaghan C. Daley, Juanita Urban-Rich, and Pia H. Moisander. Bacterial associations with the hydromedusa Nemopsis bachei and scyphomedusa Aurelia aurita from the North Atlantic Ocean. Marine Biology Research, 12(10):1088–1100, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2016.1228974.

Liu:2016:DMC

[913] Zhuhong Liu, Chang Chen, Lei Gao, Yuyang Zhang, Jingjing Pei, Chunhua Ren, and Chaoqun Hu. Differences in microbial communities between healthy and bleached coral Acropora solitaryensis from Xisha Islands, South China Sea. Marine Biology Research, 12(10):1101–1108, 2016. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2016.1236201.

Rzad:2016:TFP

[914] Izabella Rzad and Jiljí Sitko. Trematode fauna (Platyhelminthes: Digenea) of some sea ducks wintering on the southern Baltic coast (NW Poland) — a general comparison. Marine Biology Research, 12(10): 1109–1117, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1228973.

Anonymous:2016:EB

[915] Anonymous. Editorial board. Marine Biology Research, 12(10): ??, 2016. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
17451000.2016.1274551.

Odebrecht:2017:BPLa

[916] Clarisse Odebrecht, Eduardo Resende Secchi, Paulo Cesar Abreu, and José Henrique Muelbert. Biota of the Patos Lagoon Estuary and adjacent marine coast: Long-term changes induced by natural and humanrelated factors. Marine Biology Research, 13(1):1–2, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2017.1288037.

Odebrecht:2017:BPLb

[917] Clarisse Odebrecht, Eduardo Resende Secchi, Paulo Cesar Abreu, José Henrique Muelbert, and Franz Uiblein. Biota of the Patos Lagoon Estuary and adjacent marine coast: long-term changes induced by natural and human-related factors. Marine Biology Research, 13 (1):3–8, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1258714.

Abreu:2017:CFD

[918] Paulo Cesar Abreu, Juliano Marangoni, and Clarisse Odebrecht. So close, so far: differences in long-term chlorophyll a variability in three nearby estuarine-coastal stations. Marine Biology Research, 13 (1):9–21, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1189081.

Mendes:2017:PBC

[919] Carlos Rafael Borges Mendes, Clarisse Odebrecht, Virginia Maria Tavano, and Paulo Cesar Abreu. Pigment-based chemotaxonomy of phytoplankton in the Patos Lagoon estuary (Brazil) and adjacent coast. Marine Biology Research, 13(1):22–35, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1189082.

Lanari:2017:DMP

[920] Marianna Lanari and Margareth Copertino. Drift macroalgae in the Patos Lagoon Estuary (southern Brazil): effects of climate, hydrology and wind action on the onset and magnitude of blooms. Marine Biology Research, 13(1):36–47, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1225957.

Teixeira-Amaral:2017:MPL

[921] Priscila Teixeira-Amaral, Waldemar José Appolinário Amaral, Danielle Ortiz de Ortiz, Vanessa Ochi Agostini, and Erik Muxagata. The mesozooplankton of the Patos Lagoon Estuary, Brazil: trends in community structure and secondary production. Marine Biology Research, 13 (1):48–61, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1248850.

Noleto-Filho:2017:STV

[922] Eurico Mesquita Noleto-Filho, Pedro Pucciarelli, and Luiz Felipe Cestari Dumont. Spatial and temporal variation in juvenile size distribution of the pink shrimp (Penaeus paulensis) in the Patos Lagoon Estuary, Brazil. Marine Biology Research, 13(1):62–73, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1248851.

Costa:2017:LTA

[923] Micheli Duarte de Paula Costa and José Henrique Muelbert. Long-term assessment of temporal variability in spatial patterns of early life stages of fishes to facilitate estuarine conservation. Marine Biology Research, 13(1):74–87, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1213397.

Martins:2017:SMS

[924] Agnaldo Silva Martins and Manuel Haimovici. Seasonal mesoscale shifts of demersal nekton assemblages in the subtropical South-western Atlantic. Marine Biology Research, 13(1):88–97, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1217025.

Garcia:2017:TVA

[925] Alexandre M. Garcia, Marlucy Coelho Claudino, Renata Mont’Alverne, Paula Evelyn Rubira Pereyra, Margareth Copertino, and João Paes Vieira. Temporal variability in assimilation of basal food sources by an omnivorous fish at Patos Lagoon Estuary revealed by stable isotopes (2010–2014). Marine Biology Research, 13(1):98–107, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1206939.

Dias:2017:BPL

[926] Rafael Antunes Dias, Giovanni Nachtigall Maurício, and Leandro Bugoni. Birds of the Patos Lagoon Estuary and adjacent coastal waters, southern Brazil: species assemblages and conservation implications. Marine Biology Research, 13(1):108–120, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1209525.

ResendeSecchi:2017:LTG

[927] Eduardo Resende Secchi, Silvina Botta, Marcelo M. Wiegand, Luara Azevedo Lopez, Pedro Friedrich Fruet, Rodrigo Cezar Genoves, and Juliana Couto Di Tullio. Long-term and gender-related variation in the feeding ecology of common bottlenose dolphins inhabiting a subtropical estuary and the adjacent marine coast in the western South Atlantic. Marine Biology Research, 13(1):121–134, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1213398.

Haimovici:2017:LTC

[928] Manuel Haimovici and Luís Gustavo Cardoso. Long-term changes in the fisheries in the Patos Lagoon estuary and adjacent coastal waters in Southern Brazil. Marine Biology Research, 13(1):135–150, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1228978.

Abdullah:2017:IKL

[929] Mohamed I. Abdullah, Stein Fredriksen, and Hartvig Christie. The impact of the kelp (Laminaria hyperborea) forest on the organic matter content in sediment of the west coast of Norway. Marine Biology Research, 13(2):151–160, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1240369.

Kousteni:2017:DSS

[930] Vasiliki Kousteni, Paraskevi K. Karachle, and Persefoni Megalofonou. Diet of the small-spotted catshark Scyliorhinus canicula in the Aegean Sea (eastern Mediterranean). Marine Biology Research, 13(2):161– 173, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1239019.

Triay-Portella:2017:OME

[931] Raül Triay-Portella, Raquel Ruiz-Díaz, José G. Pajuelo, and José A. González. Ovarian maturity, egg development, and offspring generation of the deep-water shrimp Plesionika edwardsii (Decapoda, Pandalidae) from three isolated populations in the eastern North Atlantic. Marine Biology Research, 13(2):174–187, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1239018.

Oesterwind:2017:PPR

[932] Daniel Oesterwind, Christiane Bock, Anja Förster, Michael Gabel, Christina Henseler, Paul Kotterba, Marion Menge, Dennis Myts, and Helmut M. Winkler. Predator and prey: the role of the round goby Neogobius melanostomus in the western Baltic. Marine Biology Research, 13(2):188–197, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1241412.

Gislason:2017:PDG

[933] Óskar Sindri Gíslason, Jónas P. Jónasson, Snæbjörn Pálsson, Jörundur Svavarsson, and Halldór P. Halldórsson. Population density and growth of the newly introduced Atlantic rock crab Cancer irroratus Say, 1817 (Decapoda, Brachyura) in Iceland: a four-year mark-recapture study. Marine Biology Research, 13(2):198–209, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1240875.

Herrero-Perezrul:2017:EEF

[934] María Dinorah Herrero-Pérezrul, Héctor Reyes-Bonilla, and Rebeca Granja-Fernández. Effects of environmental factors on the abundances of the basket stars Astrocaneum spinosum and Astrodictyum panamense (Ophiuroidea: Gorgonocephalidae) in the northern Gulf of California, Mexico. Marine Biology Research, 13(2):210–219, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1236203.

Lima:2017:NHC

[935] Daniel José Marcondes Lima and William Santana. A new hermit crab of the Paguristes tortugae complex (Crustacea: Anomura: Diogenidae), with a key to the western Atlantic species. Marine Biology Research, 13(2):220–228, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1239021.

Innocenti:2017:GTI

[936] Gianna Innocenti, Gianluca Stasolla, Menachem Goren, Nir Stern, Ya’arit Levitt-Barmats, Arik Diamant, and Bella S. Galil. Going down together: invasive host, Charybdis longicollis (Decapoda: Brachyura: Portunidae) and invasive parasite, Heterosaccus dollfusi (Cirripedia: Rhizocephala: Sacculinidae) on the upper slope off the Mediterranean coast of Israel. Marine Biology Research, 13(2):229–236, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1240873.

Ondes:2017:RGS

[937] Fikret Öndes, Michel J. Kaiser, and Lee G. Murray. Relative growth and size at onset of sexual maturity of the brown crab, Cancer pagurus in the Isle of Man, Irish Sea. Marine Biology Research, 13(2): 237–245, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1248849.

Marquez:2017:ESS

[938] Federico Márquez, Antonella C. Frizzera, and Nuria Vázquez. Environmentspecific shell shape variation in the boring mytilid Leiosolenus patagonicus. Marine Biology Research, 13(2):246–252, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1248848.

Sugeha:2017:BEA

[939] Hagi Yulia Sugeha. Biology and ecology of anguillid eels, edited by takaomi arai. Marine Biology Research, 13(2):253–254, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1249488.

Sun:2017:PCG

[940] Weiming Sun and Li Gao. Phylogeny and comparative genomic analysis of Pteriomorphia (Mollusca: Bivalvia) based on complete mitochondrial genomes. Marine Biology Research, 13(3):255–268, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1257810.

BenLtaief:2017:HCT

[941] Thouraya Ben Ltaief, Zaher Drira, Jean-Luc Devenon, Asma Hamza, Habib Ayadi, and Marc Pagano. How could thermal stratification affect horizontal distribution of depth-integrated metazooplankton communities in the Gulf of Gabes (Tunisia)? Marine Biology Research, 13 (3):269–287, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1248847.

Fujimoto:2017:NGS

[942] Shinta Fujimoto and Hiroshi Yamasaki. A new genus and species of Renaudarctidae (Heterotardigrada: Arthrotardigrada) from Ryukyu Archipelago, Japan. Marine Biology Research, 13(3):288–299, 2017. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2016.1257809.

Rosa:2017:AGS

[943] Daniela Rosa, Rui Coelho, Joana Fernandez-Carvalho, and Miguel N. Santos. Age and growth of the smooth hammerhead, Sphyrna zygaena, in the Atlantic Ocean: comparison with other hammerhead species. Marine Biology Research, 13(3):300–313, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1267366.

Rusk:2017:IEM

[944] Adam B. Rusk, Andrea C. Alfaro, Tim Young, Ellie Watts, and Serean L. Adams. Investigation of early mussel (Perna canaliculus) development using histology, SEM imaging, immunochemistry and confocal microscopy. Marine Biology Research, 13(3):314–329, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1257812.

Moraes:2017:DLS

[945] Maria Cecilia Medeiros Moraes and Helena Passeri Lavrado. Distribution of loliginid squids in a eutrophicated tropical coastal bay. Marine Biology Research, 13(3):330–341, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1262041.

Ouazzani:2017:GDE

[946] Khalil Chahdi Ouazzani, Touria Benazzou, Najib Charouki, François Bonhomme, and Malika Chlaida. Genetic differentiation of European anchovy (Engraulis encrasicolus) along the Moroccan coast reveals a phylogeographic break around the 25th parallel North. Marine Biology Research, 13(3):342–350, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1257811.

Chen:2017:GPR

[947] Binbin Chen, Dinghui Zou, and Mingjun Zhu. Growth and photosynthetic responses of Ulva lactuca (Ulvales, Chlorophyta) germlings to different pH levels. Marine Biology Research, 13(3):351–357, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1267367.

Anonymous:2017:RA

[948] Anonymous. Reviewer acknowledgement 2016. Marine Biology Research, 13(3):358, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1324696.

Halvorsen:2017:IHC

[949] Kim Tallaksen Halvorsen, Torkel Larsen, Tonje Knutsen Sørdalen, Leif Asbjørn Vøllestad, Halvor Knutsen, and Esben Moland Olsen. Impact of harvesting cleaner fish for salmonid aquaculture assessed from replicated coastal marine protected areas. Marine Biology Research, 13 (4):359–369, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1262042.

Aissaoui:2017:NGG

[950] Chrifa Aissaoui, Lee Ann Galindo, Nicolas Puillandre, and Philippe Bouchet. The nassariids from the Gulf of Gabès revisited (Neogastropoda, Nassariidae). Marine Biology Research, 13(4):370–389, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1273528.

Sofawi:2017:MRC

[951] Ahmad Bakrin Sofawi, Mohamad Zakaria Rozainah, Osman Normaniza, and Hashim Roslan. Mangrove rehabilitation on Carey Island, Malaysia: an evaluation of replanting techniques and sediment properties. Marine Biology Research, 13(4):390–401, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1267365.

Aastrom:2017:NGT

[952] Emmelie K. L. Åström, P. Graham Oliver, and Michael L. Carroll. A new genus and two new species of Thyasiridae associated with methane seeps off Svalbard, Arctic Ocean. Marine Biology Research, 13(4): 402–416, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1272699.

Morales-Bojorquez:2017:MLS

[953] Enrique Morales-Bojórquez and José Luis Pacheco-Bedoya. A mantle length structured stock assessment model for the jumbo squid, Dosidicus gigas, fishery of the Ecuadorian Pacific: a limited data approach. Marine Biology Research, 13(4):417–428, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2016.1272696.

Narvarte:2017:BIO

[954] Maite Narvarte, María Cecilia Salas, María Soledad Avaca, María Andrea Roche, Federico Delucchi, Matías Tartara, and Jorge Marcovecchio. Beyond imposex: other sexual alterations in a nassariid snail from Patagonia. Marine Biology Research, 13(4):429–438, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2016.1272700.

Robinson:2017:ELH

[955] Tamara B. Robinson, Haley R. Pope, and Mhairi E. Alexander. Early lifehistory processes and their implications for the invasion of the barnacle Balanus glandula. Marine Biology Research, 13(4):439–446, 2017. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2016.1267862.

Jonsdottir:2017:PNS

[956] Ingibjörg G. Jónsdóttir. Predation on northern shrimp (Pandalus borealis) by three gadoid species. Marine Biology Research, 13(4): 447–455, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1272697.

Xu:2017:GDP

[957] Dongdong Xu, Bao Lou, Wei Zhou, Ruiyi Chen, Wei Zhan, and Feng Liu. Genetic diversity and population differentiation in the yellow drum Nibea albiflora along the coast of the China Sea. Marine Biology Research, 13 (4):456–462, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1274033.

Puce:2017:TYHa

[958] Stefania Puce. Thirty years of the Hydrozoan Society: New challenges in hydrozoan studies. Marine Biology Research, 13(5):463–464, 2017. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2017.1346338.

Puce:2017:TYHb

[959] Stefania Puce and Franz Uiblein. Thirty years of the Hydrozoan Society: new challenges in hydrozoan studies. Marine Biology Research, 13 (5):465–468, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1312454.

Pica:2017:HDE

[960] Daniela Pica, Azzurra Bastari, Claudia Francesca Vaga, Cristina Gioia Di Camillo, Simone Montano, and Stefania Puce. Hydroid diversity of Eilat Bay with the description of a new Zanclea species. Marine Biology Research, 13(5):469–479, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1236202.

Manko:2017:SGA

[961] Maciej K. Mańko, Angelika W. Słomska, and Krzysztof Jazdzewski. Siphonophora of the Gulf of Aqaba (Red Sea) and their associations with crustaceans. Marine Biology Research, 13(5):480–485, 2017. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2017.1296577.

Mapstone:2017:TDL

[962] Gillian M. Mapstone, Laure Corbari, and Lenaick Menot. Two deepliving rhodaliids (Cnidaria, Siphonophora) from the Mid-Atlantic Ridge. Marine Biology Research, 13(5):486–493, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2016.1232830.

Lindsay:2017:POB

[963] Dhugal John Lindsay, Mary Matilda Grossmann, Bastian Bentlage, Allen Gilbert Collins, Ryo Minemizu, Russell Ross Hopcroft, Hiroshi Miyake, Mitsuko Hidaka-Umetsu, and Jun Nishikawa. The perils of online biogeographic databases: a case study with the ‘monospecific’ genus Aegina (Cnidaria, Hydrozoa, Narcomedusae). Marine Biology Research, 13(5):494–512, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1268261.

Pica:2017:INS

[964] Daniela Pica and Stefania Puce. Investigation of nematocysts in stylasterids (Cnidaria: Hydrozoa: Stylasteridae). Marine Biology Research, 13(5):513–520, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1239020.

Marfenin:2017:FMH

[965] Nikolay N. Marfenin and Vitaly S. Dementyev. Functional morphology of hydrozoan stolons: stolonal growth, contractility and hydroplasmic movement in Gonothyraea loveni (Allman, 1859). Marine Biology Research, 13(5):521–537, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1276292.

Betti:2017:LTL

[966] Federico Betti, Giorgio Bavestrello, Carlo Nike Bianchi, Carla Morri, Elena Righetti, Simone Bava, and Marzia Bo. Long-term life cycle and massive blooms of the intertidal hydroid Paracoryne huvei in the North-western Mediterranean Sea. Marine Biology Research, 13(5): 538–550, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1240874.

Martell:2017:PDN

[967] Luis Martell, Enrico Ciavolino, Cinzia Gravili, Stefano Piraino, and Ferdinando Boero. Population dynamics of the non-indigenous hydrozoan Clytia hummelincki (Hydrozoa: Campanulariidae) in two contrasting Mediterranean habitats. Marine Biology Research, 13(5): 551–559, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1296160.

Bettim:2017:SRH

[968] Ariane Lima Bettim and Maria Angélica Haddad. Seasonal recruitment of the hydroid Podocoryna loyola (Hydractiniidae) in the Paranaguá Bay, South of Brazil. Marine Biology Research, 13(5):560–572, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1307990.

Morri:2017:GCC

[969] Carla Morri, Carlo Nike Bianchi, Cristina Gioia Di Camillo, Frédéric Ducarme, William R. Allison, and Giorgio Bavestrello. Global climate change and regional biotic responses: two hydrozoan tales. Marine Biology Research, 13(5):573–586, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1283419.

Gravili:2017:HRD

[970] Cinzia Gravili, Francesco Cozzoli, and Ferdinando Boero. The historical reconstruction of distribution of the genus Halecium (Hydrozoa: Haleciidae): a biological signal of ocean warming? Marine Biology Research, 13(5):587–601, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1290805.

Yorio:2017:RFF

[971] Pablo Yorio, Diego González-Zevallos, Alejandro Gatto, Oscar Biagioni, and Joanna Castillo. Relevance of forage fish in the diet of Magellanic penguins breeding in northern Patagonia, Argentina. Marine Biology Research, 13(6):603–617, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2016.1273529.

Checon:2017:FCC

[972] Helio Herminio Checon and Tânia Marcia Costa. Fiddler crab (Crustacea: Ocypodidae) distribution and the relationship between habitat occupancy and mouth appendages. Marine Biology Research, 13(6): 618–629, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2016.1273530.

Briz:2017:GZC

[973] Luciana Diaz Briz, Felisa Sánchez, Noemí Marí, Hermes Mianzan, and Gabriel Genzano. Gelatinous zooplankton (ctenophores, salps and medusae): an important food resource of fishes in the temperate SW Atlantic Ocean. Marine Biology Research, 13(6):630–644, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1274403.

He:2017:LSM

[974] Jian He, Jian Fei Qi, Yan Qiu Huang, Yan Qing Sheng, Pei Su, Dan Qing Feng, and Cai Huan Ke. Larval settlement and metamorphosis of the invasive biofouler, Mytilopsis sallei, in response to ions and neuroactive compounds. Marine Biology Research, 13(6):645–658, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2016.1274826.

Rodriguez:2017:ELC

[975] Carolina S. Rodriguez, Antonio C. Marques, Hermes W. Mianzan, Valquiria B. Tronolone, Alvaro E. Migotto, and Gabriel N. Genzano. Environment and life cycles influence distribution patterns of hydromedusae in austral South America. Marine Biology Research, 13(6): 659–670, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1280170.

Nogueira:2017:PAF

[976] João Miguel de Matos Nogueira, Kirk Fitzhugh, Pat Hutchings, and Orlemir Carrerette. Phylogenetic analysis of the family Telothelepodidae Nogueira, Fitzhugh and Hutchings, 2013 (Annelida: Terebelliformia). Marine Biology Research, 13(6):671–692, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1283526.

Barioto:2017:EDP

[977] João G. Barioto, Gilson Stanski, Raphael C. Grabowski, Rogerio C. Costa, and Antonio L. Castilho. Ecological distribution of Penaeus schmitti (Dendrobranchiata: Penaeidae) juveniles and adults on the southern coast of São Paulo state, Brazil. Marine Biology Research, 13(6):693–703, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1287923.

Wallace:2017:MFF

[978] Elizabeth M. Wallace. Marine fishes of Florida, by David b. Snyder and George h. Burgess. Marine Biology Research, 13(6):704–705, 2017. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2017.1291979.

Anonymous:2017:E

[979] Anonymous. Erratum. Marine Biology Research, 13(6):??, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1353325.

Hunt:2017:SPR

[980] Heather L. Hunt, Richard A. Wahle, John Tremblay, Michel Comeau, Angelica Silva, and Rémy Rochette. Spatial patterns of richness and abundance of benthic decapod crustaceans and fishes in the North-west Atlantic as measured by cobble-filled bio-collectors. Marine Biology Research, 13(7):707–725, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1296161.

Phillips:2017:VEP

[981] Nicole E. Phillips. Variability in early post-settlement mortality in intertidal mussels and the role of size at settlement. Marine Biology Research, 13(7):726–732, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1298806.

MacTavish:2017:SSC

[982] Andrea Liévana MacTavish and Lydia Betty Ladah. Supply-side control over the survivorship of Chthamalus spp. recruits in northern Baja California. Marine Biology Research, 13(7):733–743, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2017.1307410.

Siikavuopio:2017:FOT

[983] Sten Ivar Siikavuopio, Ragnhild Dragøy Whitaker, Bjørn Steinar Sæther, Philip James, Bjørn Ronald Olsen, Tina Thesslund, Annette Hustad, and Atle Mortensen. First observations of temperature tolerances of adult male snow crab (Chionoecetes opilio) from the Barents Sea population and the effects on the fisheries strategy. Marine Biology Research, 13 (7):744–750, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1313989.

Chuang:2017:RIN

[984] Po-Shun Chuang, Hung-An Chang, Wei-Chuan Chiang, and Jen-Chieh Shiao. Rapid identification of the north-western Pacific billfish species using quantitative real-time polymerase chain reaction techniques. Marine Biology Research, 13(7):751–763, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1296162.

Carvalho:2017:MCP

[985] Margarida Gama Carvalho, Claúdia Moreira, Joana F. M. F. Cardoso, Geert-Jan A. Brummer, Piet van Gaever, Henk W. van der Veer, Henrique Queiroga, Paulo Talhadas Santos, and Alberto Teodorico Correia. Movement, connectivity and population structure of the intertidal fish Lipophrys pholis as revealed by otolith oxygen and carbon stable isotopes. Marine Biology Research, 13(7):764–773, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1306079.

Primost:2017:HBC

[986] Mónica A. Primost, Mónica N. Gil, and Gregorio Bigatti. High bioaccumulation of cadmium and other metals in Patagonian edible gastropods. Marine Biology Research, 13(7):774–781, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1296163.

Davila-Jimenez:2017:SFD

[987] Yasmín Dávila-Jiménez, María Ana Tovar-Hernández, and Nuno Simoes. The social feather duster worm Bispira brunnea (Polychaeta: Sabellidae): aggregations, morphology and reproduction. Marine Biology Research, 13(7):782–796, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1280608.

Mincarone:2017:RHM

[988] Michael Maia Mincarone. Redescription of the hagfishes (Myxinidae: Eptatretus) from southern Africa. Marine Biology Research, 13(7): 797–810, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1287924.

Bjorkan:2017:BRSa

[989] Maiken Bjørkan. Book review: Science, information, and policy interface: for effective coastal and ocean management, edited by Bertrum H. MacDonald, Suzuette S. Soomai, Elizabeth M. De Santo and Peter G. Wells. Marine Biology Research, 13(7):811–812, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1293271.

Bjorkan:2017:BRSb

[990] Maiken Bjørkan. Book review: Science, information, and policy interface: for effective coastal and ocean management, edited by Bertrum H. MacDonald, Suzuette S. Soomai, Elizabeth M. De Santo and Peter G. Wells. Marine Biology Research, 13(7):811–812, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1293271.

Bauer:2017:MFE

[991] Arthur de Barros Bauer, Luciano Gomes Fischer, Fabio Di Dario, and Michael Maia Mincarone. Marine fishes (Elasmobranchii and Teleostei) from the Santana Archipelago, a Marine Protected Area in the southwestern Atlantic. Marine Biology Research, 13(8):813–831, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1302090.

Cepeda:2017:NRD

[992] Diego Cepeda and Patricia Lattig. New reports and description of a new species of Sabellidae (Annelida) for the Iberian Peninsula and Balearic Archipelago. Marine Biology Research, 13(8):832–853, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1303576.

Pilczynska:2017:HGD

[993] Joanna Pilczynska, Silvia Cocito, Joana Boavida, Ester A. Serrão, and Henrique Queiroga. High genetic differentiation of red gorgonian populations from the Atlantic Ocean and the Mediterranean Sea. Marine Biology Research, 13(8):854–861, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1312005.

Voight:2017:TSI

[994] Janet R. Voight and Jessica A. Kurth. Towards species identification of the bathyal octopod Graneledone (Mollusca: Cephalopoda). Marine Biology Research, 13(8):862–869, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1303166.

Bueno:2017:ISH

[995] Marilia Bueno, Gustavo M. Dias, and Fosca P. P. Leite. The importance of shore height and host identity for amphipod assemblages. Marine Biology Research, 13(8):870–877, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1306650.

Afiq-Rosli:2017:ESN

[996] Lutfi Afiq-Rosli, Daisuke Taira, Hai Xin Loke, Tai Chong Toh, Kok Ben Toh, Chin Soon Lionel Ng, Patrick Caranzo Cabaitan, Loke Ming Chou, and Tiancheng Song. In situ nurseries enhance coral transplant growth in sedimented waters. Marine Biology Research, 13(8):878–887, 2017. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2017.1307988.

Velazquez:2017:PDC

[997] Samira Nuñez Velazquez, Carlos Enrique Rumbold, and Sandra Mabel Obenat. Population dynamics of Caprella dilatata and Caprella equilibra (Peracarida: Amphipoda) in a Southwestern Atlantic harbour. Marine Biology Research, 13(8):888–898, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1317101.

DelCarmenGarcia-Rivas:2017:WCL

[998] María Del Carmen García-Rivas, Salima Machkour-M’Rabet, Gabriela Pérez-Lachaud, Juan J. Schmitter-Soto, Cassiopea Doneys, Nikolas St-Jean, Dorka Cobián, and Yann Hénaut. What are the characteristics of lionfish and other fishes that influence their association in diurnal refuges? Marine Biology Research, 13(8):899–908, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1314496.

Brewer:2017:TSD

[999] Reid S. Brewer, Brenda L. Norcross, and Ellen Chenoweth. Temperatureand size-dependent growth and movement of the North Pacific giant octopus (Enteroctopus dofleini) in the Bering Sea. Marine Biology Research, 13(8):909–918, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1309436.

Haug:2017:TLF

[1000] Tore Haug, Stig Falk-Petersen, Michael Greenacre, Haakon Hop, Ulf Lindstrøm, Sonnich Meier, Kjell T. Nilssen, and Anette Wold. Trophic level and fatty acids in harp seals compared with common minke whales in the Barents Sea. Marine Biology Research, 13(9):919–932, 2017. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2017.1313988.

Neves:2017:HIO

[1001] Vanessa Neves, Alexandra Guedes, Bruno Valentim, Joana Campos, and Vânia Freitas. High incidence of otolith abnormality in juvenile European flounder Platichthys flesus from a tidal freshwater area. Marine Biology Research, 13(9):933–941, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1331041.

Hernandez-Alcantara:2017:DMC

[1002] Pablo Hernández-Alcántara, Araceli J. Mercado-Santiago, and Vivianne Solís-Weiss. Description, morphometric characterization and distribution pattern of Paradiopatra multibranchiata, a new onuphid (Annelida: Polychaeta) from the tropical eastern Pacific. Marine Biology Research, 13(9):942–954, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1333619.

Baranano:2017:RZM

[1003] Carlota Barañano, Emilio Fernández, Gonzalo Méndez, and Jesús S. Troncoso. Resilience of Zostera marina habitats and response of the macroinvertebrate community to physical disturbance caused by clam harvesting. Marine Biology Research, 13(9):955–966, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1307989.

Smulders:2017:EFS

[1004] Fee O. H. Smulders, J. Arie Vonk, M. Sabine Engel, and Marjolijn J. A. Christianen. Expansion and fragment settlement of the non-native seagrass Halophila stipulacea in a Caribbean bay. Marine Biology Research, 13(9):967–974, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1333620.

Poquita-Du:2017:ENT

[1005] Rosa Celia Poquita-Du, Kok Ben Toh, Tai Chong Toh, Chin Soon Lionel Ng, Daisuke Taira, Hai Xin Loke, Lutfi Afiq-Rosli, Loke Ming Chou, Tiancheng Song, and Patrick Cabaitan. Effects of nursery table slope orientation on coral survival and growth. Marine Biology Research, 13 (9):975–982, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1322703.

Maggioni:2017:DTM

[1006] Davide Maggioni, Stefania Puce, Paolo Galli, Davide Seveso, and Simone Montano. Description of Turritopsoides marhei sp. nov. (Hydrozoa, Anthoathecata) from the Maldives and its phylogenetic position. Marine Biology Research, 13(9):983–992, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1317813.

Xiong:2017:ELH

[1007] Ying Xiong, Jian Yang, Tao Jiang, Hongbo Liu, Xiaming Zhong, and Jianhua Tang. Early life history of the small yellow croaker (Larimichthys polyactis) in sandy ridges of the South Yellow Sea. Marine Biology Research, 13(9):993–1002, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1319067.

Quattrocchi:2017:EDS

[1008] Federico Quattrocchi and Francesc Maynou. Environmental drivers of sardine (Sardina pilchardus) in the Catalan Sea (NW Mediterranean Sea). Marine Biology Research, 13(9):1003–1014, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1331039.

Diaz-Astudillo:2017:OLF

[1009] Macarena Díaz-Astudillo, Manuel I. Castillo, Mario A. Cáceres, Guido Plaza, and Mauricio F. Landaeta. Oceanographic and lunar forcing affects nearshore larval fish assemblages from temperate rocky reefs. Marine Biology Research, 13(10):1015–1026, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1335872.

Buhl-Mortensen:2017:CRS

[1010] Pål Buhl-Mortensen. Coral reefs in the Southern Barents Sea: habitat description and the effects of bottom fishing. Marine Biology Research, 13 (10):1027–1040, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1331040.

Carrasco:2017:SST

[1011] Sergio A. Carrasco, Loes Vandecasteele, Marcelo M. Rivadeneira, Miriam Fernández, and Alejandro Pérez-Matus. Spatial and short-term variability of larval, post-larval and macrobenthic assemblages associated with subtidal kelp forest ecosystems in Central Chile. Marine Biology Research, 13(10):1041–1058, 2017. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1322704.

Colgan:2017:FSS

[1012] Donald J. Colgan. Fine-scale spatial partitioning of genetic variation and evolutionary contestability in the invasive estuarine mussel Xenostrobus securis. Marine Biology Research, 13(10):1059–1072, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1331361.

Oliveira-Souza:2017:PST

[1013] Wilson Oliveira-Souza and Helena Passeri Lavrado. Population structure and temporal variation of the roughneck shrimp Rimapenaeus constrictus (Decapoda: Penaeoidea) in a coastal bay of the Southwestern Atlantic. Marine Biology Research, 13(10):1073–1083, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1342846.

Consoli:2017:IFV

[1014] Pierpaolo Consoli, Valentina Esposito, Manuela Falautano, Pietro Battaglia, Luca Castriota, Teresa Romeo, Mauro Sinopoli, Pietro Vivona, and Franco Andaloro. The impact of fisheries on vulnerable habitats: the case of trawling on circa-littoral grounds in the Strait of Sicily (central Mediterranean Sea). Marine Biology Research, 13(10): 1084–1094, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1348010.

daSilveira:2017:MPG

[1015] Savênia Bonoto da Silveira, Wilson Wasielesky, Ana Paula Dini Andreote, Marli Fatima Fiore, and Clarisse Odebrecht. Morphology, phylogeny, growth rate and nodularin production of Nodularia spumigena from Brazil. Marine Biology Research, 13(10):1095–1107, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1336587.

Magalhaes:2017:LSC

[1016] Tatiana Magalhães, Edvanda A. Souza-Carvalho, Renata Biagi, José A. Cuesta, and Fernando L. Mantelatto. Larval stages of the crab Mithrax tortugae (Brachyura: Mithracidae) with comparisons between all species of Mithrax. Marine Biology Research, 13(10):1108–1117, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1342035.

Coelho:2017:BRRa

[1017] Rui Coelho. Book review: Rays of the world, edited by Peter R. Last, William T. White, Marcello R. de Carvalho, Bernard Séret, Matthias F. W. Stehmann and Gavin J. P. Naylor. Marine Biology Research, 13(10):1118, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1336246.

Coelho:2017:BRRb

[1018] Rui Coelho. Book review: Rays of the world, edited by Peter R. Last, William T. White, Marcello R. de Carvalho, Bernard Séret, Matthias F. W. Stehmann and Gavin J. P. Naylor. Marine Biology Research, 13(10):1118, 2017. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1336246.

Urkmez:2018:DMS

[1019] Derya Ürkmez, Marta Ostrowska, Milena Roszkowska, Magdalena Gawlak, Krzysztof Zawierucha, Reinhardt Møbjerg Kristensen, and Lukasz Kaczmarek. Description of Megastygarctides sezginii sp. nov. (Tardigrada: Arthrotardigrada: Stygarctidae) from the Turkish Black Sea coast and a key to the genus Megastygarctides. Marine Biology Research, 14(1):1–16, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2017.1342845.

Barros-Garcia:2018:NIS

[1020] David Barros-García, Rafael Bañón, Juan Carlos Arronte, Lourdes Fernández-Peralta, Ramón García, Samuel Paco Iglésias, Daniel Y. Sellos, João Pedro Barreiros, Ángel Sebastián Comesaña, and Alejandro De Carlos. New insights into the systematics of North Atlantic Gaidropsarus (Gadiformes, Gadidae): flagging synonymies and hidden diversity. Marine Biology Research, 14(1):17–29, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1367403.

Grinyo:2018:MFG

[1021] Jordi Grinyó, Andrea Gori, Pablo J. López-González, Andreu Santín, Patricia Baena, and Josep-Maria Gili. Morphological features of the gorgonian Paramuricea macrospina on the continental shelf and shelf edge (Menorca Channel, Western Mediterranean Sea). Marine Biology Research, 14(1):30–40, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1375118.

Mercado-Molina:2018:GFO

[1022] Alex E. Mercado-Molina, Fabiola Rivera-Irizarry, Jaime Fonseca-Miranda, and Yesenia Bruno-Laureano. Growth facilitation by the octocoral Gorgonia ventalina explains spatial difference in the population size structure of the common demosponge Ircinia felix. Marine Biology Research, 14(1):41–50, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1367098.

Gerasimova:2018:LTC

[1023] Alexandra V. Gerasimova, Ekaterina V. Ushanova, Andrew A. Filippov, Nadezhda A. Filippova, Igor A. Stogov, and Nikolay V. Maximovich. Long-term changes in cohort structure of the soft-shell clam Mya arenaria in the White Sea: growth rate affects lifespan and mortality. Marine Biology Research, 14(1):51–64, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1346259.

Poza:2018:TDA

[1024] Ailen M. Poza, M. Cecilia Gauna, Juan F. Escobar, and Elisa R. Parodi. Temporal dynamics of algal epiphytes on Leathesia marina and Colpomenia sinuosa macrothalli (Phaeophyceae). Marine Biology Research, 14(1):65–75, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2017.1348606.

Christiansen:2018:HTV

[1025] Bernd Christiansen and Horst Weikert. High temporal variability of zooplankton, but only weak evidence for a near-bottom effect on composition and distribution in the deep Levantine Basin, eastern Mediterranean. Marine Biology Research, 14(1):76–86, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1350712.

duPlooy:2018:HOS

[1026] Schalk J. du Plooy, Renzo Perissinotto, and David G. Muir. Hypoosmotic shock and the role of freshwater input on the post-bloom recovery of Cyanothece sp. in Lake St Lucia, South Africa. Marine Biology Research, 14(1):87–94, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1350713.

Asanopoulos:2018:ESV

[1027] Christina H. Asanopoulos, Adrian J. Linnane, Matthew Hoare, and Charlie Huveneers. In situ video monitoring of by-catch interactions within commercial rock lobster (Jasus edwardsii) fishing traps. Marine Biology Research, 14(1):95–104, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1348605.

Yamaguchi:2018:AHS

[1028] Tadanori Yamaguchi, Yuji Kawakami, and Michiya Matsuyama. Analysis of the hatching site and migratory behaviour of the swordtip squid (Uroteuthis edulis) caught in the Japan Sea and Tsushima Strait in autumn estimated by statolith analysis. Marine Biology Research, 14 (1):105–112, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1351616.

Grzelak:2018:NSE

[1029] Katarzyna Grzelak and Martin V. Sørensen. New species of Echinoderes (Kinorhyncha: Cyclorhagida) from Spitsbergen, with additional information about known Arctic species. Marine Biology Research, 14 (2):113–147, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1367096.

Zabala:2018:EDL

[1030] Soledad Zabala, Andrés Averbuj, Pablo E. Penchaszadeh, and Gregorio Bigatti. Early development of the limpet Siphonaria lessonii Blainville, 1827 in populations affected by different physical stressors. Marine Biology Research, 14(2):148–157, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1396343.

Salvadori:2018:XXS

[1031] Susanna Salvadori, Anna Maria Deiana, Federica Deidda, Cinzia Lobina, Antonello Mulas, and Elisabetta Coluccia. XX/XY sex chromosome system and chromosome markers in the snake eel Ophisurus serpens (Anguilliformes: Ophichtidae). Marine Biology Research, 14(2): 158–164, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1406665.

Yoo:2018:MSD

[1032] Yeong Du Yoo, Kyeong Ah Seong, Jae Seong Kim, Seung Won Nam, Hae Jin Jeong, Jung-Rae Rho, Wonho Yih, and Hyung Seop Kim. Mixotrophy in the sand-dwelling dinoflagellate Thecadinium kofoidii. Marine Biology Research, 14(2):165–172, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1379604.

Klarian:2018:NIT

[1033] Sebastian A. Klarian, Carlos Canales-Cerro, Patricio Barría, Patricia Zárate, Francisco Concha, Sebastián Hernández, Maike Heidemeyer, Pauline Sallaberry-Pincheira, and Roberto Meléndez. New insights on the trophic ecology of blue (Prionace glauca) and shortfin mako sharks (Isurus oxyrinchus) from the oceanic eastern South Pacific. Marine Biology Research, 14(2):173–182, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1396344.

Dobrzycka-Krahel:2018:ESD

[1034] Aldona Dobrzycka-Krahel and Bozena Graca. Effect of salinity on the distribution of Ponto-Caspian gammarids in a non-native area — environmental and experimental study. Marine Biology Research, 14(2): 183–190, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1406666.

Osorio-Zuniga:2018:STV

[1035] Francisca Osorio-Zúñiga, Mauricio F. Landaeta, Jorge Angulo-Aros, and Fernando Balbontín. Spatio-temporal variability of ichthyoplankton and its relationship with oceanographic conditions at the shelf break off Chilean Patagonia (43^{o}S–51^{o}S). Marine Biology Research, 14(2): 191–202, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1406663.

Yaragina:2018:EPM

[1036] Natalia A. Yaragina, Yury Kovalev, and Anatoly Chetyrkin. Extrapolating predation mortalities back in time: an example from North-East Arctic cod cannibalism. Marine Biology Research, 14(2):203–216, 2018. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2017.1396342.

Nogueira:2018:SRS

[1037] João Miguel de Matos Nogueira, Orlemir Carrerette, Pat Hutchings, and Kirk Fitzhugh. Systematic review of the species of the family telothelepodidae nogueira, Fitzhugh and Hutchings, 2013 (Annelida, Terebelliformia), with descriptions of three new species. Marine Biology Research, 14(3):217–257, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1401729.

Collin:2018:SDE

[1038] Rachel Collin, Nerea Nieto, and Cynthia Peña. Seasonal differences in egg size in three species of crabs from a tropical upwelling zone. Marine Biology Research, 14(3):258–268, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2017.1406664.

Mourato:2018:RBS

[1039] Bruno L. Mourato, Mariela Narvaez, Alberto F. de Amorim, Humberto Hazin, Felipe Carvalho, Fábio Hazin, and Freddy Arocha. Reproductive biology and space-time modelling of spawning for sailfish Istiophorus platypterus in the western Atlantic Ocean. Marine Biology Research, 14 (3):269–286, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1407873.

vanOevelen:2018:TSC

[1040] Dick van Oevelen, Gerard C. A. Duineveld, Marc S. S. Lavaleye, Tina Kutti, and Karline Soetaert. Trophic structure of cold-water coral communities revealed from the analysis of tissue isotopes and fatty acid composition. Marine Biology Research, 14(3):287–306, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2017.1398404.

Zunino:2018:EMS

[1041] Serena Zunino, Valentina Pitacco, Borut Mavric, Martina Orlando-Bonaca, Petar Kruzić, and Lovrenc Lipej. The ecology of the Mediterranean stony coral Cladocora caespitosa (Linnaeus, 1767) in the Gulf of Trieste (northern Adriatic Sea): a 30-year long story. Marine Biology Research, 14(3):307–320, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2017.1408915.

Uozumi:2018:MFM

[1042] Ryosuke Uozumi, Hiroshi Yamasaki, and Euichi Hirose. Mangrove forests may serve as stable environments for the meiobenthic Echinoderes komatsui (Kinorhyncha: Cyclorhagida): distribution patterns and population dynamics in a subtropical estuary. Marine Biology Research, 14 (3):321–333, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1408916.

Xu:2018:DLA

[1043] Ran Xu, Qikang Bo, and Xiaodong Zheng. A divergent lineage among Octopus minor (Sasaki, 1920) populations in the Northwest Pacific supported by DNA barcoding. Marine Biology Research, 14(4): 335–344, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1427866.

Gao:2018:CSS

[1044] Dongkui Gao, Xiumei Zhang, Wentao Li, and Peidong Zhang. Comparative study on the surface ultrastructures of fertilized eggs of three Sebastes species. Marine Biology Research, 14(4):345–353, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2018.1425457.

Bokenhans:2018:AGD

[1045] V. Bökenhans, G. Bigatti, M. García Asorey, and A. Averbuj. Age and growth differences in two populations of the edible marine gastropod Buccinanops globulosus. Marine Biology Research, 14(4):354–365, 2018. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2017.1406667.

Murillo:2018:MGS

[1046] F. J. Murillo, B. W. MacDonald, E. Kenchington, S. E. Campana, B. Sainte-Marie, and M. Sacau. Morphometry and growth of sea pen species from dense habitats in the Gulf of St. Lawrence, eastern Canada. Marine Biology Research, 14(4):366–382, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1417604.

Rodriguez-Barreras:2018:DIP

[1047] Ruber Rodríguez-Barreras. Demographic implications of predatory wrasses on low-density Diadema antillarum populations. Marine Biology Research, 14(4):383–391, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1426861.

Xu:2018:ESM

[1048] Qiang Xu, Pei-Dong Zhang, Yan-Shan Liu, Xi-Tao Wang, and Wen-Tao Li. The effect of substrate media on the survival and growth of the eelgrass Zostera marina. Marine Biology Research, 14(4): 392–402, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2017.1417602.

Adami:2018:IMB

[1049] Mariana Adami, Evangelina Schwindt, Alejandro Tablado, Javier Calcagno, Juan Carlos Labraga, and Lobo Orensanz. Intertidal mussel beds from the South-western Atlantic show simple structure and uniform appearance: does environmental harshness explain the community? Marine Biology Research, 14(4):403–419, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2017.1417603.

Paulino:2018:SAV

[1050] Ana Isabell Paulino, Aud Larsen, Gunnar Bratbak, Dale Evens, Svein Rune Erga, Einar Bye-Ingebrigtsen, and Jorun Karin Egge. Seasonal and annual variability in the phytoplankton community of the Raunefjord, west coast of Norway from 2001–2006. Marine Biology Research, 14(5):421–435, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1426863.

deLecea:2018:ZAI

[1051] Ander M. de Lecea, Ross Coppin, Margaux Noyon, and Jenny A. Huggett. Zooplankton adrift: investigating transportation by cyclonic eddy. Marine Biology Research, 14(5):436–447, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2018.1426862.

Naik:2018:SDS

[1052] Sangeeta Mahableshwar Naik and Arga Chandrashekar Anil. Survival in the dark: strategies adopted by Tetraselmis indica (Chlorodendrophyceae, Chlorophyta). Marine Biology Research, 14(5):448– 453, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1464188.

Dong:2018:EST

[1053] Zhijun Dong, Lei Wang, Qingqing Liu, and Tingting Sun. Effects of salinity and temperature on the recruitment of Aurelia coerulea planulae. Marine Biology Research, 14(5):454–461, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2018.1459725.

Manriquez:2018:EIA

[1054] Patricio H. Manríquez, Ricardo Guiñez, Alberto Olivares, Marcela Clarke, and Juan Carlos Castilla. Effects of inter-annual temperature variability, including ENSO and post-ENSO events, on reproductive traits in the tunicate Pyura praeputialis. Marine Biology Research, 14 (5):462–477, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1425456.

Fang:2018:SPA

[1055] Jinghui Fang, Zengjie Jiang, Jianguang Fang, Bin Kang, Yaping Gao, and Meirong Du. Selectivity of Perinereis aibuhitensis (Polychaeta, Nereididae) feeding on sediment. Marine Biology Research, 14(5): 478–483, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1426864.

Kurata:2018:SCD

[1056] Kengo Kurata and Jun’ichi Hiratsuka. Spatiotemporal changes in distributions of the two dominant bivalve species in an estuarine river connecting two brackish lakes, Japan. Marine Biology Research, 14 (5):484–496, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1428754.

Goedknegt:2018:IIP

[1057] M. Anouk Goedknegt, Sarah Bedolfe, Jan Drent, Jaap van der Meer, and David W. Thieltges. Impact of the invasive parasitic copepod Mytilicola orientalis on native blue mussels Mytilus edulis in the western European Wadden Sea. Marine Biology Research, 14(5):497–507, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2018.1442579.

Xue:2018:UNF

[1058] Ying Xue, Kisei Tanaka, Huaming Yu, Yong Chen, Lisha Guan, Zengguang Li, Haiqing Yu, Binduo Xu, Yiping Ren, and Rong Wan. Using a new framework of two-phase generalized additive models to incorporate prey abundance in spatial distribution models of juvenile slender lizardfish in Haizhou Bay, China. Marine Biology Research, 14 (5):508–523, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1447673.

Medeiros:2018:DPF

[1059] Aline Paiva Morais Medeiros, Josias Henrique de Amorim Xavier, Marianna Barbosa da Silva, Luana Aires-Souza, and Ierecê Maria de Lucena Rosa. Distribution patterns of the fish assemblage in the Mamanguape River Estuary, North-eastern Brazil. Marine Biology Research, 14 (5):524–536, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1459724.

Polonia:2018:CAA

[1060] Ana Rita Moura Polónia, Daniel Francis Richard Cleary, Francisco José Riso da Costa Coelho, Leontine E. Becking, Nicole Joy de Voogd, Abdul Hamid A. Toha, and Newton Carlos Marcial Gomes. Compositional analysis of archaeal communities in high and low microbial abundance sponges in the misool coral reef system, Indonesia. Marine Biology Research, 14(6):537–550, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1498977.

Bitner:2018:RBC

[1061] Maria Aleksandra Bitner and Marco Romanin. Recent brachiopods collected during the ZhongSha 2015 expedition to the South China Sea, West Pacific. Marine Biology Research, 14(6):551–564, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2018.1502447.

Mandai:2018:PPI

[1062] Silvia Sayuri Mandai, Raquel Corrêa Buranelli, Christoph D. Schubart, and Fernando L. Mantelatto. Phylogenetic and phylogeographic inferences based on two DNA markers reveal geographic structure of the orangeclaw hermit crab Calcinus tibicen (Anomura: Diogenidae) in the western Atlantic. Marine Biology Research, 14(6):565–580, 2018. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2018.1497184.

Avila:2018:ABA

[1063] Enrique Ávila, Osmar Roberto Araujo-Leyva, María Amparo Rodríguez-Santiago, and Hugo López-Rosas. Alien barnacle Amphibalanus amphitrite epizoic on two native oyster species in the southern Gulf of Mexico: spatio-temporal variability and current status of its epibiosis. Marine Biology Research, 14(6):581–589, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2018.1459723.

Christianen:2018:RNF

[1064] M. J. A. Christianen, W. Lengkeek, J. H. Bergsma, J. W. P. Coolen, K. Didderen, M. Dorenbosch, F. M. F. Driessen, P. Kamermans, E. Reuchlin-Hugenholtz, H. Sas, A. Smaal, K. A. van den Wijngaard, and T. M. van der Have. Return of the native facilitated by the invasive? Population composition, substrate preferences and epibenthic species richness of a recently discovered shellfish reef with native European flat oysters (Ostrea edulis) in the North Sea. Marine Biology Research, 14(6):590–597, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1498520.

Hussy:2018:FSG

[1065] Karin Hüssy, Margit Eero, and Krzysztof Radtke. Faster or slower: has growth of eastern Baltic cod changed? Marine Biology Research, 14 (6):598–609, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1502446.

Pescinelli:2018:PSR

[1066] Régis Augusto Pescinelli, Alexandre Oliveira Almeida, and Rogério Caetano Costa. Population structure, relative growth and morphological sexual maturity of the snapping shrimp Alpheus brasileiro anker, 2012 (Caridea: Alpheidae) from the south-eastern coast of Brazil. Marine Biology Research, 14(6):610–620, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2018.1472383.

Mayor:2018:MSD

[1067] Ejiroghene D. Mayor and Paulinus Chigbu. Mysid shrimp dynamics in relation to abiotic and biotic factors in the coastal lagoons of Maryland, Mid-West Atlantic, USA. Marine Biology Research, 14(6): 621–636, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1472384.

Sakaji:2018:ISE

[1068] Hideo Sakaji, Kazuhisa Hamada, and Yasuhiko Naito. Identifying spawning events of greater amberjack using accelerometers. Marine Biology Research, 14(6):637–641, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1492140.

Ortega:2018:RIP

[1069] Ileana Ortega, Milena Bacaneli, Leonir André Colling, and Luiz Felipe Cestari Dumont. Responses of the infaunal peracarid Monokalliapseudes schubarti (Tanaidacea: Kalliapseudidae) to a bottom trawling disturbance and seasonality in an estuary of southern Brazil. Marine Biology Research, 14(6):642–653, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2018.1478104.

Liversage:2018:UCU

[1070] K. Liversage and J. Kotta. Unveiling commonalities in understudied habitats of boulder-reefs: life-history traits of the widespread invertebrate and algal inhabitants. Marine Biology Research, 14(7): 655–671, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1510180.

Ringvold:2018:UCS

[1071] Halldis Ringvold and Morag Taite. Using citizen science to obtain data on large, floating gelatinous spheres from NE Atlantic, attributed to egg mass of ommastrephid squid (Oegopsida, Cephalopoda, Mollusca). Marine Biology Research, 14(7):672–681, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2018.1504165.

Gastaldi:2018:MSP

[1072] Marianela Gastaldi, Thiago S. De Paula, Maite A. Narvarte, Gisele Lôbo-Hajdu, and Eduardo Hajdu. Marine sponges (Porifera) from the Bahía San Antonio (North Patagonian Gulfs, Argentina), with additions to the phylogeography of the widely distributed Cliona aff. celata and Hymeniacidon perlevis, and the description of two new species. Marine Biology Research, 14(7):682–716, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2018.1506136.

Vermeersch:2018:EEO

[1073] Xavier H. C. Vermeersch, Koen Chiers, Annemie Decostere, Stephanie Van Weyenberg, and Geertrui Vlaemynck. Evaluation of the effects of onboard sorting and experimental holding practices on the post-catch survival of brown shrimp (Crangon crangon L.). Marine Biology Research, 14(7):717–727, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1497183.

daCosta:2018:AEG

[1074] Paulo Alberto Silva da Costa, Adriana da Costa Braga, Juliana Monteiro da Silva Vieira, Raquel Rennó M. Martins, Ricardo Raphael Bastos de São-Clemente, and Bernardo Roxo Couto. Age estimation, growth and maturity of the Argentine hake (Merluccius hubbsi Marini, 1933) along the northernmost limit of its distribution in the south-western Atlantic. Marine Biology Research, 14(7):728–738, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2018.1502885.

Briz:2018:SVG

[1075] Luciana Diaz Briz, Felisa Sánchez, Noemí Marí, and Gabriel Genzano. Seasonal variation of gelatinous plankton consumption by fish in the South-western Atlantic Ocean: a question of strategy. Marine Biology Research, 14(7):739–751, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1508847.

Espinasse:2018:IPV

[1076] Marina Espinasse, Claudia Halsband, Øystein Varpe, Astthor Gislason, Kristinn Gudmundsson, Stig Falk-Petersen, and Ketil Eiane. Interannual phenological variability in two North-East Atlantic populations of Calanus finmarchicus. Marine Biology Research, 14(7): 752–767, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1506135.

Cano-Sanchez:2018:ASP

[1077] Esperanza Cano-Sánchez and Pablo J. López-González. Ammothea species (Pycnogonida: Ammotheidae) collected during the Polarstern cruise ANT-XXIX/3 to Antarctic waters, with a description of a new species. Marine Biology Research, 14(8):769–777, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2018.1521524.

Gondim:2018:RPK

[1078] Anne Isabelley Gondim, Rafael Bendayan de Moura, Martin Lindsey Christoffersen, and Thelma Lúcia Pereira Dias. Rediscovery of the poorly known sea urchin Centrostephanus longispinus rubicingulus (H. L. Clark, 1921) (Echinodermata, Echinoidea, Diadematidae) on the continental shelf of Brazil, with notes on its morphology and synonymy. Marine Biology Research, 14(8):778–789, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2018.1529414.

Ba-Akdah:2018:TSS

[1079] Mohammad Abdulaziz Ba-Akdah, Sathianeson Satheesh, Abdul Mohsin A. Al-Sofyani, Yolanda Lucas, Patricia Álvarez-Campos, and Guillermo San Martín. Taxonomy of some species of the genus Syllis (Annelida: Syllidae: Syllinae) from the Red Sea found among the first colonizers of an artificial substrate. Marine Biology Research, 14(8):790–805, 2018. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2018.1531133.

Velasquez:2018:NGS

[1080] Marcel Velásquez and J. Reuben Shipway. A new genus and species of deep-sea wood-boring shipworm (Bivalvia: Teredinidae) Nivanteredo coronata n. sp. from the Southwest Pacific. Marine Biology Research, 14(8):806–815, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1544421.

Xiao:2018:PHE

[1081] Ning Xiao, Xiaomeng Li, and Zhongli Sha. Psychropotid holothurians (Echinodermata: Holothuroidea: Elasipodida) of the tropical Western Pacific collected by the KEXUE expedition with description of one new species. Marine Biology Research, 14(8):816–826, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2018.1546012.

Turicchia:2018:CKS

[1082] Eva Turicchia, Bert W. Hoeksema, and Massimo Ponti. The coralkilling sponge Chalinula nematifera as a common substrate generalist in Komodo national park, Indonesia. Marine Biology Research, 14 (8):827–833, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1544420.

deSouza:2018:INS

[1083] Adelson Silva de Souza, Darlan de Jesus de Brito Simith, and Fernando Araújo Abrunhosa. Impacts of nutritional stress on the survival and developmental time to metamorphosis in megalopa larvae of the exploited mangrove crab, Ucides cordatus (Ucididae). Marine Biology Research, 14(8):834–845, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1547406.

Korn:2018:PBR

[1084] Olga M. Korn, Elena S. Kornienko, and Nikolai I. Selin. Population biology and reproductive characteristics of the hermit crab Pagurus minutus Hess, 1865 (Decapoda: Anomura: Paguridae) in the northern part of the species range (Peter the Great Bay, the Sea of Japan). Marine Biology Research, 14(8):846–855, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2018.1503685.

Hernaez:2018:SSS

[1085] Patricio Hernáez and Marcio Camargo Araujo João. Social structure, sexual dimorphism and relative growth in the ghost shrimp Callichirus seilacheri (Bott, 1955) (Decapoda, Axiidea, Callianassidae) from the tropical eastern Pacific. Marine Biology Research, 14(8): 856–867, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1529415.

Sokolover:2018:SEB

[1086] Noga Sokolover, Andrew N. Ostrovsky, and Micha Ilan. Schizoporella errata (Bryozoa, Cheilostomata) in the Mediterranean Sea: abundance, growth rate, and reproductive strategy. Marine Biology Research, 14 (8):868–882, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1526385.

Aguilar-Perera:2018:RSI

[1087] A. Aguilar-Perera, J. Valle-Nava, E. H. Williams, and L. Bunkley-Williams. Rocinela signata (Isopoda: Aegidae) parasitizing Indo-Pacific lionfish Pterois volitans (Scorpaeniformes: Scorpaenidae) in the Southern Gulf of Mexico. Marine Biology Research, 14(8):883–886, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2018.1528012.

Lizcano-Sandoval:2018:GSP

[1088] Luis David Lizcano-Sandoval, Edgardo Londoño-Cruz, and Fernando A. Zapata. Growth and survival of Pocillopora damicornis (Scleractinia: Pocilloporidae) coral fragments and their potential for coral reef restoration in the Tropical Eastern Pacific. Marine Biology Research, 14(8): 887–897, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1528011.

VanSteenkiste:2018:SDE

[1089] Niels W. L. Van Steenkiste and Brian S. Leander. Species diversity of eukalyptorhynch flatworms (Platyhelminthes, Rhabdocoela) from the coastal margin of British Columbia: Polycystididae, Koinocystididae and Gnathorhynchidae. Marine Biology Research, 14(9–10): 899–923, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1575514.

Bahr:2018:OWD

[1090] Keisha D. Bahr, Ku‘ulei S. Rodgers, and Paul L. Jokiel. Ocean warming drives decline in coral metabolism while acidification highlights species-specific responses. Marine Biology Research, 14(9–10): 924–935, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1551616.

Rochin-Banaga:2018:PEM

[1091] Heriberto Rochín-Bañaga, Rafael Cervantes-Duarte, Mara Y. Cortés, Felipe García-Romero, and Alejandro Ramos-Rodríguez. Parmales ecology in Magdalena Bay, Mexican Pacific coast. Marine Biology Research, 14 (9–10):936–944, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1577458.

Sotelo:2018:MDP

[1092] Inés Sotelo, Gustavo Thompson, Marina Sabatini, Raúl Reta, and María Cristina Daponte. Mesoscale distribution and population structure of the chaetognath Serratosagitta tasmanica (Thomson, 1947) from Southwestern Atlantic Ocean. Marine Biology Research, 14(9–10): 945–960, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2018.1550257.

Correia:2018:DTO

[1093] Edna Correia, José Pedro Granadeiro, Aissa Regalla, Hamilton Monteiro, and Paulo Catry. The diet of two opportunistic seabirds (Caspian and Royal Terns) confirms the importance of Sardinella spp. in West African coastal food webs. Marine Biology Research, 14(9– 10):961–972, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1572193.

Adelir-Alves:2018:OTS

[1094] Johnatas Adelir-Alves, Felippe A. L. M. Daros, Henry L. Spach, Marcelo Soeth, and Alberto T. Correia. Otoliths as a tool to study reef fish population structure from coastal islands of South Brazil. Marine Biology Research, 14(9–10):973–988, 2018. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2019.1572194.

Wiegleb:2018:PRG

[1095] Joschka Wiegleb, Paul Kotterba, Cornelius Hammer, and Daniel Oesterwind. Predation of the round goby (Neogobius melanostomus Pallas, 1814) on Atlantic herring eggs in the Western Baltic Sea. Marine Biology Research, 14(9–10):989–1003, 2018. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2019.1577977.

Lou:2019:GSS

[1096] Yu Lou, Peng Yi, Xue Wen Song, and Hong Wei Wang. Grateloupia Serra sp. nov. H. W. Wang and Y. Lou (Halymeniaceae, Rhodophyta): a new species previously confused with Grateloupia asiatica in China. Marine Biology Research, 15(1):1–12, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2019.1593461.

Athanasiadis:2019:AGF

[1097] Athanasios Athanasiadis. Amphithallia, a genus with four-celled carpogonial branches and connecting filaments in the Corallinales (Rhodophyta). Marine Biology Research, 15(1):13–25, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2019.1598559.

Lee:2019:MGR

[1098] Taekjun Lee, Yeon Jae Bae, and Sook Shin. Mitochondrial gene rearrangement and phylogenetic relationships in the Amphilepidida and Ophiacanthida (Echinodermata, Ophiuroidea). Marine Biology Research, 15(1):26–35, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1601226.

Anikeeva:2019:SWM

[1099] Oksana V. Anikeeva, Andrew J. Gooday, and Alexey N. Petrov. Softwalled monothalamids (Rhizaria: foraminifera) of the Crimean shelf (Black Sea): taxonomic composition and inter-regional patterns of species diversity and distribution. Marine Biology Research, 15(1): 36–48, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1582782.

Wang:2019:EAD

[1100] Weimin Wang, Huan Zhang, Hongju Chen, Yunyun Zhuang, Yousong Huang, and Guangxing Liu. Effects of Asian dust input on eukaryotic phytoplankton community structure in the open areas in the Northwestern Pacific Ocean. Marine Biology Research, 15(1):49–60, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1588981.

Aragon:2019:ECC

[1101] Hansel Caballero Aragón, Maickel Armenteros, Susana Perera Valderrama, Néstor Rey Villiers, Dorka Cobián Rojas, Karla Campos Verdecia, and Pedro M. Alcolado Menéndez. Ecological condition of coral reef assemblages in the Cuban Archipelago. Marine Biology Research, 15(1):61–73, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1577557.

Leal:2019:AMS

[1102] Gustavo R. Leal, Guilherme T. Nunes, Gabriela Oliveira, and Leandro Bugoni. Assortative mating, sexual size dimorphism and sex determination in a seabird with plumage polymorphism. Marine Biology Research, 15(1):74–83, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1596285.

Fernandez:2019:DCE

[1103] Santiago José Fernandez, Pablo Yorio, and Javier Ernesto Ciancio. Diet composition of expanding breeding populations of the Magellanic penguin. Marine Biology Research, 15(1):84–96, 2019. CODEN ????

ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1596286.

Acosta-Pachon:2019:TIB

[1104] Tatiana A. Acosta-Pachón and Sofía Ortega-García. Trophic interaction between striped marlin and swordfish using different timescales in waters around Baja California Sur, Mexico. Marine Biology Research, 15(1):97–112, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1578377.

Soykan:2019:DTD

[1105] Ozan Soykan, Kerem Bakir, and H. Tuncay Kinacigil. Demersal trawl discards with spatial and bathymetric emphasis in the Turkish coast of the Aegean Sea. Marine Biology Research, 15(1):113–123, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1576902.

Genc:2019:FHD

[1106] Yasar Genç, Nimet Selda Basçinar, and Murat Dagtekin. Feeding habits during migration of the Atlantic bonito Sarda sarda (Bloch, 1793) to the Black Sea. Marine Biology Research, 15(2):125–136, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1596287.

Nilssen:2019:DPC

[1107] Kjell Tormod Nilssen, Ulf Lindstrøm, Jon Ivar Westgaard, Lotta Lindblom, Taija-Riitta Blencke, and Tore Haug. Diet and prey consumption of grey seals (Halichoerus grypus) in Norway. Marine Biology Research, 15(2):137–149, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1605182.

He:2019:SMP

[1108] Yuan He, Wei Zhou, Jiaming Li, and Songdong Shen. Study of mosaicism of Pyropia yezoensis. Marine Biology Research, 15(2):150–158, 2019. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2019.1605183.

Santana-Moreno:2019:FRR

[1109] Luis Daniel Santana-Moreno and Rigoberto Moreno-Mendoza. First record and range extension of the twinspot bass Serranus flaviventris (Cuvier, 1829) (Actinopterygii: Perciformes: Serranidae) in the Gulf of Mexico. Marine Biology Research, 15(2):159–162, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2019.1607386.

Das:2019:VDO

[1110] Sourav Das and Sugata Hazra. Vertical distribution of optically active substances (OAS) in the coastal sea off Indian Sundarban during post monsoon. Marine Biology Research, 15(2):163–169, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1607387.

doCarmoSilva:2019:EAA

[1111] José Paulo do Carmo Silva, Marcus Rodrigues da Costa, and Francisco Gerson Araújo. Energy acquisition and allocation to the gonadal development of Cynoscion leiachus (Perciformes, Sciaenidae) in a tropical Brazilian bay. Marine Biology Research, 15(2):170–180, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1610777.

Rocha:2019:BAM

[1112] Carlos Alberto de Oliveira Rocha, José Jonathas Pereira Rodrigues de Lira, Julianna de Lemos Santana, Márcio Paiva Guimarães, and Tereza Cristina dos Santos Calado. Biological aspects of the marine crab Plagusia depressa (Fabricius, 1775) on the northeast coast of Brazil. Marine Biology Research, 15(2):181–190, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2019.1612070.

Jensen:2019:VJS

[1113] Arn Rytter Jensen, Mikkel Ree Schneider, Jens T. Høeg, Henrik Glenner, and Jørgen Lützen. Variation in juvenile stages and success of male acquisition in Danish and French populations of the parasitic barnacle Sacculina carcini (Cirripedia: Rhizocephala) parasitizing the shore crab Carcinus maenas. Marine Biology Research, 15(2):191–203, 2019. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2019.1612071.

Magalhaes:2019:TCM

[1114] Wagner F. Magalhães and Julie H. Bailey-Brock. On two cryptogenic maldanids (Annelida) from coral rubble habitats in Hawaii. Marine Biology Research, 15(2):204–209, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2019.1613551.

Kim:2019:NRB

[1115] Jong Guk Kim and Jimin Lee. New records of benthic harpacticoids from Korea: I. Three new species of the genus Zosime Boeck (Copepoda: Harpacticoida: Zosimeidae). Marine Biology Research, 15(2): 210–239, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1619774.

Garcia:2019:AVA

[1116] A. M. Garcia, M. C. L. M. Oliveira, C. Odebrecht, J. L. A. Colling, J. P. Vieira, F. L. Rodrigues, and R. F. Bastos. Allochthonous versus autochthonous organic matter sustaining macroconsumers in a subtropical sandy beach revealed by stable isotopes. Marine Biology Research, 15(3):241–258, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1627559.

Polonia:2019:ACS

[1117] Ana Rita Moura Polónia and Daniel Francis Richard Cleary. Archaeal communities in sponge, sediment and water from marine lakes and open water habitats. Marine Biology Research, 15(3):259–274, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1633469.

Aguilar-Camacho:2019:IGC

[1118] Jose Maria Aguilar-Camacho, Grace P. McCormack, Maria Belinda Longakit, Maria Vittoria Marra, and Liam Morrison. Insights gained from the chemical composition of spicules in Haplosclerida (Porifera, Demospongiae). Marine Biology Research, 15(3):275–281, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2019.1636282.

Vazquez:2019:PDG

[1119] María Guadalupe Vázquez and Claudia Cristina Bas. Population distribution and group structure of the invasive shrimp Palaemon macrodactylus (Caridea: Palaemonidae) living in a fouling community. Marine Biology Research, 15(3):282–289, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2019.1620286.

Padate:2019:KSN

[1120] Vinay P. Padate, Hashim Manjebrayakath, and Peter K. L. Ng. Kasagia sudhakari, a new species of deep-sea spider crab (Crustacea: Brachyura:

Majidae) from the southeastern Arabian Sea. Marine Biology Research, 15(3):290–296, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1637527.

Mincarone:2019:DSM

[1121] Michael M. Mincarone, Bárbara T. Villarins, Leandro N. Eduardo, Rodrigo A. Caires, Flávia Lucena-Frédou, Thierry Frédou, Alex S. Lira, and Arnaud Bertrand. Deep-sea manefishes (Perciformes: Caristiidae) from oceanic islands and seamounts off northeastern Brazil, with comments on the caristiids previously reported in Brazilian waters. Marine Biology Research, 15(3):297–304, 2019. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2019.1636281.

Tardin:2019:HUM

[1122] Rodrigo H. Tardin, Yongwan Chun, Sheila M. Simão, and Maria Alice S. Alves. Habitat use models of spatially auto-correlated data: a case study of the common bottlenose dolphin, Tursiops truncatus truncatus, in southeastern Brazil. Marine Biology Research, 15(4–6): 305–316, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1644773.

David:2019:DBP

[1123] Andrew A. David and Michael Krick. DNA barcoding of polychaetes collected during the 2018 rapid assessment survey of floating dock communities from New England. Marine Biology Research, 15(4–6): 317–324, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1655160.

Lucic:2019:ETC

[1124] Davor Lucić, Marijana Hure, Svjetlana Bobanović-’Colić, Jakica Njire, Olja Vidjak, Ivona Onofri, Barbara Gangai Zovko, and Mirna Batistić. The effect of temperature change and oxygen reduction on zooplankton composition and vertical distribution in a semi-enclosed marine system. Marine Biology Research, 15(4–6):325–342, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2019.1655161.

Zhang:2019:MCH

[1125] Chun Y. Zhang, Guo F. Chen, Yuan Y. Wang, Xu D. Zhang, and Cheng H. Li. Molecular characterization of heat shock protein 90 from the dinoflagellate Prorocentrum donghaiense and its transcriptional response to thermal, copper and nutrient stresses. Marine Biology Research, 15 (4–6):343–356, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1662445.

Haraguchi:2019:CNP

[1126] Lumi Haraguchi, Eero Asmala, Hans H. Jakobsen, and Jacob Carstensen. Composition of natural phytoplankton community has minor effects on autochthonous dissolved organic matter characteristics. Marine Biology Research, 15(4–6):357–375, 2019. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2019.1662449.

Omori:2019:CRR

[1127] Makoto Omori. Coral restoration research and technical developments: what we have learned so far. Marine Biology Research, 15 (7):377–409, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1662050.

Terrana:2019:SRS

[1128] Lucas Terrana and Igor Eeckhaut. Sexual reproduction of the shallowwater black coral Cirrhipathes anguina (Dana, 1846) from Madagascar. Marine Biology Research, 15(7):410–423, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2019.1662444.

Dahlet:2019:PSU

[1129] Lol Iana Dahlet, Helio Herminio Checon, Leonardo Q. Yokoyama, and Alexander Turra. Patterns of shell utilization by hermit crabs as a proxy of gastropod population dynamics. Marine Biology Research, 15 (7):424–433, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1662446.

Martins:2019:OVO

[1130] Luciana Martins. Ontogenetic variation of the ossicles in Pseudothyone belli (Ludwig, 1887) with a description of a new Sclerodactylidae (Echinodermata: Holothuroidea: Dendrochirotida) from Southwestern Atlantic. Marine Biology Research, 15(7):434–443, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2019.1662051.

Ortega-Garcia:2019:HEB

[1131] Sofía Ortega-García, Alberto Abad-Uribarren, Ulianov Jakes-Cota, and Nurenskaya Vélez-Arellano. Histological evidence of bluefin tuna (Thunnus thynnus Linnaeus, 1758) spawning in waters within the Mexican Exclusive Economic Zone in the Gulf of Mexico. Marine Biology Research, 15(7):444–450, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1662448.

Costa:2019:ICW

[1132] Leonardo Lopes Costa and Ilana Rosental Zalmon. Issues to consider when sampling the sandhopper Atlantorchestoidea brasiliensis (Dana, 1853) for impact assessments on sandy beaches. Marine Biology Research, 15(8–9):451–457, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1680848.

Llanos:2019:IBC

[1133] Elizabeth Noemi Llanos, María Lourdes Jaubet, and Rodolfo Elías. The intertidal benthic community of SW Atlantic as an environmental indicator of 50 years of human-mediated changes. Marine Biology Research, 15(8–9):458–475, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1674875.

Lucena:2019:PAC

[1134] Rudá Amorim Lucena, Joafrâncio Pereira de Araújo, and Martin Lindsey Christoffersen. Pycnogonida (Arthropoda) from coral reef environments along the Southwest Atlantic: new records and new species. Marine Biology Research, 15(8–9):476–499, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2019.1674876.

Ferreira:2019:PSC

[1135] Inês Ferreira, Diana Santos, Cláudia Moreira, Diana Feijó, Alberto Rocha, and Alberto T. Correia. Population structure of Chelidonichthys lucerna in Portugal mainland using otolith shape and elemental signatures. Marine Biology Research, 15(8–9):500–512, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2019.1673897.

Du:2019:AOC

[1136] Jianguo Du, Zehao Chen, Meiling Xie, Mingru Chen, Xinqing Zheng, Jianji Liao, and Bin Chen. Analysis of organic carbon sources in tropical seagrass fish: a case study of the east coast of Hainan Province. Marine Biology Research, 15(8–9):513–522, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2019.1673896.

Ocalewicz:2019:FDK

[1137] Konrad Ocalewicz, Lech Kirtiklis, Katarzyna Mojsa, Mariusz Sapota, and Maciej Kwiatkowski. First description of karyotypes and localization of ribosomal genes in two sand lances (Uranoscopiformes: Ammodytidae); small sand-eel (Ammodytes tobianus Linnaeus, 1758) and great sand-eel (Hyperoplus lanceolatus Le Sauvage, 1824). Marine Biology Research, 15(8–9):523–529, 2019. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2019.1662447.

Kealy:2019:HCA

[1138] Rachael A. Kealy, Thomas Busk, Josephine Goldstein, Poul S. Larsen, and Hans Ulrik Riisgård. Hydrodynamic characteristics of aquiferous modules in the demosponge Halichondria panicea. Marine Biology Research, 15(10):531–540, 2019. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2019.1694691.

Mazda:2019:MCJ

[1139] Yuichi Mazda, Tsutomu Sasagawa, Yoshiakira Iinuma, and Kaori Wakabayashi. Maternal care and juvenile feeding in a hyperiid amphipod (Oxycephalus clausi bovallius, 1887) in association with gelatinous zooplankton. Marine Biology Research, 15(10):541–547, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2019.1694692.

Theil:2019:MPT

[1140] Emma Palacios Theil and Darryl L. Felder. Molecular phylogeography of Tumidotheres maculatus (Say, 1818) and Zaops ostreus (Say, 1817) (Crustacea: Decapoda: Pinnotheridae) in the western Atlantic, with description of a new species and synonymy of Epulotheres Manning, 1993. Marine Biology Research, 15(10):548–567, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2019.1693597.

Kuguru:2019:STG

[1141] Gibbs Kuguru, Enrico Gennari, Sabine Wintner, Matthew L. Dicken, Juliana D. Klein, Clint Rhode, and Aletta E. Bester van der Merwe.

Spatio-temporal genetic variation of juvenile smooth hammerhead sharks in South Africa. Marine Biology Research, 15(10):568–579, 2019. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2019.1695058.

Bringloe:2019:DBS

[1142] Trevor T. Bringloe, Kjersti Sjøtun, and Gary W. Saunders. A DNA barcode survey of marine macroalgae from Bergen (Norway). Marine Biology Research, 15(10):580–589, 2019. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2019.1699659.

vanDuyl:2020:DCF

[1143] Fleur C. van Duyl, Sabine K. Lengger, Stefan Schouten, Tomas Lundälv, Dick van Oevelen, and Christina E. Müller. Dark CO_{2} fixation into phospholipid-derived fatty acids by the cold-water coral associated sponge Hymedesmia (Stylopus) coriacea (Tisler Reef, NE Skagerrak). Marine Biology Research, 16(1):1–17, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2019.1704019.

Cleary:2020:BCS

[1144] Daniel Francis Richard Cleary, Marina R. S. Ferreira, Nguyen K. Bat, Ana Rita Moura Polónia, Newton Carlos Marcial Gomes, and Nicole Joy de Voogd. Bacterial composition of sponges, sediment and seawater in enclosed and open marine lakes in Ha Long Bay Vietnam. Marine Biology Research, 16(1):18–31, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2019.1702214.

Wang:2020:CIB

[1145] Rui Wang, Qiaona Xue, Jiangtao Wang, and Liju Tan. Competitive interactions between two allelopathic algal species: Heterosigma akashiwo and Phaeodactylum tricornutum. Marine Biology Research, 16(1):32–43, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1702213.

Gao:2020:FRC

[1146] Dahai Gao, Zhongmin Sun, Chaohua Huang, Jianting Yao, Yongqiang Wang, Wei Tan, and Fuxiao Chen. First record of Caulerpa lentillifera J. Agardh (Bryopsidales, Chlorophyta) from China. Marine Biology Research, 16(1):44–49, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2019.1702215.

Hernaez:2020:NSG

[1147] Patricio Hernáez, Amanda M. Windsor, Caio Augusto Paula, and William Santana. A new species of the ghost shrimp genus Neocallichirus Sakai, 1988 (Decapoda: Axiidea: Callianassidae) from the southwestern Atlantic. Marine Biology Research, 16(1):50–60, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2019.1703003.

Bucher:2020:NRP

[1148] Daniel J. Bucher, Jacobus W. Mosse, B. G. Hutubessy, Matthew Tonge, and Renaud Joannes-Boyau. A new record of Pristigenys meyeri (Priacanthidae) from Maluku province, eastern Indonesia with a description of the otoliths and a first estimate of age. Marine Biology Research, 16(1):61–67, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1701195.

Athanasiadis:2020:STM

[1149] Athanasios Athanasiadis. A study of the type material of Lithophyllum hibernicum (Lithophyllaceae, Corallinales, Rhodophyta) with comments on L. bathyporum and L. incrustans. Marine Biology Research, 16(1):68–76, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2019.1703138.

Kumar:2020:NRA

[1150] K. V. Aneesh Kumar, K. P. Deepa, M. M. Muhammed Rafeeq, G. Sree Renjima, Hashim Manjebrayakath, B. R. Smitha, N. Saravanane, M. P. Prabhakaran, V. N. Sanjeevan, and M. Sudhakar. Near reef abundance and on-offshore distribution of tuna larvae around Minicoy Island in Lakshadweep Sea, India. Marine Biology Research, 16(2):77–92, 2020. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2019.1704018.

Yamaguchi:2020:RBE

[1151] Tadanori Yamaguchi, Katsumi Takayama, Naoki Hirose, and Michiya Matsuyama. Relationship between empirical water temperature and spring characteristics of swordtip squid (Uroteuthis edulis) caught in the eastern Tsushima Strait. Marine Biology Research, 16(2):93– 102, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1712420.

Santos:2020:ESR

[1152] Régis Santos, Ana Novoa-Pabon, Hélder Silva, and Mário Pinho. Elasmobranch species richness, fisheries, abundance and size composition in the Azores archipelago (NE Atlantic). Marine Biology Research, 16 (2):103–116, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1718713.

Nguyen:2020:CAR

[1153] Hai-Thanh T. Nguyen, Binh T. Dang, Henrik Glenner, and Audrey J. Geffen. Cophylogenetic analysis of the relationship between anemonefish Amphiprion (Perciformes: Pomacentridae) and their symbiotic host anemones (Anthozoa: Actiniaria). Marine Biology Research, 16(2): 117–133, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1711952.

Gopal:2020:PAA

[1154] Aiswarya Gopal, Usha Vaikom Parameswaran, Koovapurath Useph Abdul Jaleel, and Narayanane Saravanane. Prionospio atrovitta (Annelida: Spionidae), a new species from Andaman and Nicobar archipelago, northern Indian Ocean. Marine Biology Research, 16(2):134–140, 2020. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2019.1708950.

Quintela:2020:PSB

[1155] María Quintela, François Besnier, Bjørghild Seliussen, Kevin A. Glover, and Ulf Lindstrøm. Population structure of bycaught harbour porpoise (Phocoena phocoena) in Norway. Marine Biology Research, 16 (3):141–147, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1729992.

OMalley:2020:RPR

[1156] Betsy L. O’Malley and Heather L. Hunt. Role of predators in the recruitment of invertebrates in a rocky subtidal community in the southwest Bay of Fundy, NB. Marine Biology Research, 16(3):148–165, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2020.1721541.

Dautov:2020:ELD

[1157] Salim Shakirovich Dautov. The embryological and larval development of the sea urchin Diadema savignyi (Audouin, 1809) (Diadematoida: Diadematidae) from the South China Sea. Marine Biology Research, 16 (3):166–176, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1726960.

Hernandez:2020:RCG

[1158] Eleazar Hernandez, Omar A. Vázquez, André Torruco, and Md Saydur Rahman. Reproductive cycle and gonadal development of the Atlantic sea urchin Arbacia punctulata in the Gulf of Mexico: changes in nutritive phagocytes in relation to gametogenesis. Marine Biology Research, 16 (3):177–194, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1731758.

Kakui:2020:PSS

[1159] Keiichi Kakui and Yoshihisa Fujita. Paradoxapseudes shimojiensis sp. nov. (Crustacea: Tanaidacea: Apseudidae) from a submarine limestone cave in Japan, with notes on its chelipedal morphology and sexual system. Marine Biology Research, 16(3):195–207, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2020.1720249.

Padate:2020:DNS

[1160] Vinay P. Padate, Sherine Sonia Cubelio, and K. V. Jayachandran. Description of a new species of deep-water crab of the genus Homolodromia A. Milne-Edwards, 1880 from the northern Indian Ocean (Crustacea: Decapoda: Brachyura: Homolodromiidae). Marine Biology Research, 16 (4):209–218, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1735641.

Martins:2020:TBA

[1161] Luciana Martins and Camilla Souto. Taxonomy of the Brazilian Apodida (Holothuroidea), with the description of two new genera. Marine Biology Research, 16(4):219–255, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2020.1761027.

David:2020:TOC

[1162] Andrew A. David and Jordan Cahill. Tri-oceanic connectivity of the supposedly cosmopolitan polychaete, Harmothoe imbricata (Annelida:

Polynoidae): insights from the COI marker. Marine Biology Research, 16(4):256–264, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1740740.

Kenchington:2020:PAL

[1163] Trevor J. Kenchington, Daphne E. Themelis, Shannon C. DeVaney, and Theodore W. Pietsch. Pelagic anglerfishes (Lophiiformes: Ceratioidei) of The Gully Marine Protected Area. Marine Biology Research, 16 (4):265–279, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1771598.

Deli:2020:CGP

[1164] Temim Deli, Souad Guizeni, Lotfi Ben Abdallah, Khaled Said, and Noureddine Chatti. Chaotic genetic patchiness in the pelagic teleost fish Sardina pilchardus across the Siculo–Tunisian Strait. Marine Biology Research, 16(4):280–298, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2020.1752921.

Sorlie:2020:DCB

[1165] Maria Sørlie, Kjell Tormod Nilssen, Arne Bjørge, and Carla Freitas. Diet composition and biomass consumption of harbour seals in Telemark and Aust-Agder, Norwegian Skagerrak. Marine Biology Research, 16(4):299–310, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1751205.

Gil:2020:RTS

[1166] Damian G. Gil, Estela C. Lopretto, and Héctor E. Zaixso. Reproductive timing and synchronized reproduction of the sea urchin Pseudechinus magellanicus (Echinoidea: Temnopleuridae) in central Patagonia, Argentina. Marine Biology Research, 16(5):311–326, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2020.1772493.

Leon-Guzman:2020:RBD

[1167] Sairi Sarai León-Guzmán, María del Carmen Alejo-Plata, Enrique Morales-Bojórquez, and Francisco Benítez-Villalobos. Reproductive biology of the dart squid, Lolliguncula diomedeae (Cephalopoda: Loliginidae) from Gulf of Tehuantepec, Mexico. Marine Biology Research, 16 (5):327–339, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1777433.

Rumbold:2020:BIA

[1168] Carlos Rumbold, Nicolás Battini, Clara Belen Giachetti, Karen Lidia Castro, Sandra Obenat, and Evangelina Schwindt. Biological invasions in artificial habitats: factors that determine the presence of native and exotic peracarid crustacea species in Southwestern Atlantic. Marine Biology Research, 16(5):340–355, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2020.1781187.

Siciliano:2020:EBX

[1169] Salvatore Siciliano, Júlio Cardoso, Arlaine Francisco, Shirley P. De Souza, Rachel Ann Hauser-Davis, and Tammy Iwasa-Arai. Epizoic barnacle (Xenobalanus globicipitis) infestations in several cetacean species in south-eastern Brazil. Marine Biology Research, 16(5):356–368, 2020. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2020.1783450.

Ringvold:2020:ERG

[1170] Halldis Ringvold, Andreas Hatlevik, Jan Hevrøy, Marc Hughes, and Nils Aukan. Encounters with the rare genus Helicosalpa (Chordata, Thaliacea, Salpida), using citizen science data. Marine Biology Research, 16 (5):369–379, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1781189.

Sokolov:2020:PAL

[1171] Sergey G. Sokolov, Ilya I. Gordeev, and Dmirtry M. Atopkin. Phylogenetic affiliation of the lepocreadiid trematodes parasitizing some marine fishes in the North-western Pacific. Marine Biology Research, 16 (5):380–389, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1758947.

Hosoda:2020:NSP

[1172] Yushi Hosoda and Keiichi Kakui. A new species of Pantopipetta (Pycnogonida: Austrodecidae) from the North Pacific, with a note on the palp articulation. Marine Biology Research, 16(5):390–395, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2020.1758946.

Aneesh:2020:NSR

[1173] Panakkool Thamban Aneesh, Niel L. Bruce, Musaliyarakam Nashad, Kinattumkara Bineesh, and Abdulla A. Mohamed Hatha. A new species of Renocila Miers, 1880 (Crustacea: Isopoda: Cymothoidae), a fish parasitic isopod from Andaman Island, India. Marine Biology Research, 16 (5):396–410, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1761028.

Okamoto:2020:NSZ

[1174] Nobuya Okamoto, Yuki Oya, and Keiichi Kakui. A new species of Zeuxo (Crustacea: Peracarida: Tanaidacea) from Japan, with remarks on carapace pigmentation as a potentially useful taxonomic character. Marine Biology Research, 16(5):411–422, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2020.1766693.

Christophersen:2020:RCR

[1175] Gyda Christophersen, Ingebrigt Bjørkevoll, Snorre Bakke, and Margareth Kjerstad. Reproductive cycle of the red sea cucumber, Parastichopus tremulus (Gunnerus, 1767), from western Norway. Marine Biology Research, 16(6–7):423–430, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2020.1781188.

ElMghazli:2020:CSB

[1176] Hayat El Mghazli, Abdelaziz Mounir, Mohammed Znari, Mohamed Naimi, Hassan El Ouizgani, and Mohamed Aourir. A comparison of spring body condition, maturation status, and diet in the European sardine, Sardina pilchardus (Teleostei: Clupeidae) from contrasting environments off the Moroccan Atlantic coast. Marine Biology Research, 16 (6–7):431–445, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1817489.

Folkvord:2020:SFH

[1177] Arild Folkvord, Essi Lakso, Marita Laupsa, Sonnich Meier, Luciana Alves Musialak, and Svein Sundby. Swimbladder filling in herring larvae: effects of food oil on the water surface. Marine Biology Research, 16(6– 7):446–457, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1837882.

Coelho:2020:PCA

[1178] Jéssica Fernanda Ramos Coelho, Sergio Maia Queiroz Lima, and Flávia de Figueiredo Petean. Phylogenetic conservatism of abiotic niche in sympatric Southwestern Atlantic skates. Marine Biology Research, 16 (6–7):458–473, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1837883.

Tadmor-Levi:2020:ICT

[1179] Roni Tadmor-Levi, Lior David, and Daniel Golani. Indication of cryptic taxa within the flat needlefish, Ablennes hians based on analysis of the cytochrome oxidase I region on a wide range of samples. Marine Biology Research, 16(6–7):474–479, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2020.1834117.

Bascinar:2020:STV

[1180] Nimet Selda Basçinar, Ahmet Mutlu Gözler, Elnare Eldar Jafarova, Murat Erbay, Ilkay Özcan Akpinar, Elvan Atilgan, Ülgen Aytan, D. Selim Misir, and Bayram Zengin. Spatial and temporal variations in the species composition and abundance of benthic macroinvertebrates along the coast of the Eastern Black Sea. Marine Biology Research, 16(6– 7):480–493, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1805469.

Landers:2020:KCA

[1181] Stephen C. Landers, Randall D. Bassham, Jonathan M. Miller, Jeroen Ingels, Nuria Sánchez, and Martin V. Sørensen. Kinorhynch communities from Alabama coastal waters. Marine Biology Research, 16(6– 7):494–504, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1789660.

Sun:2020:DAC

[1182] Jingxian Sun, Ziqian Huang, Ruhao You, Shan Liang, Weixiao Di, Zheyan Chang, and Peng Zhang. Distribution and accumulation of Cd in Ligia cinerascens and its effect on ultrastructure of hepatopancreas. Marine Biology Research, 16(6–7):505–513, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2020.1805631.

Bitner:2020:NRR

[1183] Maria Aleksandra Bitner and Tina N. Molodtsova. New records of recent brachiopods (Terebratulida) from the northern Mid-Atlantic Ridge. Marine Biology Research, 16(6–7):514–520, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2020.1797093.

Johansen:2020:NIN

[1184] Per-Otto Johansen and Wim Vader. Nicippe isaki, a new species of Pardaliscidae (Crustacea: Amphipoda) from Svalbard. Marine Biology Research, 16(6–7):521–531, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2020.1828593.

Athanasiadis:2020:PDG

[1185] Athanasios Athanasiadis. Phragmope discrepans, gen. and comb. nov. (Mesophyllaceae, Corallinales, Rhodophyta), the species known as ‘ Mesophyllum engelhartii’ from South Africa. Marine Biology Research, 16 (6–7):532–549, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1831543.

Gonzalez-Lorenzo:2020:UPS

[1186] Gustavo Gonzalez-Lorenzo, Kilian Toledo-Guedes, Ezequiel Rodríguez-Riesco, Erika González-Méndez, Yulimar González-Rodríguez, and Cristina Piñeiro-Corbeira. The underreported presence of the subtidal seagrass Cymodocea nodosa in rocky intertidal pools of the Canary Islands. Marine Biology Research, 16(6–7):550–555, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2020.1834583.

Ahmad:2020:DBF

[1187] MuhdSyukeri Ahmad, Noraznawati Ismail, and Ibrahim M. S. Eldeen. Disruption of biofilm formation by fractions from book gills and carapace of horseshoe crab Tachypleus gigas. Marine Biology Research, 16(6– 7):556–564, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1840592.

Aneesh:2020:NGS

[1188] Panakkool Thamban Aneesh, Kerry A. Hadfield, Nico J. Smit, and Appukuttannair Biju Kumar. A new genus and species of fish parasitic cymothoid isopod (Crustacea) from Indian waters, with a key to the branchial-attaching cymothoid genera. Marine Biology Research, 16(8– 9):565–584, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2020.1851032.

Vagenas:2020:AGF

[1189] Georgios Vagenas, Paraskevi K. Karachle, Aikaterini Dogrammatzi, and Athanassios C. Tsikliras. Age, growth and feeding habits of boarfish (Capros aper) in the Aegean and Ionian Seas (Greece). Marine Biology Research, 16(8–9):585–592, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2020.1855657.

Correia:2020:REG

[1190] Edna Correia, Camilo Carneiro, and António Araújo. Reproductive ecology and growth of marine catfishes (Ariidae) supporting sustainable fisheries in Banc d’Arguin National Park, Mauritania. Marine Biology Research, 16(8–9):593–599, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2020.1855658.

Jaureguizar:2020:EII

[1191] Andrés Javier Jaureguizar, Agustín María De Wysiecki, and Martina Daniela Camiolo. Environmental influence on the inter-annual demographic variation of the narrownose smooth-hound shark (Mustelus schmitti, Springer 1939) in the Northern Argentine Coastal System (El Rincón, 38–42^{o}S). Marine Biology Research, 16(8–9):600–615, 2020. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2020.1868522.

Moraes:2020:BCS

[1192] Isabela R. R. Moraes, Alexandre Oliveira Almeida, Valter José Cobo, Douglas Fernandes Rodrigues Alves, Thiago Maia Davanso, and Antonio Leão Castilho. Biodiversity of caridean shrimps on rocky bottoms of two preserved islands on the southeastern Brazilian coast. Marine Biology Research, 16(8–9):616–631, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2020.1864830.

Dixit:2020:RPG

[1193] Sudhanshu Dixit, Hashim Manjebrayakath, and N. Saravanane. A rare polyclad genus Bulaceros (Platyhelminthes: Polycladida: Pseudocerotidae): new species and new record from Indian coral atolls. Marine Biology Research, 16(8–9):632–642, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2020.1870044.

Berg:2020:FSP

[1194] Justin T. Berg, Charlotte M. David, Melissa M. Gabriel, and Bastian Bentlage. Fluorescence signatures of persistent photosystem damage in the staghorn coral Acropora cf. pulchra (Anthozoa: Scleractinia) during bleaching and recovery. Marine Biology Research, 16(8– 9):643–655, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1875245.

Lincy:2020:CBC

[1195] Jovitha Lincy and Cathrine Sumathi Manohar. A comparison of bacterial communities from OMZ sediments in the Arabian Sea and the Bay of Bengal reveals major differences in nitrogen turnover and carbon recycling potential. Marine Biology Research, 16(8–9):656–673, 2020. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2020.1840593.

Almen:2020:TRT

[1196] Anna-Karin Almén and Tobias Tamelander. Temperature-related timing of the spring bloom and match between phytoplankton and zooplankton. Marine Biology Research, 16(8–9):674–682, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2020.1846201.

Mohan:2020:IWC

[1197] Arya P. Mohan and Abraham Biju. Influence of warm core eddy on the vertical distribution of autotrophic pico- and nanoplankton in the Bay of Bengal. Marine Biology Research, 16(10):683–694, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1872795.

Dias:2020:HMT

[1198] Cristina de Oliveira Dias, Ana Cristina Teixeira Bonecker, Márcia Salustiano de Castro, Pedro Freitas de Carvalho, Rodolfo Paranhos, and Sérgio Luiz Costa Bonecker. Holoplankton and meroplankton of three western Atlantic sedimentary basins. Marine Biology Research, 16(10): 695–713, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1894341.

Nevarez-Lopez:2020:CPG

[1199] Cintya Nevárez-López, Norma Hernández-Saavedra, Arturo Sánchez-Paz, Delia Rojas-Posadas, Adriana Muhlia-Almazán, and Juana López-Martínez. Colour polymorphism and genetic structure in the cannonball jellyfish (Stomolophus meleagris, L. Agassiz, 1860) in the Gulf of California. Marine Biology Research, 16(10):714–728, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1887495.

daCruzPires:2020:BCP

[1200] Ana Cecília da Cruz Pires, Vanessa Oliveira, Daniel Francis Richard Cleary, Ana Rita Moura Polónia, Francisco José Riso Costa Coelho, Zarinah Waheed, Nicole Joy de Voogd, and Newton Carlos Marcial Gomes. Bacterial composition and putative functions associated with sponges, sediment and seawater from the Tioman coral reef system, Peninsular Malaysia. Marine Biology Research, 16(10):729–743, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2021.1891250.

Cleary:2020:CPC

[1201] Daniel F. R. Cleary and Yusheng M. Huang. A comparison of the prokaryotic communities associated with seven seaweed species, sediment, and seawater from the Penghu archipelago, Taiwan. Marine Biology Research, 16(10):744–761, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2020.1859119.

Mincarone:2020:DSS

[1202] Michael M. Mincarone, Júlia R. Martins, Fabio Di Dario, Leandro N. Eduardo, Thierry Frédou, Flávia Lucena-Frédou, and Arnaud Bertrand. Deep-sea smelts, pencil smelts, and barreleyes (teleostei: Argentiniformes) from oceanic islands and seamounts off northeastern Brazil. Marine Biology Research, 16(10):762–773, 2020. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1891806.

Golani:2020:ZPR

[1203] Daniel Golani, Ronald Fricke, and Brenda Appelbaum-Golani. Zoogeographic patterns of Red Sea fishes — are they correlated to success in colonization of the Mediterranean via the Suez Canal? Marine Biology Research, 16(10):774–780, 2020. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.1894340.

Aneesh:2021:NSP

[1204] Panakkool Thamban Aneesh, Ameri Kottarathil Helna, Appukuttannair Biju Kumar, and Balu Alagar Venmathi Maran. A new species of parasitic copepod of the genus Lernaeenicus Lesueur, 1824 (Siphonostomatoida: Pennellidae) from the torpedo scad Megalaspis cordyla (Linnaeus) off Kerala coast of Arabian Sea, India. Marine Biology Research, 17(1):1–11, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1887498.

Khalaji-Pirbalouty:2021:NSD

[1205] Valiallah Khalaji-Pirbalouty and Jean-Marc Gagnon. A new species of Dynoides Barnard, 1914 (Crustacea, Isopoda, Sphaeromatidae) from Canada, with notes on geographic distribution of the north-eastern Pacific Ocean species. Marine Biology Research, 17(1):12–20, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1892766.

Aviz:2021:SPA

[1206] Daiane Aviz, Cleverson Ranieri Meira dos Santos, and José Souto Rosa Filho. Sabellariid (Polychaeta: Annelida) reefs as nursery ground for the hermit crab Clibanarius symmetricus (Randall, 1840) on the Amazonian coast of Brazil. Marine Biology Research, 17(1):21–30, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1887494.

Sokolov:2021:PPH

[1207] Sergey G. Sokolov, Dmitry M. Atopkin, and Ilya I. Gordeev. Phylogenetic position of the hemiuroid genus Paraccacladium Bray and Gibson, 1977 (Trematoda: Hemiuroidea) and the status of the subfamily Paraccacladiinae Bray and Gibson, 1977. Marine Biology Research, 17 (1):31–40, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1891252.

Polpetta:2021:BRS

[1208] Maila Polpetta, Francesco Piva, Stefano Gridelli, and Filippo Bargnesi. Behavioural responses in the sand tiger shark (Carcharias taurus) to permanent magnets and pulsed magnetic fields. Marine Biology Research, 17(1):41–56, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1887496.

DiMatteo:2021:NHR

[1209] Andrew DiMatteo, Gwen Lockhart, and Susan Barco. Normalizing home ranges of immature Kemp’s ridley turtles (Lepidochelys kempii) in an important estuarine foraging area to better assess their spatial distribution. Marine Biology Research, 17(1):57–71, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1896004.

Zeng:2021:OFR

[1210] Lei Zeng, Zhenzhao Tang, Pimao Chen, Jie Yu, and Guobao Chen. Optimization of fishery resources assessment methods and ecological effects evaluation of artificial reefs. Marine Biology Research, 17(1): 72–85, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1887497.

Boonsanit:2021:ESH

[1211] Phurich Boonsanit and Supanut Pairohakul. Effects of salinity on haemolymph osmolality, gill Na^{+} /K^{+} ATPase and antioxidant enzyme activities in the male mud crab Scylla olivacea (Herbst, 1796). Marine Biology Research, 17(1):86–97, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2021.1900496.

Fazhan:2021:EDM

[1212] Hanafiah Fazhan, Khor Waiho, Mohamad Jalilah, Mhd Ikhwanuddin, Ghazali Azmie, Sabri Muda, Mohd Zaidi Ibrahim, Draman Shuhaimi, Mohd Husin Norainy, Muhammad Allim Jamalludin, and Ngah Rusnita. Effect of different measuring techniques, preservation methods and storage duration on the morphometric measurements of crustacean larvae. Marine Biology Research, 17(1):98–105, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1900576.

SriHari:2021:SCG

[1213] Murugesan SriHari, Susanthi Silpa, Annam Pavan-Kumar, Shashi Bhushan, Binaya Bushan Nayak, and Zeba Jaffer Abidi. Stock characterization of greater lizardfish, Saurida tumbil (Bloch, 1795) along the west coast of India using morphological and molecular markers. Marine Biology Research, 17(2):107–119, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2021.1891251.

Bartels:2021:TIS

[1214] Paul J. Bartels, Paulo Fontoura, Diane R. Nelson, Sebastian Orozco-Cubero, Monika Mioduchowska, Magdalena Gawlak, Lukasz Kaczmarek, and Jorge Cortés. A trans-isthmus survey of marine tardigrades from Costa Rica (Central America) with descriptions of seven new species. Marine Biology Research, 17(2):120–166, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1901936.

Cirino:2021:DSS

[1215] Luca Cirino, Sujune Tsai, Fu-Wen Kuo, Zhi-Hong Wen, Pei-Jie Meng, and Chiahsin Lin. Decline of Seriatopora (Scleractinia: Pocilloporidae) fecundity in Taiwan in 2018–2019. Marine Biology Research, 17 (2):167–171, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1906904.

Gil:2021:PCH

[1216] Marta Gil and Fran Ramil. Plumularioidea (Cnidaria, Hydrozoa) from the Guinea Current Large Marine Ecosystem (GCLME). Marine Biology Research, 17(2):172–184, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.1913187.

Riyas:2021:SJB

[1217] Abdul Riyas, Neelesh Dahanukar, Krishnan Anoop Krishnan, and Appukuttannair Biju Kumar. Scyphozoan jellyfish blooms and their relationship with environmental factors along the south-eastern Arabian Sea. Marine Biology Research, 17(2):185–199, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1916034.

Hmida:2021:MGC

[1218] Leila Hmida, Chiheb Fassatoui, Sarra Missaoui, Jihen Zaghab, Mohamed Salah Romdhane, François Le Loc’h, and Frida Ben Rais Lasram. Morphological and genetic characterization of the invasive rayed pearl oyster Pinctada imbricata radiata (Mollusca: Bivalvia: Pteriidae) populations from contrasting environments along the Tunisian coast. Marine Biology Research, 17(2):200–214, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2021.1919713.

Trivellini:2021:CSS

[1219] M. Magdalena Trivellini, Silvina Van der Molen, Luis Filun, and Federico Márquez. Can shell shape be used to find the origin of South American mussels? Marine Biology Research, 17(2):215–222, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1919714.

Buhl-Mortensen:2021:DHS

[1220] Lene Buhl-Mortensen and Arne Hassel. Distribution and habitat of Scalpellidae (Cirripedia: Thoracica) in the Norwegian and Barents Seas. Marine Biology Research, 17(3):223–233, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1928220.

Linares:2021:BPB

[1221] Fátima Linares, Evaristo Pérez Rial, J. L. Rodriguez, Gema Pazos, and Blanca Álvarez-Blázquez. Biometric parameters and biochemical composition of wild wreckfish (Polyprion americanus). Marine Biology Research, 17(3):234–246, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.1923752.

Rosales:2021:RBB

[1222] José Luis Bortolini Rosales, Jesús Andrés Mejía Estrada, María del Pilar Alonso Reyes, Jesús Romero-Rodríguez, and Juan Antonio Baeza. Reproductive biology of the bopyrid isopod Robinione overstreeti, a branchial parasite of the ghost shrimp Callichirus islagrande (Decapoda: Callichiridae) in the Gulf of Mexico. Marine Biology Research, 17(3): 247–259, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1928221.

Abe:2021:IFF

[1223] Hiroshi Abé, Shizuha Mizuno, and Teruhisa Komatsu. Invertebrates found on floating Sargassum horneri (Fucales, Sargassaceae) under the Kuroshio Current in the Pacific Ocean. Marine Biology Research, 17 (3):260–271, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1931889.

Kikuchi:2021:UGR

[1224] Eidi Kikuchi, Luis Gustavo Cardoso, Delfina Canel, Juan T. Timi, and Manuel Haimovici. Using growth rates and otolith shape to identify the population structure of Umbrina canosai (Sciaenidae) from the Southwestern Atlantic. Marine Biology Research, 17(3):272–285, 2021. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2021.1938131.

Lafite:2021:FBB

[1225] Niviane Ferreira Lafite, Martin Lindsey Christoffersen, and Anne Isabelley Gondim. Feeding biology of the brittlestar Ophioderma appressum (Say, 1825) (Echinodermata, Ophiuroidea) in a tropical shallow reef from the Brazilian coast. Marine Biology Research, 17(3):286–294, 2021. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2021.1943753.

Urra:2021:DAS

[1226] J. Urra, P. Marina, T. García, and J. Baro. Damage assessment and survival estimates in the wedge clam (Donax trunculus) caught by mechanical dredging in the northern Alboran Sea. Marine Biology Research, 17(3):295–310, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1949618.

Stoltenberg:2021:RJT

[1227] Ina Stoltenberg, Jan Dierking, Dörthe C. Müller-Navarra, and Jamileh Javidpour. Review of jellyfish trophic interactions in the Baltic Sea. Marine Biology Research, 17(4):311–326, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1964532.

Sam:2021:IFS

[1228] Shu Qin Sam, Chin Soon Lionel Ng, Yuichi Preslie Kikuzawa, Tai Chong Toh, Wan Ting Sim, and Loke Ming Chou. Influence of fragment size on post transplantation growth and survival of domed scleractinian corals. Marine Biology Research, 17(4):327–340, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1957934.

ElMghazli:2021:DAH

[1229] Hayat El Mghazli, Mohammed Znari, Abdelaziz Mounir, Hassan Benaissa, and Hassan El Ouizgani. Does the Atlantic horse mackerel Trachurus trachurus (Teleostei: Carangidae) differentiate morphologically within the putative Moroccan–Saharan stock? Marine Biology Research, 17(4):341–349, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1957935.

Fang:2021:DCT

[1230] Zhou Fang, Yue Jin, Xinjun Chen, and Yan Wang. Differences in the concentrations of trace elements among different hard structures and their potential application in species identification: a case study on Loliginidae cryptic species. Marine Biology Research, 17(4):350–361, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1957936.

Tarula-Marin:2021:FRM

[1231] Alexis O. Tárula-Marín and Nancy C. Saavedra-Sotelo. First record of the mating system in the grey smoothhound shark (Mustelus californicus). Marine Biology Research, 17(4):362–368, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2021.1964533.

Andreyeva:2021:HEO

[1232] A. Y. Andreyeva, O. L. Gostyukhina, E. S. Kladchenko, D. A. Afonnikov, D. A. Rasskazov, A. O. Lantushenko, and E. A. Vodiasova. Hypoxia exerts oxidative stress and changes in expression of antioxidant enzyme genes in gills of Mytilus galloprovincialis (Lamarck, 1819). Marine Biology Research, 17(4):369–379, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2021.1967992.

King:2021:BWD

[1233] Carissa D. King, Emily Chou, Melinda L. Rekdahl, Sarah G. Trabue, and Howard C. Rosenbaum. Baleen whale distribution, behaviour and overlap with anthropogenic activity in coastal regions of the New York Bight. Marine Biology Research, 17(4):380–400, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2021.1967993.

Hune:2021:PAT

[1234] Mathias Hüne, Pablo A. Oyarzún, Pablo Reyes, Gladys Rivera, and Mario Montecinos. Phylogeographic analysis of Thyrsites atun (Perciformes: Gempylidae) reveals connectivity between fish from South Africa and Chile. Marine Biology Research, 17(4):401–413, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2021.1967994.

Wimalasiri:2021:SPK

[1235] H. B. U. G. M. Wimalasiri, W. R. W. M. A. P. Weerakoon, R. P. P. K. Jayasinghe, and P. Dalpadado. Spatial patterns and key drivers of zooplankton in the north central Indian Ocean. Marine Biology Research, 17(5–6):415–433, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1975755.

Parameswaran:2021:NNR

[1236] Usha V. Parameswaran, Aiswarya Gopal, K. U. Abdul Jaleel, and N. Saravanane. Nine new records of brittle stars (Echinodermata: Ophiuroidea) from Indian waters. Marine Biology Research, 17(5–6): 434–453, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1967995.

Sjotun:2021:FAN

[1237] Kjersti Sjøtun, Caroline S. Armitage, Mette Eilertsen, and Christiane Todt. Fauna associated with non-native Sargassum muticum (Fucales, Phaeophyceae) vary with thallus morphology and site type (sounds and bays). Marine Biology Research, 17(5–6):454–466, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1994999.

Hookabe:2021:HFD

[1238] Natsumi Hookabe, Hisanori Kohtsuka, and Hiroshi Kajihara. A histology-free description of Tetrastemma cupido sp. nov. (Nemertea: Eumonostilfera) from Sagami Bay, Japan. Marine Biology Research, 17 (5–6):467–474, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1979236.

Kunzmann:2021:FRS

[1239] Andreas Kunzmann. First record of stonefish metabolism: baseline respiration and spontaneous activity of tropical marine Synanceia verrucosa. Marine Biology Research, 17(5–6):475–485, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2021.1979237.

Signorelli:2021:BMS

[1240] Javier H. Signorelli, Javier Sellanes, and Cynthia M. Asorey. Bathymodiolinae mussels in the South-eastern Pacific: the presence of the genus Adipicola (Bivalvia: Mytilidae) at a methane seep site off central Chile. Marine Biology Research, 17(5–6):486–493, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.1975754.

Khalaji-Pirbalouty:2021:RTS

[1241] Valiallah Khalaji-Pirbalouty and Niel L. Bruce. Redescription of the type species of the genus Cassidinidea hansen, 1905 (Crustacea: Isopoda: Sphaeromatidae), with notes on geographic distribution of the New World species. Marine Biology Research, 17(5–6):494–502, 2021. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2021.1990958.

Shahdadi:2021:MND

[1242] Adnan Shahdadi, Pierre A. Mvogo Ndongo, and Christoph D. Schubart. Mito-nuclear discordance in West African mangrove crab species (Decapoda: Brachyura: Sesarmidae) suggests uni-directional mitochondrial introgression, despite prolonged evolutionary independence. Marine Biology Research, 17(5–6):503–512, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2021.1990959.

Sing:2021:PIE

[1243] Lim Chin Sing, Tan Lester Jin Xiang, Neo Mei Lin, Teo Serena Lay Ming, Beng Stephen Sui Leung, Nadarajan Saravanan, and Tun Karenne. A preliminary investigation of epibiotic macrofauna on the mangrove horseshoe crab Carcinoscorpius rotundicauda (Latreille, 1802) in Singapore. Marine Biology Research, 17(5–6):513–522, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2021.1986723.

Checon:2021:NUR

[1244] Helio H. Checon, Mariana Silva, Guilherme N. Corte, Leonardo Q. Yokoyama, M. Alexandra Teodósio, and Alexander Turra. Night underwater rides: the activity of a sandy beach gastropod is affected by interactive effects of light availability and water level. Marine Biology Research, 17(5–6):523–528, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.1974044.

Dautov:2021:NDD

[1245] S. S. Dautov. New data on distribution of muscle and serotonincontaining cells in larvae of diadematid sea urchins from nha Trang Bay (South China Sea). Marine Biology Research, 17(5–6):529– 536, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1975753.

Tejedor-Junco:2021:PMA

[1246] María Teresa Tejedor-Junco, Valeria Cubas Díaz, Margarita González-Martín, and Fernando Tuya. Presence of microplastics and antimicrobialresistant bacteria in sea cucumbers under different anthropogenic influences in Gran Canaria (Canary Islands, Spain). Marine Biology Research, 17(7–8):537–544, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1990960.

Shiyong:2021:STC

[1247] Wen Shiyong, Sun Tianli, Wang Zizhu, Li Fei, Li Dongmei, and Liu Na. Spatial and temporal characteristics of harmful algal blooms in Zhejiang Province waters during 1933–2018. Marine Biology Research, 17(7–8):545–553, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2009874.

Parvathi:2021:BGN

[1248] A. Parvathi, C. R. Vishal, N. V. Madhu, P. Anil, P. M. Mohammed Iqbal, P. Gopika, and M. F. Camey. Bioluminescence of green Noctiluca scintillans from the coastal waters of Kochi (southwest coast of India) and in vivo experiments on bioluminescence in response to mechanical stimuli and its diurnal variations. Marine Biology Research, 17 (7–8):554–568, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2012578.

Shi:2021:CMB

[1249] Xiaomeng Shi, Xing Chang, Xiaoyu Guo, Chenfei Zhao, and Shanying Tong. Calcification moderates the biochemical responses of Gephyrocapsa oceanica to ocean acidification. Marine Biology Research, 17(7– 8):569–575, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2016841.

Karunakumari:2021:SSA

[1250] R. Karunakumari, N. V. Vinithkumar, Reena Singh, D. Magesh Peter, and G. Dharani. Seaweeds of South Andaman: diversity, distribution, biochemical characterization and fatty acid profiling of some selected species. Marine Biology Research, 17(7–8):576–591, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.2016842.

Bracun:2021:STO

[1251] Sandra Bracun, Maximilian Wagner, and Stephan Koblmüller. Spatiotemporal occurrence patterns of epibiota along the leaves of the seagrass Cymodocea nodosa in the Northern Adriatic Sea. Marine Biology Research, 17(7–8):592–602, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.2015389.

Vaga:2021:IAR

[1252] Claudia F. Vaga, Alvaro E. Migotto, Maria E. A. Santos, Renato M. Honji, James Reimer, and Marcelo V. Kitahara. An integrative approach reveals a new species of Zoantharia (Cnidaria, Anthozoa), Terrazoanthus silveirai, in the South-western Atlantic Ocean. Marine Biology Research, 17(7–8):603–614, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.1998539.

Rigoberto:2021:TRP

[1253] Rosas-Luis Rigoberto and Elias-Valdez Alexei. Trophic resource partitioning among four sympatric lutjanid species in the southern Mexican Caribbean Sea. Marine Biology Research, 17(7–8):615–624, 2021. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2021.2009872.

Breedy:2021:TSG

[1254] Odalisca Breedy, Judith Camps-Castellà, G. Försterra, and Verena Häussermann. Taxonomic status and geographic distribution of Phycogorgia fucata (Valenciennes 1846) (Octocorallia: Gorgoniidae). Marine Biology Research, 17(7–8):625–633, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2021.2009873.

Ringvold:2021:FES

[1255] Halldis Ringvold and Eero J. Vesterinen. First in situ observations of the free-floating gelatinous matrix of blackbelly rosefish Helicolenus dactylopterus (Delaroche, 1809). Marine Biology Research, 17(7–8): 634–645, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2012579.

Zhai:2021:RPA

[1256] Yida Zhai, Runqi Li, Fuguo Liu, Chunyun Zhang, Yuanyuan Wang, and Guofu Chen. Recombinase polymerase amplification combined with lateral flow dipstick for the rapid detection of Prorocentrum donghaiense. Marine Biology Research, 17(7–8):646–657, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2021.2010279.

deCarvalho:2021:TAI

[1257] Cássia de Carvalho, Leandro Rabello Monteiro, Karina A. Keunecke, Helena P. Lavrado, and Ana Paula M. Di Beneditto. Trophic assessment and isotopic niches of the sympatric penaeids species Penaeus brasiliensis and P. paulensis in SW Atlantic estuarine systems. Marine Biology Research, 17(7–8):658–668, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.2011320.

Nashad:2021:NSB

[1258] Musaliyarakam Nashad, Panakkool Thamban Aneesh, Appukuttannair Biju Kumar, and Kinattumkara Bineesh. A new species of branchial fish parasitic isopod, Norileca Bruce, 1990 (Crustacea: Isopoda: Cymothoidae), from the Andaman Islands, India. Marine Biology Research, 17(7–8):669–679, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2011321.

Bruce:2021:NSM

[1259] Niel L. Bruce and Eknarin Rodcharoen. A new species of Metacirolana kussakin, 1979 (Crustacea: Isopoda: Cirolanidae) from the coral reefs of Viti Livu, Fiji, with a revised diagnosis to the genus. Marine Biology Research, 17(7–8):680–691, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.2015391.

Ceriello:2021:CBT

[1260] Hellen Ceriello, André R. Senna, Luiz F. Andrade, and Sérgio N. Stampar. Crustacea biodiversity in tubes of Ceriantharia (Cnidaria; Anthozoa), including the description of a novel species of Amphipoda from southeastern Brazil. Marine Biology Research, 17(7–8):692– 706, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2020293.

Jinno:2021:ISH

[1261] Masaki Jinno, Wataru Doi, Akira Mizutani, and Hiroyoshi Kohno. Influence of salinity on the horizontal distribution of soldier crab Mictyris guinotae in a subtropical river in Japan. Marine Biology Research, 17 (7–8):707–722, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2013502.

Colombo:2021:DTE

[1262] Julia Colombo, Ayelen Blandi, Pablo Lértora, and Martín Varisco. Differences in the timing of egg extrusion and reproductive investment between primiparous and multiparous females of the southern king crab Lithodes santolla. Marine Biology Research, 17(7–8):723–732, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.2020294.

Zhang:2021:AGC

[1263] Jun Zhang, Yan Wang, Zuozhi Chen, Yane Jiang, and Shannan Xu. Age and growth of Ceratoscopelus warmingii (Myctophidae) in the South China Sea based on sagittal otolith microstructure. Marine Biology Research, 17(7–8):733–743, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.2015390.

Albrecht:2021:LMS

[1264] Sofia Albrecht, Jean-Luc Jung, Laura Lazar, Marguerite Bathie, Delphine Lémery-Peissik, and Marianne Helene Rasmussen. Local movements and site-tenacity of minke whale in Iceland in relation to feeding and a northward distribution shift. Marine Biology Research, 17 (7–8):744–761, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2021.2016839.

Blasi:2021:CCS

[1265] Monica Francesca Blasi, Alice Rotini, Tiziano Bacci, Monica Targusi, Giusy Bonanno Ferraro, Luca Vecchioni, Rosa Alduina, and Luciana Migliore. On Caretta caretta’s shell: first spatial analysis of microand macro-epibionts on the Mediterranean loggerhead sea turtle carapace. Marine Biology Research, 17(7–8):762–774, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.2016840.

Billah:2021:IAC

[1266] Mohammad Maruf Billah and Md Saydur Rahman. Impacts of anthropogenic contaminants and elevated temperature on prevalence and proliferation of Escherichia coli in the wild-caught American oyster, Crassostrea virginica in the southern Gulf of Mexico coast. Marine Biology Research, 17(9–10):775–793, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2053161.

Hossain:2021:GAR

[1267] Md Nazmul Hossain, Tanjia Afrin Chowdhury, Mahmudul Hasan, Kazi Faizul Azim, Sayadatunnasa Jenny, Muhammad Fazle Rabbee, Md. Mahbubur Rahman, and Saneya Risa Somana. Genomic attributes and recent advances in detection of white spot syndrome virus in shrimp: a review. Marine Biology Research, 17(9–10):794–814, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2021.2023191.

Yildiz:2021:CIR

[1268] Taner Yildiz, Ugur Uzer, Emre Yemisken, F. Saadet Karakulak, Abdullah E. Kahraman, and Özgür Çanak. Conserve immatures and rebound the potential: stock status and reproduction of whiting (Merlangius merlangus [Linnaeus, 1758]) in the Western Black Sea. Marine Biology Research, 17(9–10):815–827, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2021.2019787.

Vagenas:2021:FDG

[1269] Georgios Vagenas, Charis Apostolidis, Aikaterini Dogrammatzi, and Paraskevi K. Karachle. Filling in data gaps on the phaeton dragonet [Synchiropus phaeton (Günther, 1861)]: a study on biological characteristics in the eastern Mediterranean (Aegean Sea, Greece). Marine Biology Research, 17(9–10):828–837, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2029900.

Villalba:2021:RCS

[1270] Ana Gloria Villalba Villalba, Lerma Hanaiy Chan Chan, Irlanda Lagarda Diaz, Naomi Yerith Reyes Jiménez, Christian Minjarez Osorio, Reina Castro Longoria, and Amir Maldonado Arce. Reproductive cycle of sea urchin Echinometra vanbrunti (Echinodermata: Echinoidea) from the Gulf of California. Marine Biology Research, 17(9– 10):838–852, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2029901.

Ganias:2021:FPS

[1271] Konstantinos Ganias and Irini-Fotini Kompogianni. Fecundity pattern and spawning dynamics of the common cuttlefish Sepia officinalis. Marine Biology Research, 17(9–10):853–861, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2022.2032176.

Funes-Rodriguez:2021:CHF

[1272] René Funes-Rodríguez, Rafael Cervantes-Duarte, María del Carmen Jiménez-Quiroz, Sergio Hernández-Trujillo, Rogelio González-Armas, Sofia A. Barón-Campis, Eduardo Anaya-Godinez, and Francisco Barrón-Barraza. Characterization of hydrological features and plankton during neap tides in a subtropical lagoon south of the California current. Marine Biology Research, 17(9–10):862–875, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2022.2036762.

Price:2021:SGH

[1273] Nathan William Price, Kao-Sung Chen, Chaolun Allen Chen, and Colin Kuo-Chang Wen. Scraping and grazing herbivorous/detritivorous fish display opposite feeding behaviours under different protection regulations. Marine Biology Research, 17(9–10):876–891, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2022.2037002.

Mamet:2021:SSI

[1274] Leandro Nicolás Getino Mamet, Gaspar Soria, Laura Schejter, and Federico Márquez. Shell shape as an indicator of phenotypic stocks of Tehuelche scallop in Northern Patagonia, Argentina. Marine Biology Research, 17(9–10):892–903, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2039715.

Athanasiadis:2021:CDM

[1275] Athanasios Athanasiadis. Canal differentiation in multiporate conceptacles of Mesophyllaceae (Corallinales, Rhodophyta), with recognition of Thallis gen. nov. from South Africa, and Perithallis gen. nov., Printziana gen. nov. and Sunesonia gen. nov. from southern Australia–New Zealand–Chatham. Marine Biology Research, 17(9–10): 904–945, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2043553.

Tsuyuki:2021:TNS

[1276] Aoi Tsuyuki, Yuki Oya, and Hiroshi Kajihara. Two new species of the marine flatworm Pericelis (Platyhelminthes: Polycladida) from southwestern Japan with an amendment of the generic diagnosis based on phylogenetic inference. Marine Biology Research, 17(9–10):946– 959, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2048669.

Nicolau:2021:HDW

[1277] Lídia Nicolau, Paulo Vasconcelos, André N. Carvalho, Fábio Pereira, David Piló, Laura Sordo, and Miguel B. Gaspar. Hand dredging for the wedge clam (Donax trunculus) in the Algarve coast (southern Portugal): fishing yield, bycatch, discards and damage rates. Marine Biology Research, 17(9–10):960–977, 2021. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2048670.

Flores:2021:RSY

[1278] Andrés Flores, Rodrigo Wiff, Donald I. Brown, Mauricio Ahumada, and Sebastián Larrea-Meza. Reproductive strategy of yellow squat lobster (Cervimunida johni Porter, 1903): re-evaluating the maturity criteria. Marine Biology Research, 17(9–10):978–993, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2051554.

Duan:2021:GOH

[1279] Xiaoshuo Duan, Yao Bian, Haolin Wang, and Hongwei Wang. Grateloupia oligoclora (Halymeniaceae, Rhodophyta): a new species from China based on morphological observation, life history and rbc L gene sequence analysis. Marine Biology Research, 17(9–10):994–1004, 2021. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2046279.

Mashirin:2022:MET

[1280] K. Reeha Mashirin and K. C. Chitra. Microplastics — an emerging threat in the Indian waterbodies. Marine Biology Research, 18 (1–2):1–12, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2096905.

vanGenne:2022:IAO

[1281] Barbara van Genne and Ricardo A. Scrosati. The interspecific abundance-occupancy relationship in rocky intertidal communities. Marine Biology Research, 18(1–2):13–18, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2022.2105895.

Gomes-Goncalves:2022:RBE

[1282] Rafaela de Sousa Gomes-Gonçalves and Francisco Gerson Araújo. Relationships between environmental heterogeneity and fish beta diversity in a tropical bay. Marine Biology Research, 18(1–2):19–31, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2063902.

Rodriguez-Rua:2022:FIP

[1283] Ana Rodríguez-Rúa, Marina Delgado, Luis Silva, Sara Román, and Miguel Cojan. First insights into population parameters and spatial distribution of the two morphotypes of Mactra stultorum (brownish and whitish) (Linnaeus, 1758) (Mollusca: Bivalvia) along the southwestern Spanish coast. Marine Biology Research, 18(1–2):32–47, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2064511.

Beziat:2022:UPC

[1284] Pauline Béziat and Andreas Kunzmann. Under pressure: Cassiopea andromeda jellyfish exposed to increasing water temperature or lead, cadmium and anthropogenic gadolinium contamination. Marine Biology Research, 18(1–2):48–63, 2022. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2066132.

Jonsdottir:2022:BCS

[1285] Ingibjörg G. Jónsdóttir, Chris Yesson, Kirsty M. Kemp, and Steinunn H. Ólafsdóttir. Benthic community structure on offshore northern shrimp (Pandalus borealis) grounds north of Iceland. Marine Biology Research, 18(1–2):64–78, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2080840.

Rubio-Polania:2022:MCD

[1286] J. Rubio-Polania, D. Torruco-Gómez, A. González-Solís, R. Rodríguez-Canul, M. Améndola-Pimenta, I. Mariño-Tapia, J. Herrera-Silveira, F. Árcega-Cabrera, V. Ceja-Moreno, J. Romo-Alvarado, I. Lizarraga-Castro, and C. Gutiérrez-Solís. Megabenthic communities distributed in the outer margins of the Yucatan shelf and their relationship to environmental variables. Marine Biology Research, 18(1–2):79–103, 2022. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2022.2086699.

Zagorodnyaya:2022:SIV

[1287] Yuliya A. Zagorodnyaya and Sergey A. Piontkovski. Seasonal and interannual variations of the abundance of the dinoflagellate Noctiluca scintillans in the northern Black Sea. Marine Biology Research, 18(1– 2):104–116, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2086701.

Amberkar:2022:STV

[1288] Ujwala Amberkar and Rakhee Khandeparker. Spatio-temporal variations in culturable bacterial community associated with denitrification in the Arabian Sea oxygen minimum zone. Marine Biology Research, 18(1– 2):117–130, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2086700.

Veronique:2022:UPI

[1289] Sarano Véronique, Sarano François, Girardet Justine, Preud’homme Axel, Vitry Hugues, Heuzey René, Sarano Marion, Delfour Fabienne, Glotin Hervé, Adam Olivier, Madon Bénédicte, and Jung Jean-Luc. Underwater photo-identification of sperm whales (Physeter macrocephalus) off Mauritius. Marine Biology Research, 18(1–2):131–146, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2040737.

Lahbib:2022:PPA

[1290] Youssef Lahbib, Paulo Vasconcelos, Sami Abidli, Carlos M. Barroso, Najoua Trigui El Menif, and Miguel B. Gaspar. Preferential prey and attacking strategy of the purple dye murex (Bolinus brandaris) on three common bivalve species from Tunisia (Central Mediterranean Sea). Marine Biology Research, 18(1–2):147–159, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2022.2056613.

Chen:2022:SCI

[1291] Shixuan Chen, Hao Xu, Shuo Zhang, and Wenwen Yu. Study of the characteristics of intestinal bacteria of little yellow croaker (Larimichthys polyactis) and big head croaker (Collichthys lucidus) from the Lyusi fishing ground. Marine Biology Research, 18(1–2):160–171, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2072893.

Morgan:2022:FIP

[1292] Alexia Morgan, Marina Mendes, Geoff Tingley, Susanna Wingard Brian, and Pedro Sousa. Have Fishery Improvement Projects (FIPs) adequately addressed key needs to monitor and mitigate fishery impacts on endangered, threatened, and protected (ETP) species? A review, with recommendations to the supply chain. Marine Biology Research, 18(3– 4):173–186, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2096903.

Olivares-Banuelos:2022:TRP

[1293] T. N. Olivares-Bañuelos, C. M. Gómez-Gutiérrez, L. L. García-Echauri, E. D. Moreno-Sánchez, and F. Correa-Sandoval. Temperature and reduced pH regulate stress and biomineralization gene expression in larvae and post-larvae of the sand dollar Dendraster excentricus. Marine Biology Research, 18(3–4):187–202, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2022.2105894.

mahjoubGallouli:2022:SVP

[1294] El Mahjoub Gallouli, Hassan Oulad Ali, Mohamed El Hafa, and Ahmed Aamiri. Spatiotemporal variability of phytoplankton in the Aglou-Sidi Ifni coastline (Moroccan Atlantic Coast). Marine Biology Research, 18 (3–4):203–214, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2086702.

Kasanah:2022:FRC

[1295] Noer Kasanah, Drajad S. Seto, Maria Ulfah, Zhiyuan Peng, and Triyanto. First report on chemistry of a red seaweed Croisettea sp. from the coastal area of Yogyakarta, Indonesia. Marine Biology Research, 18(3– 4):215–221, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2096904.

Bhowal:2022:FRE

[1296] Aishee Bhowal, Alfisa Siddique, Jasmine Purushothaman, Sanu V. Francis, and Chelladurai Raghunathan. First report of an epibiontic diatom Protoraphis hustedtiana R. Simonsen, 1970 (Protoraphidaceae) in association with copepod Pontella andersoni Sewell, 1912 from northern coastal waters of Bay of Bengal, Indian Ocean. Marine Biology Research, 18(3–4):222–229, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2119250.

Villegas-Hernandez:2022:LHT

[1297] H. Villegas-Hernández, E. Novelo-Ríos, K. Fuentes-Dantorie, S. Guillén-Hernández, C. González-Salas, D. Pech-Puch, R. Díaz-Gamboa, and G. Poot-López. Life-history traits facilitate the population success of the sergeant major Abudefduf saxatilis (Linnaeus, 1758, Perciformes: Pomacentridae) in the Caribbean Sea. Marine Biology Research, 18(3– 4):230–251, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2117827.

AlvesPereira:2022:IPF

[1298] Juan Alves Pereira, Alexandre da Gama Fernandes Vieira Júnior, Éden Hávila Lima Guedes, Bianca de Freitas Terra, and André Luiz Machado Pessanha. Importance of pneumatophore fringe-mudflat continuum for an estuary-dependent fish species. Marine Biology Research, 18(3–4): 252–265, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2076882.

Gonzalez-Suarez:2022:EDS

[1299] Daniel González-Suárez and José Carlos Hernández. Early developmental stages of the sea urchin Sphaerechinus granularis (Lamarck, 1816) (Echinoidea: Toxopneustidae). Marine Biology Research, 18(3– 4):266–277, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2093375.

SriHari:2022:DGD

[1300] Murugesan SriHari, Annam Pavan-Kumar, Barry C. Russell, Daniel Golani, Zeba Jaffer Abidi, and Loveson Edward. Deep genetic divergence in Randall’s threadfin bream Nemipterus randalli Russell, 1986 population from northern Bay of Bengal inferred by mtDNA sequences. Marine Biology Research, 18(3–4):278–282, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2022.2096902.

Penas-Torramilans:2022:RSC

[1301] Raquel Peñas-Torramilans, Amanda P. Horch, and Mariana Terossi. Reproduction of the shore crab Pachygrapsus transversus (Gibbes, 1850) (Decapoda: Grapsidae) in the Brazilian temperate zone. Marine Biology Research, 18(3–4):283–291, 2022. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2101123.

Kumar:2022:IES

[1302] K. V. Aneesh Kumar, M. P. Rajeeshkumar, M. Hashim, K. P. Deepa, J. L. Otero-Ferrer, N. Saravanane, Sujitha Thomas, K. K. Bineesh, and V. M. Tuset. Inferring ecological strategies of Psenopsis spp. (Teleostei: Centrolophidae) inhabiting Indian waters from morphological features. Marine Biology Research, 18(3–4):292–305, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2101124.

Sathyamoorthy:2022:GBP

[1303] Gayathri Sathyamoorthy and Thiruganasambandam Rajendran. Growth and biochemical profiling of marine microalgae Chlorella salina with response to nitrogen starvation. Marine Biology Research, 18(5–6): 307–314, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2131823.

Karadurmus:2022:RBR

[1304] Ugur Karadurmus. Recolonization behavior of rapa whelk (Rapana venosa) and common starfish (Asterias rubens) eradicated from mussel bed. Marine Biology Research, 18(5–6):315–325, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2120202. See

letter [1359].

Polo:2022:SCS

[1305] Christian Polo, Bernabé Moreno, Yessenia Arroyo, and Báslavi Cóndor-Luján. On some cryptic sponges associated with Lessonia trabeculata holdfasts in the South-eastern Pacific. Marine Biology Research, 18(5– 6):326–334, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2123521.

Eryasar:2022:GND

[1306] Ahmet Raif Eryasar. Grid-net design that successfully reduces discarded catch and damage to benthic species in the veined rapa whelk beam trawl fishery. Marine Biology Research, 18(5–6):335–346, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2123523.

Seyhan:2022:CCN

[1307] Kadir Seyhan, Umar Khan, Ahmet Sahin, and Yahya Terzi. Catch composition of non-target species escaping dredge gear of Rapana fisheries in the western coast of the Black Sea. Marine Biology Research, 18(5– 6):347–360, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2126857.

Alhajji:2022:MRA

[1308] Adnan Hussain Alhajji, Hua Hsun Hsu, Yousef Ahmed Alkhamis, Ahmed Saud Alsaqufi, Syed Ajmal Khan, and Zahid Nazeer. Maturity and reproduction in the Arabian carpet shark, Chiloscyllium arabicum from the Saudi Arabian waters of the Arabian Gulf. Marine Biology Research, 18(5–6):361–371, 2022. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2131824.

Moreno-Santos:2022:UFS

[1309] Rodrigo Augusto Moreno-Santos, Mario Vinicius Condini, Lorena Lopes Almeida, Marcelo Soeth, Áthila Andrade Bertoncini, and Mauricio Hostim-Silva. Unveiling the first steps of juvenile Atlantic goliath grouper Epinephelus itajara (Perciformes: Epinephelidae) in a tropical estuary at the southwestern Atlantic coast. Marine Biology Research, 18(5– 6):372–379, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2119251.

Hernaez:2022:IBS

[1310] Patricio Hernáez, Marcelo Antonio Amaro Pinheiro, Flávio Almeida Alves-Junior, and William Santana. Intertidal burrowing shrimps (Axiidea: Callianassidae, Callichiridae; Gebiidea: Axiannassidae, Upogebiidae) collected along the Brazilian coast. Marine Biology Research, 18 (5–6):380–408, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2105893.

Magalhaes:2022:RTP

[1311] Wagner F. Magalhães, Verônica Maria de Oliveira, and Julie H. Bailey-Brock. Redescription of two Phyllodoce (Annelida: Phyllodocidae) species from Hawaii. Marine Biology Research, 18(5–6): 409–414, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2119252.

Bjordal:2022:DJA

[1312] Vilde R. Bjørdal, Herdis L. Mørk, Kjell Rong Utne, Anders Fernö, and Leif Nøttestad. The diet of juvenile Atlantic mackerel (Scomber scombrus) feeding in new northern nursery areas along the Norwegian coast. Marine Biology Research, 18(7–8):415–425, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2147951.

Springer:2022:SCF

[1313] Karin Springer, Beatrice Brix da Costa, Sam Samsuardi, and Andreas Kunzmann. Simulating cyanide fishing: photosynthetic effects of shortterm cyanide exposure in three different hermatypic coral species. Marine Biology Research, 18(7–8):426–434, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2022.2147947.

Wang:2022:SDF

[1314] Rongfu Wang, Masayuki Chimura, Kazuhiro Sadayasu, Mitsuhiro Ishino, and Orio Yamamura. Summer diet and feeding strategy of Pacific cod (Gadus macrocephalus) inhabiting the southern coast of Hokkaido, Japan. Marine Biology Research, 18(7–8):435–447, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2022.2147948.

Garcia-Cardenas:2022:GRT

[1315] Francisco J. García-Cárdenas and Pablo J. López-González. Growth and reproductive traits of deep-sea pen Anthoptilum murrayi Kölliker, 1880, from Iceland (North Atlantic). Marine Biology Research, 18(7– 8):448–465, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2147949.

Herrera:2022:DPT

[1316] Daphine Ramiro Herrera and Rogerio Caetano da Costa. Distributional patterns of two sympatric blue crabs (Callinectes) and the implications for conservation management at the South-west Atlantic subtropical shelf. Marine Biology Research, 18(7–8):466–476, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2022.2147950.

Rubio-Polania:2022:CSM

[1317] J. C. Rubio-Polania, A. González-Solis, C. Enriquez, F. Árcega-Cabrera, V. Ceja-Moreno, L. Cetina-Euan, R. Poot-Aguayo, F. Nuñez-Chan, and D. Torruco. Community structure of megabenthos of Perdido Fold Belt (Tamaulipas, Mexico) and its relationship with the oceanographic and sediment parameters including potential pollutants. Marine Biology Research, 18(7–8):477–494, 2022. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2022.2137198.

Montalbetti:2022:STM

[1318] Enrico Montalbetti, Valerio Isa, Sara Vencato, Yohan Louis, Simone Montano, Silvia Lavorano, Davide Maggioni, Paolo Galli, and Davide Seveso. Short-term microplastic exposure triggers cellular damage through oxidative stress in the soft coral Coelogorgia palmosa. Marine Biology Research, 18(7–8):495–508, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2022.2137199.

Gao:2023:EEP

[1319] Letian Gao, Zhongjian Huang, Zhihai Zhong, Xiaomeng Shi, Xing Chang, Yuhong Sun, Nuoyan Zhang, Ning Zou, and Shanying Tong. Effects of elevated pCO_{2} on the response of coccolithophore Emiliania huxleyi to prolonged darkness. Marine Biology Research, 18(9– 10):509–519, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2169463.

Alina:2023:ELT

[1320] Dining Nika Alina, Rita Rachmawati, Neviaty Putri Zamani, and Hawis Madduppa. Exposure at low tide leads to a different microbial abundance of intertidal coral Acropora pulchra. Marine Biology Research, 18(9– 10):520–530, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2169464.

Bag:2023:DSP

[1321] Souvik Bag, Basanta Sarkar, Madhurima Seal, Arnab Chatterjee, Ayan Mondal, and Soumendranath Chatterjee. Diversity and seasonal prevalence of starch hydrolysing, phosphate solubilizing and nitrogen-fixing bacterial groups of rooted and un-rooted regions of tropical mangrove sediments of Sundarbans, West Bengal, India. Marine Biology Research, 18(9–10):531–543, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-

1019 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/17451000.2023.2174263.

Tkachenko:2022:PSS

[1322] Konstantin S. Tkachenko, Vu Viet Dung, and Vo Thi Ha. Phase shift from a stony-coral to a soft-coral community on a coral reef: a case study of an alternative state. Marine Biology Research, 18(9– 10):544–554, 2022. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2022.2153869.

Zou:2023:BSC

[1323] Xiaoxiao Zou, Tian Meng, Dandan Yao, Zuo Chen, Jun Zhu, Dan Mu, and Shixiang Bao. Benthic Sargassum composition and community characteristics in the intertidal zone of Hainan Island, China. Marine Biology Research, 18(9–10):555–565, 2023. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2023.2174262.

Micael:2023:RSN

[1324] Joana Micael, Alfonso A. Ramos-Esplá, Pedro Rodrigues, and Sindri Gíslason. Recent spread of non-indigenous ascidians (Chordata: Tunicata) in Icelandic harbours. Marine Biology Research, 18(9–10): 566–576, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2176882.

Hayeewachi:2023:BAF

[1325] Lutfee Hayeewachi, Eknarin Rodcharoen, and Kringpaka Wangkulangkul. Bioengineering by the alien false mussel Mytilopsis sallei (Bivalvia: Dreissenidae) affects small-scale distribution pattern of benthic macroinvertebrates in an estuarine system. Marine Biology Research, 18 (9–10):577–588, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2169465.

Pescinelli:2023:FGL

[1326] Régis Augusto Pescinelli and Rogerio Caetano Costa. Filling gaps in the life history of the snapping shrimp Alpheus petronioi: morphology of the first post-embryonic stages reared in laboratory. Marine Biology Research, 18(9–10):589–598, 2023. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2023.2173782.

Wang:2023:CPE

[1327] Ming-Huang Wang, Chih-Feng Chen, Frank Paolo Jay B. Albarico, Wen-Pei Tsai, Chiu-Wen Chen, and Cheng-Di Dong. Concentrations of phthalate esters on Indian Ocean silky sharks and their long-term dietary consumption risks. Marine Biology Research, 18(9–10):599–610, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2173783.

Bugota:2023:LGD

[1328] Valeli J. Bugota and Cyrus Rumisha. Low genetic diversity and lack of genetic structure among populations of the sandfish Holothuria (Metriatyla) scabra on the Tanzanian coast. Marine Biology Research, 18 (9–10):611–619, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2176883.

Grobler:2023:PPM

[1329] D. L. Grobler, J. D. Klein, J. M. Quattro, N. Bolaño-Martínez, R. H. Bennett, and A. E. Bester van der Merwe. Phylogenetic placement and molecular dating of hammerhead sharks (Sphyrnidae) based on whole mitogenomes. Marine Biology Research, 19(1):1–12, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2023.2194060.

Eskinat:2023:ARC

[1330] Deniz Eskinat, Ali Cemal Gücü, and Meltem Ok. Annual reproductive cycle of a successful Lessepsian immigrant in the Eastern Mediterranean Sea, Pomadasys stridens (Forsskål, 1775) (family: Haemulidae). Marine Biology Research, 19(1):13–24, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2023.2198243.

Jose:2023:RAP

[1331] Sheena Jose, S. Bharathi, A. T. Damodhar, K. V. Jayachandran, T. T. Ajith Kumar, and Kuldeep Kumar Lal. Reef-associated palaemonid shrimp fauna of Lakshadweep islands: special focus on Cuapetes Clark, 1919 (Decapoda; Caridea) with a description of new species and phylogenetic notes. Marine Biology Research, 19(1):25–47, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2198241.

Bahtiar:2023:RBM

[1332] Bahtiar, Muhammad Fajar Purnama, Maruf Kasim, and Ermayanti Ishak. Reproductive biology of mangrove clams Geloina expansa (Mousson, 1849) from mangrove at Kendari Bay, Southeast Sulawesi, Indonesia. Marine Biology Research, 19(1):48–58, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2023.2185639.

Andrade:2023:FRG

[1333] Luiz F. Andrade, André R. Senna, and Jesser F. Souza-Filho. First record of the genus Harpinia Boeck, 1876 (AMPHIPODA:3 Phoxocephalidae: Harpiniinae) from the southern hemisphere, with description of three new species. Marine Biology Research, 19(1):59–80, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2023.2194059.

Efimova:2023:EWT

[1334] Tatiana Efimova, Tatiana Churilova, Elena Skorokhod, Victoria Buchelnikova, Anatoly Buchelnikov, Nataliia Moiseeva, Pavel Salyuk, Igor Stepochkin, and Aleksandr Melnik. Effect of water trophic status on bio-optical properties and productive characteristics of phytoplankton of the Black Sea coastal waters near Sevastopol. Marine Biology Research, 19(2–3):81–93, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2203502.

AlBusaidi:2023:SBC

[1335] Saud AlBusaidi and Khalid Al-Hashmi. Seasonal biomass and composition of mesozooplankton communities in Sea of Oman and Arabian Sea. Marine Biology Research, 19(2–3):94–107, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2023.2206138.

Jacob:2023:SEH

[1336] Hannah Jacob, Karin Springer, Laura Belter, and Andreas Kunzmann. Symbiodiniaceae in and ex hospite have differential physiological responses under different heat stress scenarios. Marine Biology Research, 19(2–3):108–120, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2198240.

Dharmadhas:2023:PCB

[1337] Jeba Sweetly Dharmadhas, Poornima Arumugam, and Rajiv Periakaruppan. Preparation and characterization of biosorbent Sargassum myriocystum for zinc removal. Marine Biology Research, 19(2– 3):121–131, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2193897.

Delgado-Serrano:2023:AVL

[1338] Celia Delgado-Serrano and Fernando Tuya. Annual variation in leaf canopy across a seagrass distribution range is modulated by local environmental variation. Marine Biology Research, 19(2–3):132–140, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2193898.

Pinheiro:2023:APM

[1339] Marcelo A. A. Pinheiro, Fernanda V. B. Souza, Julia F. Perroca, Mauro M. T. Silva, Raimundo L. M. Sousa, Thais A. Mota, and Sérgio S. Rocha. Advances in population monitoring of the mangrove ‘uçá’-crab (Ucides cordatus): reduction of body size variance for better evaluation of population structure and extractive potential. Marine Biology Research, 19 (2–3):141–153, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2205148.

Volonterio:2023:SBA

[1340] Odile Volonterio and Rodrigo Ponce de León. Systematic and biogeographical aspects of the first Antarctic discovery of Orthoplana bregazzii Karling, 1973 (Proseriata, Otoplanidae). Marine Biology Research, 19 (2–3):154–164, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2205149.

Baran:2023:GSS

[1341] Christine C. Baran, Rhea Mae A. Luciano, Christine S. Segumalian, Darryl Anthony M. Valino, and Maria Vanessa Baria-Rodriguez. Genus and size-specific susceptibility of soft corals to 2020 bleaching event in the Philippines. Marine Biology Research, 19(2–3):165–176, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2198242.

Hassoun:2023:NRO

[1342] Mustapha Hassoun, Hanaa Moussa, Ghizlane Salhi, Loubna Benamar, Naouar Ben Ali, and Mohamed Kazzaz. New records and observations of marine algae from Morocco (Eastern Atlantic Ocean). Marine Biology Research, 19(2–3):177–190, 2023. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2023.2206139.

Buhl-Mortensen:2023:SPB

[1343] Lene Buhl-Mortensen, Trude Hauge Thangstad, Guldborg Søvik, and Henning Wehde. Sea pens and bamboo corals in Skagerrak and the Norwegian trench. Marine Biology Research, 19(2–3):191–206, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2224967.

Fazelian:2023:TIB

[1344] Nasrin Fazelian, Morteza Yousefzadi, and Ali Movafeghi. Toxicity of ironbased nanoparticles to Nannochloropsis oculata: effects of Fe_{2}O_{3}-NPs on oxidative stress and fatty acid composition. Marine Biology Research, 19 (2–3):207–217, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2214400.

Akrong:2023:VSA

[1345] M. O. Akrong, A. K. Anning, G. N. D. Addico, J. N. Hogarh, A. Adu-Gyamfi, K. A. A. deGraft Johnson, M. Ale, J. A. Ampofo, and A. S. Meyer. Variations in seaweed-associated and planktonic bacterial communities along the coast of Ghana. Marine Biology Research, 19(4– 5):219–233, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2213894.

Mohamed:2023:DII

[1346] Samia Ben Mohamed, Marwa Mejri, Abdellah Chalh, Adel A. Basyouny Shahin, Jean-Pierre Quignard, Monia Trabelsi, and Abderraouf Ben Faleh. Distinct inter and intrapopulation variation in the otolith shape and size of Mullus barbatus (Actinopterygii: Mullidae) from the Bizerte and Ghar El Melh lagoons in Tunisian waters. Marine Biology Research, 19(4–5):234–248, 2023. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2023.2203503.

Fang:2023:EIS

[1347] Jia-Feng Fang and Qi Li. The effects of inbreeding on stress resistance of the Pacific oyster Crassostrea gigas at different temperatures and salinities. Marine Biology Research, 19(4–5):249–260, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2224025.

Mgeleka:2023:PGH

[1348] Said S. Mgeleka, Mathew Ogalo Silas, Cretus Mtonga, Cyrus Rumisha, Elina Viinamäki, Patrick Polte, Mattias Sköld, Monika Winder, and Martin Gullström. Population genetics of the hound needlefish Tylosurus crocodilus (Belonidae) indicate high connectivity in Tanzanian coastal waters. Marine Biology Research, 19(4–5):261–270, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2224026.

Nunez-Gonzalez:2023:ADE

[1349] Raibel Núñez-González and Sheila M. Pauls. Abundance and distribution of Echinometra lucunter (Echinodermata, Echinoidea) at a rocky platform on the central coast of Venezuela. Marine Biology Research, 19 (4–5):271–278, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2224027.

Das:2023:RCA

[1350] Yean Das, Krishan D. Karunarathne, Mishal Roy, Mohammed Shah Nawaz Chowdhury, and SM Sharifuzzaman. Record of Crambionella annandalei Rao, 1931 from Bangladesh, with a review of the geographic distribution of the genus Crambionella (Cnidaria: Scyphozoa). Marine Biology Research, 19(4–5):279–293, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2203500.

Ceriello:2023:CBC

[1351] Hellen Ceriello and Sérgio N. Stampar. Ceriantheomorphe brasiliensis (Cnidaria; Ceriantharia): how does it behave? Marine Biology Research, 19(4–5):294–304, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2203504.

Bascinar:2023:LWR

[1352] Nimet Selda Başçinar, Devrim Selim Misir, Cemil Altuntaş, Yaşar Genç, Murat Dağtekin, Murat Erbay, Gülsüm Balçik Misir, and Erdinç Aydin. Length–weight relationship and seasonal variations in diet composition of whiting (Merlangius merlangus) in the South-eastern Black Sea. Marine Biology Research, 19(4–5):305–316, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2023.2203499.

Zagorodnyaya:2023:PEB

[1353] Yuliya A. Zagorodnyaya, Sergey A. Piontkovski, and Vladimir V. Gubanov. The pelagic ecosystem of the Black Sea goes gelatinous. Marine Biology Research, 19(6–7):317–326, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2023.2235571.

Motro:2023:BAP

[1354] Uzi Motro, Noa Simon-Blecher, Omri Bronstein, Suembikya Frumin, and Yair Achituv. Brewed in the African pot: the phylogeography of the toothed barnacle Chthamalus dentatus (Chthamaloidea: Chthamalidae). Marine Biology Research, 19(6–7):327–341, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2193899. See

correction [1360].

Paul:2023:ETC

[1355] Sourav Paul, Samya Karan, and Bhaskar Deb Bhattacharya. Effects of tropical cyclone Amphan on the copepods of the Ganges estuary. Marine Biology Research, 19(6–7):342–354, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2023.2235591.

Fernandes:2023:PEI

[1356] Abigail Fernandes, Avelyno D’Costa, Akhil Nair, Santosh Jathar, Anoop Kumar Yadav, Pamela Jha, Vinothkannan Ravichandran, and Renitta Jobby. Protective effects of indigenous lactic acid bacteria in Artemia salina challenged with Vibrio parahaemolyticus: an in vitro and in silico approach. Marine Biology Research, 19(6–7):355–365, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2235590.

Kumar:2023:PVS

[1357] K. V. Aneesh Kumar, V. P. Padate, S. S. Cubelio, and N. Saravanae. Population variation and sexual dimorphism in the deep-sea crab Parilia alcocki (Brachyura, Leucosiidae). Marine Biology Research, 19(6– 7):366–380, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2246473.

Chunga-Llauce:2023:LDS

[1358] Juan A. Chunga-Llauce, Liliana Tapia-Ugaz, and Aldo S. Pacheco. Latitudinal distribution of skeleton shrimps (Amphipoda, Caprellidae) in artificial habitats throughout the eastern South Pacific coast of Peru. Marine Biology Research, 19(6–7):381–394, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17451000.2023.2246488.

Cinar:2023:LEM

[1359] Melih Ertan Çinar and Baki Yokes. Letter to editor: Misidentification of species in the experimental study carried out in the Sea of Marmara by Karadurmus (Marine Biology Research, 2022:18, 315–325). Marine Biology Research, 19(6–7):395–396, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2023.2235579. See [1304].

Anonymous:2023:C

[1360] Anonymous. Correction. Marine Biology Research, 19(6–7):397, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2259770. See [1354].

Bordbar:2023:FET

[1361] Leila Bordbar, Evangelia Strogyloudi, Ioannis Hatzianestis, Vasiliki Paraskevopoulou, and Archontia Chatzispyrou. First evidence of trace metals and persistent organic contaminants from an endangered marine species, Mobula mobular (Bonattere, 1788) caught in Hellenic waters (Saronikos Gulf). Marine Biology Research, 19(8–9):399–406, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2256343.

Jeong:2023:MPB

[1362] Dong Hyuk Jeong, Hyeon Been Lee, Aaron A. Heiss, Byung Cheol Cho, and Jong Soo Park. Morphological, phylogenetic and biogeographic characterizations of three heterotrophic nanoflagellates isolated from coastal areas of Korea. Marine Biology Research, 19(8–9):407–418, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2271939.

Gil:2023:SCH

[1363] Marta Gil and Fran Ramil. Sertularioidea (Cnidaria, Hydrozoa) from the Guinea Current Large Marine Ecosystem (GCLME). Marine Biology Research, 19(8–9):419–437, 2023. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2023.2253823.

Ganias:2023:ICM

[1364] Kostas Ganias, Alexandra Karatza, and Dimitrios Lachouvaris. Investigating the capture mechanism of scavenging gastropods in hanging-net fisheries. Marine Biology Research, 19(8–9):438–446, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2256345.

Rezende-Gois:2023:MAM

[1365] Gabriel Vieira Martins Rezende-Gois, Caio Santos Nogueira, Isabela Ribeiro Rocha Moraes, Rafael Carvalho Santos, and Rogerio Caetano Costa. Morphometric analyses of Mithraculus forceps (Brachyura:

Mithracidae): linear and geometric approaches in a no-take marine reserve population. Marine Biology Research, 19(8–9):447–459, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2262482.

Colombo:2023:APJ

[1366] Julia Colombo, Vinuesa Julio, and Martín Varisco. Aggregation patterns in juvenile southern king crab Lithodes santolla (Molina, 1782). are the aggregations related to moulting activity? Marine Biology Research, 19 (8–9):460–465, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2262485.

Castro-Gutixerrez:2023:AGF

[1367] J. Castro-Gutiérrez, A. García-González, C. Rodríguez-García, Á. R. Domínguez-Bustos, and R. Cabrera-Castro. Age, growth, and feeding of boarfish, Capros aper (Linnaeus, 1758) in the Southwest of the Iberian Peninsula. Marine Biology Research, 19(8–9):466–479, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2267566.

Boonsanit:2023:ESP

[1368] Phurich Boonsanit and Supanut Pairohakul. Effects of salinity on physiological responses in the sesarmid crab Episesarma mederi (H. Milne Edwards, 1853). Marine Biology Research, 19(8–9):480–490, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2282425.

Gu:2023:DPR

[1369] Lide Gu, Wanli Yan, Xinli Yue, Haowen Zhong, and Deli Wang. Distribution and potential roles of microbial protoporphyrin IX in marine sediments. Marine Biology Research, 19(8–9):491–506, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2282429.

Warneke-Cremer:2023:CSN

[1370] Cornelia Warneke-Cremer and Juergen Guerrero-Kommritz. Cephalopods sampled near Azores–Biscay rise and along the mid-Atlantic Ridge north of the Azores by the FRV Walther Herwig II in 1982. Marine Biology Research, 19(8–9):507–533, 2023. CODEN ???? ISSN 1745-1000 (print),

1745-1019 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17451000.2023.2282435.

Lxfcskow:2023:STP

[1371] Florian Lüskow, Gudjon Mar Sigurdsson, Jörundur Svavarsson, and Astthor Gislason. Spatial and temporal patterns of scyphozoan jellyfish abundance and growth in Icelandic coastal waters — a climate change perspective. Marine Biology Research, 19(10):535–548, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2291569.

Kennedy:2023:KCM

[1372] Madison C. Kennedy, Martin V. Sørensen, Nuria Sánchez, and Stephen C. Landers. Kinorhynch communities of Mobile Bay, Alabama. Marine Biology Research, 19(10):549–555, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/17451000.2023.2291572.

Chen:2023:DAN

[1373] Chong Chen, Noriyuki Isobe, and Hidetaka Nomaki. A deep abyssal natural wood fall in the Northwestern Pacific and its associated fauna. Marine Biology Research, 19(10):556–563, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17451000.2023.2291579.

Vagenas:2023:BAR

[1374] Georgios Vagenas, Aikaterini Dogrammatzi, Gerasimos Kondylatos, and Paraskevi K. Karachle. On the biology of the alien Red Sea goatfish, Parupeneus forsskali (Fourmanoir and Guézé, 1976) in the Aegean Sea, eastern Mediterranean. Marine Biology Research, 19(10): 564–573, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
17451000.2023.2299978.

Feyzioglu:2023:FCI

[1375] Ali Muzaffer Feyzioglu, Yahya Terzi, Rafet Çagri Öztürk, Ersan Basar, Ilknur Yildiz, and Ertugrul Agirbas. Flow cytometric investigation and comparison of Synechococcus spp. pico-, nano- and microplankton in the Arctic and Antarctic polar regions during the summer period of 2019. Marine Biology Research, 19(10):574–581, 2023. CODEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17451000.2023.2299995.

Karadurmus:2023:IIV

[1376] Ugur Karadurmus, Dilek Ustaoglu, and Mehmet Aydin. Intra- and interspecific variability in external morphology among three Symphodus species (Labridae). Marine Biology Research, 19(10):582–591, 2023. CO-DEN ???? ISSN 1745-1000 (print), 1745-1019 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17451000.2023.2299996.